A Complete Bibliography of Publications in Annales de l’Institut Henri Poincaré (B)

Probability and Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

06 April 2022 Version 2.12

Title word cross-reference

(0 < p ≤ 1) [Lon81]. (1, 2) [NRS10]. (1 − ∥t∥_{2n+1})_{+}^{n+1} [LR86]. (α, d, β) [MV07]. (Δ1) [Mil83]. (R^{+})^{d} [Pei92]. (n + 1) [AP05]. (N, d) [Imk84]. {−1, 1}^{n} [Lar93]. 1 [FGS18, HMNP98]. 2 [AKSS07, Dey20, Fou01, HMNP98, PZ21, TW18, Yao18]. 3 [FP18a]. 4 + 1 [vdHS03, vdHdHS07]. 6 − 8 log 2 [CH12]. ^{3} [Shi10]. ^{4} [Shi10]. ^{d} [Kur10]. _{κ} [BLM21]. α [Hey11, Zha13]. β [DV13, TV98]. C(0, 1)Z^{d} [RZ93, RZ93]. C(R_{+}^{∗}; R) [Yor75]. C^{∞} [BN90]. c_{0} [Sch83]. D(R_{+}^{∗}; R) [Yor75]. D[0, 1] [BF92]. CAR(p) [SG02]. d [ČTW11, Enr96, Val12]. D = 1 [BDP95]. δ [Pov98]. ℓ_{p} [Sod08]. ℓ_{p}^{n} [NR03]. ϵ [GBR73]. F [Zhe12]. G [BP85, Son13, PX05]. γ [Viv21]. γ ∈ (0, 2) [GM21]. H [GL89]. H^{p} [Lon81]. H_{p} [KBR71]. H ∈ (0, \frac{1}{2}) [CN05]. K [CG68, DL96, BHS12]. κ [BLM21]. L^{1} [Bru73]. L^{2} [FHL12, GKMMR20, JS20, PX05, Sch83, CCS92]. L^{p} [BJ20, BCK66, Tal89].

L^{p} (1 ≤ p < ∞) [Hac82]. L^{p}, 0 ≤ p < 1 [Sch72b]. L_{E}^{p} [Ass87]. L_{0}^{∞} [VW04]. L_{1} [AJ76b, AJ76a, BHL93, Émi85, Kip74]. L_{∞} [AJ76b, Ast76, Émi81a]. L_{p} [CH90, Kol92, Mat93, TV98]. l_{p}(B)(1 ≤ p < ∞) [Led83]. L_{p} (1 < p < ∞) [Émi81b]. Λ [BEV13, BBL14, LZ15, JS20]. [0, ∞[^{s} [Saa88]. L log L [Led81]. M [BT08, Fer15, Gue99, CLP14, GNRV05, Mat98c]. F_{q}[X] [MN83]. GL_{N} [CK22]. R [LR14, Kur10, Rou80]. R^{∗+} [LP97]. R^{2} [CX21]. R^{d} [BBC78, LS91, LP97, Leg89, Pei92, Roy74, Leg89, Pei92]. R_{+}^{d} [All78]. R^{N} [Sch85, Lar93, Mic92b, PPH75, SP76]. R^{n}(n ≥ 3) [Yor85]. R_{+} [Leg89]. R_{+}^{m} × R^{n} [CGB17]. Z [DV21, Shi10, Shi10]. Z/qZ [NN20]. Z^{2} [GZ98, GM20]. Z^{d} [MST20, Mou11a, PS14, vdBvE22]. Z × Γ [KSS21]. R^{k} [Ber70]. L [HK09]. μ [Rou72, Tor78b]. N [CGL21, AP05, Gra84]. ∇ϕ [CD12]. o(n) [EN20]. Sine_{β} [HV17]. P [Jac71b, Alt87, Bas82, Ben79, DS96, KL05a, Mus06, Tal89]. q [TV81]. R [BNNS18, CGR79, GL89]. R^{1} [Ros10]. R^{2} [Doz83]. R^{d} [Cun03]. R^{N} [Sch89, LM83, LLO06]. R^{N} ∪ {∞} = S_{N} [Sch89, Sch89]. R^{r} [Smy75]. ρ [Sar21]. s [BK95]. S^{p} [Mac87]. S_{N} [Sch85]. σ [Cal81, Loo77a, Yan13, vW83]. SL(2, Z) [DK79]. SL(d, R)/SL(d, Z) [Le 02]. Sp(4) [Ras11a]. T [FSP71, BW01]. τ [Ano77, Tor76]. Ã_{r} [Sch09]. U [DDP87, Eic01, ES02, Har05, PSV17, dIS02]. V [Her05, PSV17, TV81]. Ã_{1} [Cha12]. X [Tor78b]. X_{n} = A_{n}X_{n−1} + B_{n} [BBD12]. X log X [Aïd20]. Z [Bré04, Loo77b]. Z^{2} [CRR98, MR94a]. Z^{4} [MR97]. Z^{d} [Bla02, Lan91]. Z^{n} [Dep65]. Z_{2} [Lem88].

-adic [Mus06]. -adiques [Mus06]. -algebras [vW83]. -algèbre [Loo77a]. -ary [CLP14]. -Banach [Bas82]. -block [HMNP98]. -bornés [TV81]. -circulants [BHS12]. -coalescents [BBL14]. -colorings [FP18a]. -conjecture [Har05]. -continue [Rou72]. -convergence [BP85]. -dependence [Mat98c]. -dependent [BCHJ21, HMNP98]. -diagonals [BNNS18]. -dimensional [AP05, Enr96, Fou01, TW18, Val12]. -dimensionnels [TV81]. -empirical [ES02]. -entropy [GBR73]. -estimation [Sar21]. -estimations [Fer15]. -estimators [BT08]. -expectation [Son13]. -extensions [Gue99, Lem88]. -finies [Cal81]. -finite [Yan13]. -geometrically [Her05]. -géométriquement [Her05]. -harmonic [HK09]. -hereditary [Sch83]. -level [AP05]. -Lyapounov [Pov98]. -matrices [Zhe12]. -order [GNRV05]. -player [CGL21]. -positivity [JS20]. -processes [Eic01]. -processus [Jac71b]. -regular [ČTW11]. -régularité [Ano77, Tor76]. -spaces [Kol92]. -spectral [FHL12]. -spin [KL05a]. -stable [Hey11, Tal89, Viv21, Zha13]. -statistic [DDP87]. -statistics [PSV17, dIS02]. -supermartingale [PX05]. -superprocess [MV07]. -systèmes [DL96]. -systems [DL96]. -transformations [BW01]. -valued [CCS92]. -variations [GL89]. -Wiener [Imk84].

1 [CG68, Coh68, Coh72, Fli81].

2-point [GM20]. 2d [MP19].

33 [Mat98b]. 39 [Bal04, She04]. 3D [Fla02].

41 [BW06]. 465 [Mat98b].

685 [Bal04].

876 [She04].

être [dlRd03].

A. [Ano77, Rou79, Tor76, Yor93]. Aalen [Lem16]. ABC [Mis16]. Abel [GRVY99]. abelian [Fog75, CL13b, DS07]. abéliens [Tor69]. aberrantes [Gar78, Gar79]. above [MM18, vdHS03, vdHdHS07]. Absence [CDL20, Hir16]. absolute [BS10b, Hu02]. absolutely [DMR95]. absorbed [Mai13]. absorption [BBS15, Jaf12]. absorptions [CGL21]. abstract [Gua77, KKK85]. abstraits [Gho75]. accroissements [Gal76, Hai87, Jac71c, JKM82]. actifs [AS94]. acting [AABV16, Her08a]. action [BL12, Cha91]. actions [Bew71, DK79, DS07]. Activated [BGHR19, RT18]. Adaptation [GJTZ08a]. adaptative [BM89, Mac78]. adaptés [GLS73a, GLS73b]. Adaptive [BLR11, BLR16, BEK19, BK21, BMP09, CM21, CGG11, FLRB11, HMSH12, Lac07, CGC10, CL13a, GNS13, GG01, IL06, Pol02, TV98]. Addendum [She04]. additif [MO83, Nev83]. additifs [Émi81a, Hac82]. Additions [AG91]. Additive [Émi86, Bas08, BCR09, Bax96, Ber89, CDLR17, Dev16, Émi85, Fer15, Fou95, GG79, HMSH12, KHT12, LL08, Wu99]. additives [Ber89, Fou95, RR70a]. adic [FP08, JdlR04, Mél06, Mus06]. adiques [Mus06]. admissible [EHJ07]. admitting [DK08]. advection [CX21]. affecté [Mac78]. Affine [CGB17, Cha12, Bro03, Cun03, GGL15, GLL18, GL16, JJK^{+}14]. affines [Con72a]. affinités [Has90]. affinities [Has90]. against [Hol12]. age [RDL20]. Ageing [MOS11, AD11, OR18]. aggregate [Ami17]. Aggregated [Aud04]. aggregation [Ami17, BDCKL20, Bla02, FNS00, LM13, ST20a]. Aging [DD07]. agree [Bor77]. Agrégation [Bla02]. Airy [Dim18, LS19a]. alcoves [Def16]. Aldous [FM09, HP17, HLT21]. aléatoire [AJ76b, Con65, Cun03, Deg82, Deh74, Dep99, FK93, Fis73, FLP06, Jan70, Leg89, Nev69, Pei92, Saa95, Sun87]. aléatoires [Ahm65, Bas89, Bec89, Ber70, BLT69, BDC67, BDC68, BK95, BH80b, BH80a, CD01, Dav74, DV98, Del87, Del64, Der76, EV03, Fer90, Fis76, For68, For73, FK75, GS97, Gre08, Gue99, Hai81, Han73, Hen91, Hil78, Hög74a, Hög74b, Kah87, Le 89, Led92, Loo77b, Mat93, Mén78, Mus06, Pay67, Ray68, Rei82, Roy80, Sch00, Sun87, Tor69, Tor71, Val89, Van72, lJ87]. algebra [Ben67a]. algebraic [BW15, Pou08]. algebras [GSS05, Rab94, vW83]. algèbre [Ben67a, Loo77a]. algèbres [Dho72, Rab94]. algorithm [HLT21, KOS19]. algorithme [Col78, Pel90]. algorithmes [BD96a]. algorithms [BH08, BD96a, Cat91a, Cer96, DG01, EM83, Pel90, Tro96]. allelic [Han04]. Allen [Web10]. Almost [BS10b, BDS21, CL13b, JRdlR10, RAS09, Wan07, DGM10, DDP87, Del64, Éme05, EM83, FR03, Pem97, RY95, RR94, DST91, Sen90b]. Almost-sure [JRdlR10, Wan07, EM83, Sen90b]. along [DW17, KRS04]. alpha [ACRC17]. alpha-shape [ACRC17]. alternative [Bir06]. alternatives [Kyp04]. amarts [Gho77]. among [BBHK08, BGN19, Mis16, Mou11a]. Analyse [Rou97, Web80, Bar81, Col78, DP75, Heu79, Hey68]. Analysis [Dia05, BDN20, Čer04, FGV16, FFJ19, FV05, HHK06, Hey68, Imk84, Kyp04, Rou97, Heu79]. Analytic [KKK85, AHK77, Per92]. Analytical [Pic64]. analytique [Pic64]. analytiques [TTL74]. ancillarité [Pet70]. Andersen [MST20]. Anderson [AD14, CCP12, CK12, Che17a, Che19, Che20, DM09, EdHM14, GS15a, GK09, HL22, HLN17, MP19, MOS11]. angle [BAC21]. Anisotropic [CL13a, FJR21, VST12]. Ann [BW06]. Annales [Bal04, Mat98b, She04, Ano05a]. Annealed [Ass11, BS10a, CGGvdH22, GKS07, GS15a, Poi13, Pov98]. annealing [Cat91a, Con94, Mic92a, Mic92b, Tro96, You88]. Anomalous [BBHK08, Bas14, MSC19, ST20a]. anon [KLO95]. anon-gradient [KLO95]. anti [GLS05]. anti-voter [GLS05]. anticipating [GNSS94, OP89]. Anticipation [FI93]. Anticipative [DS06, Nic03]. anticommutator [And15]. Antisymmetric [Gol95]. Any [GLTZ17, GNRV05]. apart [Bra09]. aperiodic [Kre76]. apparaissent [Lap86]. apparentés [Gil66b]. Application [GZ19, Pel72, Ber68, Ber88, CGPP06, CP04, CL09, DDT04, Hai87, KW20, LPP15, Mal11, MØP04, ÖÇE17, Sch00, SV08, Tor17, Tra15, Zhe85, de 95, dlP94, Bas69, Ber91a, Fer15, Hel99, Hil73, Lem73a, Lem73b, Mic92a, OP89, dZ81]. Applications [Ber89, Cat91b, Cha70, Ben67b, BCR20, Bos75, BR74, CG04, Coq93, CP11, DP75, Deg82, DG01, DW11, ES02, FGR09, FV05, Gas02, GW18, GH88, Hey68, HHM21, Ist92, Kip74, KRS18, Le 89, Lim12, MS83a, NRS10, Nua84, RV13, Röl13, Sam07, Ust82, Ber89, Der76, Fou00, Her94, Her05]. Applying [Tal96]. approach [Bai10, BG85, BCR19, BPTZ18, CDP09, CG96, CF17, ESYY12, FMP14, FGJ17, GOVW09, GB16, HM16, Pou08, WY16, WZ90, Woo16]. approximate [ÁEdBCAM11, JKMP05, BCG15]. approximated [CCFR09]. Approximating [Tal92]. Approximation [All78, Aza89, Bou88b, DSS13, MRZ98, Rec86, BCV21, CI03, CDM21, EM83, HLOT^{+}19, Hof15, HMP20b, LP90, Le 85, NPR10, Pel10, Tor17, Zai02, Aza89]. Approximations [BP04, FV05, BH85, CLB98, CE00, Kas20, SZ97]. Arbitrage [Str90, DS97]. arbitraires [Geo78]. arbitrary [AG89, AG91, DGR16, FWY00, MRZ98, Yan21a]. Arbitration [Str90]. Arbres [Cha86a, Nev86]. arcsin [Hob94]. Arcsine [ELS10]. area [BCK21, BJ20, DNT12, FLS06]. arises [Zen16]. arising [BDEG88, BDEG89, Bel08]. arithmetical [Gal70]. arm [Yao18]. ARMA [Ber91a]. ARMAX [Ber91a]. arms [vdBvE22]. Aronszajn [TV81]. arrêtées [Mat93]. array [CAM98, Tra02, dBMCA99]. article [Ano77, Bod71, Cha71, Jac74, Rou79, She04]. ary [CLP14]. Askey [LR86]. assets [AS94]. associated [AD01, BDT11, BL05, CW13, Cro18, DGM10, DTMS98, Gar99, GW18, KM14, Lou00, MR90, RW17]. associé [BKRR71]. associée [CL84, Tor69, dZ81]. associées [LM76, SM79]. associés [DTMS98, Wei84]. assumptions [Cas07, Gil65a]. asymétrique [KL95]. Asymmetric [CCEL13, AAD19, AI05, BMRS17, BF88, BW19b, Che06, DPS89, FFV00, FGM09, JLT21, KPS16, KL95, Kom98, Lam00, Lan91, LM97, Lou05, Rez02, Set07, Var95a]. Asymptotic [AD08, Ang16, BBL14, Bud03, IL06, Kom92, MCRT06, Mik88, R.83, RSH19, Sim07, Tou15, Uek99, Ako93, AI99, ACA02, Bag85, Ber85a, Bra09, BM89, CF17, CM16, CAM98, DY11, Dos80, DST91, GGS15, GO13, HMP20b, Jia12, LL15, MRW20, Mic09, MR89, MT08, NNRT09, NZ13, Sen90b, Sta89, Tuy81, Web10, Bab91, FLP06, Fra91, LS91, Mey91, Mic92a]. Asymptotically [JM93, KM02, CG96, Sin07]. Asymptotics [BK18, DLS02, FHSX21, GHS11, Gui00, Hob94, HL22, KL16, LW20b, MNR21, AD13, Bar15, CK20, CR05b, CY07, CHSX15, Che17a, DE05, FW09, FR03, Gan02, GL16, HV17, HLN17, MW02, Riv12, XZ09]. asymptotique [Ako93, ACA02, Bab91, DL84, DL78, FLP06, Fra91, LS91, Loo77a, Mey91, Mic92a, Mic09, MR89, PPH75, Rev83, Sen90b, Web80]. asymptotiques [DP84a, Dos80, DR69, DST91, Hei78, MV69, Rao70]. asynchronous [Fis01]. Atlas [DJO19]. atom [AP05]. Atomes [Tor78a]. attached [FT05]. Attainable [DS96]. attenuated [NdS20]. attracting [BY19]. attraction [FT12, Gin80, Gin81, MP84, MR94a]. attractive [BGRS14, BRT14, Lan91]. attractiveness [GS10a]. attractors [MR94b]. au-dessus [Val89]. auto [WY16]. auto-covariance [WY16]. automata [AMS95, BT00, MM14]. automorphismes [CM70]. autonomous [GI80, SSV03]. autorégressif [DST91, Gas90, MR89, Tou96]. autorégressifs [Mok90]. autoregression [DST91]. autoregressive [Gas90, MR89, Mok90, PD14, Tou96]. autres [Cal81]. aux [Ber88, Ber89, Coq93, Del87, Der76, Fou00, GH88, Hai87, Her94, Her05, dZ81]. Average [Che97, Har15, DGW06, GS97, Mör01, Tes18, Dep99]. averageable [Mus06]. Averaged [Yil10, SY11]. averaging [CP17, HKPG16, HP17, Kif14]. avoiding [BST20, DCKY14, GM20, Hey11, Zer05]. away [EO22, PPR18]. axes [Pei92]. axiomes [Gil65a]. axioms [Gil65a]. ayant [FSP71, Hög74b]. Azéma [Enr07, PP98]. Aztec [BK18].

B [Mat98b]. backtracking [Her19]. backward [CR20, Wu99, Ham96]. Backward-stochastic [Ham96]. Baird [Duh97]. Bakry [CDP09, Joh17]. balanced [CCKW21, DGR18]. balayage [GS87, Mal90]. Balkema [Wüt04]. ball [HPT02, KM02]. Ballistic [CP12, JP17, KK03, Pet09a, RAS09, Sch06, Sep00, She03, She04]. balls [Bor77, Sod08]. Banach [ALM20, Bas82, BH80a, BH81, DCS74, DS97, Gho77, GP74b, GL80, Hei74b, NN81, Rao72, Sch83, Wer80, Yor74b, ZBH17]. band [BGP14]. Barak [MR21]. Barrier [BRZ19, Jaf12, Mai13]. barrières [Dep65]. barriers [De 18, Dep65]. Barycentres [Pic94]. based [HLNV13, Hir16, MNR21, VW15]. Basic [EdHM14]. basis [GLTZ17]. Baxter [GM20, Rei82]. Bayesian [BCG15, FFJ19, FT79, GvdVY20]. bayésien [CM72, MPL73]. bayésiennes [Flo82, Lan82]. be [dlRd03]. Beboutoff [Bru71]. behave [BRR02]. Behavior [GH10, Ako93, Ang16, AI99, Bab91, BMRS17, BBS08, CGG10, CM16, DS90, FGN13, Gre08, Kes86, Lin17, MCRT06, Mic92a, Mou96, NZ13, Sch06, Sin07]. behaviour [BEV13, CM15, CPV22, GP14, GS15a, GO13, MM13, Mer06, R.83, Tuy81]. Bellman [Cha86a]. Bennett [Pin14a, Rio02]. Bernoulli [BS17, BS10b, Bou80, EN20, JdlR04, Tes06]. Bernstein [Dub75]. Berry [Gou05, KRT16, Tol00]. Besicovitch [LT06]. Besov [Lia07]. Bessel [GRVY99, LLO06, LR96]. beta [BBS08, DE05, LS19a, RS18]. beta-Dyson [RS18]. Between [Fle10, BG10, BGJS18, BFK20, BC19, DV08, DT03, Fou95, FX19, GS15a, Gil66b, Hue16, Mal65, Pan14, Pra07, Pru97a, STW00]. beyond [JS20]. Bhatnagar [Hof15]. bialgebras [Hud05]. bias [Jia12, MSZ15]. Biased [FP18b, AHOZ13, BGN19, Jir78, MSZ15, RT18, SSS^{+}21]. bibliographie [Ano77]. bidimensionnel [BG92, Cai82, Led81]. bifurcating [PD14]. bifurcations [CDLR17]. big [Hu16]. Biggins [IKM20]. Biharmonic [Gol89]. biharmoniques [Gol89]. bilinear [HLNV13, KL16]. billiard [Pèn19]. billiards [CP12]. Bin [Sta90, MR21]. Bin-packing [Sta90]. binary [BRZ19, DRZ21]. Bingham [CPV22]. Biographie [Bre02, DC02]. biography [Ibr02, Bre02, DC02]. bipartite [Abr16, BL20]. Birkhoff [Bew71]. birth [CCM19, CD19b, LM09]. birth-and-death [CCM19]. birth-death [CD19b]. births [Mou96]. Bismut [Bañ18]. Bismut-Elworthy-Li [Bañ18]. bistochastic [BQZ20]. bivalentes [Han72]. Bivariate [Wüt04, Yin96, CLB98, GvdLW95, Vig03]. bivarié [Vig03]. black [EF16]. blanc [Enr96]. block [HMNP98, MS21]. blocking [BB18]. Blow [Thi91]. Blow-up [Thi91]. BMO [Eme85]. Board [Ano05d, Ano16a, Ano16b, Ano17a, Ano97, Ano99, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05b, Ano05c, Ano05e, Ano05f, Ano05g, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f]. Bochner [CDP09, HS96, NN81]. Boguslavsky [Die82]. Boltzmann [Car15, Fou01, GM19, RT97]. Boltzmann-grad [RT97]. bond [CI03, EFG^{+}20, Gol18, GW18, JL08]. bonds [FGN13, GLS05]. bones [Bou00]. Boolean [AS17]. booléens [Bod70, Bod71]. bootstrap [AMS95, AG89, AG91, Bre83, CAM98, DMS93, GR08, dBMCA99]. bord [Pri68]. Borel [DGR16]. borélien [ERR71]. boréliennes [SP76]. bornée [SP75]. Bornes [Mus06, Léo86, TV81]. Bose [Dia05]. both [Zho07]. Böttcher [FW09]. bound [Ber85a, DP11, Lac12a, Lac12b, Mej04, Yur08, de 95, dlP94, vdBvE22]. boundaries [AD01, BW19b, De 18, GM13, McG89, Ove94]. Boundary [Gra85, AC20, AB02, AR06, AG21, AT20, Ass20, BDCH16, Bet15, BP95, BCT11, BW05, BW06, Col04, DR22, FRS22, Fun99, Gra88a, GM19, HPT02, HP19, LS19b, PW17, Pri68, Qia19, Sło93]. boundary-weighed [DR22]. bounded [And14, BCV21, BBHK08, BEK19, BCQ21, De 98, DMPU09, Dep65, DK08, NRR20]. boundedness [CKK17, DHJ^{+}21, KMR06]. boundedness-by-entropy [DHJ^{+}21]. Bounds [Cou20, DGM08, Kol03, AC99, BKS12a, BC20, BOT14, BGM20, CKS87, Cha21, CFP19, DHHX20, Gui02, KL16, KRT16, LLL11, MRVZ16, Mus06, Rio09, SS10b, Tra15, Vay03]. box [RT10]. boxes [Abe15]. bracketing [Pol02]. Brailovsky [Bar81]. branched [BC19]. branchement [Cha86b, Hae72, Rou87a, Rou87b]. branchements [Lot85]. Branching [BP15, HH09, HLOT^{+}19, Nak15, AD09, ABKV14, ABK15, AD11, BB14a, Bar00, BBS15, Ber13, BHR17, BD16, BRT07, BDZ16, CH14, CLP14, CH20, CG14, CR96, DFM^{+}03, DGdH^{+}08, Eng08, EHK10, Eva91, FS04, FM19, FR03, GHS11, GKV03, GK17, HR12, Her77, HH81, HL05, HT04, Hu16, Jaf12, Jiř70, JS20, KM20, Kli09, KLMSR12, Lab14, LS16, Le 91, Liu96, Liu01, Löc02, Mai13, MZ16, Mar20, OR18, Ove94, ÖÇE17, ÖE19, Pai18, Riz15, SBM21, Zho07, Zhu21]. branching-selection [CG14]. Brascamp [CCEL13]. Bravais [MP19]. breaking [BZ12]. breeding [HH09]. Bretagnolle [Bre02]. Breuer [NZ21]. bridge [LR96]. bridges [Con18, GP11, NR17, PZ04, Vig11]. Broder [HLT21]. Brownian [BB02, Abr16, AT20, Arc21, ABK15, AWX08, ASX11, Bab91, BBD19, BB89, BH97, BB01, BBC05, BO11, BOT14, BOZ16, Bef03, BSC04, BB11, BBS15, BK21, Ber89, BL08b, Bet16, BL03, BJ20, BDP95, Bra09, BÉL09, BL12, BM93, CCM03, CPY04, Car85, CK22, CG79, Cha94, CDH00, Che15, CF09, CK12, CN05, Chu84, CGPP06, CV14, CP11, CP04, ClJ89, Cro08, CCFR09, CLM13, Dah17, De 99, DS90, DV05, DPRZ00, DNT12, DMY91, DN01, Dub04, EH03, Eld14, EF16, Éme05, Eng08, EPT17, FGV16, FFW05, FGR09, FWY00, FINR06, Fra91, FV05, Gan67, GS18, Gol89, Gra99, GNRV05, GMV93, GM19, HP11, HH09, HR12, HNPS19, HLLS22, HP19, Jac18, Jac10, LR14, JN19, KPS16, KL96, KRS04, KM17, KLMSR12, Lac91, Lac97b]. Brownian [Lac12a, Lac12b, LS19a, LTT18, LSW02, LW06, Le 19, LY19, LR93, Lou68, Mai13, MZ16, MM18, Mal90, McG89, Mer06, Mey91, MSS06, Mör01, MT08, Nak15, NdS20, NRS10, NR17, NNT10, Osa98, ÖÇE17, ÖE19, Pec01, Pem97, Per89, PT98, Pia95, Qia19, Rai99, RS18, RSS16, RVY09, SY05, SV07, Sar17, Sch09, STW00, SBM21, Wan07, Wer93, WZ90, Wüt98, Wüt99, vdB07, vdHS03]. brownien [All78, Bab91, Ber89, Bia86, Fra91, Gol89, Lac91, Lac97b, Lou68, Mey91, Nev76, Rec86, Sch85, Wer93, Yor85]. browniennes [BL12, LR93, Mal90, Pia95]. browniens [Rai99]. bruit [Bro83, Enr96, Lap86]. bruits [BL79]. Brunel [FG79]. Brunn [Ehr86]. BSDE [PX05]. BSDEs [BPTZ18, CPX17, DHR11, DKGT16, HKM14, MPR17]. building [Sch09]. Bulk [LWZ16, Péc12]. bundles [Mal11]. Burgers [Gir02, Gou07, JLT21, Mor99]. Burgers-type [Mor99]. Burkholder [RW17, dIS02]. Burkholder/Rosenthal [RW17]. burnt [Mar16]. busy [CG68]. bypass [BD96a].

C [Duh97]. cône [Mey91]. cactus [CLM13]. cactuses [CLM13]. Cahn [ABK18, DDT04, Web10]. Calcul [CG92, Gra88b, Bec74, Ber88, Der90, Fis73, Lon69, Mic09, Pel90, RZ93, TTL69]. calculating [RZ93]. calculation [Mic09, Pel90]. calculus [CG92, DOS18, DS06, Der90, FHSX21, Gra85, Hud05, Maj19, NPR10, Pic05, KHT12]. Cameron [KKK85]. can [BT18a, Ber95, dlRd03]. canal [Mac78]. cancelled [FI93]. canonical [Coq93, TTL68]. canonique [Coq93, DP75]. canoniques [TTL68]. Cantor [Bat00, Kig13]. Capacity [FS89, BP92, GS87, LT21, vdB07]. Caractérisation [Bal81, FSP71, RZ93]. caractérisations [Sch72b]. caractéristique [Hen91, Le 89]. Caractéristiques [Arn93a, Cup69]. Carlo [HLOT^{+}19, MNR21]. carpet [BB89]. carré [Gal76]. carrés [Ber91a, BD92, CPP71, Deg82, DST90, DST91, Tou96]. Cas [Tou96, BLT69, Gal76, Gil65b, GLS73b, Ham96, Mal65, MR89, Leg89, Les87, Mac78, Pei92]. cascades [BDK19]. case [AFGJ16, BBD12, CI03, CF18, Eng09, EG08, FW09, Gil65b, GG08, GH10, Gor17, GNRV05, Ham96, HJY11, Kob13, Mal65, MR89, Pèn19, Riv12, Smo83, Tou96, Var06, dT07, Leg89, Pei92]. Castelle [DC02]. catalytic [AN09, CH14, DFM^{+}03, DV05, Kli09]. Catching [ES79]. catégorie [Ben67b]. category [Ben67b]. Cauchy [Dav05, MR94a]. causal [Gas90]. Cavity [Pan09]. Cayley [Mar16]. Cech [IY20]. cells [DLP07, GC03]. cellular [AMS95, BT00, HKPG16, MM14]. censored [CG14, GG01]. Central [BY08, Bay15, CDMF12, DVNZ20, LSX21, NNT10, Zhe12, BBC78, BF92, BG99a, BG92, BU18, CDGR20, CR78, De 98, DR00, DR08, DMR94, EV03, Gri04, Gui02, HJY11, Her08b, JL06, JL08, Led83, Mas86, Ouc05, RAS09, Rez02, Rio09, RR94, Roy74, BFK20, Bli14, Bru93]. central-limite [Roy74]. centrale [BF92, Bru93, JKM82, Mai78, Tou83]. centred [Bro03]. centrifugal [FLP06]. centrifuge [FLP06]. Centroids [Pic94]. cercle [Car81]. certain [BLMZ15, CPY04, DGZ03, Éme05, Gan67, Joh12, Val89]. certaines [Bre83, Fer90, GMP78, Hen82, Pet79, Pic64, Tor82, Val89]. certains [Ant82, BDC69, Car81, Dep65, Die81, GLS73b, Hei74a, Rot90, Yor77]. Cervonenkis [Pis84]. CGL [Kuk13]. Chaîne [Rou72, CG92, Gir70, Gra88b, LS91, Loo77a, Mai78]. Chaînes [Gal96, Bab88, Geo78, GH88, Hen82, Her05, Her08b, Jac71a, Jac74, Mét72, Mic99, Nev72b, Sun75]. Chacon [Egg80]. chain [BH20, BB17b, CY07, CG92, Dia05, GZ98, Lac07, LS91, LPS03, MNR21, Sar14, SJ14, SZ97]. chains [AP98, AKU20, AP05, AMZ21, BFG15, Bor08, CG04, CL09, DGM10, DMS05, DE02, DGM08, DFM16, Gal96, Her05, Her08b, MRZ98, McD75, Mic99, NRR20, Pel10, Sud21]. Challenging [Cat12]. chamber [Gra99, Ras11a, Sch09]. champ [GLS73a, Val78]. Champs [DAP76, Del77b, EV03]. Change [GJTZ08b, GJTZ08a, Kom98]. Change-point [GJTZ08b, GJTZ08a]. changed [BT18b, CF09]. changement [Bou84]. changements [Arn93a]. changent [DS84]. changes [Arn93a, ALT99, CGB17, VA89]. Changing [AD09, NS81, CP11]. channel [Szn82]. chaos [AZ99, Car15, DK19, GQ15, GK93, Har01, JM98, MØP04, Pec01, ST94]. chaoses [AL12]. characteristic [Har15, Hen91, Le 89, LR86]. characteristics [Arn93a, Con18]. Characterization [Che87, GRS^{+}18, Har05, SS10a, BM08, KK75, RZ93, Ros01, SS18]. Characterizations [Son13, Egg80]. charged [Ass11, Che08, GPdS16]. charged-polymer [GPdS16]. Chatard [Cha71]. Chemins [Mer83]. Chervonenkis [Vay03]. choc [Leg89]. chocs [Pei92]. choice [LW20a]. Choquet [Gla76, Rau83, Rau04]. chordal [Zha10]. Chung [Deh00, KL96]. Chung-type [Deh00]. Ciesielski [KP11, Yor91]. circle [BS13, ClJ89, Lac17, MOW09]. circulants [BHS12]. Circular [Coo19]. claims [DS96, DS97]. clairvoyant [BT18a]. Clark [Ané01]. class [Bal03, Bal04, BKS12a, BO11, BDC67, Bur07, CD12, DVW00, El 00, EPP06, EW06, FK95, FGN19, GSS05, GRS^{+}18, Gri77, KM14, KPR18, LPP20, LZ15, Mar15, Mél06, OP89, Pri18, SSV03, Sch83, SS18, dlR02]. class-one [BO11]. Classe [Led81, BB78, BDC67, Dav74, GH88, RLT81, Rei82, Sun75, dlR02]. Classes [Del87, Bag85, GP07, KPT20, Pis84, RV20]. Classical [AB18, ADL16, Cue18, Dah17, Kel17, NW18]. classification [DNN73, Jac71a, Jac74, Pet79]. classique [DP84a]. Claude [SL95]. CLE [SWW17]. clients [Gra84]. closed [Del64, Per89]. closure [KL02]. CLT [ADL16, CL09]. cluster [ELS10, GMSS18, Kes86, PSS09, Qia19]. clusters [Abe15, VST12, Yao18]. coagulation [NO88]. coalescence [BLNR11, KPR18, Sep20, STW00]. coalescent [BL05, FFJ19, KZ17, Sen17]. coalescents [Bas08, BBS08, EP98, Lim12, BBL14]. Coalescing [Eva97, FLT04, Zho07]. Coarsening [FINR06]. Coboundaries [VW04]. cocycles [BU18, Der00, Len04, LS05b, Tia17]. coefficient [GCJ93, Hof01, HMSH12, MGY20, Nag05, Sue05]. coefficients [BT15, BP04, Def16, FX19, FL20, Hen91, Ist92, Léo86, XZ20]. Coexistence [CS21, CH12, CP01]. cohomologous [Der00]. coin [AKU20]. collection [FK93]. collections [Hut20]. colliding [Gra99]. Collision [FGM09]. Collisions [BPS12, KPS16, Pei92]. coloring [ABGG^{+}12, Hol17]. colorings [FP18a]. comb [AR16]. combinatoires [Laf72]. combined [CCFR09]. Combining [Mat98c]. combustion [CQR07]. comme [Led70]. common [JP17, Rou21]. commutatif [Bru73]. commutative [Bli14, Gui02, Hey68]. commuting [KT97, Ros89]. Compacité [Tor69, BR74, For78]. compact [BT15, Bor21, Cha94, Cha70, CL13b, Dah17, DS07, EP06, FK10, Fra91, Gre08, Haz93, Her94, Her05, Hög74a, Hög74b, MR94b, Pri68, Rau92, Yan21a]. compact-support [FK10]. compacte [Fra91, Her05, Pri68]. compactes [BBC78]. compactness [Her08a, Lin75, Wu99]. compacts [BR79, Con72b, Con72a, FSP71, Her94]. Comparaison [Bou88a, Fer84b, Fli81, MPPL71, Vig03, Tor78b]. Comparing [AK20, BHLP19, BG85]. Comparison [DV08, Pan14, CK17, Fli81, Vig03].

Comparisons [Pru97a]. compensated [ZBH17]. Compensation [BLT69]. compensators [Tuy81]. competing [CS21, Hol12, Sar17]. Complément [Ano77, Bod71, Tor77a, Tor78c]. complementary [HNPS19]. compléments [ACTR85, Jac74]. complet [Sun75]. complete [RN21]. Completely [Lam00, Ber68, Tor65]. completely-connected [Ber68]. complètement [Hei74b, Lon69, TTL69, Tor65]. completeness [HHM21]. complètes [Ber68]. complex [Hos18, IKM20, Oda06, Péc12]. complexes [IY20]. complexité [Die76, Die78]. complexity [Aud04, GJTZ08b]. component [ČTW11, CD12, GKMMR20, Gan13, KKW15, LMMR18, Val06, Was15]. Components [Ste74, Ben67a, HNPS19, Mar16, Vig03]. Comportement [Ako93, Bab91, Mic92a, Mou96, DL84, DL78, Gre08, Rev83]. composantes [Vig03]. composite [BB14b]. computation [ÁEdBCAM11, BCG15]. Computing [PP98]. concave [AL12, CCEL13, DT03, Hei74a]. concavity [BKS12a]. Concentration [AWS21, GP07, NRR20, NS22, ORB20, BR08, CM10, Con18, Fle10, Goz10, Haf20, KL16, LM09, Sam07]. concept [Gui90, Jan70]. concernant [Ant82, Bar81, FK69, Loo77b]. concerning [JV07, PS92]. Condensation [BY19, RCG18]. Condition [For78, Aïd20, BCP19, BDMBP18, BD97, CMR02, Cat86, DV05, DK19, FGN19, Joh17, LLL14, Mic09, Pig18, Riv07, Yos01, dZ81]. Conditional [ABKV14, BCQ21, CKL14, LM83, Mat03, ÖÇE17, BLR16, Chu84, CGG11, GR11, GGS15, Gru66a, Mat12, RF64, Car80, Str88]. conditionally [LMD03]. Conditioned [GMV93, LW06, CC08, Eva91, GLL18, HH81, Kor15, Pin90, Ric13, Riz15, Véc19]. Conditioning [BCQ21, Qia19, BCK21, FLS06]. conditionnelle [Car80, MPL73, Nev76, Str88, Val80]. conditionnelles [BH85, Gru66a, Ngh70, RF64, Thi81]. conditionnement [Bod70, Bod71]. Conditions [AJ76a, Bla75, AB02, BDCH16, BCT11, DHR11, Egg80, Gra88a, HSS21, Her77, Joh12, LS19b, Sło93, dBMCA99]. conductance [And14, FHS19]. conductances [And14, BBHK08, BGN19, Val12]. conduction [Bor96]. Cone [GMS19, KRS21, Mey91]. confiance [Hil77]. configuration [AvdHH19, CGGvdH22, DvdHvLS20]. configurations [BBS15, Mis16]. confined [Lam00]. confinement [GK09]. Conformal [BDCH16, GM21, HP21]. conformally [Bef03]. conjecture [Fle10, Har05, HP17, Loo81, Pia95, RV20]. Conjugaison [CM70]. connected [Ber68, GMV93, SWW17]. Connectedness [BT16]. Connectivity [SS10b]. connexes [Bal81, Ben67b, Ben67a]. consecutive [HP17]. conservatifs [KRR70]. Conservation [MU11, BF95, Hof15]. conservative [GS10a, LPY02, Set01]. consistency [GG01, GG02]. consistent [ABCF21]. Constant [COBZ21, CMR02, Cou20, Mal65, PRS11]. constantes [Mic09]. constants [Mic09]. constituants [Ben67b, Ben67a]. constituents [Ben67b]. constrained [BT18c, JM93, KM02, PS19]. constraint [CPX17, DF20]. constraints [BAC21]. Construction [BDMBP18, CE20, EP98, FGV16, Wat87, BB89, BW15, DE02, GN02, Szn80, VW15, GLS73a]. Constructive [DSS13, BLM21]. contact [BBCH21, GW18, LS06, dHdS14]. contains [Pem97]. contamination [GKMMR20]. content [HLLS22]. Contents [Ano16c, Ano16d, Ano17b]. Context [Fri67]. Context-Sensitive [Fri67, Fri67]. contingence [Col78]. contingency [NN20]. contingent [AS94, DS97]. contingents [AS94]. continu [Ben67c, Cup69, Lot85]. continuation [Per92]. continue [Mal65, Rou72]. continued [CT11]. continues [Ber88, Del64, Nua84]. continuité [Bec89]. Continuity [BLM21, Émi85, AC04, AWX08, ASX11, Bat00, Bec89, BS10b, CCS92, DLZ20, Hu02, HHM21, HK09, KMR06, MP84, Ros96, Ros99]. Continuous [FM19, JYZ05, Ros10, AD09, BCC02, BFG15, BD16, BK04b, BDM11, Che17b, Del64, Der00, Eva97, GG79, HJY11, KL05b, Lab14, Lin77, MP84, Nak86, Smo83, Var06, Wu04, Mal65]. Continuous-state [FM19, BD16, Lab14]. Continuous-time [JYZ05]. continuously [Ben67c]. continuously-deformable [Ben67c]. continus [GP74b, Thi81]. Continuum [GM18b, Čer04, CW13, Cro08, DH21, RT97]. contour [Che76]. contournent [BD96a]. contournent-ils [BD96a]. Contrôle [Geo78, LM77]. contractible [Haz93]. contraction [Ast76, Lon69, TTL69]. contractions [Bru73, Kip74, Leg89, Pei92]. Contre [EV03]. Contre-exemple [EV03]. converge [Abr16, FFW05]. Convergence [Ber68, BDZ16, CV14, CR96, Cro08, Dar85, DJO19, DMGM18, Fis76, FR14, GL89, GM19, Jac79, Kun86, LR20, Mét84, Nev72a, Ngh70, OTT02, vdHS03, AJ76b, AJ76a, ABW17, AAP09, BDS87, BP85, Bag85, BH08, BKK10, BR03, BHR17, BCL15, BL14, CCC16, CL13b, DFM16, EM15, EM83, FP17, Fou84, GP01, GB16, Her08b, JKM82, Jac98, KMS97, Lam20, LS14, Lou00, Mas86, MZ15, Mic99, MR89, Mil83, MO83, Nak86, Ouc05, Pra92, RY95, RSX21, Sar18, Sla22, Tes18, Tou96, Tro96, WZ90, dZ81, Ber68, GL89, Pra92]. convergentes [Jac82]. converges [Hey11]. Convex [CDP09, ACJ20, And09, BDT11, BGKL17, CD12, DDT04, DHR11, DKGT16, Har05, LM13, LR19, LM09, Mas93, RW17, Sen90b, SS18, Zer05, DV98, Dav02]. convexe [Mas93, Val80]. convexes [DV98, Dav02, SP76, Sen90b, Tor77b, Van72]. convexity [Bor96, Cas07, Mat12]. Convolution [LW96, CL13b, Cou20, DL84, Fog75, LS05b, Riv12, Sam07, Sie84, WW16]. Convolutional [MR94b]. convolutions [AN09, Bel06, BG10, BS10b, BW01, KMR06, OV20, Tor66, ZBH17]. cookie [KP17, Pin10]. cookies [Bau13, KOS16]. cooling [Cat91b]. Copolymer [Pet09b]. cordes [Ber89]. core [PS14]. cores [ABW17]. corner [Del08]. corners [DR22]. Correction [Bal04, CL95, de 95, AG91]. Corrections [Jac74, MS21]. correlated [ACLS19, BT13, EM83, Gri04, HLN17, KK03, Mad15]. Correlation [KMS14, Gil65b, CD19a, FMP14, Gil65b, Har05, Lac12a, Lac12b, Bas69, Bas89, Lar93, Bas89, Lar93]. correlations [BMP94a, FRS22, Hel99, Poi13, BLT69]. cosmology [Mal11]. costs [BH08]. cotype [Bas82]. Coulomb [KM17]. countable [Tor66, vW83]. Counterexample [EV03]. counting [CGG11]. couplage [CTR95]. Coupled [Fis01, ACLS19, Par05]. coupling [Geo97, Jui16, Maj19, Pet09b, SW11, CTR95]. Couplings [BJ20, GS10a, EZ19]. courant [Dep99]. courbure [Bab91]. Courte [Nev83]. Couverture [AS94]. Covariance [GBB69a, Rio93, CCEL13, Deg82, Gil66b, GM21, Heu79, KL16, LSX21, Pan14, Péc12, WY16]. covariances [CD12, DE02]. covariate [BK21]. covariation [DY11]. cover [BRZ19, DRZ21, KW20]. Coverage [AS94]. covering [BE95, FK93, FW04, JJK^{+}14]. Cox [Sen90a]. creep [Ber95]. criteria [DV08, MZ84]. Critical [BBS15, Čer04, HH81, dT16, AC18, BS17, BRZ19, BCL15, BS10a, BBD12, BDK19, CGN16, Che19, DE16, GLT11, GM20, GMSS18, GMS19, Han17, HP21, Jaf12, LS16, Lin17, Pai18, PSS09, Poi13, Zhu21, vdBvE22, vdHS03, vdHdHS07]. criticality [CGK20, DvdHvLS20]. Croissance [Roy80]. croissants [Deh74, Mer83]. cross [GM13]. Crossing [BS17, BDCH16, BCL15, Dub04, Wüt99]. crossings [CG68]. cube [Ané01, Cau69]. cumulants [HS11]. Curie [CGK20, EM15, Gor17]. Curie-Weiss [CGK20]. curlicue [Cel11]. current [Dep99]. currents [FGR09]. curtain [Jui16]. curvature [Bab91, Dar85]. curve [Bef03]. curved [GM13]. curves [BB01, BB02, DF20, Poi13, RZ17]. customer [FM09]. customers [CP78]. Cut [Bla03, BSZ03, ABDG19, Ber15b, CH17, HLLS22]. cut-offs [CH17]. cut-tree [ABDG19, Ber15b]. Cutoff [EN20, ST20b, Yan21b, Fou01, Lac17]. cutset [DT20]. cutsets [PS14]. Cycle [vdHS14, BGHR19, CG68, DK08, RZ20, SZ16]. cycles [NZ13]. cyclic [LLO06, PT08]. cylinder [Gua77, Smo83]. cylinders [BT16]. cylindrical [PW79].

D [Gra84, FGS18, AKSS07, Dey20, PZ21, Yao18]. D/D/1 [Gra84]. Dacunha [DC02]. Dacunha-Castelle [DC02]. Dade [Thi91]. d’aide [Col78]. d’algorithmes [Bou88b]. damped [MN18]. d’analyse [JT76]. d’Anosov [GH88]. dans [Ahm65, Alt87, Arn93b, Ass87, BF92, BLT69, BG92, BH80a, Cau69, Dep65, Dep99, DP84a, Dho72, ERR71, EV03, Émi81a, Émi81b, FK93, Fou84, Gar79, Gil65b, Gil66a, GBB69a, GP74b, Gra84, GLS73b, Hac82, Hei71, Hei74a, Hei74b, Hei83, Her08b, Hög74b, Jum80, Kip87, LS91, Lar93, Led83, LM77, Mal65, Mar88, MPL73, Mét84, Mey91, MR89, Ouv73, Pay67, PPH75, Rou87b, Saa88, Sch85, Sen90a, Szp78, Tor65, Tor66, Tor69, Tor77b, Tor78a, Tor78b, Tor81, Tor82, Val78, Yor74a, Yor74b, Yor85, dlRd03]. Dantzig [BLR11]. d’apprentissage [Ber68]. d’arêtes [Bou00]. d’association [CGR79]. data [CPX17, FK10, GG01, Gir02, JM98, PW17]. d’attente [Fli81, Gra84, Jai80, Pel79, dZ81]. d’auto [CF87]. d’auto-organisation [CF87]. d’Aven [Die82]. Davie [Pri18]. de-Poissonization [Zai02]. death [CCM19, CD19b, LM09]. deaths [Mou96]. Decay [CD12, AMS95, BBHK08, CDP09, Hel99, Mou11b, BMP94a]. decaying [KL20]. d’échange [Lic70, Mal96]. d’échantillonnage [Lac73]. d’échantillons [Deh74]. décision [Lan82, Rom69b]. Decoherence [Reb05]. Decomposition [DW17, Cup69, AS92, Ber91b, ÖÇE17, Rau09, Ros99, Wu99, AS92, Ber91b, Pel79]. decompositions [Duq03, LS14, Yar19]. deconvolution [ATV20, CL13a, Tra15]. decorrelation [CFP12, Pèn19]. decrease [SV07]. decreasing [CD18, Gar99, Pru97b]. d’EDP [GP01]. Deep [KRS18]. default [SJ14]. d’effectif [Gil66a, GBB69b, Mal65]. defined [BC17]. définie [PPH75]. defining [Kig13]. definite [Gan67]. Definition [Ber13, Eme85, Bos75, Eme85, Lem73a, Lem73b]. defocusing [PSS16]. deformable [Ben67c, Ben67c]. deformations [CDMF12, PRS13]. deformed [Ben20, LSX21]. Degenerate [Kli09, CD16, DG19, DPSZ16, FHS19, GM18a, GLT09, HM16, KMM10, LPP20, Rho08, Rho09, dR17b]. Degeneration [KP96]. dégénéré [BKRR71]. dégénérés [Léa85, Léa88]. degree [MS21, Riz15, dIS02]. Degrees [MH18]. Dehn [SY11]. delay [SvG11]. delays [KMB92]. delocalization [BGP14]. delta [MØP04]. demande [Lic69]. demi [Hög74a, Hög74b, Lar72a, Lar72b, Lar72c]. demi-groupe [Hög74a, Hög74b]. demi-groupes [Lar72a, Lar72b, Lar72c]. Demiconvergence [MS83b]. démographie [Pic65]. demography [Pic65]. démonstration [Nev83, Rau83]. dendrites [FTV20]. dénombrable [Gru66a]. dénombrables [Tor66]. d’ensembles [Mal96]. Densité [BL12, AJ76b, AJ76a, BB78, Con94, CG92, Lot85]. densités [AS92]. densities [AS92, Cas17, CM21, CCC16, FJR21, GL20, Ish97, KHT12, LPP20, RSH19, Tsi02]. density [AD00, BH97, BOT14, BG20, BLR11, BLR16, BEK19, BM89, BW19b, CM16, Con94, CG92, Dev89, ES17, FL20, GNS13, GG01, GG02, HL22, Lac07, Ler12, Mik88, Mor99, Mör01, PS19, R.83, RSH19, Sab01, Sar14, Sar21, TV98, BL12]. d’entrée [ERR71, Gir70]. d’entropie [DCG90, Del87]. Deny [Gla76, Rau83, Rau04]. dépend [Fis73]. dépendance [Gil66c]. dépendant [Lot85, Lot91]. dépendante [Hai87]. dépendantes [Hai81]. dependence [CGL21, CM08, DLS02, FF18, Gil66c, HMNP98, Mat98c]. dependent [BDEG88, BDEG89, BCHJ21, CGG11, DR17c, DGR18, ES17, FL21, GPP10, HMNP98, HSS21, Hol17, LR17, MRTZ16, RDL20, Rho08, de 95, dlP94]. depending [FL20, Lot91]. déplacements [Bal76, Roy74]. deposition [Bal03, Bal04, BKS12a, Sep00]. d’équations [DL78, Oga74, Szp78, Thi91]. d’équilibre [KMS97]. Dérivabilité [Hen91]. Derivation [JLT21, LNS89, LNS89]. Derivative [SBM21]. derivatives [BRR02, BCT11, FG92]. dérive [Gal96]. derived [AP98, Sud21]. d’erreur [Bar81]. Derrida [BK04a, BK04b]. descentes [Ber88]. describing [Tsi02]. Description [Car81, Fla02, HKPG16, Sam07]. design [Ver10]. Désintégration [Gav75, SP75, Cal81, Pel72]. d’espaces [Del77b, Szn80, Val78]. dessus [Val89]. d’estimateurs [BB78, DLP86]. d’estimation [Ant82, Jum80]. destruction [Loo77b]. d’état [Jac73]. Détection [Gar79, AC20, ACA02, BG16, Gua77, LMMR18, Gar78]. Determinantal [ARRS16, DW08, Bor08, BD11, Def10, ER13, Har15]. détermination [Bro75, Dub75]. Determinism [BDC69]. déterminisme [BDC69]. deterministic [BLMZ15, DGM21, KV22, LL08]. deux [Bec74, Ben67c, CL84, CGR79, Der76, Fou95, Hil73, Len77, Nua84, RR70a, Vig03]. Deviation [PD14, BM16, Cat03, Cat12, CH20, Dos91, FK94, FW07, FV07, GZ19, Gou07, HP00, Ish97, R.83, Rab94, Tro96, Wu00, DLP86]. Deviations [Kob13, Aid10, AB97, Aug18, BESY08, BCR09, BA09, BFG15, BD96b, Cas04, Cat91b, Cat91a, CR05a, CD19a, CG04, CGK20, DO20, DMPU09, DGZ03, DMS05, DGW06, Eic01, ES02, ERS15, FG12, FS17, GZ98, GKS07, Gin83, GHR01, Gui02, Haf20, HK02, IR98, KPT20, KL95, LT18, Led92, Léo95, LS05a, Lot91, MRTZ16, MM09, MN18, MSS06, MO07, PD14, RAS11b, RT10, Tra02, Wu91, Yil10, Zhu14, Dos91, KL95, Lap86, LR83, Led92, Lot91, Rab94, Wu91]. d’évolution [Oga74]. d’évolutions [Szn80]. d’exclusion [Gui97]. d’existence [BDNT74]. d’expérience [GLS73a, GLS73b, JT76, Páz77]. d’expériences [Col78]. d’expérimentation [Rom69a]. d’explosion [Thi91]. d’hypothèses [GLS73a, LOR69]. Diaconis [FM09, NN20]. diagonal [BSC04, BNNS18, Le 02]. diagonales [Les87]. diagonals [BNNS18]. diagonaux [Le 02]. diameter [DW17, MS19]. diamonds [BK18]. dichotomy [CDLR17, Oka85]. d’identification [Sen90a]. d’identité [Gil65a, Gil65b]. Didier [DC02]. Diestel [BW06, BW05]. difféomorphismes [GH88, lJ87]. diffeomorphisms [Kun86]. difference [BBD12, Ouc05]. different [CH17, DE02, EZ19, RS18]. Differentiability [Bar00, Hen91, And09, LR19, MPR17, Ros10]. differentiable [BDMBP18]. différentiables [Yor77]. Differential [FV10, AB02, Bañ18, BOT14, BOZ16, BL05, CL21, CG96, CR20, Coq93, CKL14, Dab14, DG19, Dar21, DVW00, FP17, Fou13, FJR21, Gas21, GNSS94, Ham96, JKMP05, Kli09, KPP95, LM11, Mac87, MS96, NNRT09, OP89, PPR97, Pie09, RSX21, SSV03, ST20a, Sch89, Thi91, Wu10, XZ20, Zha13]. differentiation [Émi85]. différentiel [LM76, LM77]. différentielle [Sch89]. différentielles [Coq93, Dos77, DL78, Ham96, LM76, Szp78, SM79, Thi91]. diffusing [Gra85]. Diffusion [Ber15a, KOS19, Ami17, Ang16, Bas14, BT18b, BCV21, BDCKL20, BN90, BGJS18, Bla02, BT18c, BKRR71, Bru93, CI03, CL94, CL95, CR05a, Che06, CCC16, CT11, CPX17, CFS87, Del08, DD07, DZ05, FR03, GQ16, GCJ93, Gra92, Han04, HLOT^{+}19, Hof01, HMSH12, HMP20b, HQZ17, Jac98, KMM10, KP96, KMB92, LT18, Mar21, Nag05, Pel10, Pic64, Pig18, Pin90, PZ04, Rho09, Roy73, ST20a, Sin07, Sud21, Sue05, Var95a, Yan18, Zhe85, dZ81]. diffusion-aggregation [ST20a]. diffusion-limited [Ami17, BDCKL20]. Diffusions [Sch06, AWS21, AEFdR20, AGL94, Ass20, Aza89, BP85, Bai10, BCP19, BGM20, BR03, BS01, BDN20, Cas04, Cat86, CS02, CD16, DVW00, EZ19, EHK10, Feh19, FZ88, FT05, GP14, GG08, Gob02, GO13, HH81, HL05, Kob13, Kur10, LPP20, LS19b, LPP15, Léo86, LNS89, LR19, Löc02, LLL11, LR17, LLS11, LL08, MRTZ16, Mic17, PW94, PW02, Pet96, PS92, Pin14b, Rho09, Sar18, She03, She04, SZ97, Yor74a]. Diffusive [MST20, BGJS22, BSZ03, Lac17]. diffusivity [Set07]. digraphs [Coo19, Cos21]. dilatations [TV81]. dilaté [Mas93]. Dilation [GSS05, Wer80]. dilations [BP95]. dimension [Aïd20, Bat00, BS11, CN21, Dho72, Enr96, GLTZ17, Goz10, Heu98, JJK^{+}14, Jin14, Kel17, Mic92a, Sab01, Shi19, Spo17, SBM21, Van72, Zhu21]. dimensional [AKSS07, ADCKS15, ADG21, ABO16, Ami17, ASS90, AI99, AP05, BBCH21, Bar15, BT18b, BSC04, BL19, Bod97, BCT11, BG92, Bré09, BG16, CGPP13, CD19a, Che06, Che18, Cue18, CPV22, DR17c, DW11, Eld14, EF16, Enr96, FG12, FS17, FGS19, FFW05, Fou01, GCJ93, Gri04, IPP08, JLS13, Jin14, KK03, LS19b, Lem16, LY19, LLL11, LL08, MS70, Mar00, Mar21, MWRZ04, Mic17, PS09, Pet09a, PS13, PS19, Pin10, Pov98, DL15, Pru97b, STW00, TW18, Val12, Ver10, XZ09, Zhe12, vdHS14]. dimensionnels [TV81]. dimensions [AB97, BL15, Ber13, DE16, Gin80, Gin81, Gol18, GM18b, HMP20a, Kah87, Liu01, MM09, RS18, Röl13, Shi19, Tal92, Yar19, vdHS03, vdHdHS07]. dimer [dT07]. dimers [GMT17, Ken97]. d’intégrales [All78]. d’intégrandes [Thi81]. d’intersection [Wer93]. d’introduction [Jum80]. d’invariance [DP84b, Le 02, Mas93, Mét84]. Diophantine [BH08]. direct [LP97, Nic03]. Directed [Viv21, BA09, CN21, LZ22, Mej04, NS22, QR15, RSS16, Sep20, Var06]. direction [Sim07, Tou15]. directional [ABO16, ST11]. directions [MM14]. Dirichlet [AB02, Ben79, Coq93, DE02, Li19, LY19, LS20, MRZ98, OTT02, Ros01, ST11, Tou15]. discernable [Hae72]. disciplines [Fli81]. discontinues [Est92]. discontinuous [CG21, Est92, Mac87, MGY20, PSV17]. Discounted [Bax96]. discrepancy [CFP19]. discret [Der73, Pel90, SG02]. Discrete [LT20, Ané01, BR15, BL19, Bla03, CCP12, Cro08, CLM13, FP18a, Gob02, Joh17, Mar20, PZ21, Pel90, Pru97b, SG02, Var06]. discrete-time [CCP12]. discretely [CDM21, Gug12]. Discretisation [EH19]. discretization [GS18]. discrets [Lar72a, Lar72b, Lar72c]. discriminante [Bar81]. disintegration [Bra09]. disjointness [FPW95]. disk [GM19, HRV18]. disks [Le 19]. d’isomorphisme [Bla75]. d’isomorphismes [DCS74]. Disorder [BS11, GLT11, BMRS17, CI03, CP17, JL08, LP17, Zyg13]. disorder-induced [LP17]. dispersion [KR16, dlR02]. displacement [GP11, LS16]. dissipation [DD10]. dissipative [Maj19]. distance [BL14, BJ20, BDC68, CDM21, CG09, DFM16, DT03, Hei74a, Jac79, SZ16, Tes18]. distances [BFK20, Rio09]. distaux [Les87]. distinct [BFK20]. distorted [FGV16]. distributed [GLTZ17, Sta89]. Distribution [CG68, Gal70, And15, AFGJ16, BLS22, Ben20, Ber85a, Bor87, BGS10, Coh68, Coh72, Con65, CFS87, CAM98, DV13, GP01, GBR73, HLN06, Has90, HP00, HMP20b, KLO95, PZ21, Phu81, QR15, Rhe84, Tak70, WY16, Yan21a]. Distributional [AW18]. Distributions [Der90, AKSS07, BCV21, BNNS18, BFK20, BKS12b, Cas17, CLP14, CCM19, Che87, CKL14, CH90, DHHX20, Dep65, EHJ07, EG08, For68, GPdS16, Han04, HSW20, Haz93, HT04, JV07, Kol03, MP19, MSZ15, Nak20, PS13, Phu81, PT02, SZ17, Tor65, Der90]. d’Itô [Nua84]. d’itération [MM71]. Divergence [Nak20, CN05]. divisibility [BNNS18, Pie09]. divisible [AD15a, Phu81]. divisibles [Cup69, Tor77a, Tor77b, Tor78a, Tor78c, Tor79, Tor82]. DLA [AR16, BQR03, ELS10]. do [Sch83, BD96a]. d’occupation [Ako93, Kip87].

Doeblin [For78, Haz93, Thu81]. does [AS17, JdlR12]. Doléans [Thi91]. domain [BCV21, BCQ21, MP84, MR94a, Yur08]. domaine [TV81]. Domains [Gin80, Gin81, Per92, BBC05, BEK19, CF17, DF87b, Feh19, GMV93, LR19, DL15, SWW17]. dominance [Gil65b]. Dominant [Din92]. dominantes [Din92]. domination [AN09, Len77, LS06]. domino [BK18]. données [Mac78]. Donsker [Del87, MØP04]. dont [Con65, Fis73]. Doob [Dau73, Ros99]. d’opérateurs [Her94]. d’ordre [Ben79]. d’Orlicz [DC73]. d’Ornstein [Fer90]. double [Dev89, DMY91, DMGM18, FPW95, KW20, PX05]. double-well [DMGM18]. Doubly [BRR17, CDH00, SWW17]. downwards [Ber95]. dozen [DF87a]. drainage [CV14]. drift [ANV00, BDMBP18, Cas04, DS90, DR17c, Gal96, KR16, LL08, MGY20, Pin14b, PRS11, SY05, dR17b]. driftless [AI99]. drifts [EZ19, FMM92, Hol12, KHT12, Zha13]. driven [BLNR11, BOT14, BOZ16, BL08b, CL21, CG19, DNT12, FRS22, FP17, FNS00, Fou13, FJR21, FV10, GS18, HP11, HLT16, HM16, HLN17, JKMP05, KP19, KPP95, MS96, Zha13, ZBH17]. driving [Mac87]. droite [CR79, CR80, Jan70, Mic09]. Dual [DKGT16]. dualité [Pic96, RR70b]. Duality [VAG86, Zha10, FRS22, Pic96]. duals [Cue18]. Dubins [Loo81]. dummies [BvdHK18]. d’un [Ant82, Bar81, Ber91a, Ber68, Ber89, Ber91b, Bru93, Cai82, Cha70, CR79, CF87, CSC90, Deg82, Dep99, DR82, DST90, DST91, Duh97, Enr96, Gas90, GL89, Kip87, KMB92, LS91, Mac78, Mal65, Mas93, Mél78, MR89, Ngh70, Pel72, Pel90, Rau83, Rau92, Rec86, Rou97, SG02, Tou96, Val78, Vig03]. d’une [BN90, BB78, Che76, CG92, Deh74, DL84, Fis73, FLP06, Gir70, Gra88b, Hai87, Jai80, Let72, Lon69, Loo77a, LR96, Lus78, Nev69, Rou97, Roy73, SP75, TTL69, Tor78b, Tou69, Val89, dZ81]. Dunkl [Cha12, Chy08]. duplicated [LPP15]. dust [Haa04]. Dvoretzky [BLM99]. dynamic [And14, EdHM14, GL20, Gui00]. Dynamical [FT12, HPS97, PSS09, BP95, BMP94a, BMP94b, BF95, CDLR17, DNN73, Dos80, FFV17, Gor17, GSS05, IR98, Kom92, LT18, MZ15, PT21, ST20b]. dynamics [AI05, Bay15, BL15, Ber12a, BCC02, BJM10, CMR02, Cer96, De 02, GP20, GJ19, GLP20, HRSS21, KMS97, KL05b, LP17, Sab01, Set07, Yan21b]. dynamique [Bou88b, Bru71, LR83]. dynamiques [Con72b, Dos80, KMS97]. Dyson [RS18].

eat [Pin10]. Eaton [SJ14]. échangeables [BK95]. échantillon [Gar79, Mas93]. échelle [Rou97]. Économie [Lic70]. économique [Lic70]. Edgar [Egg80]. Edgar-type [Egg80]. Edge [ASS90, LS19a, CK20, LWZ16]. edges [AvdHH19]. Edgeworth [FL21, Fer15, GK17]. Editorial [Ano05d, Ano16a, Ano16b, Ano17a, Ano97, Ano99, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05b, Ano05c, Ano05e, Ano05f, Ano05g, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f]. EDP [GP01]. EDS [Sch89]. EDSR [GP01].

Edwards [AD14, LZ22]. effect [BOZ16, Gil66a, Enr96]. effectif [Lot91]. Effective [Abe15, BP04, CI03, KP96, Mal65, Spo17]. Effet [Enr96]. Efficient [KZ21, KP12, CG96]. eigenfunctions [KM02]. eigenvalue [BCHJ21, HP00, LSX21, Yur08]. Eigenvalues [DE05, AD15b, AAP09, BY08, BG85, CDMF12, Gus05, HLN06, Lou68]. Eigenvectors [Ben20]. Eikonal [MRTZ16]. Einstein [Dia05, AHOZ13, FGS19, Lou05]. elastic [Leg89, Pei92]. élastique [Leg89]. élastiques [Pei92]. electric [Dep99]. électrique [Dep99]. élémentaire [Duh97]. elementary [Duh97]. elements [Ahm65, Har01, Har01, Ahm65, Ehr86]. elliptic [ADG21, AC99, BL95, Mar00]. elliptique [BKRR71]. Elworthy [Bañ18]. embedding [AEFdR20, CW13, De 18, LTT18]. Embeddings [CG79, Eld16, Kol92]. Emery [CDP09, Joh17]. Empirical [LW20a, And15, Aud04, BFG15, BL14, CD01, CLB98, Cat12, Deh00, DDP87, DSS13, DGW06, DMR95, DLS02, DT03, ES02, GZ19, HLN06, HP00, Kol09, LM13, Lou00, Mas88, Mas86, Rio02]. empirique [CD01]. empiriques [Rio02]. end [DGKP21]. endomorphismes [Att95, Con72b, DL96]. endomorphisms [Att95, DL96]. endpoint [QR15]. ends [Mal96]. énergie [Mic92a, Mic92b]. Energy [BK04a, BK04b, Ass11, BL20, Che15, Che08, FS89, GS87, HRSS21, MW02, Mic92a, Mic92b, Nag05]. engendré [Bru73]. engendrées [Gal76]. enhanced [FV07]. Ensemble [NR17, BMW21, Led09]. Ensembles [FK75, ACLS19, BCR15, Car81, DE05, KK19, LS19a]. entier [Has90]. entrance [AR06]. entre [Bia86, DC73, DCS74, Dos77, Fou95, Gil66b, Len77, Mal65]. entropic [ACLZ20]. Entropie [Con72a, MS83a, Mic99, Mou96]. Entropy [BKYY10, CHV20, Mkr14, BS01, CDP09, DCG90, DHJ^{+}21, Erb14, GBR73, GRS^{+}18, HU09, JdlR12, JV07, KT97, Leh13, Mic99, Mou96]. entry [Yan21a]. envelope [Vig03]. Enveloppe [Cai82, Mas93, Vig03]. Enveloppes [Dav02]. environment [ABKV14, Aid10, AB18, AFJV15, BGRS14, BB14a, BC17, Bau13, BQR03, BT13, Ber12a, BMP94a, BMP94b, BSZ03, Che06, CG21, CP12, DGR18, EdHM14, FHSX21, GPS10, GP11, GKV03, Guo14, IR98, JP17, Liu96, Mej04, Mou11b, Nak15, NS95, OR18, Pet09a, PS13, Pin10, RAS09, RAS11b, ST11, Sch06, She03, She04, Sin07, Tou15, Yil10, Zer00, BMP94b]. environments [CCKW21, DK08, GM18a, MWRZ04, Sim07, Viv21]. equality [Har05, Pra07]. equation [ANV00, AR08, ABK18, BDT11, BRR17, Bas14, BL08a, BDP95, Bré09, BBD12, CG96, CD19a, CKK17, Che17a, CJKS14, DDT04, Dab14, DT21, DVNZ20, DV05, Dey20, DMGM18, Erb10, FK10, Fou01, Gou07, HHK06, HT04, Hos18, HW22, Kyp04, MRTZ16, MN18, Mar21, Pic64, Sch89, TW18, Web10, Bal74, Mac78, Pic64, Sch89]. equations [AKSS07, AB02, AB18, Bañ18, BRT03, BOT14, BOZ16, BL05, CL21, CG96, CR20, CK17, Coq93, CPX17, CKL14, DG19, DGG21, Dav75, DVW00, DW11, DZ05, FP17, Fou13, FJR21, FV10, Gas21, GNSS94, GB16, Ham96, HQZ17, IKPY18, JKMP05, JLT21, Kal20, KR16, Kli09, Kuk13, KP19, KPP95, LM11, Mac87, MS96, NNRT09, OP89, Oda06, PPR97, Pel10, Pie09, RSX21, SSV03, ST20a, SvG11, Sud21, Thi91, Wu10, XZ20, XZ09, Zha13, dS03, Coq93, Dos77, LM76, SM79, Ham96]. equicontinuity [BT00]. équidistribuées [Bir84]. Equidistribution [Bor21, Sta89]. equilateral [Val89]. équilatère [Val89]. Équilibre [lJ87]. Equilibrium [BEM03, BGJS22, CG09, DJO19, DMGM18, EFG^{+}20, FPV88, FRS22, FP17, JL08, KLO95, Val06, Var95a]. équitabilité [Lem73a, Lem73b]. Equivalence [AMS95, SSV03, Bas08, FX19, MRW20, Yos01, Szp78]. Equivalent [Oka85, LL08, Riv12, dlR98, dlR98]. Equivalent-singular [Oka85]. erased [LPS03, Shi19]. Erdos [BDS87, HL20, KLS21, MR21, ST20b, SZ16]. Ergodic [Aar78, Boi98, DK79, Lin77, MM13, Smy75, AW18, AWS21, ABK15, Ass98, AD97, Ast76, Bew71, CD16, CD12, CL09, DL96, Eng09, Fit98, FHS19, Fur97, GS97, Gob02, Gri77, Her05, Her08b, LPWR94, Len04, LLL11, MR86, MS96, MW05, Oka85, Rau09, Ros89, Rot90, Sch00, Ste89, Wu99]. ergodicité [Mic92a, dlR98]. Ergodicity [AAD19, HP11, HLT16, LR17, Oda06, XZ20, Ben20, Bud03, FP08, Her08a, HL05, LMW21, Lin75, Mic92a, FMM92, dlR98]. ergodique [Bru73, Cha70, DL96, Émi81b, MO83, Nev83, Pet79, Sch72a, Sch00]. ergodiques [Ass87, BH81, CR80, CM70, GS97, Hac82, Her05, Her08b, Les87, Reg70, Rot90]. Errata [Ano66, CW20, Rou79, Ano66]. Erratum [BDEG89, BB02, BW06, CL95, Cha71, Gin81, Mat98b, Mat98a, de 95, Cha71]. error [HMSH12]. errors [BG20, BT08]. escape [BT18a, Tho13]. ESD [Woo16]. espace [Att95, Bal76, Coq93, Dep65, Der90, Fou84, Gui79, Hei74b, Hil78, Jac73, Lon81, May68, Ngh70, Pay67, PPH75, Pic96, Rau92, Sun75, TTL74, Tor65, Tor78a, Tor78b, Tor82, Yor74a, Yor74b]. espace-temps [May68]. Espaces [For73, Ahm65, Alt87, Arn93b, Ass87, BCK66, BDC69, BH80a, BH81, Con72a, DC73, DCS74, Geo78, GP74b, GL80, Hei74a, Hei83, Led83, Mar88, Mét84, Ouv73, RF64, Sch72b, Tor77b]. espérance [BH80b, BH80a, DR82, Nev76, Val80]. Espérances [Rou87a, Thi81, Ngh70]. Esseen [Gou05, KRT16, Tol00]. est [Led78]. Estimate [Wu04, BM89, DW11, Ish97, SG02]. Estimates [Zai02, AL12, AB97, BCP19, BH97, BMW21, BRZ19, BPTZ18, BW05, BW06, CCEL13, Cat91b, Cat91a, Cat03, Che17b, DG19, DP83, FL20, HMP20b, HQZ17, HK09, MSC19, NZ21, Pig18, R.83, Tro96, Heu98]. estimateur [AJ76b, Ber91a, Bru93, DST90, DST91, MR89, Tou96]. estimateurs [DP84a]. Estimating [BB14b, CG19, FF18, Sar21, ACRC17, HPT02, Hof01, MW07]. Estimation [BL08b, Che76, Con94, DR82, Gas90, KMB92, Lac73, Sar14, SG02, You88, Alq08, ABCF21, ADGP13, BT18b, BG20, BLR11, BLR16, BEK19, Cas07, CGC10, CGG11, CPP71, Dev89, GNS13, GCJ93, GJTZ08b, GJTZ08a, HMSH12, IL97, JM93, KZ21, KP12, KM02, Lac07, Ler12, LL08, MW05, RSH19, Sar21, Spo17, Tou69, TV98, Con94]. Estimations [Heu98, Lap86, Fer15, FK69, Páz77]. Estimator [BGH14, Ber91a, BLM99, Bla99, Bru93, DY11, DST90, DST91, GGS15, MR89, Tou96]. estimators [Aud04, Bir06, BT08, GvdLW95, GG01, GG02, MNR21, MH18, PD14]. Eta [BNNS18]. Eta-diagonal [BNNS18]. étalés [GL80]. étapes [Con65]. état [KMS97]. eternal [Bas08]. étude [DCS74, Gol89, Vig03, BB78, CGR79, CF87, Dep65, FLP06, GBB69b, Hai87, Her94, LR96, Mey91, Mic92b, MR89, Pic64, Yor75, Yor77]. Euclidean [ADG21, BH08, BT16, GJN20, Pin85, SSS^{+}21]. Euler [BGRS14, FP08, JKMP05, MGY20]. Evaluating [SJ14]. Eve [Lab14]. even [Che14]. even-spin [Che14]. event [VW15, Yan21a]. event-based [VW15]. events [AD97]. everywhere [CL13b]. évolutif [Mél78]. Evolution [Mis16, BRT03, Mal65, Mic92b, Mic92b, Jai80, Mal65]. Evolutionary [PT02]. évolutions [CGR79]. evolving [DHMP17, FVY06, vEEIK12]. Exact [BDS87, Vay03, BL02, DL96, HKPG16, Led78, LL15]. exacts [DL96]. example [McG89]. examples [CP17, LR14, PS92, Tes06]. exceedances [JD18]. Exceptional [AGJ^{+}20]. excess [Ass10]. excessifs [Alv73]. exchange [ANT17, Mal96]. Exchangeable [ANS06, Aus15, BQZ20, BK95, CW13, CKL14, Kas20, Lim12, Tra02]. Exchanging [vW83]. excité [BL79, Bro83]. Excited [Hol12, KP17, KOS16, ABO16, KM11]. exclusion [AFJV15, BT18a, BGJS22, BESS20, BW19b, DP83, DPS89, FGL95, FFV00, FGM09, FGN13, Gui97, JL06, JL08, JLT21, Kip87, Kom98, Lac17, Lou05, Rez02, Sue05, Val12, Var95a]. excursion [Bia86, HS99, LS14, PPR18, RVY09, Wat87]. Excursions [CT11, EO22, Jac10, LTT18]. exemple [EV03]. Exemples [Bou83, Tes06]. exhaustive [Lus78]. Exhaustivité [Pet70, LPPV69]. Existence [BE95, CK77, CFP19, FJR21, GL20, GO13, Len04, Sep20, Yor74a, BCR19, BK95, Bro20, DR17c, Der00, DL78, Haz93, LM77, Lus78, Pel72, Sło93, dR17b]. Exit [DW15, AD13, Feh19, KM99, iSW04, iSW05]. expanding [Gou05]. expansion [BvdHK18, CHV20, DY11, Fer15, Mas93, MØP04, Sla22, Uek99, vdHdHS07]. expansions [FL21, GK17, Jia12, Mik88, NNRT09]. expansive [EG08]. expectation [PP98, Son13]. Expériences [Flo82, MPPL71]. explanation [Yor91]. explication [Yor91]. Explicit [KMM10]. Exploitation [Tou69]. explosion [KRS04]. exponent [ABL08, DV13, EdHM14, Hen91, Lac12a, Lac12b, Le 89, Mej04, Per92, Pov98, Tro96, vdBvE22]. Exponential [AZ97, BO11, BR08, CR05b, CHSX15, Dal20, FGL95, LMW21, Véc19, AMS95, CE20, CG21, Con65, ERS16, Est92, Haf20, LLS11, Mic99, Pie09, Riv12]. Exponentially [GLP20, Gar99]. exponentials [BHS12]. Exponentielle [Est92, Con65, Mic99, Roy80]. exponents [AMZ21, LSW02, MS96, Tia17, Zer00]. exposant [Hen91, Le 89]. exposants [Bou88a]. extended [BGJS22, PS09]. Extension [Bew71, BCR20, CF09, DP75, Gui90, Rot90, Wüt04]. Extensions [Con72b, BBC78, Bou80, Gue99, Lem88, LY19]. external [EM15, FGS18]. Extinction [MP00, ERS16, Eva91, GH10]. extrêmes [Hai87]. extrapolation [LPP15]. Extremal [Eld14, Par05, ABK15, FF18, FM19, Set01]. extrémales [Deh74, Hai81]. extrémaux [Ehr86]. Extreme [BCHJ21, Coh68, FFV17]. extremes [DMS93].

faces [IY20]. facilitated [BESS20]. facteurs [Jum80]. factor [BV17, Tal95]. factorielle [JT76]. factorisation [KRS18, LR86]. Factorising [McG89]. factors [CD19b, GJ11, Hol17]. faible [Eme85, Fis76, Mét84, Tes06]. faiblement [Bru76]. fails [Lac97a]. falling [Bra09]. famille [Lot91]. familles [GMP78, Le 85]. family [EH03, Eva87, Lot91, Sud21]. fan [FK95, FGM09]. Fast [BF95, CD18, CX21, HQZ17, RSX21]. Fatou [BE82]. favoring [Dal20]. feature [Alq08]. Feller [Ass20, DPSZ16, HM18, HK09]. fellerien [Hae72]. fermé [Del64]. fermés [Van72]. Fermi [GR11]. fermions [LW20b]. Fernique [CCS92, Tal96]. ferromagnet [AK21]. ferromagnetic [Che14]. few [HR17]. Feynman [AD08, CDG11, CZ02, KKK85, Yin96]. fiabilité [CTR95]. Fibonacci [JRdlR10]. field [AS18, AK21, AZ99, Bañ18, BA09, BL19, CGN16, CGL21, EM15, FGS18, GP20, GLP20, HRSS21, Jac18, KM17, Mad15, ÖE19, PZ21, PW17, RDL20, Sep19]. fields [BT15, BZ12, Che18, CH17, DE02, DLS02, DLZ20, EV03, Gar99, HL22, MM14, Mal11, MV20, You88]. file [Gra84, Jai80]. files [Fli81, Pel79, dZ81]. Fill [HP17]. filling [DCKY14, FGL95]. filter [Bud03]. filtered [DS06]. filtering [AZ97, DG01, Dos91, FPW95, Rab94]. filtrage [Dos91, Rab94, Szp78, SM79]. filtration [DTMS98, Mal90]. Filtrations [Mal90, BPTZ18, Éme05]. financial [AI05]. fine [AKDR66, CL84, CM10, AKDR66]. finely [ALT99]. Finetti [DF87a]. Finetti-style [DF87a]. fini [Mal65]. finie [BDC68, Dho72, Jac79, Jeu93, Van72]. finies [Cal81, Laf72]. finiment [Bru73]. finitarily [MW07]. finitary [Hol17]. Finite [ACLS19, AI05, CI03, Hoc19, KL02, AKSS07, Bré09, BZ12, CCC16, Jac79, Jeu93, Lam00, LW20b, Mal65, MS19, PRS13, Poi13, Rou21, Yan13, BDC68]. finite-diameter [MS19]. finite-dimensional [AKSS07]. finite-distance [Jac79, BDC68]. Finite-time [Hoc19]. finitely [BK04a, Tal92]. fins [Mal96]. fireproof [Mar16]. Fires [Ber12b, ADCKS15, Mar16]. First [AD00, BW19a, BR08, BG85, CS21, CT16, DHHX20, DS90, DT20, Fer15, Lac91, Nak20, RT10, Tes18]. first-order [Fer15]. first-passage [DHHX20, Nak20]. fish [Bou00]. Fisher [BS01, CCC16, HHK06, NN16]. fit [BMP09, GR08, MS17, Ste74]. fixation [RT18]. fixe [FSP71, RR70b]. fixed [AR08, Woo16]. FKG [LS06]. Fleming [BEV13, LZ15, VW15, AFGJ16, CDGR20, KL20]. Fleming-Viot [CDGR20]. flip [AS17]. flipping [CD18]. flocks [BRT14]. Flot [Del77a, Led78, dlRd03]. Flots [Rai99, Bla75, Con72a, Han71, Le 02]. flow [ANT17, BFG15, Cas04, CP11, Del77a, Dep99, Erb10, ESYY12, FGHV16, Hoc19, LMX09, dlRd03, vdB07]. flows [BW15, BL05, Eld16, ES79, Erb14, ES17, HKPG16, LR14, Kre76, Kun86, Lab14, MT20, Pro86, Rai99, Rho08, RT10]. Fluctuation [BC20, Wüt98, AD00, BKS12a, DD10, FS17, Nak20]. Fluctuations [AR16, CD01, CK22, CQR07, Con18, IKM20, KLS21, Kip87, KL05a, AD15b, BW19a, Bal03, Bal04, BEM03, BGJS22, BDP95, BW19b, BDK19, CN21, DHHX20, EFG^{+}20, FPV88, GMT17, Gus05, JLS13, JM98, LT21, Liu18, LZ22, NW18, SZ17, CD01]. fluids [ACLZ20, CX21, Fla02]. Flux [Dep99]. Fokker [DMGM18]. Földes [Yor93]. Föllmer [AJ76c, AS92, ELS20]. follows [Con65]. fonction [Del87, Lic69, Nev69, PPH75]. fonctionnel [EV03, Lon69, Mai78, TTL69]. fonctionnelle [JKM82, LR96]. Fonctionnelles [Jac71b, Ber89, Bre83, Dav74, Fou95, Gav75, LR93, Páz77, RR70a]. fonctionnels [Tou83]. Fonctions [Bas69, Bas89, Del64, Pay67, Bec89, Ber70, Ber88, BLT69, BDC69, Cup69, DS84, Fer90, For73, Lac73, Bas69]. fondamental [Cau69]. fondamentaux [Bec74]. force [Zha22]. forced [AKSS07, FGHV16, Ors82]. forest [ADCKS15, Hut20, RSS16]. forests [KKW15, LPS20, SSS^{+}21]. form [Bas82]. formation [Haa04]. Forme [Coq93]. forms [Cro18, Dar21, HLNV13, KL16, LS20, OTT02]. formula [AD15a, Bañ18, FG12, GNRV05, KF98, KKK85, Led09, Leh13, OP89, Yin96]. formulae [AD08, BBL14]. Formulas [BK93, Ané01, Dah17, Dos80, Pic96]. Formule [GP74b, Pel79, BY87, Nua84]. Formules [Bou84, Pic96, Dos80]. forte [Ngh70, RE79]. fortement [Ben79, Her08b]. fortes [Led81]. Fortuin [GMS19]. Forward [MW05, Wu99]. Forward-backward [Wu99]. Fourier [BT15, BP06, GBR73, Hey68]. fourth [DK19, NN16]. fractal [Bro20, LT06]. fractals [Ber13]. Fractional [BM09, DT21, AWX08, BOT14, BOZ16, BGJS18, BL08b, CCM03, CK17, Che17a, Che20, CN05, CGPP06, DVNZ20, DNT12, Dey20, FG92, FGR09, FP17, GNRV05, GN09, HP11, JN19, Jin14, LS05a, MSS06, NNRT09, NNT10, PT08, Sud21, Wan07]. fractions [CT11]. fragment [Bra09]. fragmentation [BCK21, Bra09, HKK10, KK14, KPR18, Shi20]. fragmentation-coalescence [KPR18]. fragmentations [Ber02, Duc20, GH10, Haa04]. framework [DT03, TV98]. Frèchet [Phu81]. Free [BL20, BS01, Fun99, AS18, ATV20, Bel06, BG10, BDCH16, BL19, Bor08, Che18, CH17, Dab14, DvdHH13, DMGM18, DGKP21, GP20, Goz10, GM19, HS17, HSW20, HU09, Jac18, KLL21, LW20b, Mic92a, Mic92b, PZ21, PW17, Sep19]. freedom [MH18]. French [AJ76b, AJ76a, Ahm65, Ako93, All78, Alt87, Alv73, ACTR85, Ano66, Ano77, AS92, AS94, Ant82, Arn93a, Arn93b, Ass87, Att95, Aza89, AKDR66, ADR67, AJ76c, Bab88, Bab91, Bal74, Bal76, BBC78, Bal81, Bar81, Bas69, Bas89, Bec74, Bec89, Ben79, Ben67b, Ben67a, Ben67c, Ber91a, BD92, Ber70, BN90, Ber68, Ber88, Ber89, Ber91b, BF92, Bia86, BY87, Bir84, Bla02, BLT69, Bla73, Bla75, Bod70, Bod71, BKRR71, Bos75, BB78, Bou83, Bou88a, BG92, Bou80, Bou84, Bou00, Bou07, Bou88b, Bra73, BD96a, BCK66, BDC67, BDC68, BDC69, Bre83, Bre87, BK95, Bre02, BL79, Bro83, Bro75, BL12, BH80b, BH80a, BH81, BH85, Bru93, Bru73, BDNT74, BR74, BR79, Bru76, Bru71, CD01, Cab79, CM72, Cai82, CL84, Cal81]. French [Car80, Car81, Cat86, Cau65, Cau69, Cha70, Cha71, Cha86a, Cha86b, Che76, CTR95, CK77, CGR79, CR79, CR80, Col78, Con94, Con65, CM70, Con72b, Con72a, CG92, Coq93, CF87, CSC90, CPP71, CR78, Cun03, Cup69, DC73, DCS74, DCG90, DC02, Dau73, DP75, Dav74, DV98, Dav02, Deg79a, Deg79b, Deg82, Deh74, Del77a, Del87, DTMS98, DAP76, Del77b, Del64, Dep65, Dep99, Der90, DL96, Der73, Der76, DL84, DP84a, DP84b, Dho72, Die81, Die76, Die78, Die82, Din92, Dos77, DL78, Dos80, Dos91, DLP86, Doz83, Dub75, DR82, DR69, DST90, DST91, Duh97, DS84, Ehr86, ERR71, EV03, Eme85, Émi81a, Émi81b, Enr96, Est92, FK93, FS03, Fer84a, Fer85, Fer84b, Fer90, Fis73, Fis76, Fli75, Fli81, FZ88]. French [Flo82, For68, FK69, For73, FK75, For78, Fou00, FLP06, Fou84, Fou95, Fou05, FlJ84, Fra91, FSP71, Fri67, Gal96, Gal76, GS97, GMP78, Gar78, Gar79, Gas90, GMP02, GP01, Gav75, Geo78, Gho75, Gil65a, Gil65b, Gil66a, Gil66b, Gil66c, GBB69a, GBB69b, Gir70, GP74a, Gol89, GP74b, Gra88b, Gra84, Gre08, Gru66a, Gru66b, Gue99, GL80, Gui79, Gui97, GH88, Gui90, GLS73a, GLS73b, GL89, Hac82, Hae72, Hai81, Hai87, Ham96, Han71, Han73, Han72, Har01, Has90, Hei71, Hei74a, Hei74b, Hei78, Hei83, Hei90, Hen82, Hen91, Her94, Her05, Her08b, Heu79, Heu98, Hey68, Hil73, Hil78, Hil77, Hög74a, Hög74b, JT76, Jac79, Jac82, Jac71b, Jac71c, Jac71a, Jac73, Jac74, JKM82, Jac90, Jai80, Jan70, Jeu93, Jum80, Kah87, KRR70, Kip74]. French [Kip87, KL95, KMS97, KMB92, Lac73, Lac91, Lac97b, LS91, Laf72, LP90, Lan82, Lap86, LR83, Lar72a, Lar72b, Lar72c, Lar93, LOR69, Le 02, Le 85, Le 89, Léa85, Léa88, Led81, Led83, Led92, Led70, Led78, Leg89, Lem73a, Lem73b, Len77, Léo86, LM76, LM77, LNS89, Les87, Let72, Lic69, Lic70, LPPV69, Lon69, Lon81, Loo77a, Loo77b, Loo77c, Loo78, Loo81, LR93, LR96, Lot85, Lot91, Lou68, Lus78, MM71, Mac78, Mai78, Mal65, Mal90, Mal96, MS83a, Mar88, MV69, MPPL71, MPL73, Mas93, Mat93, May68, Mél78, Mén78, Mer83, Mét72, Mét84, Mey91, Mic92a, Mic92b, Mic99, Mic09, MN83, MR89, Mok90, MO83, Mou96, Mus06, Nev69, Nev72a, Nev72b, Nev76, Nev83, Nev86, Ngh70, Nua84, Oga74, Ouv73, Pay67]. French [Páz77, Pei92, Pel72, Pel79, Pel90, Pet70, Pet79, Pet96, PPH75, Phi71, Pia95, Pic64, Pic65, Pic94, Pic96, Pis84, Pra92, Pri68, Rab94, Rai99, RLT81, Rao70, Rau83, Rau92, Ray68, Rec86, Reg70, RR70a, RR70b, Rei82, RF64, Rev83, Rio02, Rob87, RZ93, Rom69a, Rom69b, Rot90, Rou78, Rou79, Rou87a, Rou87b, Rou72, RE79, Rou80, Rou86, Rou97, Roy80, Roy73, Roy74, Ruc80, Saa88, Saa90, Saa95, SL95, SP75, SP76, Sch72a, Sch00, Sch72b, Sch85, Sch89, Sen90a, Sen90b, Sie84, SG02, Str88, Str90, Sun75, Sun87, Szn80, Szp78, SM79, Tes06, Thi81, Thi91, TTL74, TTL68, TTL69, Tor65, Tor66, Tor69, Tor71, Tor76, Tor77a, Tor77b, Tor78a, Tor78b, Tor78c, Tor79, Tor81, Tor82, Tou83, Tou96]. French [Tou69, TV81, Val78, Val80, Val89, Van72, Vig03, Web80, Wei84, Wer93, Wu91, Yor74a, Yor74b, Yor75, Yor77, Yor85, Yor91, dZ81, dlR98, dlR02, dlRd03, lJ87]. fréquences [Mal65]. frequencies [Mal65]. frog [JJ18]. frogs [KM20]. front [CQR07]. frontière [Cat86, Der73, dZ81]. frontières [Wei84]. Fubini [DMY91]. Fuchsian [Lal98]. fugacity [GM20]. full [BW05, BW06, FR14, LZ22]. fully [FP08]. function [AS17, BDEG88, BDEG89, BLR16, BM89, CN21, GvdLW95, GM20, GC03, KLS21, Kel17, KW20, LR86, MØP04, Nev69, SY11, de 95, dlP94]. Functional [CRR98, JN19, Kas20, WW16, AKU17, AKU20, BRT03, BL02, CG04, DR00, Deh00, DMR94, Eic01, EV03, GS87, GK93, GN02, KM17, Leh13, LR96, LT06, MS96, RAS09]. functionals [BO11, Bax96, Ber89, BCR15, Cha91, Dev16, DGW06, ERS15, FG12, Fer15, Fou95, FL20, Gol95, Gui02, GG79, KHT12, KZ21, KRT16, LR93, LL08, Mou11b, Pie09, Riv12, SY05, SY13, Smo83, Véc19, Wu99]. functions [Aar78, BB14b, Bas89, BO11, Bec89, BA09, BE95, BDC69, CKS87, CG19, CGPP06, DGM10, Del64, Fer90, Gal70, GBR73, Gau20, Glo91, GLM18, Har05, HK09, KR16, LLO06, MR86, Mim14, Pay67, RW17, Ras11a, dT16]. Further [HHK06]. Fusion [Def16]. fuzzy [Häg03].

G [CG68, Coh68, Coh72, Fli81, Coh68, Fli81]. G/G/1 [Fli81]. G/M/1 [Coh68]. Galton [ADH12, Aid10, Aïd20, AP98, AR06, AC18, AHOZ13, AFGJ16, BC17, BRZ19, Che97, CG14, FW07, FW09, Kor15, Lin17, Liu01, Loo77a, Lot91, MSZ15, Mie08, Nev86, Riz15, Stu19, dR17a]. games [CGL21, COBZ21, De 99]. gamma [Has90]. Gangolli [AD15a, NN20]. gap [BY16, BCC02, BQZ20, Cos21, CFP19, FHL12, GGL15, GJ19, GL16, LPY02, LM09, Mat97, Mat98b, TW18, Wu04]. gaps [LLS11]. gas [BL08a, BCC02, Fun99, Nag05, Wu04]. gases [BEM03, BT18c]. gaskets [BH97]. gauche [Cun03, Fer85]. Gauss [Bor77, Phi71]. Gaussian [AS18, AC99, BGH14, BP06, BM00, Ber85a, Ber85b, BL19, CDM21, Cas04, Cha91, CP17, Che18, Che19, Che20, CH17, CM08, CN21, Dav02, DE02, Eld16, Fou00, FV07, FV10, FR14, GP20, GK93, Gor17, Gua77, Gui02, Gus05, GL89, HMP20b, Ist92, IL97, Jac18, Jeu93, KK75, KZ21, KP19, LMMR18, Led09, Lif94, Mad15, MV20, NW18, PZ21, Pet09a, PW17, RSH19, Rhe84, Rou97, SO75, Sep19, Tal88, TTL68, Ver10]. gaussien [Ant82, BL79, Bro83, Gar79, GL89, Lap86, Rou97]. gaussienne [Bou88b, Fou00]. gaussiennes [Ehr86, Fer84b, Le 85]. gaussiens [Ben79, Dav02, Die81, Jeu93, TTL68, Web80]. gene [ACA02, Mal65, Mal65]. genealogical [BD16]. Genealogies [Lim12]. Genealogy [Lab14]. General [BC17, Ber83, BPTZ18, DHHX20, GS18, Her77, HW22, KPP95, KLMSR12, LW20a, Nak20, Num78, PW17, Rau04, Bod70, Bod71, Led78, SM79]. généralisation [Jan70, TV81]. généralisations [Pic64]. généralisé [Ber91a]. Generalised [BK04a, AD15a]. généralisés [Cau65]. generalization [CV14, SJ14]. generalizations [Pic64]. Generalized [BE82, BK04b, Kol92, BP15, Ber91a, El 00, GNRV05, JRdlR10, KR16, Lia07, NZ13, OP89, Tro96, Cau65]. generated [AL12, ELS10, FFV17]. Générateurs [Jac71c]. generating [de 95, dlP94]. generations [AC18]. generators [BP85, BW15, DHR11, GSS05, HKM14]. generic [Kor15, dlRd03]. générique [dlRd03]. genes [Pic64, ACA02, Pic64]. genetic [DG01, Gil65a, Gil65b]. genetics [Eva97, Gil66c]. génétique [Gil65a, Gil65b, Gil66c]. genotypic [Gil66b]. génotypique [GBB69a]. génotypiques [Gil66b]. genus [Ray15]. Geodesic [BGKL17]. geodesicity [Tes18]. Geodesics [Bet16, Wüt99, Bud20, Sep20]. Geometric [HK15, KMS14, BESY08, Ben67c, BC19, ERS15, FMP14, Gan13, SY13]. geometrically [Her05]. Géométries [Laf72]. géométrique [Ben67c, Ruc80]. géométriquement [Her05]. Geometry [GH12, BHS21, GMSS18]. Giant [ČTW11, Gan13, Was15]. Gibbs [AKKR01, BL03, DKP16, EG08, GZ19, Lam21, Pai18, RZ93, SY13, Wu91, vEEIK12]. Gibbsian [Gar99, Häg03, You88]. gigue [Mac78]. Gillespie [Han04]. Ginibre [LWZ16]. Ginzburg [BDP95, Hos18, Oda06]. Girsanov [CZ02, FI93, Ouv73]. given [HLN06, Riz15]. glass [AD14, CH06, GJ19]. glasses [De 02]. Glauber [Ber12a, BCC02, De 02, KL05b]. Glauber-type [BCC02]. Global [Ahn20, DHJ^{+}21, DLN21, Hos18, ERS16, FK10]. globale [GLS73a]. globally [Maj19]. goes [Bro03]. goodness [BMP09, GR08, MS17, Ste74]. goodness-of-fit [BMP09, GR08, MS17, Ste74]. governed [DLZ20]. gPAM [CF18]. grad [RT97]. Gradient [BGM20, Erb14, HQZ17, DDT04, Erb10, GP14, KLO95, DL15]. Gram [HLN06]. grand [Hen91, Le 89]. Grandes [KL95, LR83, Lot91, Wu91, CD01, Dos91, Lap86, Rab94]. grands [Alt87, DST90, Gal96, Léo86, RE79]. graph [AKM21, Ass20, ČTW11, Gan13, HL20, Kel17, RN21, ST20b, Tim14, Was15]. Graphes [Pel90, Ray68]. Graphs [CH67, AGJ^{+}20, BV17, BHLP19, BW05, BW06, CT18, CS21, CF17, CX21, CY07, Che17b, CDL20, DP11, GP20, Hua19, KLS21, KMP14, KLL21, LPS20, MR21, Ray68, Tel08, Pel90]. Grassmannian [AMD17]. Gravitation [BBD19]. gravity [AG21, Ber15a, GM21, HP21, HRV18, MS19]. greater [Heu98]. greatest [Le 89]. Green [BW06, BW05, KW20, Ras11a]. grid [KW20]. Gronwall [HHM21]. Gross [Hof15]. ground [AD14, MW02]. group [BP06, Bla03, Bro03, Cun03, Fog75, Gre08, Kun86, Leg89, NZ13, Pei92, Tor65, Tor66]. Groupe [Let72, Bal76, Bru73, Cha70, Cun03, Gre08, Hei71, Hög74a, Hög74b, Leg89, Pei92, Roy74, Tor65, Tor66, Tor81]. groupes [Bal81, Bou83, BR79, Cau65, Con72b, CR78, FSP71, Jac73, Kip74, Lar72a, Lar72b, Lar72c, Mal65, Mus06, RLT81, Tor69]. Groups [Hor74, AABV16, AD15a, BT15, Bor21, BE95, BW05, BW06, Bur07, CL13b, Cue18, Dah17, DS07, Der00, Émi85, Haz93, Lal98, Lia16, Mal65, MR94b, Mus06, RY89, Rau04, Sie84, Tho13, Tol00, Var95b]. growing [DHMP17, Hua19, SZ16]. Growth [Bal03, Bal04, ASS90, BC20, BCK21, CK20, CE20, CFP12, ERS16, JRdlR10, LT21, Sep20, Shi20, Tel08, Tim14, vdBPSV08]. growth-fragmentation [BCK21, Shi20]. guarantees [LC20]. GUE [DR22, Gus05]. Gundy [Yar19].

H [BW06]. Haan [Wüt04]. Haar [GNS13]. Half [AC15, Aar78]. Half-plane [AC15]. halfplane [KM99]. Hall [Dim18]. Hamilton [CPX17].

Hamiltonians [CHSX15]. Hammersley [FM09, SZ17]. Hamming [AKM21]. Hanson [ALM20]. Hanson-Wright [ALM20]. Hard [CK20, PS14]. hard-core [PS14]. Hard-edge [CK20]. Hardy [Var95b]. Harmonic [BY16, ALT99, BM08, Bat00, BE95, Enr96, Hoc19, HK09, KF98, Lin17, Mim14, Pin85, Sud21]. Harmonicité [Nev69]. Harmonicity [Nev69]. harmonics [MRW20]. harmonique [Enr96]. harmonisables [RE79]. harmonizable [Lac04]. harness [SZ17]. Harris [Bra73, BR79, Cha86a, CG04, Gho75, Kyp04, Mai78, Mét72]. hasard [Dep65, Lar72a, Lar72b, Lar72c, Lem73a, Lem73b]. Hastings [VST12]. Hausdorff [Aïd20, JJK^{+}14, Kah87, LT06, Per89, Shi19]. having [Ric13]. Hawkes [RDL20, Tor17, Zhu14]. hazard [Sar21]. Heat [MSC19, Shi10, vdB07, ANV00, BKK10, BBHK08, Bor96, CKK17, CK17, CJKS14, DT21, DV05, Erb10, FK10, HW22, MRVZ16]. heat-kernel [BBHK08]. Heavy [DvdHvLS20, AAP09, BCHJ21, BGP14, CCP12, MSC19, Soh13]. Heavy-tailed [DvdHvLS20, BCHJ21, CCP12, MSC19]. Height [GMT17, Liu18, Bra09, Ric13, SZ17]. heights [FVY06]. Heisenberg [AKM21, AK21, BP06, Bla03]. helix [Tal92]. Henri [Ano05a, Bal04, Mat98b, She04]. hereditary [Sch83]. Hermite [DE05]. Hermitian [Joh12, OW17]. heteropolymers [CGG10]. hexagonal [BS13]. Hidden [BRR02, GG08]. Hiding [PRS11]. Hierarchical [BT13]. hierarchies [BK04a, BK04b, CW13]. High [Ver10, BT18b, BG16, Che18, CPV22, De 02, DE16, Eld14, GLTZ17, GM18b, HMP20a, Lam21, Lem16, MSZ15, Ray15, Röl13, vdHS14]. High-dimensional [Ver10, BT18b, BG16, Che18, CPV22, Eld14, Lem16, vdHS14]. Higher [RZ17, Gol18, Sab01]. highly [GB16]. Hilbert [GNSS94, AHK77, BDT11, Cau69, KKK85, KK75, Ngh70, Ouv73, Pay67, TTL74, Yor74a]. Hilbert-valued [GNSS94]. Hilliard [ABK18, DDT04]. histogramme [AJ76a]. Hitting [Del08, Mat97, Mat98b, CGZ13, LLL11, LLL14, LLS11]. Hodgkin [HLT16]. Hoeffding [Pin14a, Tal95]. Hölder [AGL94, BB01, BB02, DLZ20, IL97, LT06, SSV03, dR17b]. Hölderian [KMS97]. höldériennes [KMS97]. holding [GG01]. hole [Mic99]. holes [LMD03]. holomorphes [Yor77]. holomorphic [Sab01]. homogames [GBB69a, GBB69b]. homogamie [Gil66a]. Homogamy [Gil66a]. homogène [Bab91, Mal65]. Homogénéisation [Wei84, GP01]. homogeneity [FLRB11]. Homogeneous [Vig11, Arn93b, ADGP13, Bab91, BT15, Fou01, Gan67, KK14, Mal65, RCG18]. homogènes [Arn93b, Con72a]. Homogenization [BT17, Ber12a, FHS19, IPP08, Rho09, BP04, Dar21, Feh19, GP01, GM18a, HKPG16, Rho08]. honneur [SL95]. honor [SL95]. Hopf [GJ11, Mac78]. hopping [CFP13, FGS18, FGS19]. Hörmander [BCP19, Pig18]. Horton [KZ17]. Hua [Ass20]. hubs [KLL21]. hull [Mas93, Zer05]. hulls [Dav02]. Hungarian [GN02]. Hunt [AHK77, KRR70]. Hurst [CN05, GNRV05]. Huxley [HLT16].

Hydrodynamic [AFJV15, BESS20, GLT09, JP17, Kom98, LM97, Val12, Val06, DLP07, GOVW09, KL95, Mou96]. Hydrodynamical [BF88, FGN13, Lan91, BLNR11]. hydrodynamics [BGRS14, GS10a]. hydrodynamique [KL95, Mou96]. Hydrostatics [LT18]. hyper [LLO06]. hyper-Bessel [LLO06]. hyperbola [Val89]. hyperbole [Val89]. hyperbolic [BM08, Bud20, CS21, DM09, MZ15, Val06]. Hypercontractivité [Fou00]. Hypercontractivity [Fou00]. hypergroup [BG92]. hypergroupe [BG92]. hypergroupes [Gal96]. hypergroups [Gal96]. hypoelliptic [HP11]. Hypoellipticité [Cat86]. hypothèse [RR70b]. hypothèses [GLS73b]. Hypothesis [GJN20]. hysteresis [Car13].

i.i.d [BV17, BRR02, MM14]. Ibragimov [Ibr02]. idéal [Hög74b]. identically [Sta89]. Identification [BL79, PT08, Ber91a, Sen90a]. Identifications [Bro83]. Identités [Rei82]. identity [Gil65a, Gil65b, KP11, Rog84, Yor93]. II [BDT11, BR03, BL05, BMP94b, CGN16, Duc20, Gil65b, Kan72a, KMS14, KRS18, Lac12b, Lar72b, LPPV69, Loo78, Rho08, TTL69, Zha10, vdHdHS07]. IID [BA09]. III [Ami17, Dav75, Hel99, Kan72b, Lar72c, LSW02, Rom69b, Zhu21]. ill [MS17]. ill-posed [MS17]. ils [BD96a]. immigration [AD09, HL05, Löc02]. impatient [CP78]. implications [AWS21]. incipient [Yao18]. incomplete [ÁEdBCAM11]. incomplets [Col78]. Incompressible [CX21, ACLZ20, BL08a]. inconnues [Tou69]. increase [Ber95, Mic99, SV07]. Increasing [AZ99]. increment [Sep20]. increment-stationary [Sep20]. increments [SS10a, She18, Son13]. indéfiniment [Cup69, Tor77a, Tor77b, Tor78a, Tor78c, Tor79, Tor82]. Indefinite [Cha91, Lia07]. Indépendance [Car80, MPL73, Bos75]. indépendantes [Bir84, BK95, Hen91, KL95, Le 89, Led92]. indépendants [Gal76, JKM82, Tor65, lJ87]. independence [Car80, Mat12, MW11, RV19]. independent [And15, AD00, CI03, COW20, CCFR09, GN02, Hen91, JP17, KL95, Le 89, Led92, Pru97a, RR94, SS10a, Son13, STW00, Sta89, Stu17, Tor65, vdB10]. independently [BK95]. index [FF18, GNRV05, IL97]. indexé [All78]. indexed [DMS05, MS83b]. indexés [Car81]. indices [Fis76, MS83b, Nua84]. indirect [BMP09, BG16, GJTZ08b, GJTZ08a]. Indistinguishability [Hut20]. individuals [FM19, Mai13]. individuels [Reg70]. Induced [LPS20, LT06, LP17]. induces [dlR98]. induction [PT21]. induit [dlR98]. Inefficient [GvdLW95]. inégalité [FS03, Rio02]. Inégalités [Lar93, Mat93, Ehr86, Hil77, Mat93]. Inequalities [HS96, Ané01, BE82, BR08, BCT11, CCEL13, Cat03, CGZ13, CM10, Che87, CCS92, Dar96, Egg80, FMP14, FGJ17, FI88, Gas02, Goz10, GRS^{+}18, GOVW09, Haf20, Hel99, KMS14, Lar93, Leh13, Lem16, LM09, Maj19, Mat97, Mat98b, ORB20, PD14, Pol02, RW17, Rio93, SS18, Tal95, WW16, Wu10, ZBH17, dIS02, Mat93]. inequality [ALM20, BLM99, ELS20, FS03, Fle10, Har05, HU09, HHM21, Joh17, LPY02, LLL14, Pin14a, Rio02, Sod08, Völ16, Wu00, Yos01]. infection [BT20]. Inference [PW02, CM21, Gug12]. inférieure [Heu98]. inférieures [Mus06]. Infimum [Sam07, Ber91b, Pra07]. Infimum-convolution [Sam07]. infini [Rou72]. infinie [Mic92a]. infinis [CK77, Tor65]. Infinite [BJM10, Bud20, KM20, Mar21, MW02, Pie09, Sar17, AZ14, AW18, ABL16, BSC04, BNNS18, Ber13, CT16, Cue18, DJO19, DR17c, FS04, GPS10, Gri04, HMP20a, Lac97a, LM97, MR21, MS70, MT20, Mén10, Mic92a, Pèn19, PSS09, DL15, Tor65, Yao18, Yar19, dIS02, Rou80, Saa88]. infinite-bin [MR21]. infinite-degree [dIS02]. Infinite-dimensional [Mar21, Cue18, DR17c, MS70]. infinitely [AD15a, BGJS22, Eva97, Phu81, dR17a]. infinitesimal [CAM98]. infinitésimaux [Jac71c]. infinity [Bro03, Rou80]. influence [DMS93]. influences [KMS14]. Information [Die76, Die78, Tra15, AI05, BS01, BZ12, CL94, CL95, MO83, NN16]. inhomogeneous [AEFdR20, CS02, GO13, HH09, MZ16, NS95]. Initial [CJKS14, FGN19, FK10, Gir02, JM98]. inner [Aar78]. input [HLT16]. insérée [dlRd03]. inserted [dlRd03]. insights [BCG15]. instabilité [Bir84]. instant [Lac91]. Institut [Bal04, Mat98b, She04, Ano05a]. integer [EG08, HK02, Has90]. integers [KM11]. integrability [BDMBP18, LLL11, LLS11]. integrable [NS22, Gal76]. integral [BB11, CN05, EH03, Est92, IKPY18, Jac10, Kal20, Kan72a, Kan72b, KR16, Lac91, Lac97b, Lia07, SY05]. intégrale [Est92, Lac91, Lac97b]. intégrales [All78, Jac82, Yor74b]. Integrals [PW79, CŚ07, DMY91, GS18, GK93, GRVY99, GNRV05, KKK85, Kaz71, Mar00, Uek99]. integrand [Mar00]. Integrated [Sab01, AD13, DW15, GG08, Vys14]. Integration [Dah17, Hei71, Hei74a, CCM03, CL21, Dec05, Ust82]. integro [Pie09, ST20a, LM76, LM77]. integro-differential [Pie09, ST20a]. intégro-différentiel [LM76, LM77]. intensity [CGG11, DRS14, HPT02, Lem16, Sar21, Tor17]. Interacting [BST20, BMP94a, BMP94b, AAD19, AD97, BR03, DVW00, DG01, DD07, DW08, FL16, FTV20, GMT17, Gra88a, KLO95, Kur10, Léo95, Löc02, LM22, NO88, PSS16, Völ16]. interaction [AAD19, BP15, BRT14, BJM10, Che14, Hil73, KL05a, KM17, Léo86, MO07, Oga74, Osa98, Rou86]. interactions [Sud21]. interchange [AKM21]. Interface [Bod97, ABK18, BDP95, HLS22, Sep00]. interfaces [Pet09b]. Interlaced [MOW09, Def10]. interlacement [FP18b]. interlacements [DRS14, DE16, MRT21, SS10b]. intermediate [BL19]. intermediately [ABKV14]. Intermittency [AD11, DM09]. intermittent [DGM10]. Internal [BQR03, BDCKL20, AR16, ELS10, Bla02]. Internal-diffusion-limited [Bla02]. interne [Bla02]. Interpolation [BGJS18, ACLZ20]. interpretation [PPR97]. intersecting [BCR15, NR17]. Intersection [BK93, CR05b, JN19, LSW02, MV07, Ros96, Ros99, Ros01, Ros10, Wer93, vdB10]. intersections [Ass10, LPS03, MR97, vW83]. intertwining [Dub04]. Intertwinings [CD19b, RS18]. interval [ANT17, BP92, Del64, FGL95, Lam00, Pac00]. intervalle [Del64, MS83a]. intervalles [FK93]. intervals [FK93, FW04, GG01]. intervenant [Szp78]. intrinsèques [FlJ84]. intrinsic [BCK21, Ros01]. Introduction [LOR69, LPPV69, MV69, Rom69a, Rom69b]. Invariance [And14, CFP13, CC08, CD16, CM08, DMR95, Mie08, Pri12, BDCH16, CCKW21, CP04, DMP14, DDP87, DK08, DGR18, DV08, Enr07, FT79, GI05, Mas88, MR90, Osa98, Pet70, Uek99, Le 02, Mas93]. Invariant [BDEG88, BDEG89, LPP15, Mal11, SZ17, SvG11, ADR67, BT15, Bef03, BCR19, Bré09, BBD12, Bro20, CR05a, CG92, DK79, DGM21, FG12, HMP20b, HL05, LMD03, Mik88, Oka85]. invariante [ADR67, BDNT74, CG92, Gra88b]. invariantes [Flo82, Jac73, Saa88, Yor75]. invariants [KT97]. invasion [NS95]. inverse [GvdVY20, MS17, NdS20, Tra15, Yor85]. inverse-square [NdS20]. inversion [GQ16]. Inverting [ABDG19]. Inviscid [FGHV16]. involving [YYY10, Zho07]. irréductibilité [Nev72b]. Irregular [DDT04, HJY11, AKU17, BH97, KHT12, LR93]. irreversible [BB17b]. Ising [BDCH16, CGN16, CGGvdH22, CP01, CD18, DP11, GW18, KLS21, LS06, MR22, PS19, dT16, vEEIK12]. Isolated [IY20]. isolée [Tou69]. isometric [Kol92]. Isométries [BH81]. isomorphism [BHS21, BC19]. isomorphismes [DC73]. isoperimetric [Sod08]. isopérimétrie [Fou00]. Isoperimetry [Gol18, BM00, CT18, Fou00]. isoradial [dT07]. isotropes [Rai99]. isotropic [Feh19, Rai99]. Issue [Mat98b, SL95]. Itô [GK93, GNRV05, GP74b, Hud05, OP89, PSV17, Tor79]. iterated [AN09, BDEG88, BDEG89, BL02, CRR98, Deh00, DDG13, GK93, KL96, MR94a, MR97, Zen15]. Iterates [Fog75]. itératifs [Her05]. Iterative [Alq08, Her05]. itinéraires [MS83a]. Itzykson [CŚ07]. IV [MV69].

J [Cha71]. J. [Cha71, Duh97]. Jacobi [BB18, CK20, CPX17]. Jean [Bre02]. jeux [Lem73a, Lem73b]. Joint [ABW17, AWX08, JD18, Ros96, Ros99, ASX11, CFS87]. Jump [HK09, ADGP13, BKK10, BKS12b, CGC10, Erb14, Ish97, LR17, MU11, MRT21, Pel10, RN21, Rou80, Yan18, Zen16]. Jumping [BRT14, JS96, FX19]. jumps [BGJS22, CDM21, CP12, DPSZ16, FHS19, FJR21, Gra92, Kun86, KM14, Léo95, LM11, Maj19, Mou96, Pri18, PZ04, Wu10, XZ20].

K-P-P [Kyp04]. Kac [AD08, Bod97, Car15, CDG11, CZ02, MO07, Yin96]. Kakutani [Loo77c, Loo78]. Kantorovich [Pra07]. Kaplan [BLM99]. Kasteleyn [GMS19]. Kawasaki [BL15, CMR02]. Keller [GQ15]. kernel [BBHK08, BM89, BW05, BW06, CL13a, Dev89, GG01, GG02, MRVZ16, MSC19, Shi10]. kernels [BKK10, BN90, CFP19, DTMS98, DW08, Gan67, Her08a]. Khintchine [AD15a]. Kiefer [BLM99]. kill [JdlR12]. killed [AD01, GHS11, Ras11a]. killing [KK14]. kinetic [Bas14]. Kinetically [BT18c]. Kingman [FFJ19, KZ17, Ste89]. Kipnis [SL95]. Kirkpatrick [BL20, CH06, Che14]. Kneading [MV91]. Knight [BP15]. Knudsen [CP12]. Kob [MST20].

Kogan [Die82]. Kolmogorov [HHK06, BDT11, BGM20, Gru66b, TV81, Zen15]. Kostlan [BMW21]. KPP [HT04]. KPZ [CFP12, Dim18]. Krengel [JdlR12]. Kronecker [AEKN19]. Krook [Hof15]. Kullback [CL95, CL94]. Kuramoto [LP17].

L. [Loo81]. labelled [Eva97]. labyrinths [BG99a]. Lace [BvdHK18, Sla22]. Laguerre [DE05, NR17]. Lambda [VW15]. laminations [BM08]. LAMN [GG08, Löc02]. lamplighter [BW05, BW06, KMP14]. LAN [Gob02, Löc02]. Landau [BDP95, Hos18, Oda06]. landscape [Che15, HRSS21]. langages [Fri67]. Langevin [AB18, HRSS21, KP19]. language [Ben67b]. languages [Fri67]. Laplace [BDC69, Cat03, EN20, Lif94]. Laplacian [KM02, Lou68]. Laplacians [Hel99, HL20]. laplacien [Lou68]. laplaciennes [BDC69]. Large [Aid10, BCR09, BA09, BEV13, BFG15, BD96b, CGG10, CR05a, CD19a, DMS05, ES02, EHK10, FS17, FK94, FV07, GZ98, Gin83, Gou07, Gui02, HK02, HLN17, IR98, Ish97, Kob13, Léo95, MRTZ16, MM09, MSS06, MO07, Ray15, Tra02, AD08, AD00, Aug18, Ber15b, BM16, CD01, Cat91b, Cat91a, Cer96, DO20, DLN21, Dos91, DST90, DE05, EW06, Eng08, Eng09, FG12, Feh19, FGL95, Gal96, GZ19, GHR01, GM13, GB16, Gui00, HLNV13, HKK10, HP00, JS20, KPT20, KK03, LT18, MN18, Mar16, Pan14, RAS11b, Ric13, RT10, Sep00, Stu19, Tro96, WY16, Yil10, Zhe12, Zhu14, KL95, Lot91, Rab94, Wu91]. Large-deviation [Rab94]. largest [AAP09, Hen91]. Lasota [LMD03]. Lasso [Lem16, ORB20]. Last [CG21, iSW05, CH12, iSW04]. later [Mas02]. lattice [BL08a, BEM03, BL95, BT18c, Fun99, GK17, Ken97, MP00, Nag05, Yos01]. lattices [AKKR01, FS17, Fis01, Gho77, KK03, MP19, SSS^{+}21]. Law [DLN21, EW06, Eng09, EHK10, ABO16, And15, BR03, BG10, BL02, BDZ16, CD01, CPY04, Con65, DRZ21, DST90, Eng08, ELS10, FGN19, FW09, Gal96, Gui90, GL89, HKK10, KL96, KK03, Led92, Mén10, Pra92, Sep00, SZ16, Tou96, Yor93, Zen15, Coo19, Gui00]. Laws [FFV17, MR94a, MR97, AvdHH19, AS92, BF95, BCK66, Che08, CRR98, Deg79b, Deh00, GGL15, GS15a, GK93, Hof15, JS20, KM99, KM11, LT06, Mar00, MZ84, NO88, RT10, Str90, Tor66, Zhe85]. Leader [BW06, BW05]. leading [DV08]. learning [Ber68, Kol03, LC20]. least [Alq08, Aud04, Ber91a, BD92, Cha91, DST90, DST91, Tou96]. least-action [Cha91]. least-squares [DST90]. leaves [MSZ15]. Lebesgue [All78, dlR98]. lecture [Bre87]. left [AC04, ČTW11, Cun03, FW09, Jui16, Was15]. left-curtain [Jui16]. Lemma [Mas02, HR17, Zai02]. length [BD16, Con65, DF20, MRW20, Szn82]. lengths [HS99, RVY09]. leurs [FK69]. Level [AS18, PW17, AP05, BL19, CG68, MV20, RAS11b, Zhu14]. Level-set [AS18]. Levitov [VST12]. Lévy [AD15a, Ber95, Duq03, AD15a, BT18b, Ber91b, BD97, Ber20, BY87, BJ20, Bre87, BL12, CDM21, CM16, CG19, CGC10, Der90, DNT12, DM04, DW17, EO22, Fou05, Fou13, FJR21, FLS06, Gan67, Gol89, GM13, Gug12, HS17, JKMP05, Kol92, KP19, Kun86, KL10, KRS21, Lac04, Lam00, LMW21, MR94a, MØP04, MZ15, Mim14, PPR18, Pie09, Ric13, Riv07, Riv12, Rog84, iSW04, SW11, SV08, Tra15, Véc19, Vig03, YYY10, Yan13]. Li [Bañ18]. liaisons [Ber68]. librairies [Die82]. libre [Mic92a, Mic92b]. Lie [AD15a, Bal81, BT15, Cue18, Dah17, Der00, Hud05, LCO09, Lia16, Rab94, Tol00]. Lieb [CCEL13]. liée [LR96]. liées [Deg79a]. Liens [Dos77]. liés [Dho72]. lifetime [Chu84]. lifetimes [Pin90]. like [AMS95, BRR02]. Likelihood [Gas02, Bir06, Jac90, MNR21, MR89, Spo17]. likelihoods [BRR02]. lilypond [Hir16]. Limit [AZ14, AvdHH19, BB17a, BFK20, Bli14, Bru93, BW19b, CGPP13, CLP14, Che08, DVNZ20, DRZ21, EG08, FGN19, FHL12, FG92, FNS00, GKV03, GLL18, GPP10, Kor15, KM11, MWRZ04, NO88, SY13, ABKV14, Ami17, And15, ADHV19, AKU17, AKU20, ABK18, BY08, BH08, BLNR11, Bay15, BF88, Bet15, BF92, BG99a, BESS20, BG92, BDC68, BU18, CDMF12, Car16, CDGR20, Coh72, CD12, DLP07, De 98, DR00, DR08, DGM08, DMR94, EV03, ELS10, FW09, GS15a, GLT09, Gou05, Gri04, GOVW09, Gui02, HJY11, Her77, HH81, Her08b, JL06, JL08, JP17, JN19, KM99, Kif14, Kom98, KMM10, KV22, KRT16, KP19, KOS19, Lan91, LM97, LP97, LSX21, Lin76, LT20, LT06, Mas86, Mis16, MO07, Mou11a, NNT10, Num78, OR18, Ouc05]. limit [RAS09, RT97, Rez02, Rio09, RR94, Shi19, Val12, Val06, Var06, Zhe12, dR17a, dT07, Har01]. limite [BBC78, BF92, BG92, Bru93, CR78, EV03, Fis73, Her08b, JKM82, Led83, Mai78, Mét72, Roy74, Tou83, Dep65, Gil66a, GBB69b]. limited [Ami17, BB11, BDCKL20, Bla02, Gil66a]. limited-effect [Gil66a]. limitée [Bla02]. limites [BDC68, Bre83, Fis73, GH88, Har01]. Limiting [BLS22, BGS10, Cel11, Sin07, CM15, GPdS16, Guo14, PS13]. Limits [BD11, GS15b, AW18, AS17, AFJV15, CGN16, CDH00, Cro18, CLM13, CL17, Dar96, Dim18, FGHV16, GMS19, KL20, Le 91, LT21, Mar20, PWB20, Pet09a, Pet09b, RDL20, Riz15, Sen17, Soh13, Stu17, Stu19, VST12, Yao18]. line [DW08, GRS^{+}18, Hir16, KM20, LTT18, Pem97, SS18]. line-segment [Hir16]. linéaire [Ber68, Bro83, CPP71, Dos91, Oga74, Rab94, Szp78, SM79]. linéaires [Cab79, GP01, Páz77, Saa88]. Linear [Smo83, Bay15, Ber12a, Ber68, BRT07, CFS87, DV05, Dey20, DGW06, DLS02, EF05, FGS19, GP01, GvdVY20, GL16, KP19, LW20a, LLS11, MS96, PPR97, Zhe12]. lines [FLT04, GI05, KRS04, PW17, RV19]. linguistics [Ben67a]. Linguistique [Ben67a, Ben67b]. link [Stu17]. link-weights [Stu17]. Liouville [AABV16, AG21, Ber15a, GM21, HP21, HRV18, Jac18, MRVZ16, MS19, Sab01]. lip [BBC05]. Lipschitz [CR05a, EO22, GH12, Ham96, HHM21, Kob13, MH18, SvG11]. Lipschitzian [DW11]. lipschitzien [Ham96]. little [Dos80]. Littlewood [Dim18, FS03, Var95b]. Local [ASX11, BK93, Ber83, BDK19, DRS14, Ken97, KV22, LPP20, Liu01, MN18, Mer06, RY95, Stu19, And15, ADHV19, AWX08, Aza89, BH08, BBC78, BCP19, BB01, BB02, BB11, Bor87, BB17b, CR05b, CW18, CW20, CRR98, CCFR09, DOS18, Émi81b, Émi86, ESYY12, FF18, FG92, FZ88, FX19, GZ98, Gou05, GRVY99, HMNP98, Her05, HHM21, HP19, Imk84, IL97, Jac98, JN19, KRS04, KMM10, LR20, LP97, Lia07, MR94a, MØP04, MV07, Nak86, OTT02, Ros96, Ros99, Ros01, Ros10, Sep19, Var06, Wer93, Arn93a, Her05]. local-time [Wer93]. localement [BR79, Cha70, Fer84a, FSP71, Gre08, Ham96, Tor77b]. locales [Arn93a, Ber88, Ouv73, Rao70]. localised [GNS13]. Localization [BGP14, BKK10]. locally [DS07, Gre08, Ham96, Haz93, Rho09, SvG11]. Location [AEKN19, She18, CFS87]. locaux [Bou83, Wer93]. log [CCEL13, HU09, Joh17, Mad15, Mat12, SS18, Völ16, Yos01]. log-concave [CCEL13]. log-convexity [Mat12]. log-correlated [Mad15]. log-Sobolev [HU09, Joh17, SS18, Völ16, Yos01]. logarithm [BL02, CRR98, Deh00, GK93, KL96, MR94a, MR97, Zen15]. logarithmic [CMR02, ELS20, GOVW09, Hel99, LPY02, Mic09, RZ20]. logarithmically [AL12]. logarithmiques [Mic09]. logistic [Pic65]. logistiques [Pic65]. loi [Alt87, CD01, Che76, Con65, Die81, DR82, DST90, DS84, Fou84, Gal96, GP01, Gui90, GL89, Has90, Led92, Léo86, Lic69, Pra92, RLT81, RE79, Tou96]. Lois [AS92, BCK66, Bre83, Deg79a, Der76, DP84a, Rev83, Rou78, Tor65, Tor66, Tor79, Tor81, Ano77, Deg79b, FK69, For73, Gir70, Str90, Tor76, Tor77a, Tor77b, Tor78a, Tor78c, Tor82]. Long [GP14, Hey11, KL20, LP17, Ami17, BGJS22, CCKW21, CM08, Com87, DLS02, FR03, Gui97, Lac12a, Lac12b, Léo95, Sud21]. Long-range [Hey11, Ami17, CM08, DLS02, Gui97, Sud21]. Long-time [KL20]. Longest [Rou21, BB17a]. longue [Gui97]. longueur [Con65]. lookdown [VW15]. Loop [Bor08, LPS03, Qia19, Shi19]. loop-erased [LPS03, Shi19]. Loop-free [Bor08]. loop-soup [Qia19]. loops [AvdHH19, Car13]. looptrees [Arc21]. loosely [JdlR04]. Lorentz [Pèn09]. lorsque [Wu91]. lose [Pin14b]. Low [BT18b, BL15, Dal20, De 02, PS19]. Low-rank [BT18b]. low-rankness [Dal20]. Lower [CH20, DHHX20, FW07, Lac12a, RT10, AC99, BC20, Cha21, CDH00, DP11, Heu98, HKM14, Yur08, Mus06]. Lusin [Smo83]. Lusin-measurable [Smo83]. Lyapounov [Bou88a, Pov98, Zer00]. Lyapunov [Bré04, EdHM14, HSS21, MS96, Per92, Tia17].

M [Coh68, BCHJ21, Hai81, Hai87, CG68, Coh68, Coh72]. m-dépendante [Hai87]. m-dépendantes [Hai81]. M. [Ano90]. M/G/1 [CG68, Coh68, Coh72]. même [Led92]. Macdonald [Ahn20]. Machlup [BRT03]. Macroscopic [KLO95, MO07]. magic [CD19b]. magnetic [Com87]. magnetization [CGN16, MCRT06]. main [Ber89]. Major [NZ21]. majorant [DT03]. Majoration [Kip74, Mic99]. majorations [Rou87a]. majorization [AN09]. malignant [DLP07]. Malliavin [BDMBP18, CG92, Gra88b, KHT12, Maj19, MPR17, NPR10]. Mallows [BB17a]. malthusiens [LR83]. Mandelbrot [BDK19]. Manifold [ALT99, CE00]. Manifold-valued [ALT99, CE00]. manifolds [BBCH21, BE95, Dar85, Dar96, Erb10, GZ19, Yan21a]. many [AB97, BK04a, Eva97, HR17, Tal92, dR17a]. many-to-few [HR17]. map [Abr16, ABGG^{+}12, Fis01, GMS19, Hoc19]. mapping [CY07]. maps [Abr16, ANT17, AC15, ALT99, Bet16, Car16, CL17, DGM10, Gou05, LMD03, MV92, MV91, Pac00, Ray15, Stu17]. Marche [Con65, Leg89, Cun03, Dep65, FLP06, Pei92]. Marchenko [BG10]. Marches [BR79, Hög74a, Lar72a, Lar72b, Lar72c, Pei92, BDC67, BDC68, Dav74, DV98, Der76, Gre08, Hög74b, Mus06, Sch00, Sun87, Val89]. margin [Kol03]. marginal [Mar00]. marginality [GLT11]. marked [ADGP13, FINR06, Smy75]. marker [CGG11]. marker-dependent [CGG11]. markets [AI05]. Marking [NRS10]. Markov [BDEG89, ADH12, AP98, AKU20, AMZ21, AKDR66, ADR67, BDEG88, Bax96, BLMZ15, BH20, BFG15, BCR20, BP95, BRR02, Bor08, BB17b, CKS87, CZ02, CG04, CY07, Chy08, CG92, Cue18, CL09, DGM10, DMS05, Der73, Der76, DE02, Dia05, DGM08, DR69, DFM16, DGM21, ERR71, ET90, Eva87, Eva91, Eva97, FHL12, Fer15, FS89, FlJ84, Gal96, GZ98, Gav75, Geo78, Gir70, Glo91, Gol95, GLL18, Gra88b, GH88, Her77, Her05, Her08a, Her08b, HS96, Hor74, JS96, JS20, KMS97, KP11, Lac07, LS91, Le 91, Lin75, Lin76, LPS03, MRZ98, Mai78, MWRZ04, May68, McD75, MT20, Mét72, MNR21, Mic99, Mik88, NRR20, Nev72b, Num78, Oga74, Osa98, Pei92, Pel90, Pel10, Phi71, Pri68, Reb05, RR70a, RR70b, Rev83, Riz15, Ros86, Rou72, Rou97]. Markov [Sar14, ST20a, SJ14, Soh13, SZ97, Sun75, Tou83, VA89, Wu99, Wu00]. Markovian [BCR19, Dav75, DD10, EP98, Kal20, KP17, MRTZ16, MW07, Pin14b, PZ04, Vig11]. markovien [Lot85, Pel90, Rou87a, Rou87b, Rou97, Szp78]. markovienne [Bou88b, Ruc80]. markoviennes [Bab88, Bro75, CGR79, Doz83, FK69, Hen82, Jac71a, Jac74, KMS97, Rei82, Rou78]. markoviens [Ben79, Bou88a, CL84, Car81, CK77, Jac71c, Jac73, Pei92, Roy80, dZ81]. Markovness [Mat98c]. marquée [Saa90]. marqués [Bla73, Hil73]. Martin [BW06, AD01, AR06, BW05, Col04, Der73, KKK85, Ove94]. Martingale [PSS16, AS92, De 98, DHJ^{+}21, Gra88a, HQZ17, Kaz71, Ouc05, Sch85, SBM21, Str90, Wu99, Yar19, AS92]. Martingales [Dar96, Hei78, Ouv73, Pia95, Van72, Arn93a, Arn93b, ALT99, Att95, Bag85, BS03, Ber88, Dar85, DST90, Enr07, FI88, Gal76, Gri04, IKM20, JKM82, Led81, LM76, Lon81, Mat93, Mét84, Nak86, Nev72a, NS81, Nua84, PW79, Pic94, Pic05, RY95, Rau83, Zen15, dZ81, dlR02]. Maruyama [MGY20]. mass [AKKR01, ABCF21, CE20, ERS16, FNS00, GP07, Pra07]. master [Dav75, Pel10]. matching [AAH21, HPPS09]. matchings [HMP20a]. mathematical [Ben67b, Ben67a]. mathématique [Ben67b, Ben67a, LOR69, LPPV69, MV69, Rom69a, Rom69b]. matings [AG21, MS19]. matrice [Deg82]. matrices [ARRS16, ACLS19, AEKN19, And15, AD15b, AAP09, Aug18, BCHJ21, BGP14, BGBK20, Ben20, BQZ20, BGS10, BJM10, BM16, CD01, CDMF12, Cau65, COW20, DGZ03, ESYY12, Gra99, Gui02, HLNV13, Hen91, HP00, Joh12, KK19, Lam20, Le 89, LSX21, LWZ16, LV20, OW17, Pan14, Péc12, PRS13, Roy80, Zhe12]. Matrix [Li19, BT18b, BDN20, BG16, CI03, Def10, DMSY94, ES17, HLN06, LW20b, Nag05, Rho09, Tia17, WY16, Woo16]. Maurey [Bas82]. maxima [AR08, Rio02]. Maximal [BÉL09, GP11, LS16, Pol02, DL15, ZBH17, Bau13, Dar96, De 98, RT10, dlR98]. Maximum [DCG90, Mad15, AS94, Ber85a, Ber85b, Bir06, BDZ16, CH20, CFS87, Die81, FK10, JV07, MNR21, MR89, SV07, Spo17, dlR98, DCG90]. McKean [HSS21, LMW21, Mar21, RSX21]. Mean [BZ12, HRSS21, KM17, AZ14, AK21, AB97, AG89, AG91, AZ99, Bañ18, BL14, CGL21, Cat12, CAM98, DR08, EM83, FS04, GLP20, Her08a, JYZ05, Kom98, LM97, Lou05, RDL20, Sue05, Var95a, dBMCA99]. Mean-field [HRSS21, KM17, AK21, Bañ18, CGL21, GLP20, RDL20]. measurability [Ber13]. measurable [BB21, Sie84, Smo83]. measure [Aïd20, ADR67, BP06, Bat00, BY16, Bré09, BBD12, CM10, Con18, DVW00, DGZ03, EP06, Eva91, Gou07, Goz10, Gui79, HSS21, HL05, KT97, Lac97a, LPP15, Lin17, Liu01, LT06, MT20, NR03, NZ13, Pai18, PPR18, Pèn19, Per89, Pin85, Pro86, R.83, Rhe84, SO75, Smo83, Yan13, dlR98]. measure-preserving [KT97, Lac97a]. measure-valued [DVW00, EP06, Eva91]. measurement [BG20]. measurements [HMSH12]. Measures [GH73, ANS06, AKKR01, ACJ20, AD15a, Aus15, BV17, BM08, BB18, BDEG88, BDEG89, BESY08, BL03, BCR19, BL14, Bor77, BW01, BHL93, BCQ21, Bur07, CCEL13, Cel11, CR05a, CJKS14, Cou20, DK79, Def10, Del77a, DSS13, Din92, DGM21, ES02, Eld16, FG12, Fou95, GZ19, GR11, Gla76, Gua77, Heu98, Hue16, Jac79, KBR71, KK75, KM17, Lam21, LTT18, LS06, LM83, LMD03, MN18, MR94b, Mkr14, Oka85, Pri12, RY89, RV13, RZ93, Sam07, Set01, SvG11, Tes06, WW16, Wu91, Yin96, ZBH17]. mechanics [Zhe85]. mechanism [AD09, KLMSR12]. media [BRR17, CS02, DGG21, Saa95, Var06]. Médianes [Hei90, BH85]. medians [Hei90]. medium [Ben67c, Bré04, Bré09, CD16, Eng08, FNS00, HQZ17]. Meier [BLM99]. Meilleures [BH85]. mélange [Mok90]. mélangeant [Del77a]. memoriam [Ano90]. memory [BKS12b, GP14, MW07]. Mendelian [Gil65a, Mal65]. mendélien [Mal65]. mendéliens [Gil65a]. mesoscopic [LSX21]. mesurable [Val78]. mesurables [DAP76, Del77b]. Mesure [PPH75, ADR67, AJ76c, BDNT74, DLP86, Han73, Hei71, SP75, Tor69, Tor78b, dlR98]. Mesures [Cal81, Din92, Han73, Hei74b, Hei83, Hil78, Saa88, Tes06, Del77a, Del87, Fer84b, Fou95, Gui79, Han72, Heu98, Jac79, Jac73, Le 85, Led70, Les87, Pel72, RLT81, RZ93, Szp78, Tor69, Tor71, Wu91, Yor75]. metabelian [Tho13]. Metastability [LS19b, LM22]. Metastable [Che15, Bar15]. method [AMD17, BLM21, Dev89, DHJ^{+}21, Gau20, GW18, HPT02, HQZ17, JKMP05, Kas20, KHT12, Kur10, NPR10, Pan09, Röl13, Sar21, Wat87]. méthode [CPP71, JKM82, Jum80, dZ81]. Méthodes [Bro75, Dub75, JT76, CGR79]. methods [ATV20, LW20a, MNR21]. Metivier [FG79, Ano90]. metric [GM21, Rau92]. metrics [Ben67c, CP11, RY89]. métrique [Rau92, Sun75]. métriques [Ben67c, Bod70, Bod71, Val78]. Metropolis [KOS19]. Meyer [BB21, Ros99, Yar19]. Meyer-measurable [BB21]. Microscopic [BKS12a]. migration [Mal65]. milieu [Ben67c, Saa95, Sun87]. Milstein [DNT12, GN09]. Milstein-type [DNT12]. min [BLNR11]. min-driven [BLNR11]. minima [AR08]. Minimal [Her77, HKM14, LLO06, Szn82, ABL08, AFGJ16, Boi98, DT20, DKGT16, Rio09]. Minimax [MS17, Bir84, JM93, KM02, GJTZ08b]. minimaxity [BM89]. Minimization [CL94, CL95, Kol09, LM13, LW20a]. minimum [Hu16, Pra07]. Minkowski [Ehr86, HLLS22]. minorant [EO22]. misanthropes [ACTR85]. mise [Heu79]. Mises [Deg79a, Deg79b]. missing [ABCF21, Tal95]. mixed [Che14, Tou96]. Mixing [DS07, DP11, Pèn19, PS19, Ass00, BHLP19, BD96b, CMR02, DMR94, El 00, Eld14, GLP20, Guo14, HP17, Her18, KMP14, Lin76, MT20, NN20, Rio93, TV98, Yos01, Del77a, Mok90]. mixte [Tou96]. mixture [Bla99]. Mixtures [Gua77, CM10, GKMMR20, Gas02, LMMR18]. Mme [Cha71]. Mobility [Lou05]. Model [Bir06, And14, AvdHH19, AD14, AD11, BY08, BL20, BW19a, BB14b, BL08a, BDCH16, Ben67c, Ber91a, BT13, BRR02, BS10a, BS11, BB17b, CGGvdH22, CCP12, CH06, CD18, CK12, Che14, Che19, Che20, CV14, CGK20, Com87, CQR07, CH12, CP04, DLP07, DV21, DH21, DP11, DST90, DST91, EM15, EdHM14, EN20, Eva97, FFV00, FGR09, FHS19, FL16, FGHV16, Fun99, GLS05, GS15a, Gas90, GQ15, Gor17, GK09, GPdS16, GMS19, Häg03, Hir16, HLT16, HLN17, IL06, JJ18, JM98, KM20, KL05a, KP12, LS91, LT18, Lem16, Ler12, LW20b, LP17, MM09, MR21, MCRT06, MP19, MST20, Mis16, MOS11, NO88, NW18, Pan09, PS14, Poi13, Saa95, Sch83, Sen90a, Shi20, Tou96, Ver10, dT16]. model-super-Brownian [CP04]. Modèle [SM79, Ben67c, Ber91a, CM72, DP84a, DST90, DST91, Gas90, LS91, Lic70, MV69, MPL73, Saa95, Sen90a, Tou96, Rom69a]. modèles [Ber91a, GL80, Her05, Pic65, Sen90b]. Modeling [BRT14]. models [AKM21, ABL08, ASS90, AC15, AZ99, Bal03, Bal04, BC20, BL08b, BK04a, BK04b, Bro20, CGGvdH22, CP01, CDG11, CCC16, CP17, DGdH^{+}08, DGR16, DvdHvLS20, DR22, GR08, GvdLW95, GW18, Gua77, Han04, Her05, KLS21, KZ21, LS06, MM13, MS21, MR22, NS22, ORB20, PD14, PWB20, Pic65, Ros89, Sen90b, Tra15, dT07, vEEIK12]. Moderate [BESY08, Cas04, DMPU09, DGZ03, DGW06, Eic01, ERS15, CG04, CH20, CGK20, GZ98, Haf20, JM98, Led92, PD14]. modérées [Led92]. modified [BR08]. moduli [CCS92]. modulo [Cab79]. Moindres [BD92, Ber91a, CPP71, Deg82, DST90, DST91, Tou96]. Moment [Che17a, WY16, AL12, DCG90, DK19, Her77, Joh12, NN16, de 95, dlP94]. Moments [DM04, iSW04, DCG90, DT21, DS96, HHM21, Led09, NRR20, Nev69, RZ17]. Monge [Pra07]. monotone [AS17, BB17a, Bel08, DT03, HS11]. monotonic [MSZ15]. Monotonicity [MV92, RCG18]. Monte [HLOT^{+}19, MNR21]. montées [Ber88]. mort [CR79, CR80, Rob87]. morts [Mou96]. Moshe [LW20b]. motion [AT20, Arc21, ABK15, ASX11, Bab91, BBD19, BB89, BH97, BBC05, BO11, BBS15, BL08b, BL03, BM93, CCM03, Car85, CG79, Cha94, CDH00, CK12, CN05, Chu84, CP11, CP04, ClJ89, Cro08, CCFR09, Dah17, DS90, DV05, DPRZ00, DNT12, Eld14, Eng08, EPT17, FGV16, FGR09, FS04, Fra91, Gan67, Gra99, GNRV05, GMV93, HP11, HH09, HR12, HNPS19, HP19, Jac18, Jac10, JN19, KL96, KLMSR12, Lac91, Lac97b, LLO06, Lac12a, Lac12b, LTT18, LY19, Lou68, Mai13, MZ16, MM18, McG89, Mer06, Mey91, MSS06, Mör01, Nak15, NdS20, NNT10, ÖÇE17, ÖE19, Pem97, Per89, PT98, SY05, SV07, Sch09, SBM21, Wer93, WZ90, Wüt98, Wüt99, vdB07, vdHS03]. motions [BOT14, BOZ16, BSC04, BJ20, BÉL09, CPY04, CK22, CF09, Dub04, FWY00, Lac04, LS19a, PT08, RS18]. Mott [CFP13, FGS18, FGS19]. Mouvement [Lou68, All78, Bab91, Bia86, Fra91, Gol89, Lac91, Lac97b, Mey91, Nev76, Rec86, Sch85, Wer93, Yor85]. mouvements [Loo77b]. movement [Gol89]. moving [DGW06]. moyen [Dep99]. moyennables [Mus06]. moyenne [Ant82, Cha70, GS97, MO83]. moyennes [Lac73]. Multi [Yao18, GCJ93, Han04, KK03, Rou97, Szn82, Rou97]. multi-allelic [Han04]. Multi-arm [Yao18]. multi-channel [Szn82]. multi-dimensional [GCJ93, KK03]. multi-échelle [Rou97]. Multi-scale [Rou97]. Multiapplications [SP76]. Multiclass [FM09]. multiclass-customer [FM09]. Multicolour [AAH21]. Multidimensional [LMMR18, Bru93, Chy08, EZ19, GQ16, LR17, Sch00, Sło93, GS97]. multidimensionnel [Gar79]. multidimensionnelle [Bru93, MM71]. multidimensionnelles [Gar78, Sch00]. multidimensionnels [GS97]. multifractal [PT98, RV13]. Multifractality [Yan18]. multifractional [ASX11, CM08, Lac04]. multiindex [Gas90]. multiindice [Gas90]. multilevel [GS15b]. Multiparameter [ASX11, CGB17, FS89]. Multiple [Ass00, AvdHH19, FFV00, HR17, Zha22]. multiple-force-point [Zha22]. multiplicatifs [Bou88a]. Multiplications [Kah87]. Multiplicative [BCR15, ST94, Bar00, BM09, BG20, CR05a, DGG21, Est92, FP17, Fur97, Jin14, Lem16, MS96]. multiplicativement [Dho72]. multiplicatives [Jac71b]. Multiplicities [Eva94]. multiscale [GLM18]. multisections [DAP76]. multitype [CLP14, CCM19, Mie08, dR17a]. multivalued [Bag85]. Multivariate [NPR10, Dal20, Gau20, GG02, KRT16, R.83]. multivoque [Dau73, Val80]. multivoques [Nev72a]. mutation [Cer96]. mutation-selection [Cer96]. mutual [BZ12]. Mutually [DFM^{+}03].

naissance [CR79, CR80]. naissances [Mou96]. narrow [BL95, CF17]. Nash [LLL14]. Natural [HLS22, BCR20, RZ17]. naturally [Zen16]. Navier [AKSS07, BL08a, CD19a, XZ09]. Near [ABL08, BD11, CGN16, DE16, GH10, Pai18, Sen17]. near-critical [CGN16, DE16, Pai18]. Near-minimal [ABL08]. nearly [OW17]. nécessairement [Gas90]. nécessaires [AJ76a]. necessarily [Gas90]. Necessary [dBMCA99]. needlets [CM15]. Negative [Kop82, Bab91, Con65, Dar85, PPR18, Bab91, Con65]. Nested [Duc20, Mar15]. network [CV14, Dep99, FTV20]. networks [ADHV19, HRSS21, Kli09]. Neuberger [LW20b]. Neuberger-Shapiro [LW20b]. Neumann [Bew71, BCT11, DV05, DL15]. neural [HRSS21]. neurons [FL16, FTV20, LM22]. never [Ber95]. Newtonian [vdB07]. Nicodym [Ano77]. Nikodym [Tor76]. nilpotent [Rab94, Rau04]. nilpotentes [Rab94]. nilpotents [CR78]. nilvariétés [CM70]. Nisio [Tor79]. no [Bou00, Tim14]. nodal [MRW20, RV19]. nodes [Gan02]. Noise [Ber20, ANV00, AB02, BDP95, CDLR17, CR05a, CD19a, CX21, Che17a, Che19, Che20, DDT04, DGG21, DVNZ20, EF16, Enr96, FP17, Gir02, HT04, Hos18, HW22, HLN17, KMS14, KP19, LMW21, RSH19, Woo16]. noises [AP05, Tsi02]. nombre [Ber88, Gra84]. nombres [Alt87, DST90, Gal96, Léo86, RE79]. Non [BM93, CE00, DV05, EFG^{+}20, FPV88, Gas21, HMP20b, KP19, NR17, RT18, ADGP13, Ben67b, Ben67a, Bli14, BRT07, BCR15, BB78, Bru93, Cal81, CR05a, CX21, CW18, CW20, CD12, CPV22, DY11, DR17c, Dey20, DGW06, DN01, Dos91, DLP86, DR82, EF05, ET90, Eng09, Eva91, FGN19, FRS22, Fog75, FX19, FFV17, Gas90, Gou05, Gra99, GP74b, Gru66a, Gui02, HK15, Her19, Hey68, JL08, KR16, KLO95, KMS97, Kor15, LS19b, Lap86, Len04, Léo86, Lon69, LR93, LL08, MRTZ16, Maj19, May68, Mél78, MT20, NNT10, OTT02, Páz77, Pin85, Pin14b, Rab94, SP75, SSV03, Szp78, SM79, TTL69, Tor69, Val80, Wu00, May68, vEEIK12]. Non-asymptotic [HMP20b]. non-autonomous [SSV03]. non-backtracking [Her19]. non-central [NNT10]. non-colliding [Gra99]. non-commutative [Bli14, Gui02, Hey68]. non-connexes [Ben67a]. non-convex [CD12]. Non-equilibrium [EFG^{+}20, FRS22, JL08, KLO95]. non-ergodic [Eng09]. non-Euclidean [Pin85]. non-extinction [Eva91]. Non-fixation [RT18]. Non-Gaussian [KP19]. non-generic [Kor15]. non-geometric [HK15]. non-Gibbs [vEEIK12]. non-globally [Maj19]. non-Hölderian [KMS97]. non-homogeneous [ADGP13]. Non-intersecting [NR17, BCR15]. non-linéaires [Páz77]. Non-linear [DV05, BRT07, Dey20, DGW06, EF05, KP19]. non-Lipschitz [CR05a]. non-local [CW18, CW20, FX19, OTT02]. non-Markov [ET90]. non-Markovian [MRTZ16, Pin14b]. non-null [CPV22]. non-parametric [Bru93, LL08]. non-paramétriques [BB78]. non-perturbed [DN01]. Non-polar [BM93]. non-random [FGN19]. non-regular [DR17c]. non-related [Ben67a]. non-reversible [LS19b, Wu00]. non-smooth [KR16]. Non-stationary [May68]. Non-symmetric [CE00]. non-synchronous [DY11]. non-uniform [Len04]. Non-uniqueness [Gas21]. nonamenable [Per00]. nonasymptotic [CDG11]. noncoalescence [ClJ89]. Noncommutative [RW17, Zen15]. noncompact [Dar96, Kig13]. nonconservative [BP95]. nonconstant [Ber85b]. Nonconventional [Haf20, Kif14]. nondeterminism [ASX11, Ber83]. Nonequilibrium [JL06, JLS13]. nonequivalence [RSH19]. Nonexistence [Tia17]. nonlattice [BDZ16]. Nonlinear [Gra92, AZ97, Bud03, CK17, Dos91, GS10b, Kuk13, MN18, Rab94, Tor17, Zhu14]. Nonparametric [ADGP13, BG20, CGC10, Gug12, FFJ19, GGS15, GLM18, Hof01, KP12]. nonsingular [Kop82]. nonstationary [MR90, Mél06]. nonuniformly [MZ15]. norm [DW11, Mat93, Rhe84]. normal [Gau20, NPR10]. normalité [MR89]. normality [MR89]. normalization [HH81]. norme [Mat93]. normed [Her08a]. normées [Mén78]. norms [AGL94, CH90, Kol92, LT06]. Note [ADR67, Gil66b, AC07, Bel06, Bor77, Bor96, MS04, ADR67, Gil66b]. Notes [Bre87]. notion [Gui90]. nouveau [Dos91]. nouvelle [Jum80, Lem73a, Lem73b]. noyau [For73]. Noyaux [DTMS98, Alv73, BN90]. nucléaire [Fou84]. nucléaires [Mar88]. nuclear [Ust82]. Nucleation [Com87, FINR06]. null [AMD17, CPV22, FMM92, Yan21a]. null-space [AMD17]. nulle [Bab91]. number [AD14, BL95, Mai13, Riz15]. numbers [DLN21, DST90, EW06, Eng08, Eng09, Gal96, Gui00, HKK10, JS20, KK03, RT10, Sep00, Yao18, EHK10]. numéro [SL95].

oblique [And09, Kob13]. Obliquely [CR20, Sar18]. observables [DLN21]. observation [Tou69]. observations [BG20, BMP09, BG16, EM83, Gob02, GJTZ08b, GJTZ08a]. observé [SG02]. observed [CDM21, SG02]. obstacles [ÖÇE17, PX05, vdBPSV08]. obtained [Lin77, Yor93]. Occupation [Cha94, GG79, Ako93, BL14, Gou07, HL05, KM17, KL20, MM09, MN18]. Occurrence [AD97, AD00]. Ocone [Ané01]. octopi [AK20]. Odd [PS14]. ODE [Kur10]. œuvre [Bre87]. off [Tsi02]. off-white [Tsi02]. offs [CH17]. Once [KSS21]. One [Ami17, Bré09, Che06, Hol17, Pru97b, ACLS19, ABO16, ASS90, AD13, BLS22, Bar15, BO11, Bod97, BCT11, CGPP13, CDL20, DS07, DW11, FG12, FS17, FGS19, Gan02, HHK06, IPP08, JLS13, LS19b, LY19, LLL11, LL08, Mar00, MWRZ04, Mic17, OTT02, Pet09a, PS13, Pin10, Pov98, STW00, Véc19, YYY10]. One-dependent [Hol17]. One-dimensional [Ami17, Bré09, Che06, Pru97b, Bod97, BCT11, DW11, FGS19, IPP08, JLS13, LS19b, LY19, LLL11, LL08, Mar00, MWRZ04, Mic17, Pet09a, PS13, Pin10, STW00]. one-point [BLS22]. one-sided [AD13, Véc19, YYY10]. one-way [Gan02]. ones [BRR02]. Only [Has90]. Onsager [BRT03]. open [BW19b, Koz11]. opérateur [BKRR71, CSC90, LM76, Rau92]. opérateurs [Dho72, Her94, TTL74]. operator [Ast76, CSC90, Kan72a, Kan72b, Lin77, Rau92, Yur08]. operator-valued [Kan72a, Kan72b]. Operators [Hor74, BHL93, CW18, ER13, Her94, Lin75, Lin76, MS17, Sab01, CW20]. Optimal [GS18, GS10b, Hue16, Ler12, ÖE19, RV13, ÁEdBCAM11, ABCF21, Cat91b, De 18, DGKP21, Pra07, Tro96, Deg79b]. optimales [Bro83, LM77]. optimality [IL06, LM13]. optimaux [Deg79b]. optional [Mal96]. optionnels [Mal96, Mer83]. Oracle [Lem16]. Orbit [Def10, Geo97]. orbites [BL12]. orbits [BL12]. order [ACJ20, DY11, Fer15, FWY00, GNRV05, KM02, Pay67, RSX21, Wer80, vW83]. order-constrained [KM02]. orders [JJ18]. ordinaires [Dos77]. ordinary [CG96]. ordre [BLT69, Hei78, Pay67]. organisation [CF87]. organized [CGK20]. oriented [vdHS03, vdHdHS07]. Ornstein [CCS92, El 00, EPP06, Fer90, FG79, HLT16, Shi20]. Orthogonal [FRS22, NR17, Led09]. orthonormal [GLTZ17]. oscillateur [Enr96]. oscillating [BF95]. oscillator [Enr96]. oscillatory [GB16, Uek99]. other [Bud03, CFP13, GR11, KM99]. out-degree [Riz15]. outcome [AS17]. outdegree [CDL20]. outdegree-one [CDL20]. outerplanar [Car16, Stu17]. Outliers [COW20]. Overcrowding [HV17]. overlap [PZ21].

P [Kyp04, Kyp04]. P. [Der90, Duh97, Gol89, Mat98b, Yor93]. packing [BS13, Sta90]. Pages [Mat98b]. paire [Lus78]. Pairing [Han17]. pairs [Kas20]. Palm [Del77a, GH73, Han73, Tor69]. panmictiques [GBB69b]. Paouris [Fle10]. paper [Bal04]. Parabolic [Che19, Che20, Bor96, BCT11, CCP12, CK12, Che17a, EdHM14, GS15a, GP01, GK09, GB16, HL22, HLN17, MP19, MOS11, MP00, PPR97]. paraboliques [GP01]. Paracontrolled [MP19, FHSX21]. paradox [FI93]. paradoxe [Lem73a, Lem73b]. paralyzing [vdBPSV08]. Parameter [GKMMR20, Bli14, CN05, Has90, HS96, LW20a, Mik88, Mil83, NS81]. parameterization [RZ17]. parameters [Deg79b, Gan67, Gol89, IKM20, KM02, MR22, PD14]. paramètre [Cai82, Has90, Led81]. paramètres [BL79, Bro83, Deg79b, Gol89]. Parametric [Fer15, Bru93, BT08, Gas90, GR08, Hof01, LL08, KMB92]. paramétrique [Bru93, DR82, Gas90, KMB92]. paramétriques [BB78, DLP86]. parametrix [HM16, KMM10]. parametrization [HLS22]. Pareto [GL16, OR18, PT02]. parfaites [Gru66b]. part [AB18, MS96]. partial [Ako93, AB02, CKL14, Dab14, DG19, DVW00, DDG13, Gin80, Gin81, Jac90, LL15, PPR97, SSV03]. partially [MS17]. Particle [BBS15, BGRS14, BB18, BT18a, BT18c, BMP94b, CDG11, CDGR20, CKL14, DKP16, FGN19, FRS22, FNS00, GS10a, GLT09, JL06, JL08, JLS13, KLO95, KL05b, Lan91, Léo95, Lou05, MO07, Mou11b, RCG18, RT97, Saa90, Set01, Set07, Var95a, Völ16]. particles [BRT14, DW08, FK95, Gra99, Gra88a, KPS16, KL95, NO88, Osa98, Rou80, Sar17]. particular [For68]. particule [Saa90]. particules [CK77, KL95, Loo77b, Rob87, Rou80, Rou86, Saa88]. particulier [For68]. particuliers [GLS73b]. partielle [Cat86, Jac90]. partielles [Ako93]. parties [Gil66b]. partition [AJ76b, BA09, CN21, KLS21, dT16]. partitions [Ahn20, CW13, Dim18, EG08, Eva97, Lab14]. Pascal [JdlR04]. passage [BW19a, BR08, CS21, CT16, CG21, CH12, DHHX20, DS90, DT20, DM04, Lac91, Lac97b, Nak20, RT10, Tes18, Val89]. passé [Mey91]. past [CGL21, Zer05]. Pastur [BG10]. Path [BLM21, CGK20, DR17c, Duq03, Bel08, BB11, BK21, Čer04, DS90, DMY91, FL20, FV05, LS14, MRTZ16, She18, Wan07]. Path-dependent [DR17c, MRTZ16]. path-functionals [FL20]. Path-space [CGK20]. paths [BSC04, BAC21, BC19, BCR15, Cha91, FR14, GQ16, HK15, HR12, LRZ19, LT20, MSS06, PW02, STW00]. Pathwise [And09, DOS18, KR16, LR19, Fou13]. Paul [BY87, Bre87]. payoff [COBZ21]. PCA [BGKL17]. PDE [FHSX21, IPP08, MP00]. PDEs [BRT07, DLZ20, FX19, HM18, HLOT^{+}19]. peeling [CL17]. Penalisation [YYY10]. penalisations [RVY09, Yan13]. Penalising [KLL21]. Penalized [Spo17, Bir06, Kol09]. perceptron [MCRT06]. Percolation [CT18, HPS97, Lal98, Per00, AS18, Abe15, AMS95, BW19a, BS17, BR08, BS13, CS21, CP01, CT16, Čer04, CG21, CH12, CDL20, DHHX20, DvdHH13, DT20, Gol18, GW18, GMSS18, GM18b, GH12, Hir16, HLS22, MV20, Nak20, NS95, PSS09, RT10, Tes18, vdBvE22, vdHS03, vdHdHS07, vdHS14]. Percolations [AC15]. percolative [DRS14]. Perfect [FPW95, Gru66b]. Performance [LC20, MGY20]. perimeter [ACRC17]. periodic [BLS22, KP96, KOS16, LS19b, Pèn19, PT08]. periodicity [Kom92]. periodogram [DGW06]. perméables [Wei84]. permutable [Car80]. permutahedron [Mat03]. Permutation [CH67, AD15b]. permutations [BB17a, BDS21, FP08, RZ20, SZ16]. perpetual [SY05]. Persistence [AMZ21, DDG13, Dev16]. perspective [BS13]. Perturbation [CW18, BDN20, CS02, DD10, Val06, CW20]. Perturbations [Szn80, ACLS19, Bat00, BF95, Dos80, EHJ07, HHM21]. Perturbed [DN01, DZ05, ER13, CPY04, CDH00, DDT04, Kom92]. Perturbing [Bau13, BS13]. Pétersbourg [Lem73a, Lem73b]. petite [Lus78]. petites [Dos80]. Petrovskii [HHK06]. peut [dlRd03]. Pfaffian [BCQ21]. Phase [AKM21, CP01, Was15, Bro20, DH21, GI05, MV20, Pin14b, Yan21b]. phénomènes [LR83]. phenomenon [GHR01]. phénotypiques [GBB69b]. physics [Hut19]. Pickands [Wüt04]. Pickrell [Ass20]. pieces [Wat87]. Piecewise [DGM21, BLMZ15, MH18]. pièges [BD96a]. pinning [BT13, BS10a, BS11, Pet09b, Poi13, Yan21b]. Pinsker [Ros86]. Piscounov [HHK06]. Pisier [Bas82]. pitchfork [CDLR17]. pivotal [HLS22]. place [BDEG88, BDEG89]. place-dependent [BDEG88, BDEG89]. placed [Gan02]. plan [Wer93]. Planar [CGN16, CH67, Pèn09, Abr16, BC20, Bef03, Ber15a, Bet15, Cha94, Dub04, GMV93, GMS19, HNPS19, MV20, SY11, Sep20, Val89, Wer93, Zer05]. Planck [DMGM18]. plane [Aar78, ARRS16, Ahn20, AC15, Chu84, DFM^{+}03, Dim18, FMM92, FMP14, Mén10]. planes [Val89]. Plans [GLS73a, Páz77, ÁEdBCAM11, Col78, GLS73b, JT76]. player [CGL21]. plongement [Val78]. plus [Hen91, Le 89, Lus78, Rou86, Woo16]. plusieurs [Cup69, Gar79, Mat93, Tou69]. poids [Mat93]. Poincaré [Ano05a, Bal04, BW06, Mat98b, She04, Ané01, BR08, CGZ13, Che87, Cou20, Goz10]. Poincaré-type [Che87]. Point [BHR17, ARRS16, AR08, BLS22, BESY08, BR15, BMP94b, Bor08, BCQ21, CFP13, ClJ89, ET90, FK93, For68, FSP71, FT05, GLT11, GM20, GJTZ08b, GJTZ08a, HPT02, Har15, SY13, Smy75, Tor17, Val06, Wat87, Zha22, Zhu14]. point-process [HPT02]. Points [Roy73, BDCKL20, BS10a, Bla73, BT00, BL95, BSZ03, BM93, Che18, CH17, Dav05, DPRZ00, Eld14, Gir02, Han17, HLLS22, HLS22, Jac18, NRS10].

Pointwise [Che17b, CL09, Ass98, BLR16]. poisson [Bou00, AAH21, AT20, AAP09, Bal74, BT16, Bro20, CK12, Coq93, CDL20, Cup69, DS06, DW11, FLT04, FFW05, For68, FLRB11, FT05, GC03, HPPS09, HMP20a, JdlR12, JM93, KF98, Lam21, NdS20, Nic03, Pic96, Pra92, Pri12, Tor17, Tor69, ZBH17]. Poisson-weighted [HMP20a]. Poissonian [Lac12a, Lac12b, MW02, ÖE19, Wüt98, Wüt99]. poissonien [BL79]. Poissonization [Zai02]. polaires [Roy73]. polar [BM93]. policy [LC20]. Polish [GZ19]. politiques [LM77]. Pólya [Cas17, CM21, Pou08]. polycyclic [Tho13]. polydisc [JM93]. polygons [BST20]. polyhedral [LR19]. polyhedron [And09]. polymer [Ass11, Che08, CN21, GPdS16, LZ22, NS22, Viv21, Yan21b]. polymers [BA09, Mej04, MT08, QR15, Var06, Zyg13]. Polynômes [Sch00, Dub75]. polynômial [DL96]. Polynomial [LLL11, DL96, FRS22, KK19, Mok90]. polynomial-time [DL96]. polynomially [SZ16]. polynomials [ADL16, BMW21, Han17, Har15, Sch00]. polynomiaux [Mok90]. polyspectra [CM15]. ponctuel [Bru73, DL96, Sch72a]. ponctuelles [For68, FK69, FK75, Szn80]. ponctuels [CGR79, Han71, Hil73, Les87]. pondérés [BD92]. pont [Bia86, LR96, Rec86]. population [BY08, Cer96, Gil66a, Lot91]. populations [GBB69a, GBB69b, LR83]. Porous [DGG21, BRR17, HQZ17]. portée [Gui97]. portfolio [JYZ05]. portion [Qia19]. posed [MS17]. posedness [Hos18]. positif [Ber91b]. positifs [Émi81a, Hen91, Her94]. position [LR17]. Positive [Gan67, AZ14, AW18, Ast76, BB14a, Ber91b, BE95, BHL93, CC08, Cau65, GLL18, Hen91, Her94, Hög74b, KP11, Mai78, Véc19]. positively [GW18]. positives [BH81, Cau65, Fli75]. Positivity [Vys14, CK17, Fou01, JS20]. possibly [Rho09]. posterior [Cas17, CM21]. Postulats [Gil65a]. Potential [Eva87, GK09, AHK77, Ast76, CCP12, CK12, CD12, DMGM18, EF05, GB16, HH09, KP96, Lac12a, Lac12b, MO07, NdS20, PT21, Wüt98, Wüt99, DTMS98, MW02]. Potentials [KRS18, BCR09, Bod97, Gar99, Pet09b]. potentiels [DTMS98]. Potts [CP01, EM15, Häg03]. pour [AJ76b, AJ76a, ACA02, Bab88, Bir84, BDC68, BL79, Bro83, Bro75, Bru93, Bru73, Bru76, Cat86, CK77, CR80, Cun03, DCS74, DL96, Der76, Dos80, Dub75, DST90, Ehr86, EV03, Fou84, Fra91, GP74a, Gol89, GP74b, Gui97, GH88, Har01, Hei71, Hei78, Her05, Her08b, Jac71a, Jac73, Jac74, KL95, KMS97, LS91, Lap86, Le 02, Lem73a, Lem73b, Léo86, Les87, LR93, Lot91, Mai78, Mas93, Mat93, Mét72, Mét84, Mic99, Mic09, Mus06, Nua84, Páz77, Pet96, RLT81, Rei82, Rev83, Rio02, Rou87a, RE79, Roy74, Saa88, Sch72a, Tou83, Tou69, Wei84, Wu91, dZ81, dlR02, lJ87]. poursuite [BD92]. power [HJY11, HP17, NNT10]. powers [CL13b, LW96, CSC90]. PR [BW06]. Precise [BM16]. Précision [GLS73a]. Précisions [AJ76c]. Prediction [Rao72, Bru76]. Préface [SL95, GMP02, SL95, GMP02]. premier [Lac91]. prescribed [Dar96]. presence [BY16, CX21, KP96, Rou87b]. preserve [Att95, Has90]. préservent [Att95, Has90]. preserves [BB17b]. preserving [Hu02, KT97, Lac97a, Pro86]. presque [Del64, DST91, ERR71, MO83, Nev72a, Sen90b]. presque-borélien [ERR71]. presque-sûre [Sen90b]. preuve [Duh97]. previsible [Tuy81]. price [AI05, AS94]. prices [BRT14]. principal [DF20, Yur08]. principales [Ber89]. Principe [DP84b, Rab94, Dos91, Mét84]. Principes [Le 02, Mas93]. Principle [Kob13, And14, BDMBP18, CFP13, CK17, CCKW21, CD16, CM08, CP04, DMP14, DDP87, DK08, DGR18, Dos91, DV08, Enr07, FV07, GZ19, Gou07, GI05, MN18, MR90, Osa98, Rab94]. principles [CC08, DMR95, Le 02, Mas93, Mie08]. prior [Dal20, EHJ07]. priori [BPTZ18, GBB69a, GLS73a, GLS73b]. priorité [Jai80]. priors [SJ14]. prix [AS94, Lic69]. probabiliste [Duh97]. probabilistes [Bro75, BR74, DCS74]. Probabilistic [BG85, MM14, PPR97, PT21, BRR17, BRT07, DV05, Duh97, EF05, Fla02, HHK06, Kyp04, Rau09]. probabilité [AJ76a, Bar81, Che76, CG92, DL84, FK69, Gra88b, Her05, Hög74b, Let72, Mas93, Tor65, Yor75]. Probabilités [Ano05a, Bal04, Gru66a, Gru66b, Rou87b, She04, Arn93a, Bec74, Bod70, Bod71, DR69, Fis73, For73, Hey68, Pel90, RF64, Rev83]. Probabilities [Haz93, ALT99, BDEG88, BDEG89, BS17, BDCH16, BCL15, Che17b, CH20, Dub04, FGM09, FK94, FW07, Gru66a, Gru66b, Lif94, LW96, Mat03, Pel90, Pru97a]. Probability [DP83, Hu02, KBR71, Mat98b, ABL08, ACJ20, Arn93a, BESY08, Bel08, Che87, CG92, CH12, ERS15, GHS11, GBR73, Her05, Hey68, Liu96, LMD03, Mas93, NZ13, Pem97, RF64, Sam07, WW16, vdHdHS07, Tor65]. problem [AEFdR20, ACLZ20, AD13, BDT11, DCG90, DM09, Fun99, Gra88a, LM13, Sen90a]. Problème [LM76, Sen90a, Ben79, BDNT74, DCG90]. Problèmes [Loo77b, Ant82, DCS74, FK75, Jum80]. problems [GvdVY20, Kol03, LS20, MS17, DL15, Sta90, Tra15]. procedure [LM13]. Process [RAS11b, Zhu14, AD09, Ade76, Ako93, AR06, ABGM21, ABK15, BMRS17, BT18a, BT20, BLNR11, BRZ19, BF88, BEV13, Ber85a, Ber85b, BGJS18, Ber68, Ber91b, BHR17, BESS20, Bru93, CK20, CG14, CL17, Dav05, DPS89, DS90, DDP87, DJO19, Dep65, DR22, DT03, ELS20, Eva91, EO22, FPV88, FGL95, FM09, FG92, FLRB11, GW18, GM13, Gui97, GL89, HPT02, HV17, Imk84, Ist92, IL97, Jac90, JL08, KMB92, KRS18, Lac17, Le 91, LS06, Liu96, LT06, Mas88, MØP04, Mik88, MR89, Mou96, Pel90, Pèn09, PS19, RN21, Rez02, Ric13, Rou80, Rou97, Saa90, SZ17, Sep20, Soh13, Sta90, Sue05, Tal88, Val12, VW15, Vig03, Yan21a, YYY10, Yan13, Zen16, dHdS14]. Process-level [RAS11b, Zhu14]. Processes [CCS92, JS96, AZ14, AW18, ADH12, ARRS16, ABKV14, AHK77, AP98, ASS90, AD15a, Ass20, AD13, ADGP13, AKDR66, ADR67, BP15, BV17, BDS87, BKS12a, BB21, BB14a, BBCH21, BDEG88, BDEG89, Bar00, BM09, BCR09, BKK10, Bax96, BT18b, BLMZ15, BCV21, BR15, Ber83, Ber95, BD97, BL05, Ber20, BCR20, BD16, Bor08, BD11, BW19b, BDM11, BCQ21, BKS12b, CGB17, CFP13, CPY04, CDM21, CL21, CLB98, CL94, CL95, Cha91, Cha12, CM16, CLP14, CCM19, CZ02, CR05b, Chy08, CG19, CD19b, CGPP06, CM08, CGC10, CGG11, CT11, CG14, Cro18, CH90, Cue18, Dav02, Dec05, DS06, DR00, Def10, Deh00, DG01, DD07, DGW06, DN01, DM04, DZ05, DMR94, DMR95, DLS02, Duq03, DGM21, DF87b, Egg80, ER13, Eic01, ET90, Émi85]. processes [EP06, Erb14, Eva87, FGM09, FHL12, FT12, FS89, FW07, FW09, FM19, FR03, Fou13, FFV17, FJR21, FV07, FT05, Gan67, GKV03, GCJ93, GG08, Gol95, GS15b, GJ11, Gra85, Gug12, Har15, HKK10, HS17, Her77, Hey11, HS96, HK09, Ish97, JLS13, JLT21, Jeu93, Jin14, Jiř70, JS20, KK14, Kom98, KMM10, KL20, Kun86, KL10, KP11, KPR18, KRS21, Lab14, LS16, Lam00, LS14, LM97, LL15, Li19, LMW21, LS05a, LM09, LZ15, Lot91, Lou00, MRZ98, MS70, MP84, MR94a, MU11, Mas86, May68, MZ15, MRT21, MOW09, MS83b, Mim14, Mok90, MW07, MW11, Num78, Osa98, Ove94, PPR18, PW79, PS09, Pie09, Pig18, Pin90, Pou08, Pri68, PZ04, RDL20, Ric13, Rio93, Rio02, Riv07, Riv12, Rog84, Ros01, Ros86]. processes [SS10a, iSW04, iSW05, ST20a, SW11, SY13, Sen17, She18, Shi20, Smy75, SV08, Son13, Str88, Tak70, Tal89, TTL68, Tor17, Véc19, VAG86, VA89, Wat87, Wer80, Wu99, Wu00, Yan18, Zha13, Zhe85, Zhu14]. Processus [Ben79, Bla73, BKRR71, ERR71, Émi81a, Gho75, Han71, Hil73, Jeu93, KRR70, Lot85, May68, Oga74, Pri68, Rou86, Saa90, TV81, Ako93, ACTR85, Ant82, AKDR66, ADR67, Ber68, Ber91b, Bou88a, Bru93, Bru76, Cai82, CL84, Car81, Cau69, Cha86a, CK77, CGR79, CR79, CR80, CF87, Dav74, Dav02, Dep65, Der73, Der76, DP84b, Die81, Doz83, DR82, DR69, Fli75, Fou84, FlJ84, Gal76, GP74a, GP74b, Gui97, GL89, Hae72, Hai87, Jac71b, Jac71c, Jac73, JKM82, Jac90, KMB92, Léa85, Léa88, Len77, Lot91, Mar88, Mél78, MR89, Mok90, Mou96, Nev86, Pel90, Rec86, RR70a, RR70b, Rev83, Rio02, Rob87, Rou87a, Rou87b, RE79, Rou80, Rou97, Str88, TTL68, Tou83, Vig03, Web80, Wei84, Yor77, dZ81]. proches [Rou86]. Product [BB18, BHS12, Chy08, GR11, Hut19, LP97, Mat93, Sam07]. Products [KK19, ARRS16, ACLS19, BJM10, BM16, COW20, Gui79, Hen91, JD18, Le 89, LWZ16, Per00, Tor65]. produit [CL84, Fer84b, Hil78, Mat93, Pel79]. Produits [Gui79, Hen91, Le 89, Led70, Roy80, Tor65, lJ87]. profile [HLN06]. profiles [GK17]. progeny [MSZ15]. progressive [Bla99]. projecteurs [Ngh70]. Projecting [NR03]. projection [ACJ20, Car85, Jui16]. projections [AKSS07, GNS13, HU09, KF98]. projective [Le 91]. projectors [KL16]. Prokhorov [Alt87]. prolongements [KRR70]. promenade [Jan70]. promenades [Rei82]. Proof [HP17, Sep20, AMD17, BB18, Duh97, Jiř70, Kel17, Sla22]. Propagation [GQ15, JM98, AZ99]. proper [HH81]. Properties [SY05, AI05, Ast76, BLMZ15, BT00, Bud03, CM10, Che97, DM09, DPSZ16, DRS14, DST91, EdHM14, FG92, GGL15, GGS15, GL16, HSW20, Kaz71, KLO95, KM14, LMX09, Lin77, LZ15, Mal96, MV92, Mok90, Sam07, Smy75, Sue05, Wan07, vEEIK12, CGN16]. property [AABV16, AMD17, Bat00, BF92, El 00, EP06, GG08, Gob02, GK09, HM18, HK09, KM17, Lab14, LS91, MU11, NN81, Ros86, SW11]. propos [Bar81, Phi71, Yor85]. propres [Lou68, TTL74]. Propriété [LS91, Ano77, BF92, Cha86b, FSP71, Tor76]. Propriétés [Deg82, Doz83, DR69, DST91, Mal96, Mél78, Mok90, Rao70, TTL74, Ber70, Bra73, BH81, Cal81, Cha70, Die82, GL80, MV69]. pruned [BRT07]. Pruning [ADH12]. pseudo [Bas89]. pseudo-aléatoires [Bas89]. pseudorandom [Bas89]. puis [Vig03]. Puissances [CSC90, DL84]. pure [BS03, CGC10, Wu10]. purement [Reg70].

QFTs [ADG21]. quadrangulation [GM19]. quadrangulations [ABW17, Bet15, Mén10]. Quadratic [EM83, IL97, BD16, Cha91, DHR11, Gar99, LW20b, R.83]. quadratique [DLP86]. quadri [dT07]. quadri-tilings [dT07]. Qualitative [BLMZ15, Bos75, Car15]. quantile [CH90]. quantisation [Hud05]. quantitatifs [ACA02]. Quantitative [Car15, Dar21, GM18a, PS09, ACA02]. quantization [ADG21, DSS13, TW18]. Quantum [AC04, LS05b, AKM21, AKKR01, AK21, AG21, AB18, BW15, Ben20, Ber15a, BP95, Col04, GSS05, GM21, HP21, HRV18, MS19, Par05, Reb05]. quarter [FMM92]. quartic [DGR16]. Quasi [DO20, Her08a, Lin75, RV19, AFGJ16, AC04, BCV21, BR74, CCM19, FS17, Fit98, For78, Her94, Her05, Oka85, Tes06, Yor75, Tes06]. quasi-Bernoulli [Tes06, Tes06]. quasi-compacité [BR74, For78]. quasi-compact [Her94, Her05]. quasi-compacte [Her05]. Quasi-compactness [Her08a, Lin75]. quasi-compacts [Her94]. Quasi-independence [RV19]. quasi-invariant [Oka85]. quasi-left [AC04]. Quasi-static [DO20]. quasi-stationary [AFGJ16, BCV21, CCM19]. quasi-sure [Fit98]. quasilinear [DG19, DHJ^{+}21]. que [Sch85]. quelconque [Jac73]. Quelles [DS84]. Quelques [ACTR85, Ass87, Ber70, Bra73, BR74, Die82, Dos80, Fis73, GL80, Hac82, Hil77, Sch72b, Bou80, BH81, Cal81, Nua84, DP75]. Quenched [CCKW21, DGR18, Eng08, JL08, Pet09a, BS10a, CP17, DMP14, DK08, EdHM14, GS15a, GPdS16, MM18, PS13, Poi13, RAS11b]. question [Heu79]. questioned [Heu79]. Queue [Fli81, CG68, Coh72, GI80]. queueing [Coh68]. queues [CP78, FM09]. qui [Att95]. quotient [Mét72].

Rademacher [DK19, KRT16]. radial [ER13]. radii [BGBK20]. radius [BHS12]. Radon [Ano77, SO75, Tor76]. ramification [DR82]. ramifiés [Bla73]. Random [ABL16, BP92, BR15, BK04a, BK04b, Bré04, Car13, CCFR09, FS17, FMM92, FP08, KKW15, LPWR94, MRT21, NS95, RZ20, RT18, RSS16, Sch09, Tel08, vdBPSV08, AL12, AD08, ABW17, Ade76, ARRS16, ABKV14, Aid10, ACLS19, AD01, AEKN19, ABO16, AAD19, AC18, And14, AC15, ADHV19, AHOZ13, AD15b, AGJ^{+}20, AI99, AD00, AC07, Ass10, AAP09, Aug18, AC99, Aus15, AFJV15, ADL16, BGRS14, BT15, BB14a, BC17, BPS12, Bar00, BB17a, BGHR19, BC20, Bau13, BHS21, Bec89, BMW21, BHLP19, BQR03, BGBK20, BKYY10, BAC21, BY19, BBHK08, BT13, BGN19, Ber13, Ber12a, BD97, Bet15, BG99a, BHR17, BS10a, BS11, Bla03, BL95, Bod97, Boi98, BMP94a, BMP94b, BSZ03, Bor21, Bor87, BF95, BHS12, BDZ16, Bré09, BDC67]. random [BDC68, BK95, Bro03, BW05, BW06, BBD12, Bro20, BJM10, Bud20, BK18, BM16, BZ12, CD01, CDLR17, Car16, CC08, CH14, Cas04, CGPP13, ČTW11, Che06, Che97, CR05b, CY07, CW13, Che17b, CH20, CCKW21, CD16, Col04, CP12, Coo19, Cos21, COW20, Cro08, CRR98, Cun03, CL17, DV98, DP83, DMPU09, Def10, Def16, DGZ03, DMS05, DRZ21, DW15, Dep65, Dep99, DK08, DGR18, Dim18, DMSY94, DRS14, DE16, DLZ20, DF87b, EV03, Eld14, Eng08, ESYY12, ES17, EdHM14, Eva94, FK93, FW04, FL21, Fer90, FLT04, FGN19, FHS19, FNS00, FVY06, For68, FLP06, FFV17, FP18b, FHSX21, GS97, Gan13, GP20, Gan02, GKS07, GPS10, GHS11, GP11, GGL15, Gau20, GKV03, GM18a, GS10b, GW18, Gra99, GN02, Gre08, GK17, GB16, Gue99]. random [GPP10, Gui00, Gui02, GL16, Guo14, Gur14, GMS19, HLNV13, HK02, Han17, Hen91, Her19, HP00, Hol12, Hu16, HL22, Hua19, Hue16, IPP08, IR98, IY20, Jaf12, JD18, JRdlR10, JP17, JJK^{+}14, KLS21, Kes86, KM99, KK19, Kig13, KSS21, KMP14, KLL21, KK03, KM11, KP17, KOS16, KOS19, LLO06, Lam20, LTT18, Le 89, Led92, Leg89, LT21, LW20b, Liu96, LWZ16, Lou05, LV20, MM14, MM18, Mal11, MS04, MR94a, MR97, MRW20, Mat93, MR90, MWRZ04, Mej04, MR94b, MO07, Mou11a, MSC19, Mus06, Nak15, NO88, NN20, Nev69, OW17, OR18, ÖÇE17, Pai18, Pay67, Pei92, Pèn09, PSS16, Pet09a, PS13, PS92, Pin10, Poi13, Pri12, Pru97a, Pru97b, Ras11a, RAS09, RAS11b, Ray68, Rho08, RV13, RR94, Rou21, Saa95, ST11]. random [ST94, SY11, Sch06, Sch00, She03, She04, Shi10, Shi19, SS10b, Sim07, ST20b, Sta89, SZ16, Stu17, Stu19, Tho13, Tim14, Tou15, Tra02, Tuy81, Val89, Var95a, Var06, Viv21, Vys14, Was15, Woo16, Yil10, Yur08, Zer00, Zer05, ZBH17, Zhu21, Zyg13, de 95, dlP94, dHdS14, Ahm65, Con65, Del64, Kah87, Leg89]. random-walk [Dep65]. Randomly [Ors82, AKSS07, Gan02, GG01]. rang [Fer84a, Fer85]. Range [AC18, Ami17, ACTR85, BMRS17, BT20, BF88, BKYY10, Bré09, CFP13, CCKW21, CM08, Com87, DLS02, FGS18, FGS19, FPV88, FHS19, Gui97, HK02, Hey11, JLS13, Lac12a, Lac12b, LM97, Léo95, Poi13, Set07, Sud21, Saa90]. ranges [Mar15]. rank [ACLS19, BT18b, DS07, KL02, PRS13]. rank-one [DS07]. ranked [Saa90]. rankness [Dal20]. rapport [Nev76]. Rare [Yan21a, AD97]. rarefaction [FK95, FGM09]. Rareté [FK93]. Rate [FP17, ABCF21, ADGP13, BCL15, CL09, FG12, JRdlR10, KM20, LR17, Ouc05, Sar21, Tho13, Tim14, Zai02]. Rates [GG02, Jac98, Mas86, AMS95, BDS87, BL02, CG21, DFM16, EM15, FR14, GG01, GLT09, KL20, Vay03]. ratio [Fit98, Gas02, Lin76, Num78]. rationnelles [Fli75]. Ray [BP15]. RBM [LRZ19]. reaction [CR05a, LT18]. reaction-diffusion [CR05a, LT18]. Real [Lac04, BU18, Duq03, EV03, KM20, Riv07, Rog84, SS18]. real-valued [BU18]. realization [KT97]. Réarrangements [DV98, DV98, Pru97b]. Recherche [GLS73b]. reciprocal [Con18]. Recollement [RR70a]. reconstruction [Cha21]. Records [HMNP98]. recouvrant [FK93]. recouvrement [FK93, FK75]. recovery [GKMMR20]. rectangles [ABL16]. rectangular [BG10, BK18]. Recuit [Mic92b, Con94, Mic92a]. Recurrence [ANT17, BH20, CP78, AKDR66, AKDR66, FMM92, PS09, PS92, Pin10, Pin14b, Loo77b, Sun87]. recurrent [ADR67, CG04, CRR98, GZ98, GPS10, LL08, MR94a]. récurrente [Mai78]. récurrentes [Bra73, Mét72, Sun75]. récurrents [ADR67, Bal81, DR69]. recursion [CR96, Led09]. recursions [GL16]. recursive [Ber15b, HMP20b]. reductions [LP90]. réduites [LP90]. réel [Bia86]. réelle [CR79, CR80, Deh74, Jan70]. réelles [Pet79]. réels [EV03, GP74a, Ouv73]. réfléchi [Ber89]. réfléchies [LNS89, Pet96]. Reflected [Dub04, AI99, Ber89, BM93, CR20, Del08, DZ05, FGV16, Gas21, LS19a, LNS89, LR19, PW94, PX05, Pet96, Sar18]. reflecting [BBC05, CF09, ClJ89, Sło93, WZ90]. Reflection [STW00, And09, BDT11, Gra88a]. reflections [Kob13]. Reformulation [Rot90, Rot90]. Refracted [KL10]. regeneration [DGM08]. regenerative [LS14, Num78]. regime [AKKR01, Che17a, CP12, GP14]. régions [Hil77]. Régression [CPP71, Alq08, Aud04, BK21, Bla99, BT08, Dal20, DST90, DT03, GGS15, GLM18, IL06, KP12, DST90]. Regular [LY19, ČTW11, Coo19, DR17c, DMR95, GP20, Gir02, Lim12, Tel08, Tor65, Gru66a]. régularisant [CSC90]. Régularité [BN90, Le 89, Léa85, Léa88, Ano77, CG92, Fer90, Tor76]. Regularity [Haa04, MRVZ16, Nag05, Sue05, AC99, Bel06, CG92, EH19, Fer90, FGR09, GvdVY20, Lia07, Mil83, OV20, DL15, SSV03, BN90, Le 89]. regularization [LW20a, Mar21]. regularly [BHR17]. regulated [MS04]. régulier [Tor65]. régulières [Gru66a, Lon81, LR93]. reinforced [KSS21, MRT21, RN21, Zen16]. reinforcement [Ber20]. related [Bar00, Ben67a, CCEL13, CPY04, CG14, Gan67, Gil66b, LR96, PW79, RVY09, Shi20]. Relation [Gil65a, Gil65b, Len77, AHOZ13, BG10, FGS19, Gil66c, Lou05]. Relations [Bia86, Dub04]. relationship [Gil65a, Gil65b]. relative [BL03, Hai87, Mou96]. relatives [All78, Del87]. relativistic [Ang16, AT20, Bai10]. relaxation [ESYY12, Liu18]. relevance [BS11, GLT11]. relèvement [Han72, Pel72]. reliability [CTR95]. remark [Bou80]. remarkable [BS03, Yan13]. Remarks [BW01, GH73, Hel99, HL05, BM00, PS92]. Remarque [Mén78, Bou80, RR70b]. Remarques [DC73, Hög74b, Pis84, Sch72b]. Renewal [McD75, MT20, dS03, ADGP13, BDS87, Num78, Soh13, Sta90]. renormalization [DGdH^{+}08]. renormalized [CK12, FR03, NdS20, Ros96, Ros99, Ros01, Ros10]. renouvellement [Bab88, Jac71a, Jac74]. Rényi [BDS87, HL20, JV07, KLS21, ST20b]. répartition [Con65, Lac73, Loo77c, Loo78]. répartitions [For68, FK69, FK75]. repeated [BJM10, De 99]. replacement [de 95, dlP94]. Representation [Leh13, Alv73, Att95, Fou95, Gal76, AK21, BB21, BRR17, BRT07, DV05, DK08, DKGT16, FG12, HLOT^{+}19, Kol92, Kre76, Lin77, Pec01, VW15, Bas69, Ruc80, Tor69, Att95, Fou95]. Représentations [Jac82, TTL68, BCT11, BDM11, Chy08, EH03, TTL68]. representative [GS97, LPWR94]. représentatives [GS97]. reproduction [DR82]. reproduisant [For73]. Repulsion [FVY06]. repulsive [Yan21b]. rerooted [Stu19]. Rescaled [Abr16, FK94]. rescaling [LT21]. réseau [Dep99]. réseaux [Pel79]. reservoirs [BGJS22]. residues [Fra91]. résidus [Fra91]. resistance [Cro18, Tel08]. resistances [Abe15, Dep99]. resistor [Dep99]. résolubles [Bou83]. Résolution [Mac78]. résolvantes [Bra73, Bro75]. resolvent [HLNV13]. respect [CCM03, CM10, CN05, CP11, Dec05, DS06, MSZ15]. response [DD10, FGS19]. respresentation [EF05]. Restricted [CW13, Bas82]. result [DR17c, Duh97, Gui97, Hob94, Jin14, MT08, Pec01, Rau09]. résultat [Bar81, Duh97, Gui97]. résultats [Tes06]. results [Ber12a, BM16, DGM10, DMS93, DF87a, JV07, KMR06, MP84, Tes06, Wu99, Wüt98, Zhe85]. retardants [Bou07]. retardateurs [Bou07]. retards [KMB92]. réticulé [Hei74b]. réticulés [BH80a, Hei83]. Retournement [RR70b, ERR71]. rétrogrades [Ham96]. return [Ass00, Lac97a, Pac00]. reversal [ET90, Zha22]. reverse [HLT21]. Reversed [ST11, De 18]. Reversibility [Her19]. Reversible [Han04, Gol95, LS19b, PZ04, Wu00]. Révész [Yor93]. revisited [CMR02, FG79]. Revuz [Yin96]. Riccati [BDN20]. Ricci [CP11]. rice [AMD17]. Richardson [LPP15]. Riemann [BM08]. Riemannian [BBCH21, CP11, Fra91]. riemannienne [Fra91]. rien [Rev83]. Riesz [BCR09, EF05, GP74a, Hei74a, Led70]. rigged [AHK77]. right [DV13, Mic09]. Rigidity [FP18a]. ring [Liu18]. Rips [IY20]. risk [JYZ05, Kol09, LM13, LW20a, SV08]. risque [Bir84]. Robertson [Ang16]. robust [KP12]. Rolling [HPT02]. Rolling-ball [HPT02]. Romberg [LPP15]. rooted [AABV16]. Rosenthal [RW17, dIS02]. Rosenthal-type [dIS02]. Rost [De 18]. rotation [Uek99]. rotations [BU18]. Rouault [Rou79]. Rough [Tro96, BOZ16, BC19, CL21, Che19, Che20, FX19, FV05, FR14, Gas21, HK15, HW22, LT20, MSS06, BLM21]. roughly [CT18]. roughness [LRZ19]. roulette [Jir78]. round [HNPS19]. Rowlinson [DH21]. rows [BQZ20]. Rudolph [Kre76]. rule [LW20a]. rules [LCO09]. rupture [DP84a]. Russo [BMW21]. RW [DLN21].

s [Bag85]. sûre [Nev72a, Sen90b]. sûrement [Del64]. sûres [DST91]. Saint [Lem73a, Lem73b]. Saint-Pétersbourg [Lem73a, Lem73b]. Sakai [Pia95]. Salem [FS03]. same [Led92, Mén10]. sample [AB97, AG89, AG91, BSC04, CAM98, DS90, KL16, LSX21, MRW20, Mas93, Pan14, Péc12]. sampled [Gug12]. samplers [DKP16]. Sampling [ACJ20, GR11, BBL14, HJY11, KL20, Lim12, de 95, dlP94]. Sanov [KPT20, Din92]. Sanov-type [KPT20]. satisfying [Völ16]. Sato [Han04]. sausage [Eva94]. sausages [vdB10]. saut [Rou80, Rou86]. sauts [Léa85, Léa88, Mou96, dZ81]. scalaire [Bru76]. scalaires [Gui79]. scalar [AWS21, Hof15, Gui79]. Scale [DvdHH13, BH97, BEV13, Bro20, CGG10, Fla02, GOVW09, KLL21, Liu18, Rou97]. Scale-free [DvdHH13, KLL21]. scaled [MW02]. scales [CCM19, DLP07, GvdVY20]. Scaling [AS17, Bet15, Cro18, CL17, GMS19, LT21, Mar20, Mou11a, OR18, Riz15, Sen17, Stu17, VST12, Yao18, dR17a, dT07, dHdS14, ABL08, Bas14, CGN16, Car16, CŚ07, CD12, MST20, MO07, Shi19, Soh13, CLM13]. scanning [Mal90]. Scarcity [FK93]. scenery [AC07, CGPP13, GKS07, GPP10, Gui00, Pèn09]. Schatten [GP07, KPT20, RV20]. schedules [Cat91b]. Scheffers [LCO09]. Schelling [DV21]. schéma [Bou80]. scheme [Bou80, DNT12, MGY20]. schemes [CG96, GN09]. Schilder [LR93]. Schur [Mkr14]. Schweizer [AS92]. SDE [CG19, DR17c, KHT12, Sch89, Sło93, dR17b]. SDEs [And09, AKU17, BDMBP18, BCR20, DNT12, DPSZ16, FX19, FL20, GN09, HP11, HSS21, HM16, KM14, LMW21, Maj19, MGY20, Pri18]. search [CLP14, DF87a]. Second [DY11, FK95, Tor78c, Wer80, BLT69, FGN19, Pay67, Pay67]. Second-order [DY11, Pay67]. section [Mer83]. Segel [GQ15]. segment [Hir16]. segmented [Ben67b]. segmentés [Ben67b]. segregated [Mis16]. selection [Alq08, BB14b, BGH14, Bir06, Cer96, CG14, JYZ05, Ler12, Ver10]. selective [Pet09b, Hil73]. selects [AFGJ16]. Self [BR03, Ber02, GM20, MV07, Var95a, AvdHH19, Ass10, BY19, BCK21, BST20, CP17, CGK20, DCKY14, GH10, Haa04, Hey11, JN19, KZ17, Kur10, KP11, PSS16, Pie09, Sab01, She18, VAG86, VA89]. self-attracting [BY19]. self-averaging [CP17]. Self-avoiding [GM20, BST20, DCKY14, Hey11]. Self-interacting [BR03, Kur10, PSS16]. Self-intersection [MV07, JN19]. self-intersections [Ass10]. self-loops [AvdHH19]. self-organized [CGK20]. Self-similar [Ber02, BCK21, GH10, Haa04, KP11, Pie09, Sab01, She18, VAG86, VA89]. self-similarity [KZ17]. selon [Rou86]. Semi [Arn93b, Cau65, Hor74, Jac73, ST20a, AD15a, Arn93a, Bab88, Bru73, BT08, DDT04, Émi85, Gas90, GP01, Gui90, Hen82, Jac71c, Jac71a, Jac74, Kip74, LP97, Led78, Leg89, Num78, Oga74, PPR97, Pei92, RY95, Rei82, Sch85, Thi81, Arn93b, Gas90]. semi-continus [Thi81]. semi-convex [DDT04]. semi-direct [LP97]. semi-flot [Led78]. semi-group [Leg89, Pei92]. semi-groupe [Bru73, Leg89, Pei92]. Semi-groupes [Cau65, Jac73, Kip74]. Semi-Groups [Hor74, Émi85]. semi-linéaire [Oga74]. semi-linéaires [GP01]. semi-linear [GP01, PPR97]. Semi-Markov [ST20a]. semi-markoviennes [Bab88, Hen82, Jac71a, Jac74, Rei82]. semi-markoviens [Jac71c, Jac73]. semi-martingale [Sch85]. Semi-martingales [Arn93b, Arn93a, RY95, Arn93b]. semi-parametric [BT08, Gas90]. semi-paramétrique [Gas90]. semi-regenerative [Num78]. semi-simple [AD15a]. semi-stable [Gui90]. semiamarts [Edg79]. semicontinuous [HKM14]. semidirected [Zyg13]. semiflexible [CGG10]. semigroup [Bew71]. semigroups [BCR19, BP95, Cau65, GSS05, KM14, LS19a, Lam20, Reb05, Sie84]. semilinear [HLOT^{+}19]. semimartingale [Bel08, GS18, KP12]. Semimartingales [Pro86, CE00, IKPY18, KPP95, Mac87, MZ84, MS96, PSV17, Wat87]. semiparametric [GR08]. semisimple [Tol00]. Semistable [Sie84, iSW05]. sens [Ano77, Bra73, Mai78, Mét72, Tes06, Tor76]. sense [Haz93, Tes06]. Sensitive [Fri67, Fri67]. sensitivity [Her18, KMS14, PSS09]. séparabilité [Bec89]. Separabilities [SO75]. separability [Bec89]. séparables [Ouv73]. separation [GI05, GJN20, Ano77, Tor76]. sequence [BDC69]. sequences [BK95, DR08, DMPU09, GS97, HMNP98, JRdlR10, KRT16, LPWR94, Lou00, MR90, MR94b, MV91, Nak86, Ouc05, Rao72, Tuy81]. sequential [FFV17, KKK85]. Séries [Fli75, Lia16, MW05]. server [CP78, GI80]. service [Fli81, GI80]. ses [Ben67b, Bos75, Pic64, Tor78b]. Set [Cas07, AS18, BDT11, DRS14, DE16, EG08, FP18b, Gua77, HP19, Kig13, LMD03, MV20, Riz15, SS10b, Was15]. sets [Bat00, BL19, CD18, Cha94, DHMP17, ES79, GBR73, GG79, JJK^{+}14, Mal96, Sep19, Yan21a]. setting [BGH14]. Seules [Has90]. Sevastyanov [Jiř70]. Seven [ADCKS15]. Seven-dimensional [ADCKS15]. several [Eva87, Gan67, Gin80, Gin81, Mat93]. Seymour [BMW21]. shadow [Jui16]. Shape [Ass10, ACRC17, CT16, Coq93, Nak20]. shapes [EG08]. Shapiro [LW20b]. Sharp [AB97, Bar15, BG16, Cat91a, DH21, ABK18, Cat91b, MV20, dIS02]. shear [Cas04]. sheet [CGPP06, CCFR09, KRS04, Wan07]. sheets [AWX08]. shell [FGHV16, RV20]. Sherrington [BL20, CH06, Che14]. shift [GLT11, KZ21]. Shih [FlJ84]. shock [Leg89]. shocks [FFV00]. short [Web10]. Shortest [HS99]. sided [AD13, HHK06, LSW02, MS04, Véc19, YYY10]. Sierpiński [BB89, BH97]. signals [BG16, FV10]. signatures [GQ16]. signed [Jac79, MM13]. signées [Jac79]. similar [Ber02, BCK21, GH10, GS15b, Haa04, KP11, Pie09, Sab01, She18, VAG86, VA89]. similarity [KZ17]. Simon [Kyp04]. Simple [SWW17, AD15a, BW19b, Cro08, DP83, DPS89, FGL95, FFV00, GM19, JL06, Kip87, Lac17, Lou05, LV20, RF64, Rez02, Sla22, Var95a, de 95, dlP94]. simpler [Kel17]. simples [RF64]. simplexes [Ben67b]. simplices [Ben67b]. simplified [Jiř70]. simply [GMV93]. simulated [Cat91a, Con94, Mic92a, Tro96, Mic92b]. simulé [Con94, Mic92a, Mic92b]. Sinaï [Riv07, Dev16, ELS10, Pèn19]. single [CP78, PX05]. singular [BCR20, CF18, EPP06, GZ19, GHR01, HM18, IPP08, LWZ16, Oka85, Osa98, WY16, XZ20]. singularité [Rou97]. singularités [Wer93]. singularities [Wer93]. Singularity [CGPP06, Hoc19, Rou97]. site [Kip87]. sites [Eva97, KK03]. sitewise [BMRS17]. six [BB17b, DR22]. six-vertex [DR22]. Size [DT20, Gan13, AG89, Lot91]. sizes [Cer96, Mar16]. SK [Pan09]. skeleton [ÖÇE17]. Skew [Chy08, MV92, MV91]. Skew-product [Chy08]. Skorohod [CL21, DZ05, Lia07]. Skorokhod [AEFdR20, Che17a, De 18, Eld16, Jak86]. slab [CM21]. slabs [BS17]. SLE [BLM21, RZ17, Zha10]. Slow [KT97, CFP12, EFG^{+}20, FGN13, GLP20, RSX21]. slow-fast [RSX21]. Slowdown [MZ16]. Small [BB14a, BBS08, CS02, GM13, LS05a, Tal88, AKKR01, Bor77, CDM21, DRS14, ES79, FW04, Fla02, Mik88, RSH19, Woo16, XZ09]. Small-time [BBS08]. smallest [PX05]. Smooth [GLM18, BDT11, BG16, CGL21, CX21, JM98, KR16, KT97, KZ21, MV20, Rao72]. Smoothing [BOZ16, RY89, CSC90, KM14, PW02]. smoothness [ALT99]. snake [Le 19]. snakes [Mar20]. Snell [Cai82]. Sobolev [CMR02, CM10, ELS20, GOVW09, Hel99, HU09, Joh17, LPY02, Mét84, Mic09, DL15, SSV03, SS18, Völ16, Yos01, Zha13]. SOC [Gor17]. soft [LWZ16, Wüt99]. soft-edge [LWZ16]. solution [CG96, FK10, Fou01, Mor99]. Solutions [Lan82, AKSS07, BDMBP18, BOT14, BRT07, BCT11, CK17, DHR11, DHJ^{+}21, DL78, Kob13, Kyp04, LM11, Mac87, Pie09, SSV03, Sło93]. solvability [KR16]. solvable [Bur07]. Some [Ade76, BM00, BS03, BAC21, BL02, BT00, CP17, DGM10, DMS93, Egg80, FI88, HSW20, JV07, Kaz71, LMX09, LZ15, MP84, PS92, AZ14, AWS21, ADG21, Ast76, BESY08, Bas08, BRT07, Bou80, BDC69, Col04, Dep65, Der00, Dev16, DW08, Fer90, FK94, FP17, FGJ17, FX19, FL20, GO13, HM16, Kom92, KMM10, Kur10, LS14, MV92, Ove94, Pic64, Pie09, Rot90, Zer00, Dos80]. sommabilité [Mén78]. sommes [Ako93, Led92]. son [Ber91b]. sortie [Gir70]. sorting [ADHV19]. soup [Qia19]. sources [Rou78]. sous [Bro75, BL12, Car81, Jac71b, Lus78, Mic99, MO83, Yor75]. sous-additif [MO83]. sous-ensembles [Car81]. sous-exponentielle [Mic99]. sous-markoviennes [Bro75]. sous-processus [Jac71b]. sous-tribu [Lus78]. sousliniens [Del77b]. space [ANT17, Ang16, Att95, AMD17, BDT11, BGKL17, BT16, CHSX15, Che17a, CGK20, Coq93, DGG21, DS97, Dep65, Der90, DCKY14, Eld16, Erb10, FI93, Gui79, Hos18, KK75, May68, Nic03, Num78, ORB20, Pay67, Pec01, Phu81, Pic96, Rau92, Rho08, Tor65]. space-filling [DCKY14]. space-time [DGG21, Hos18, ORB20]. spaces [ALM20, Ahm65, AHK77, AD15a, Arn93b, BT15, Bas82, BCK66, BDC69, Gan67, GZ19, Gin83, Gua77, Her08a, KKK85, KBR71, Kol92, NN81, Rao72, RF64, Sch83, Ust82, Wer80, ZBH17]. spanning [ABL08, Hut20, KKW15, KW20, LPS20, SSS^{+}21, dT16]. spans [EPT17]. Sparse [LV20, ORB20, BHLP19, BGBK20, BQZ20, BG16, Cos21, HL20]. Sparsity [Kol09]. spatial [BEV13, KLL21, Mie08, Rou87a, Rou87b, VW15, vdBPSV08]. Spatially [GNS13, Fou01, HLN17]. spatio [Bas69]. spatio-temporelle [Bas69]. SPDE [CF18, Mor99]. SPDEs [Bar15, CF17, CX21, DHJ^{+}21]. special [Mar15, SL95, Led78, SL95]. spéciaux [Bla75, Han71]. specific [Mic92a]. spécifique [Mic92a]. Spectra [OW17, Ass20, Gue99]. Spectral [BV17, BGBK20, BQZ20, GL16, HL20, LPY02, LM09, LLL14, LLS11, TW18, Tsi02, And15, BY16, BCC02, Ber89, BGS10, BHS12, Cos21, CFP19, DGZ03, FHL12, GGL15, GJ19, GC03, KL16, Mat97, Mat98b, Mic99, TV81, Wu04, Zhe12, dlR98, Rau92]. spectrale [Ber89, Rau92]. spectralement [Ber91b]. spectrales [Led70]. spectrally [Ber91b, PPR18, Véc19]. spectrally-positive [Ber91b]. spectre [Cup69, Mél78]. Spectres [Gue99]. spectrum [AEKN19, Bré04, CDLR17, COW20, EPP06, LV20, Péc12]. Speed [Tes18, ABGM21, BH08, BGN19, BL14, Cun03, Kom98, MSZ15, MR89, ÖÇE17, Pin10, Pin14b, Tro96, Zer05, Her08b, KMS97, Tou96]. spent [Mey91]. sphere [Car15, Eld14, GLTZ17, NR03, Rai99, Rai99]. spheres [CPV22, MS19]. spherical [BL20, GJ19, MRW20, Pan09]. Spike [CM21]. spiked [BY08, LS19a, PWB20]. spin [AD14, AD97, CH06, Che14, De 02, Fou00, GJ19, Gri77, KL05a, LPY02, Yos01]. spines [HR17]. spins [CD18, Fou00]. Spitzer [BD97, Rei82]. Spread [BT20, CH14, CE20, LW96, vdHdHS07]. spread-out [LW96, vdHdHS07]. spreading [Sch83]. sprinkling [BT17]. square [Alq08, Aud04, BG10, BDS21, Del08, LL15, MR86, NdS20]. squares [Ber91a, BD92, DST90, DST91, Tou96]. squarings [ABL16]. SSEP [EFG^{+}20]. Stabilité [Bir84, Gre08]. Stability [ELS20, Jui16, KM99, RDL20, AZ97, Bud03, DG01, GM13, Mac87, Sło93, SvG11, Woo16, Gre08]. stabilizing [ERS15]. Stable [GGL15, GS15a, HMP20a, KRS21, Arc21, CR05b, CG19, DMSY94, FT12, FG92, FS04, Fou13, Gal96, Gin83, Gir02, GH10, GJ11, Gui90, HSW20, Hey11, HM16, KL20, KRS18, LS16, LS05a, LM83, MP84, Mar20, MZ15, PT08, Sin07, Tal89, Viv21, YYY10, Yan13, Zha13, BCK66, Gal96]. Stable-drift [Gal96]. stables [Ber70, BCK66, Dav74, Tor81]. stacks [KP17]. stage [Con65]. stage-length [Con65]. Stam [Let72]. standard [BGJS18]. standards [KRR70]. starting [BDCKL20]. state [AD09, BD16, FM19, KMS97, Lab14, MW02, Num78]. states [AD14, BG99b, Che15, Par05, Sab01]. static [DO20]. stationarity [KOS19]. Stationary [ABGG^{+}12, BKS12b, AZ14, AW18, AFGJ16, BCV21, BD16, BG99b, BK95, Bur07, CCM19, CD16, Coh72, DR00, DR08, DMPU09, Dep99, FHL12, FFV17, GP14, Gas90, GL89, HMNP98, HT04, KLO95, Liu18, MS70, May68, Mic17, MR94b, MW11, PW79, Pel90, Rho09, SS10a, Sep20, She18, Son13, Tim14, Wu99]. stationnaire [Dep99, Gas90, GL89, Hai87, Mél78, Rec86]. stationnaires [Bas69, BK95, Bru76, Hai81, Han71, Han73, Hil73, Hil78, May68, Pel90, Szn80, TV81, Web80]. statistic [DDP87]. Statistical [Hut19, PWB20, AWS21, LS91]. statistically [Sen90b]. statistically-convex [Sen90b]. Statistics [Mat98b, Pac00, ANT17, BCHJ21, ESYY12, FT79, GR11, HL20, Ken97, Lam21, LSX21, PSV17, Sen90b, Ste74, Zhe12, dIS02, MN83]. Statistique [Sen90b, DP84a, Jum80, LS91, Lan82, LOR69, LPPV69, MV69, MPL73, Rom69a, Rom69b, lJ87]. Statistiques [FZ88, MN83, Bal04, Deg82, Sen90b, She04, Ano05a]. stay [CC08, GLL18, Véc19]. steady [KMS97]. Stein [CD19b, CFP19, Gau20, GW18, Kas20, NPR10, Röl13]. Steiner [Pru97b]. steps [BHR17, HP17, Pei92]. stereographic [Car85]. Sticky [EZ19, FGV16, Gra88a]. Stieltjes [All78, GBR73]. stiff [LS20]. Stochastic [ANV00, AB02, AN09, BBCH21, BCV21, BL05, BCT11, CCM03, CD18, CN05, DVW00, Dec05, DLZ20, DGKP21, DF87b, HW22, IKPY18, Imk84, JJ18, LS06, Pic05, Ust82, Zha13, ADG21, AR08, ASS90, ABK18, Bai10, BB21, Bañ18, BDT11, BRR17, BRT03, BLNR11, BOT14, Ber83, BDN20, BP04, BD96a, CL21, CG96, COBZ21, CR05a, CD19a, CR20, CKK17, CK17, CQR07, CJKS14, Coq93, CKL14, Cro18, DDT04, Dab14, DG19, DT21, DOS18, DVNZ20, Der90, Egg80, Eld16, EM83, Feh19, FGR09, FP17, FK10, Fou13, FJR21, FV05, Fun99, Gan67, GS18, Gou07, GNSS94, GL16, HM18, Ham96, Hoc19, Hof15, HLT16, HHM21, JKMP05, Kan72a, Kan72b, KR16, Kli09, Kuk13, Kun86, KPP95, KMR06, LS19a, Lam20, LCO09, LMX09, LM11, LS05b, LP17, Mac87]. stochastic [MS70, Mar00, MN18, MS96, MP00, MS21, NNRT09, OP89, Oda06, OV20, Pel10, Pic65, RCG18, RSX21, RZ93, SSV03, SvG11, Str88, Sud21, Tak70, Thi91, Tor17, TW18, Uek99, Web10, Wer80, Wu10, XZ20, XZ09, Zhe85, ZBH17, Est92, LNS89, Sch89]. stochastically [FGHV16, Kom92]. stochastique [Ben79, Der90, Est92, LNS89, MM71, Sch89]. stochastiquement [Yor77]. stochastiques [All78, Bou88b, BD96a, Coq93, Dos77, DL78, Dub75, Fli75, Ham96, Jac82, LM76, Pic65, RZ93, Str88, Szp78, SM79, Thi91, Yor74b]. Stokes [AKSS07, BL08a, CD19a, XZ09, Yur08]. stopped [Mat93]. stopping [AC04, De 18, PP98]. Strained [Tor66]. strategy [GKMMR20, ÖE19]. Stratonovich [GQ16, KPP95]. strength [Bau13]. Strict [Fou01, CK17]. strictly [Ros89]. string [Ors82]. strings [Ber89]. strip [BL95, KM99]. Strong [CLB98, DPSZ16, HKK10, LM11, Lin76, Mic17, RSX21, Sep00, Zyg13, dR17b, BCP19, BDMBP18, BD96b, ES02, FI88, GG02, HM18, HHM21, Mas88, NS81, EHK10]. Strongly [RN21, DMR94, EHJ07, Her08b, NS95, Rio93]. strongly-ergodic [Her08b]. structure [AI05, Bel08, DDP87, GMP78, Hu02, PT98, vdHS14]. structures [EH19, Fla02]. study [Cat12, FLP06, Gol89, KHT12, Mey91, Pic64, Vig03, Dep65, Her94, LR96, Mic92b, MR89]. Sturm [Sab01]. style [DF87a]. sub [BBCH21, JP17, Mic99, Pet09a]. sub-ballistic [JP17]. sub-exponential [Mic99]. sub-Gaussian [Pet09a]. sub-Riemannian [BBCH21]. subadditive [Ste89]. subcritical [ABKV14, GKV03, GQ15]. Subdiffusive [Kes86]. Subexponential [Gan02]. Subgeometric [DFM16, DGM08]. subgraphs [DHMP17]. subinvariant [BHL93]. subjectivité [Jum80]. sublinear [JLS13]. submartingales [Mil83]. subordinates [HS96]. subordinateur [Vig03]. Subordination [ATV20, MS70]. subordinator [Vig03]. subordinators [Mar15]. subsequences [BB17a]. subsets [Bef03, Tal89]. subshifts [ANS06, Rou21]. substring [Rou21]. successifs [Lac97b]. succession [FLT04]. successive [Lac97b]. sufficient [FT79]. suffisantes [AJ76a]. suit [Con65]. suite [Hai87, dZ81]. suites [BLT69, BDC69, BK95, GS97, Hai81, Lac73, Loo77c, Loo78, Mén78, Sch72a]. sum [COBZ21, LL15, de 95, dlP94]. summability [DGR16]. sums [Ako93, Cel11, DDG13, GN02, Led92, Pru97a, Sta89]. super [CE20, CP04, DV05, FS04, FM19, HP19, KLMSR12, Mer06, Mör01, Nak15, Per89, PT98, vdHS03]. super-Brownian [DV05, HP19, KLMSR12, Mer06, Mör01, Nak15, Per89, PT98, vdHS03]. super-exponential [CE20]. super-individuals [FM19]. super-stable [FS04]. superadditive [Émi86]. Supercritical [BMRS17, DCKY14, KLMSR12, Abe15, BB14a, DT20, FW07, FW09, Gol18, JS20, dHdS14]. Superdiffusions [CE20, EW06, Eng09, ERS16]. Superdiffusivity [Lac12a, Lac12b]. supérieure [Heu98, Vig03]. supermartingale [PX05]. supermartingales [LR20]. Superposition [LCO09]. superprocess [MV07, Zho07]. Superprocesses [Le 91, FK94, LMX09]. supersolutions [BPTZ18, DKGT16, HKM14]. Support [CF18, AGL94, Che76, EP06, FK10, HPT02, LZ15, Per89, Pet96, Tor77a, Tor77b, Tor78c, Tor82]. supported [FP08]. supports [GG79, Mar88]. suprema [Lif94, Mat93, SV08, vW83]. Supremum [DG19, BB01, BB02, CM16, Her08a, She18, Tak70, Tal88, YYY10, Yan13]. sur-additif [Nev83]. sur-additifs [Hac82]. sure [BS10b, DGM10, DDP87, DST91, EM83, Fit98, FR03, JRdlR10, RY95, RAS09, RR94, Sen90b, Wan07]. surely [Del64]. surface [FVY06, Fra91, NR03]. surfaces [ANT17, AG21, BM08, BBCH21, Bet16, Boi98]. surprising [BG10]. survey [Gal70]. Survival [KK14, GHS11, GvdLW95, Jaf12, Liu96, ÖE19, vdHdHS07]. survivor [LMD03]. swaps [EN20]. Swendsen [GLP20]. symbiotic [AD11, KM20]. symétrique [Kip87]. Symmetric [AFJV15, BKK10, PW94, AL12, AD15a, BGJS22, CKS87, CZ02, CD16, CE00, DP83, DMSY94, FGL95, FGN13, FT05, HS96, JL06, LS16, Mac87, MU11, Mim14, NZ13, Pru97a, SZ97]. symmetries [Ade76, LM83]. symmetrised [AD08]. symmetrization [Pru97b]. symmetry [BZ12]. synchronous [DY11]. system [AAD19, AD00, BB18, DLZ20, GLT09, Gra88a, JP17, KL95, KLO95, PPR97, RT97, dS03]. système [BL79, Bro83, Bru71, KL95, Oga74]. Systèmes [Fer84a, Fer85, CK77, Con72b, DL96, Dos80, Fou00, Lap86, Léo86, Les87, Saa88, dZ81]. Systems [KPS16, AD08, Ass00, AD97, BGRS14, BDEG88, BDEG89, BG99b, BF95, CDLR17, CDGR20, CR05a, Coh68, CD12, CKL14, DDT04, DNN73, DL96, Dos80, FRS22, Fou00, FFV17, Fur97, GS10a, Gri77, Kom92, KL05b, KV22, Lac97a, Lan91, LPY02, LCO09, Len04, Léo95, Löc02, LM22, MZ15, MO07, Par05, PT21, DL15, RCG18, Sar17, Set01, Val06, Völ16, Yos01].

Table [Ano16c, Ano16d, Ano17b]. tables [Col78, NN20]. taboo [DN01]. tacnode [BD11]. Tagged [BT18c, Bra09, JL06, JL08, JLS13, Lou05, Saa90, Set07, Var95a]. Tail [AL12, Lif94, Riv12, Ber85a, Coh72, Deh00, DV13, FW09, Gan02, Mas88, Pru97a]. tailed [AAP09, BCHJ21, BGP14, CCP12, DvdHvLS20, MSC19, Soh13]. Tails [QR15, AL12, Ass11, Bur07, GGS15, Tal88]. taken [GM13]. taking [vW83]. Talagrand [Völ16]. Tanaka [BK93]. tant [Sch85]. TASEP [BLS22, FGN19, GS15b, Liu18]. Taylor [KP11, Yor91]. TAZRP [ABGM21]. TCL [Cun03]. Techniques [CTR95, DCS74, Gol89]. Teichmüller [ANT17]. temperature [CD18, De 02, Lam21, LW20b, Mis16, Wu91, Wu91]. temperatures [BL15, PZ21]. temporal [BU18, OV20]. temporelle [Bas69]. Temps [Gra84, Thi91, Ako93, Aza89, Der73, ERR71, FZ88, Fou95, Kip87, Lac97b, Lot85, Mat93, May68, Mey91, RR70b, SG02, Wer93]. tend [Wu91]. tends [Wu91]. tendues [Tor66]. tensor [DGR16, PWB20]. tensors [Gur14]. tent [MV92, MV91]. term [CR05a]. termes [Tor65]. terminal [DHR11]. terms [DNT12, Tor65]. test [Ant82, ACA02, BB11, CPV22, IL06]. Testing [CPV22, MW11, MS21, Bir06, BMP09, GR08, GJN20, MS17, MR22]. Tests [Deg79b, Gar78, DP84a, FK69, FLRB11, Rao70, LOR69]. ’th [DS96]. their [ABW17, BRR02, BJ20, DGM10, DW17, DGM21, FTV20, GBR73, Tuy81]. Theorem [BP15, DVNZ20, AKU17, AKU20, ABK15, AGL94, Ass98, BH08, BB21, Bas82, BG99a, BG92, BU18, CT16, CDG11, CDGR20, CF18, De 98, DR00, DD10, DL96, Din92, DMY91, DMR94, EV03, Fit98, FG79, Fra91, Fur97, GQ16, Gou05, Gri77, Gri04, Gru66b, GPP10, Her05, Her08b, HP00, JL06, JL08, JN19, Jiř70, KV22, KP19, KL02, Lac97a, LP97, LSX21, Lia16, LLL11, LR93, Mas88, Mas86, McD75, NN16, NZ21, Ouc05, RAS09, Rau04, Rez02, Rio09, RR94, Ste89, SvG11, Tal96, Thu81, Tol00, Var06, Wüt04, Yor91, dS03, BFK20, Bli14, Bru93, Pet96]. Théorème [BBC78, Bru93, Bru73, CR78, Émi81b, Fra91, GP74a, Han72, Her05, Jac71a, Jac74, JKM82, Led83, Mai78, MO83, Pet96, Roy74, Sch72a, BG92, Dau73, DL96, Din92, EV03, FlJ84, Her08b, LR93, MM71, Mer83, Nev83, Ouv73, Pel72, Phi71, Rau83, TV81, Yor91, Gru66b]. Théorèmes [BDC68, CR80, GH88, Les87, Mét72, Reg70, Tou83, Ass87, Bec74, Bou83, Fis73, Hac82, Rot90, Tor79, GS97]. theorems [ABKV14, Ass00, BDS87, BY08, BB17a, BHS21, Bay15, Bew71, Boi98, BDC68, CDMF12, CGPP13, Coh72, CL09, DR08, DGM08, Émi86, FS17, FHL12, FG92, GS97, GKV03, GLL18, Gui02, Haf20, HJY11, Kif14, KMM10, Kor15, KRT16, LPWR94, Lin76, LT20, MWRZ04, MT20, NNT10, Num78, Rot90, SY13, Zhe12]. Théorie [Bab88, Pet79, Rau92, Ruc80, Ben67c, Ber89, Cha70, Dos91, Hey68, JT76, Lan82, Rab94, Sch00]. theory [Aar78, AZ14, AHK77, AK21, BLM21, Bel08, Ben67c, Ber89, BP95, Col04, Def10, DS97, DF87a, Dos91, EF05, Eva87, FHS19, Her77, Hey68, MS96, PT21, Rab94, Rau92, Sch00, SV08, Var95b, VAG86, Wer80, BLM21]. thermodynamique [Lic70]. theta [Cel11]. Thick [Che18, CH17, Dav05, Jac18]. Thin [DPRZ00, RV20, Sep19]. thin-shell [RV20]. Three [AI99, LR14, BBCH21, GvdLW95, Gol18, Gol89, Lia16, Shi19, vdB10]. Three-dimensional [AI99, BBCH21]. three-series [Lia16]. threshold [AS17, Per00]. Thresholds [IY20]. tide [Hol12]. Tightness [MZ84, Zhe85]. tilings [BK18, dT07]. Time [ET90, VA89, Zha22, Ako93, AD00, BRZ19, BT18b, BB11, BBS08, BFG15, Bor87, BDM11, CGB17, CCP12, CCM19, CF09, CHSX15, Che17a, Che17b, Che20, Coh72, CP11, CCFR09, DGG21, DS90, Del08, DRZ21, DL96, DGR18, DP11, DGKP21, ES17, Feh19, FGL95, FG92, FR03, GP14, GZ98, GH10, GO13, GM13, GG79, HP17, Hoc19, Hos18, HLN17, HP19, Jac98, JN19, JYZ05, KMP14, KL20, Lac97a, Lac91, Lia07, Liu18, LP17, MØP04, Mat93, May68, Mey91, MW05, MV07, NN20, NS81, ORB20, Pec01, Rho08, Sen17, SG02, Web10, Wer93, XZ09]. time-changed [BT18b]. Time-changes [VA89]. time-changing [CP11]. time-dependent [DGR18, ES17]. time-inhomogeneous [GO13]. Time-reversal [Zha22]. time-space [CHSX15, Che17a, Pec01]. Times [BK93, Ang16, Ass00, AC04, AWX08, ASX11, Aza89, Bar15, BB01, BB02, BHLP19, Ber83, Bla03, CGZ13, CT16, Cha94, CR05b, CRR98, DOS18, DW15, DM04, DGM08, Eld14, Fou95, GRVY99, GMS19, Her18, Imk84, KM99, KRS04, KL20, Lac97b, LLL11, LLL14, LLS11, MM09, MR94a, Mat97, Mat98b, Mer06, Mic17, Pac00, PP98, PS19, RY95, Ros96, Ros99, Ros01, Ros10, iSW04, iSW05, Thi91, Yan21a]. tired [KM20]. Toeplitz [ER13, Lem88]. top [Per92]. topological [Ahm65, Sie84, Tor65, Tor66]. topologie [CL84]. topologies [ES02]. topologique [CM70, Hei71, PPH75, Tor65, Tor66]. topologiques [Ahm65, Reg70]. topology [Jak86]. tore [Die81]. tori [vdHS14]. Tortrat [Ano77]. torus [BSC04, BB17b, FG12, FP18a, JJK^{+}14, PS19, TW18]. tossing [AKU20]. Total [BD16, CDM21, NZ21, SZ16, CG09, ERS16]. tout [Rev83]. toute [DS84]. toy [FL16, FTV20]. trace [Cha21, Shi10]. traces [Aug18, BH20]. Tracy [DV13]. traitement [Ben67b]. trajectoire [Ber91b]. trajectoires [Aza89, Ber70, BL12, ERR71, FlJ84]. trajectories [Aza89, BL12]. trajectory [Ber91b]. transfer [Her94]. transfert [Her94]. transform [ABDG19, BP06, Cat03, GRVY99, Lif94]. transformation [BL12, DGdH^{+}08, FI93, FP08, JdlR04, Jia12, KL02, dlRd03]. transformations [AW18, BW01, Cab79, CZ02, Con72a, El 00, EPP06, GS10b, KT97, Kop82, Mél06, Nic03, Pet79, Pri12, Ros89]. transforms [GBR73, JM93]. Transience [DHMP17, DE16, Hen82, Pin10, Pin14b, ABO16, Bal76, BB11, PSS16, PS92, ST11, Sun87]. Transience/recurrence [Pin10]. transient [AS18, Aid10, Ass10, Bau13, KM11, MWRZ04, Pet09a, PS13, iSW05]. transientes [Bab88]. Transition [BH97, AKM21, Ass10, Bar15, Bro20, CKS87, Che17b, DH21, DW08, DR69, HKPG16, Her05, Ish97, KM14, Lac07, MV20, Rau92, Rev83, Sar14, Sar21, Was15]. Transitions [Kig13]. transitive [BV17, CT18]. translatées [Tor78b]. translation [ANT17, ST94]. translations [CM70, Yor75]. transmission [KLL21]. transport [DGKP21, FGJ17, GRS^{+}18, Hue16]. transport-entropy [GRS^{+}18]. Transportation [Maj19, Wu10, ÁEdBCAM11, Pra07, RV13]. Transporting [LTT18]. transposition [DMSY94, GP20]. Transversal [BW19a]. trap [ÖE19]. Trapping [Eva91]. traps [BD96a, Mou11a]. traveling [LP17]. Travelling [Kyp04, HHK06, KLMSR12]. travelling-waves [HHK06]. treatment [Ben67b]. Tree [AP98, AS18, ADH12, ABDG19, Aid10, BRZ19, BP92, Ber15b, BD16, Cas17, CM21, Cro08, DRZ21, HMP20a, KW20, Kig13, KZ17, Liu01, Mar16, MSZ15]. Tree-valued [AP98, ADH12]. Trees [Duc20, FFW05, NNRT09, ADH12, Aïd20, ABL08, AABV16, AC18, AG21, AHOZ13, BC17, Ber12b, Ber15b, Bla99, BRT07, CLP14, Che97, CW13, Cro08, DMS05, DW17, FLT04, Hut19, Hut20, Kor15, LW06, Lin17, Mie08, MS19, Pic05, Riz15, Stu19, dR17a, dT16, vEEIK12]. treillis [Bod70, Bod71]. Trends [CG09]. triangle [Koz11]. triangular [Tra02, dT07, dBMCA99]. triangulations [Bud20]. tribes [Gru66a]. tribu [Lus78]. tribus [GMP78, Gru66a]. trichotomy [CKK17]. tridiagonal [Lam20]. trigonometric [ADL16, Sch00]. trigonométriques [Sch00]. triple [BB18]. trispectrum [MRW20]. trivariate [CPY04]. trivial [Val78]. trois [Gol89]. trou [Mic99]. troughs [AI99]. truncation [Pol02]. trunk [KW20]. Tube [BCP19, Pig18, KM17]. tués [KRR70]. Tulcea [Ano77, Tor76]. tumor [DLP07]. Tunneling [BL15]. Turán [SZ16]. turbulence [Bas69, FGR09, Gir02]. turbulent [FGHV16]. Tusnady [Mas02]. Two [Bli14, DLP07, FFW05, KMR06, LSW02, PZ21, ACLS19, ADG21, BL15, Ben67c, BFK20, BL19, BJ20, BG92, CGPP13, CD19a, ClJ89, DGdH^{+}08, EF16, Fou95, GKMMR20, GOVW09, HP17, HU09, HS96, Jin14, KKW15, LMMR18, MM09, MS04, McG89, Mén10, Mil83, MS83b, MP00, NS81, Pan14, PS09, PS19, Pin14b, Val06, Vig03, XZ09, vdBvE22]. two-arms [vdBvE22]. two-component [GKMMR20, KKW15, LMMR18, Val06]. Two-dimensional [FFW05, BL19, BG92, CGPP13, CD19a, EF16, Jin14, PS09, PS19, XZ09]. Two-parameter [Bli14, HS96, Mil83, NS81]. two-phase [Pin14b]. two-scale [GOVW09]. Two-sided [LSW02, MS04]. Two-temperatures [PZ21]. two-type [DGdH^{+}08]. type [Alt87, BDS87, BRR17, BM00, BGM20, Ber12a, BFK20, BCC02, BLM99, CCEL13, Cas07, CM10, Cha12, Che87, CZ02, CN05, CR96, CCS92, DGdH^{+}08, Deh00, DNT12, Egg80, FSP71, GI80, Gin83, GN09, Gru66a, Has90, HU09, Hob94, HLT16, Joh17, KPT20, KT97, KP11, Mor99, OTT02, Pet96, Pri18, Rei82, Ros99, Rou21, Sch09, Shi20, Thi91, dIS02, dlR98]. types [Eva97, Gru66a, Pan14, dR17a]. Typical [Lin17]. typically [BDS21].

UHF [GSS05]. Uhlenbeck [CCS92, Fer90, HLT16, Shi20]. UIPT [GMSS18]. unbounded [BW15, CD16, CP12, CPX17, DGM10, DHR11, EM83, FHS19, GSS05, LPY02, Yos01]. unbounded-range [FHS19]. unconnected [Ben67b]. uncountable [Gru66a]. unicellular [Ray15]. Unicité [CR79, DL78, Yor74a]. unified [BPTZ18]. Uniform [Edg79, GGS15, KMP14, KM17, LL08, Num78, SSS^{+}21, Aïd20, BDCKL20, CLB98, DP11, GG02, Her18, Hut20, Jin14, KW20, Len04, Lin75, LPS20, Mén10, Jac79]. uniforme [DS84, Jac79, Mén78]. uniformes [Sch72a]. uniformity [CPV22]. uniformly [GG01, GLTZ17, Gou05, Mar00]. unifying [Yan13]. unimodales [MS83a]. Unimodality [HS17]. unimodular [Var95b]. unique [Ben20, CF09, FP08]. uniquely [Fur97, Len04]. Uniqueness [AKKR01, ÁEdBCAM11, BBC05, BL03, CD12, DFM^{+}03, DHR11, Fou13, FL20, Gas21, Per00, Pri18, Sep20, Sło93, dR17b]. unit [HRV18]. unitaire [Lon69, TTL69]. unitary [HP00, SS10a]. univariate [BFK20]. Universal [Phu81, Haz93]. Universality [AD13, CH06, CFP12, DGR16, ES17, GM20, Gur14, Joh12, KPR18, Péc12, Woo16, Ahn20, ESYY12, LWZ16, MP19]. universally [GS97, LPWR94]. universellement [GS97]. unknown [MS17]. unnormalized [CDG11]. unscaled [HR12]. updatings [Fis01]. Upper [BOT14, CKS87, CDH00, Lac12b, Mej04, Rio09, Aar78, Ass11, Che17b, Gui02, Vig03, vdBvE22]. urn [EN20]. use [Ade76, Pin14b]. using [AD09, ACRC17, BLM21, BRT07, DV05, LS14, NPR10]. utility [AI05].

V [BR79, LOR69]. vacant [ČTW11, DRS14, DE16, SS10b, Was15]. valeur [Fis73]. Valeurs [Deh74, Hai81, Har01, BH80a, Cau69, Fou84, Gar78, Gar79, GP74b, Hei71, Hei74b, LS91, Lou68, Mar88, Mén78, Mét84, Pay67, PPH75, Saa88, SP76, Szp78, TTL74, Yor74a, Yor74b]. Validity [GR08]. valley [GPS10]. Value [FFV17, Coh68]. valued [ADH12, AP98, ALT99, BU18, CE00, CCS92, DVW00, EP06, Eva91, Gho77, GNSS94, HK02, Kan72a, Kan72b, Riv07]. Values [LSW02, BB14a, Der00, Har01, LS91, LWZ16, MSZ15, Pay67, WY16]. vanishing [CD19a, IY20]. Vapnik [Pis84, Vay03]. variable [CFP13, Deh74, FGS18, FGS19, Fis73]. variable-range [FGS18, FGS19]. variables [AL12, Bir84, BLT69, Bou84, BK95, BH80b, BH80a, BT08, Cup69, DMPU09, DMSY94, EH03, FL21, Gau20, GS10b, GW18, GN02, Hai81, Led92, MR94a, MR90, Mén78, Pru97a, RR94, Sta89, Tra02, Tuy81, de 95, dlP94]. Variance [BL95, Mou11b, Ber85b, Bru93, Cat12, DMY91, DR82, Fle10, Gil66b, HLN06, MZ16, PW02]. Variation [Str88, CDM21, CG09, De 98, FG92, Jeu93, LL15, NZ21, SZ16]. Variational [BDM11, GLM18, BCT11, FGJ17]. Variations [BY87, GL89, HJY11, IL97, NNT10, RZ93]. variété [Bab91, Pic94, Pri68]. variety [Bab91, Pic94, Pri68]. various [Gas02]. varying [BHR17]. vecteur [Deg82]. vecteurs [TTL74]. vector [Ahm65, DST90, DST91, JM93, Led92, Mal11, Hei90]. vectoriel [DST91, PPH75, Tor78a, Tor78b, Tor82]. vectorielle [Bru76]. vectorielles [BH80b, DST90, Hei83, Hei90, Led92, Mén78]. vectoriels [Ahm65, Tor77b]. Velocity [Zer00, FGS18, FNS00, Guo14, dlR02]. Ventcell [Pet96]. Ventcell-type [Pet96]. Ventzell [OP89]. vers [JKM82, KMS97, Pra92, Wu91, dZ81]. Version [Dau73, Thu81]. versus [BBD19, ERS16, HK15, Hof01]. vertex [BV17, BB17b, DR22, MRT21, RN21, Stu19, Zen16]. vertex-reinforced [MRT21, RN21]. vertex-transitive [BV17]. vertices [Riz15]. Vervaat [Fou05]. via [AK21, BT17, Bir06, CDP09, DHMP17, DHJ^{+}21, Eld16, ELS20, GJN20, HMP20a, JM93, Kas20, Lab14, LS20, Maj19, Pol02, Sar21, Sep20]. vibrantes [Ber89]. vibrating [Ber89]. vibrations [Ors82]. vicinity [Rou97]. vie [Rob87]. Vietoris [IY20]. view [BMP94b]. viewed [Mou11b]. VII [RVY09]. Viot [AFGJ16, BEV13, CDGR20, KL20, LZ15, VW15]. viscosity [Kob13]. viscous [ACLZ20, CX21, CPX17, JLT21]. Vitali [Egg80]. Vitali-conditions [Egg80]. Vitesse [Her08b, KMS97, Tou96, dlR02, Cun03, MR89]. Vlasov [HSS21, LMW21, Mar21, RSX21]. voisinage [Rou97]. Volterra [CL21, Dec05]. Volume [Mat98b, CI03, Lac12a, Lac12b, LM97, Mej04, MW02, Tel08, vdB10]. Voronoi [GC03]. vortex [BG99b]. votant [Saa95]. voter [CCC16, CP04, GLS05, GW18, MM09, MM13]. voting [Saa95]. vraisemblance [DP84b, Jac90, MR89]. vs [BS10a].

W. [Haz93]. waiting [Coh72, FGL95]. walk [AD01, AC18, AHOZ13, AGJ^{+}20, AC07, Ass10, BC17, BGHR19, Bau13, BHS21, BKYY10, BY19, BBHK08, BGN19, BS10a, BS11, BMP94a, BMP94b, Bor87, BDZ16, Bré09, Bro03, CH14, ČTW11, CH20, Con65, CCFR09, Cun03, DRZ21, Dep65, DGR18, Dev16, ELS10, FLP06, Gan02, GKS07, GPS10, GHS11, GM20, GPP10, Gui00, HK02, Her19, Hey11, Hol12, Hu16, Hua19, IR98, Jaf12, Kes86, KSS21, KMP14, KOS16, KOS19, Leg89, Lou05, LPS03, MM18, MS04, MSZ15, Mou11a, NN20, NS95, OR18, Pai18, Pei92, PSS16, Pet09a, PS13, Pin10, RAS09, RAS11b, Sch09, Shi10, Shi19, ST20b, Tel08, Var95a, Was15, Yil10, Zer05, Zhu21, dHdS14]. Walker [Ang16, SG02]. walking [GI80]. walks [AD08, Aid10, ABO16, AAD19, AI99, AD00, AC99, BPS12, Bar00, BP92, BR15, BD97, BG99a, BHR17, Bla03, BSZ03, Bor21, Bré04, BDC67, BDC68, BW05, BW06, CFP13, CC08, CGPP13, Che97, CR05b, Che17b, CCKW21, Col04, CP12, Cro08, CRR98, DV98, DP83, Def16, DW15, DK08, DCKY14, Eld14, FS17, FMM92, FLT04, FP18b, GP11, GGL15, GLL18, Gre08, GK17, GL16, Guo14, JP17, KM99, Kig13, KK03, KM11, KP17, Leg89, MR94a, MR97, MWRZ04, MSC19, Mus06, Nak15, Pei92, Pru97b, Ras11a, ST11, SY11, Sch00, Sim07, Tho13, Tou15, Val89, Vys14, Zer00, RT18]. wall [MM18]. walls [FVY06]. Walsh [IKPY18]. Walton [CG14]. Wang [GLP20]. Wasserstein [ACJ20, BFK20, BGKL17, BL14, DFM16, Erb10, Mar21]. Watanabe [BDMBP18]. Watson [ADH12, Aid10, Aïd20, AP98, AR06, AC18, AHOZ13, AFGJ16, BC17, BRZ19, Che97, FW07, FW09, Kor15, Lin17, Liu01, Loo77a, Lot91, MSZ15, Mie08, Nev86, Riz15, Stu19, dR17a]. wave [DVNZ20, Dey20, Kyp04, MN18]. Wavelet [Ist92, HMSH12]. waves [HHK06, KLMSR12, Le 02, LP17]. way [Gan02]. Weak [CT16, ERS16, FL20, GB16, Lou00, MZ15, PS13, Ros86, Sar18, BP85, DMP14, DV08, El 00, Eme85, FWY00, GRS^{+}18, Joh12, KL02, LS14, MP19, Nak86, Pet09b, Pig18, Tes06, WZ90]. Weakly [Kuk13, Ass00, DPS89, FL21, GPP10, JLT21]. web [EF16, FFW05, FINR06, NRS10, RSS16, CV14]. wedge [LRZ19]. Weierstrass [Kel17]. weighed [DR22]. weight [Mat93]. weighted [AR08, Ass98, CH90, DLS02, Her08a, HMP20a, NZ13, NNT10, Pac00, BD92]. Weights [MR86, Boi98, Dal20, FHS19, GM20, RZ20, SZ16, Stu17]. Weiss [EM15, CGK20, Gor17]. welding [HP21]. well [DMGM18, Hos18, LW20b]. well-posedness [Hos18]. Welsh [BMW21]. Weyl [Gra99, Mkr14, Ras11a, Sch09]. which [Bor77, Sch83]. white [ANV00, AB02, DGG21, Enr96, Hos18, RSH19, Tsi02]. white-noise [ANV00, AB02]. Whittaker [BO11]. whose [Con65]. Widom [DH21, DV13]. Wiener [Att95, BN90, DMY91, FI93, GK93, GJ11, Gua77, Hai87, Imk84, KKK85, LT06, Mac78, vdB10]. Wigner [And15, Ben20, CDMF12, Joh12, LSX21, PRS13]. Wilkinson [LZ22]. winding [Yao18]. Windings [SY11]. window [AS17, Lac17]. Wishart [ACLS19]. within [Duc20]. without [Aïd20, BDMBP18, DNT12, Fou01, MSZ15, de 95, dlP94]. Witten [Hel99]. Wolfowitz [BLM99]. Wood [Duh97]. workable [DS97]. Wright [ALM20, CCC16].

X [Tor66]. XX’ [BGS10]. XY [NW18].

Yamada [BDMBP18]. Yang [GM20]. years [Mas02]. Yoeurp [Yar19]. Yor [PP98]. Yorke [LMD03]. Yule [SG02]. Yule-Walker [SG02]. Yuriskii [Tor79].

Zero [ABO16, JdlR12, Jia12, Bab91, BMRS17, BT20, BF88, COBZ21, Émi85, FMM92, FPV88, HP19, JLS13, Kom98, LM97, Leg89, Lou05, Mis16, PPR18, Sen17, Set07, Sue05, Var95a, Wu91, ACTR85, Der76, Leg89, RLT81, Saa90, Wu91, Saa90]. Zero-one [ABO16]. zero-range [BMRS17, BF88, JLS13, Set07, ACTR85, Saa90]. Zero-ranked [Saa90]. zero-sum [COBZ21]. zerotemperature [Mis16]. zéro-un [RLT81]. zeros [ADL16, Han17]. Zipf [Rou78]. Zuber [CŚ07].

References

Amir:2016:LPG

[AABV16] Gideon Amir, Omer Angel, Nicolás Matte Bon, and Bálint Virág. The Liouville property for groups acting on rooted trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1763–1783, November 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

              http://projecteuclid.org/euclid.aihp/1479373247.

Andreis:2019:ESI

[AAD19] Luisa Andreis, Amine Asselah, and Paolo Dai Pra. Ergodicity of a system of interacting random walks with asymmetric interaction. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):590–606, February 2019. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

              https://projecteuclid.org/euclid.aihp/1547802410.

Amir:2021:MPM

[AAH21] Gideon Amir, Omer Angel, and Alexander E. Holroyd. Multicolour Poisson matching. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):1811–1833, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/Multicolour-Poisson-
matching/10.1214/21-AIHP1147.full.

Auffinger:2009:PCL

[AAP09] Antonio Auffinger, Gérard Ben Arous, and Sandrine Péché. Poisson convergence for the largest eigenvalues of heavy tailed random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):589–610, August 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1249391376;   http://www.numdam.org/item?id=AIHPB_2009_
_45_3_589_0.

Aaronson:1978:ETI

[Aar78] Jon Aaronson. Ergodic theory for inner functions of the upper half plane. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):233– 253, July–September 1978. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1978__
14_3_233_0.

Andriani:1997:SED

[AB97] Cristina Andriani and Paolo Baldi. Sharp estimates of deviations of the sample mean in many dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33 (3):371–385, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1997__33_3_371_0;   http://
www.sciencedirect.com/science/article/pii/S0246020397800970.

Alos:2002:SPD

[AB02] Elisa Alòs and Stefano Bonaccorsi. Stochastic partial differential equations with Dirichlet white-noise boundary conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(2):125–154, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_2_125_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010974.

Attal:2018:CQP

[AB18] Stéphane Attal and Ivan Bardet. Classical and quantum part of the environment for quantum Langevin equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54 (4):2159–2176, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1539849795.

Ayed:2021:CRO

[ABCF21] Fadhel Ayed, Marco Battiston, Federico Camerlenghi, and Stefano Favaro. On consistent and rate optimal estimation of the missing mass. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1476–1494, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/On-consistent-and-rate-
optimal-estimation-of-the-missing-mass/10.1214/20-AIHP1126.
full.

Addario-Berry:2019:ICT

[ABDG19] Louigi Addario-Berry, Daphné Dieuleveut, and Christina Goldschmidt. Inverting the cut-tree transform. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1349–1376, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1569398872.

Abe:2015:ERS

[Abe15] Yoshihiro Abe. Effective resistances for supercritical percolation clusters in boxes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):935–946, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1435759235.

Angel:2012:SMC

[ABGG^{+}12] Omer Angel, Itai Benjamini, Ori Gurel-Gurevich, Tom Meyerovitch, and Ron Peled. Stationary map coloring. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):327–342, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1334148201.

Amir:2021:TSP

[ABGM21] Gideon Amir, Ofer Busani, Patrícia Gonçalves, and James B. Martin. The TAZRP speed process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1281–1305, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/The-TAZRP-speed-process/
10.1214/20-AIHP1117.full.

Arguin:2015:ETE

[ABK15] Louis-Pierre Arguin, Anton Bovier, and Nicola Kistler. An ergodic theorem for the extremal process of branching Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):557–569, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1428672682.

Antonopoulou:2018:SIL

[ABK18] D. C. Antonopoulou, D. Blömker, and G. D. Karali. The sharp interface limit for the stochastic Cahn–Hilliard equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):280–298, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1519030829.

Afanasyev:2014:CLT

[ABKV14] V. I. Afanasyev, Ch. Böinghoff, G. Kersting, and V. A. Vatutin. Conditional limit theorems for intermediately subcritical branching processes in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (2):602–627, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1395856142.

Aldous:2008:NMS

[ABL08] David J. Aldous, Charles Bordenave, and Marc Lelarge. Nearminimal spanning trees: a scaling exponent in probability models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5):962–976, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1222261920;   http://
www.numdam.org/item?id=AIHPB_2008__44_5_962_0.

Addario-Berry:2016:RIS

[ABL16] Louigi Addario-Berry and Nicholas Leavitt. Random infinite squarings of rectangles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):596–611, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1462367886.

Amir:2016:ZOL

[ABO16] Gideon Amir, Noam Berger, and Tal Orenshtein. Zero-one law for directional transience of one dimensional excited random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):47–57, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1452089259.

Abraham:2016:RBP

[Abr16] Céline Abraham. Rescaled bipartite planar maps converge to the Brownian map. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):575–595, May 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1462367885.

Addario-Berry:2017:JCR

[ABW17] Louigi Addario-Berry and Yuting Wen. Joint convergence of random quadrangulations and their cores. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1890–1920, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773730.

Auscher:1999:GLB

[AC99] Pascal Auscher and Thierry Coulhon. Gaussian lower bounds for random walks from elliptic regularity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35 (5):605–630, September 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_5_605_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399001090.

Attal:2004:QST

[AC04] Stéphane Attal and Agnès Coquio. Quantum stopping times and quasi-left continuity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(4):497–512, July/August 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_4_497_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000238.

Asselah:2007:NRW

[AC07] Amine Asselah and Fabienne Castell. A note on random walk in random scenery. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(2):163–173, March/April 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_2_163_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000409.

Angel:2015:PRM

[AC15] Omer Angel and Nicolas Curien. Percolations on random maps I: Half-plane models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):405–431, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1428672675.

Andreoletti:2018:RCG

[AC18] Pierre Andreoletti and Xinxin Chen. Range and critical generations of a random walk on Galton–Watson trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):466–513, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1519030836.

Aaron:2020:BD

[AC20] Catherine Aaron and Alejandro Cholaquidis. On boundary detection. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2028–2050, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/

volume-56/issue-3/On-boundary-detection/10.1214/19-
AIHP1027.full.

Azais:2002:ATQ

[ACA02] Jean-Marc Azaïs and Christine Cierco-Ayrolles. An asymptotic test for quantitative gene detection [Un test asymptotique pour la detection de gènes quantitatifs]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):1087–1092, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_1087_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020302011366.

Alfonsi:2020:SPM

[ACJ20] Aurélien Alfonsi, Jacopo Corbetta, and Benjamin Jourdain. Sampling of probability measures in the convex order by Wasserstein projection. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1706–1729, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Sampling-of-probability-
measures-in-the-convex-order-by-Wasserstein/10.1214/
19-AIHP1014.full.

Akemann:2019:FRP

[ACLS19] Gernot Akemann, Tomasz Checinski, Dang-Zheng Liu, and Eugene Strahov. Finite rank perturbations in products of coupled random matrices: From one correlated to two Wishart ensembles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):441–479, February 2019. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1547802406.

Arnaudon:2020:EIP

[ACLZ20] Marc Arnaudon, Ana Bela Cruzeiro, Christian Léonard, and Jean-Claude Zambrini. An entropic interpolation problem for incompressible viscous fluids. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2211–2235, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/An-entropic-interpolation-

problem-for-incompressible-viscous-fluids/10.1214/19-
AIHP1036.full.

Arias-Castro:2017:EPU

[ACRC17] Ery Arias-Castro and Alberto Rodríguez-Casal. On estimating the perimeter using the alpha-shape. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1051–1068, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1500624030.

Andjel:1985:QCP

[ACTR85] E. Andjel, C. Cocozza-Thivent, and M. Roussignol. Quelques compléments sur le processus des misanthropes et le processus^{≪} zéro-range ^{≫}. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(4):363–382, ???? 1985. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1985_
_21_4_363_0.

Asselah:1997:ORE

[AD97] Amine Asselah and Paolo Dai Pra. Occurrence of rare events in ergodic interacting spin systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(6):727–751, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1997__33_6_727_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397801112.

Asselah:2000:FOT

[AD00] Amine Asselah and Paolo Dai Pra. First occurrence time of a large density fluctuation for a system of independent random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(3):367–393, May 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_3_367_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001291.

Alili:2001:MBA

[AD01] L. Alili and R. A. Doney. Martin boundaries associated with a killed random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(3):313–338, May/June 2001.

CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_3_313_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010694.

Adams:2008:AFK

[AD08] Stefan Adams and Tony Dorlas. Asymptotic Feynman–Kac formulae for large symmetrised systems of random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5): 837–875, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1222261915;   http://www.numdam.org/
item?id=AIHPB_2008__44_5_837_0.

Abraham:2009:CBM

[AD09] Romain Abraham and Jean-François Delmas. Changing the branching mechanism of a continuous state branching process using immigration. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):226–238, February 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1234469979;   http://www.numdam.org/item?id=AIHPB_2009_
_45_1_226_0.

Aurzada:2011:IAS

[AD11] Frank Aurzada and Leif Döring. Intermittency and ageing for the symbiotic branching model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2):376–394, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1300887274;   http://www.numdam.org/item?id=AIHPB_2011_
_47_2_376_0.

Aurzada:2013:UAO

[AD13] Frank Aurzada and Steffen Dereich. Universality of the asymptotics of the one-sided exit problem for integrated processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):236–251, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1359470133.

Arguin:2014:NGS

[AD14] Louis-Pierre Arguin and Michael Damron. On the number of ground states of the Edwards–Anderson spin glass model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):28–62, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1388545264.

Applebaum:2015:GGL

[AD15a] David Applebaum and Anthony Dooley. A generalised Gangolli–Lévy–Khintchine formula for infinitely divisible measures and Lévy processes on semi-simple Lie groups and symmetric spaces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):599–619, May 2015. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1428672684.

Arous:2015:FER

[AD15b] Gérard Ben Arous and Kim Dang. On fluctuations of eigenvalues of random permutation matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):620–647, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1428672685.

Ahlberg:2015:SDF

[ADCKS15] Daniel Ahlberg, Hugo Duminil-Copin, Gady Kozma, and Vladas Sidoravicius. Seven-dimensional forest fires. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3): 862–866, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1435759231.

Adelman:1976:SUS

[Ade76] Omer Adelman. Some use of some “symmetries” of some random process. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(2): 193–197, April–June 1976. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1976__
12_2_193_0.

Albeverio:2021:ESQ

[ADG21] Sergio Albeverio, Francesco C. De Vecchi, and Massimiliano Gubinelli. The elliptic stochastic quantization of some two dimensional Euclidean QFTs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2372–2414, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/The-elliptic-stochastic-
quantization-of-some-two-dimensional-Euclidean-QFTs/
10.1214/20-AIHP1145.full.

Azais:2013:NEJ

[ADGP13] Romain Azaïs, François Dufour, and Anne Gégout-Petit. Nonparametric estimation of the jump rate for non-homogeneous marked renewal processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1204–1231, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1380718744.

Abraham:2012:PGW

[ADH12] Romain Abraham, Jean-François Delmas, and Hui He. Pruning Galton–Watson trees and tree-valued Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):688–705, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1340714868.

Angel:2019:LLR

[ADHV19] Omer Angel, Duncan Dauvergne, Alexander E. Holroyd, and Bálint Virág. The local limit of random sorting networks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):412–440, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1547802405.

Azais:2016:CZC

[ADL16] Jean-Marc Azaïs, Federico Dalmao, and José R. León. CLT for the zeros of classical random trigonometric polynomials. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):804–820, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1462367894.

Azema:1967:NMI

[ADR67] Jacques Azema, Marie Duflo, and Daniel Revuz. Note sur la mesure invariante des processus de Markov récurrents. (French) [Note on the invariant measure of recurrent Markov processes]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(4):397–402, October– December 1967. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1967__3_4_397_0.

Alvarez-Esteban:2011:UAC

[ÁEdBCAM11] P. C. Álvarez-Esteban, E. del Barrio, J. A. Cuesta-Albertos, and C. Matrán. Uniqueness and approximate computation of optimal incomplete transportation plans. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2): 358–375, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887273;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_358_0.

Ankirchner:2020:SEP

[AEFdR20] Stefan Ankirchner, Stefan Engelhardt, Alexander Fromm, and Gonçalo dos Reis. The Skorokhod embedding problem for inhomogeneous diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1606–1640, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/The-Skorokhod-embedding-
problem-for-inhomogeneous-diffusions/10.1214/19-AIHP1012.
full.

Alt:2019:LSK

[AEKN19] Johannes Alt, László Erdős, Torben Krüger, and Yuriy Nemish. Location of the spectrum of Kronecker random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):661–696, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1557820827.

Asselah:2016:FVS

[AFGJ16] Amine Asselah, Pablo A. Ferrari, Pablo Groisman, and Matthieu Jonckheere. Fleming–Viot selects the minimal quasistationary distribution: The Galton–Watson case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52 (2):647–668, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1462367888.

Avena:2015:SER

[AFJV15] Luca Avena, Tertuliano Franco, Milton Jara, and Florian Völlering. Symmetric exclusion as a random environment: Hydrodynamic limits. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):901–916, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1435759233.

Arcones:1989:BMA

[AG89] Miguel A. Arcones and Evarist Giné. The bootstrap of the mean with arbitrary bootstrap sample size. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):457–481, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?

id=AIHPB_1991__27_4_583_0. See additions and correction [AG91].

Arcones:1991:ACB

[AG91] Miguel A. Arcones and Evarist Giné. Additions and correction to “The bootstrap of the mean with arbitrary bootstrap sample”. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(4):583–595, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//www.numdam.org/item?id=AIHPB_1989__25_4_457_0. See

[AG89].

Ang:2021:LQG

[AG21] Morris Ang and Ewain Gwynne. Liouville quantum gravity surfaces with boundary as matings of trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):1– 53, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-1/Liouville-quantum-
gravity-surfaces-with-boundary-as-matings-of-trees/
10.1214/20-AIHP1068.full.

Aru:2020:EGR

[AGJ^{+}20] Juhan Aru, Carla Groenland, Tom Johnston, Bhargav Narayanan, Alex Roberts, and Alex Scott. Exceptional graphs for the random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2017–2027, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Exceptional-graphs-for-
the-random-walk/10.1214/19-AIHP1026.full.

Arous:1994:HNS

[AGL94] Gérard Ben Arous, Mihai Gradinaru, and Michel Ledoux. Hölder norms and the support theorem for diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(3): 415–436, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1994__30_3_415_0.

Albeverio:1977:HPA

[AHK77] Sergio Albeverio and Raphael Høegh-Krohn. Hunt processes and analytic potential theory on rigged Hilbert spaces. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(3):269–291, July–September 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_3_269_0.

Ahmad:1965:EAE

[Ahm65] Salah Ahmad. Éléments aléatoires dans les espaces vectoriels topologiques. (French) [Random elements in topological vector spaces]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(2):95– 135, October–December 1965. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1965__
2_2_95_0.

Ahn:2020:GUM

[Ahn20] Andrew Ahn. Global universality of Macdonald plane partitions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1641–1705, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/Global-universality-of-Macdonald-
plane-partitions/10.1214/19-AIHP1013.full.

Arous:2013:ERB

[AHOZ13] Gerard Ben Arous, Yueyun Hu, Stefano Olla, and Ofer Zeitouni. Einstein relation for biased random walk on Galton– Watson trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):698–721, August 2013. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1372772641.

Aspandiiarov:1999:TDR

[AI99] Sanjar Aspandiiarov and Roudolf Iasnogorodski. Threedimensional reflected driftless random walks in troughs: new asymptotic behavior. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(1):49–83, January/February 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_1_49_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399800053.

Ankirchner:2005:FUF

[AI05] Stefan Ankirchner and Peter Imkeller. Finite utility on financial markets with asymmetric information and structure properties of the price dynamics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):479–503, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_479_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000117.

Aidekon:2010:LDT

[Aid10] Elie Aidékon. Large deviations for transient random walks in random environment on a Galton–Watson tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1): 159–189, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1267454113;   http://www.numdam.org/
item?id=AIHPB_2010__46_1_159_0.

Aidekon:2020:HDU

[Aïd20] Elie Aïdékon. Hausdorff dimension of the uniform measure of Galton–Watson trees without the X log X condition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2301–2306, November 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-4/Hausdorff-dimension-of-the-uniform-
measure-of-GaltonWatson-trees-without/10.1214/19-AIHP1031.
full.

Abou-Jaoude:1976:CNS

[AJ76a] Saab Abou-Jaoude. Conditions nécessaires et suffisantes de convergence L_{1} en probabilité de l’histogramme pour une densité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12 (3):213–231, July–September 1976. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1976__12_3_213_0.

Abou-Jaoude:1976:CEP

[AJ76b] Saab Abou-Jaoude. Sur la convergence L_{1} et L_{∞} de l’estimateur de la partition aléatoire pour une densité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(4):299–317, October–December 1976. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1976__
12_4_299_0.

Azema:1976:PMF

[AJ76c] J. Azéma and T. Jeulin. Précisions sur la mesure de föllmer. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(3): 257–283, July–September 1976. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1976__12_3_257_0.

Alon:2020:CO

[AK20] Gil Alon and Gady Kozma. Comparing with octopi. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2672–2685, November 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-4/Comparing-with-octopi/10.1214/20-
AIHP1054.full.

Alon:2021:MFQ

[AK21] Gil Alon and Gady Kozma. The mean-field quantum Heisenberg ferromagnet via representation theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1203–1228, August 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-3/The-mean-field-quantum-Heisenberg-
ferromagnet-via-representation-theory/10.1214/20-AIHP1067.
full.

Azema:1966:RFP

[AKDR66] Jacques Azéma, Marie Kaplan-Duflo, and Daniel Revuz. Récurrence fine des processus de Markov. (French) [Fine recurrence of Markov processes]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(3):185–220, January–March 1966. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1966__2_3_185_0.

Albeverio:2001:UGM

[AKKR01] Sergio Albeverio, Yuri Kondratiev, Yuri Kozitsky, and Michael Röckner. Uniqueness for Gibbs measures of quantum lattices in small mass regime. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(1):43–69, January 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_1_43_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010578.

Adamczak:2021:PTI

[AKM21] Radosław Adamczak, Michał Kotowski, and Piotr Miłoś. Phase transition for the interchange and quantum Heisenberg models on the Hamming graph. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):273–325, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-

statistiques/volume-57/issue-1/Phase-transition-for-
the-interchange-and-quantum-Heisenberg-models-on/10.
1214/20-AIHP1079.full.

Akonom:1993:CAD

[Ako93] Jacques Akonom. Comportement asymptotique du temps d’occupation du processus des sommes partielles. (French) [The asymptotic behavior of the occupation time of the process of partial sums]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(1):57–81, ???? 1993. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1993__29_1_57_0.

Agrachev:2007:FDP

[AKSS07] A. Agrachev, S. Kuksin, A. Sarychev, and A. Shirikyan. On finite-dimensional projections of distributions for solutions of randomly forced 2D Navier–Stokes equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43 (4):399–415, July/August 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_4_399_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000732.

Ankirchner:2017:FLT

[AKU17] Stefan Ankirchner, Thomas Kruse, and Mikhail Urusov. A functional limit theorem for irregular SDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3): 1438–1457, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1500624047.

Ankirchner:2020:FLT

[AKU20] Stefan Ankirchner, Thomas Kruse, and Mikhail Urusov. A functional limit theorem for coin tossing Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2996–3019, November 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-4/A-functional-limit-theorem-for-coin-
tossing-Markov-chains/10.1214/20-AIHP1066.full.

Adamczak:2012:TME

[AL12] Radosław Adamczak and Rafał Latała. Tail and moment estimates for chaoses generated by symmetric random variables with logarithmically concave tails. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1103–1136, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1353098442.

Allain:1978:API

[All78] Marie-France Allain. Approximation par des intégrales de Stieltjes–Lebesgue d’intégrales stochastiques relatives au mouvement brownien indexé par R_{+}^{d}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4):441–464, October–December 1978. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1978__14_4_441_0.

Adamczak:2020:HWI

[ALM20] Radosław Adamczak, Rafał Latała, and Rafał Meller. Hanson-Wright inequality in Banach spaces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2356–2376, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/HansonWright-inequality-
in-Banach-spaces/10.1214/19-AIHP1041.full.

Alquier:2008:IFS

[Alq08] Pierre Alquier. Iterative feature selection in least square regression estimation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1):47–88, February 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1203969868;   http://www.numdam.org/item?id=AIHPB_2008_
_44_1_47_0.

Alt:1987:LGN

[Alt87] Jean-Christian Alt. La loi des grands nombres de Prokhorov dans les espaces de type p. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(4):561–574, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1987__23_4_561_0.

Arnaudon:1999:MVM

[ALT99] Marc Arnaudon, Xue-Mei Li, and Anton Thalmaier. Manifoldvalued martingales, changes of probabilities, and smoothness of finely harmonic maps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(6):765–791, November 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_6_765_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001144.

Alvarez:1973:RNE

[Alv73] Miguel Angel Garcia Alvarez. Représentation des noyaux excessifs. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(3):277– 283, July–September 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_3_277_0.

Azais:2017:RMP

[AMD17] Jean-Marc Azaïs, Stéphane Mourareau, and Yohann De Castro. A rice method proof of the null-space property over the Grassmannian. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1821–1838, November 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1511773727.

Amir:2017:ODL

[Ami17] Gideon Amir. One-dimensional long-range diffusion-limited aggregation III — the limit aggregate. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1513–1527, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773716.

Andjel:1995:EED

[AMS95] Enrique D. Andjel, Thomas S. Mountford, and Roberto H. Schonmann. Equivalence of exponential decay rates for bootstrap percolation like cellular automata. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):13–25, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1995__31_1_13_0.

Aurzada:2021:PEM

[AMZ21] Frank Aurzada, Sumit Mukherjee, and Ofer Zeitouni. Persistence exponents in Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1411–1441, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Persistence-exponents-
in-Markov-chains/10.1214/20-AIHP1114.full.

Aubrun:2009:SDI

[AN09] Guillaume Aubrun and Ion Nechita. Stochastic domination for iterated convolutions and catalytic majorization. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3): 611–625, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1249391377;   http://www.numdam.org/
item?id=AIHPB_2009__45_3_611_0.

Andres:2009:PDS

[And09] Sebastian Andres. Pathwise differentiability for SDEs in a convex polyhedron with oblique reflection. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1): 104–116, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469973;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_104_0.

Andres:2014:IPR

[And14] Sebastian Andres. Invariance principle for the random conductance model with dynamic bounded conductances. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):352–374, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1395856131.

Anderson:2015:LLL

[And15] Greg W. Anderson. A local limit law for the empirical spectral distribution of the anticommutator of independent Wigner matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):809–841, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1435759229.

Ane:2001:COF

[Ané01] Cécile Ané. Clark–Ocone formulas and Poincaré inequalities on the discrete cube. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(1):101–137, January 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_1_101_0;   http://www.sciencedirect.com/science/
article/pii/S024602030001061X.

Angst:2016:ABR

[Ang16] Jürgen Angst. Asymptotic behavior of a relativistic diffusion in Robertson–Walker space–times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):376–411, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089273.

Anonymous:1966:EF

[Ano66] Anonymous. Errata. (French) [Errata]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(4):0, April–June 1966. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1966__2_4_279_0. See [Tor66].

Anonymous:1977:CBA

[Ano77] Anonymous. Complément à la bibliographie de l’article de A. Tortrat: ^{≪} τ-régularité des lois, séparation au sens de A. Tulcea et propriété de Radon–Nicodym ^{≫}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(1):43, January–March 1977. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1976__12_2_131_0.

Anonymous:1990:MMM

[Ano90] Anonymous. In memoriam M. Métivier. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):1–4, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1990__26_1_1_0.

Anonymous:1997:EB

[Ano97] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(1):ifc, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020397900005.

Anonymous:1999:EB

[Ano99] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(1):ifc, January/ February 1999. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020399900006.

Anonymous:2004:EBa

[Ano04a] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):ifc, January/ February 2004. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020304000032.

Anonymous:2004:EBb

[Ano04b] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2):ifc, March/April 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020304000135.

Anonymous:2004:EBc

[Ano04c] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):ifc, May/June 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020304000329.

Anonymous:2004:EBd

[Ano04d] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(4):ifc, July/August 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S024602030400041X.

Anonymous:2004:EBe

[Ano04e] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(5):ifc, September/ October 2004. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020304000688.

Anonymous:2004:EBf

[Ano04f] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(6):ifc, November/ December 2004. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020304000779.

Anonymous:2005:AIH

[Ano05a] Anonymous. Annales de l’institut Henri Poincaré. Probabilités et statistiques. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):255–, 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0246020305000099.

Ce numéro spécial est dedié à la mémoire de Paul-André Meyer (1934–2003).

Anonymous:2005:EBa

[Ano05b] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(1):co2, January/ February 2005. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020305000038.

Anonymous:2005:EBb

[Ano05c] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(2):co2, March/April 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020305000282.

Anonymous:2005:EBc

[Ano05d] Anonymous. Editorial board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):co2, 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.sciencedirect.com/science/

article/pii/S0246020305000373. En hommage a Paul André

Meyer.

Anonymous:2005:EBd

[Ano05e] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):co2, July/August 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020305000506.

Anonymous:2005:EBe

[Ano05f] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5):co2, September/ October 2005. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020305000713.

Anonymous:2005:EBf

[Ano05g] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(6):co2, November/ December 2005. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020305000786.

Anonymous:2006:EBa

[Ano06a] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1):co2, January/ February 2006. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020305001019.

Anonymous:2006:EBb

[Ano06b] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(2):co2, March/April 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020306000070.

Anonymous:2006:EBc

[Ano06c] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(3):co2, May/June 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020306000239.

Anonymous:2006:EBd

[Ano06d] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(4):co2, July/August 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S024602030600032X.

Anonymous:2006:EBe

[Ano06e] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5):co2, September/ October 2006. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020306000483.

Anonymous:2006:EBf

[Ano06f] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6):co2, November/ December 2006. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020306000598.

Anonymous:2007:EBa

[Ano07a] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1):co2, January/ February 2007. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S024602030600077X.

Anonymous:2007:EBb

[Ano07b] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(2):ifc, March/April 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S024602030700012X.

Anonymous:2007:EBc

[Ano07c] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(3):ifc, May/June 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020307000192.

Anonymous:2007:EBd

[Ano07d] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(4):ifc, July/August 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S024602030700026X.

Anonymous:2007:EBe

[Ano07e] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(5):ifc, September/ October 2007. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020307000337.

Anonymous:2007:EBf

[Ano07f] Anonymous. Editorial Board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6):ifc, November/ December 2007. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020307000404.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):??, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723507.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):??, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1479373237.

Anonymous:2016:TCa

[Ano16c] Anonymous. Table of contents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):??, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723506.

Anonymous:2016:TCb

[Ano16d] Anonymous. Table of contents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):??, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1479373236.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial board. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):??, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544881.

Anonymous:2017:TCa

[Ano17b] Anonymous. Table of contents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):??, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544880.

Aaronson:2006:EMS

[ANS06] J. Aaronson, H. Nakada, and O. Sarig. Exchangeable measures for subshifts. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6):727–751, November/December 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_6_727_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000136.

Antoniadis:1982:CPE

[Ant82] Anestis Antoniadis. Sur certains problèmes d’estimation et de test concernant la moyenne d’un processus gaussien. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(3):223–236, July–September 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1982__18_3_223_0.

Aimino:2017:RSS

[ANT17] Romain Aimino, Matthew Nicol, and Mike Todd. Recurrence statistics for the space of interval exchange maps and the Teichmüller flow on the space of translation surfaces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):

1371–1401, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1500624045.

Alos:2000:SHE

[ANV00] E. Alòs, D. Nualart, and F. Viens. Stochastic heat equation with white-noise drift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(2):181–218, March 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_2_181_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001229.

Aldous:1998:TVM

[AP98] David Aldous and Jim Pitman. Tree-valued Markov chains derived from Galton–Watson processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(5):637–686, September/October 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_5_637_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800034.

Attal:2005:LAC

[AP05] Stéphane Attal and Yan Pautrat. From (n+1)-level atom chains to n-dimensional noises. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):391–407, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_391_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000129.

Alsmeyer:2006:MEB

[AR06] Gerold Alsmeyer and Uwe Rösler. The Martin entrance boundary of the Galton–Watson process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5):591–606, September/October 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_5_591_0;   http://
www.sciencedirect.com/science/article/pii/S024602030500097X.

Alsmeyer:2008:SFP

[AR08] Gerold Alsmeyer and Uwe Rösler. A stochastic fixed point equation for weighted minima and maxima. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1): 89–103, February 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1203969869;   http://www.numdam.org/
item?id=AIHPB_2008__44_1_89_0.

Asselah:2016:FID

[AR16] Amine Asselah and Houda Rahmani. Fluctuations for internal DLA on the comb. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):58–83, February 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1452089260.

Archer:2021:BMS

[Arc21] Eleanor Archer. Brownian motion on stable looptrees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):940–979, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-
2/Brownian-motion-on-stable-looptrees/10.1214/20-AIHP1103.
full.

Arnaudon:1993:CLS

[Arn93a] Marc Arnaudon. Caractéristiques locales des semi-martingales et changements de probabilités. (French) [Local characteristics of semi-martingales and probability changes]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(2): 251–267, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_2_251_0.

Arnaudon:1993:SME

[Arn93b] Marc Arnaudon. Semi-martingales dans les espaces homogènes. (French) [Semi-martingales in homogeneous spaces]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(2): 269–288, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_2_269_0.

Adhikari:2016:DPP

[ARRS16] Kartick Adhikari, Nanda Kishore Reddy, Tulasi Ram Reddy, and Koushik Saha. Determinantal point processes in the plane from products of random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):16–46, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089258.

Ansel:1992:LMD

[AS92] Jean-Pascal Ansel and Christophe Stricker. Lois de martingale, densités et décomposition de Föllmer Schweizer. (French) [Martingale laws, densities, and Föllmer Schweizer decomposition]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(3):375–392, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1992__28_3_375_0.

Ansel:1994:CAC

[AS94] Jean-Pascal Ansel and Christophe Stricker. Couverture des actifs contingents et prix maximum. (French) [Coverage of contingent assets and maximum price]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):303–315, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1994__30_2_303_0.

Ahlberg:2017:SLT

[AS17] Daniel Ahlberg and Jeffrey E. Steif. Scaling limits for the threshold window: When does a monotone Boolean function flip its outcome? Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):2135–2161, November 2017. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1511773741.

Abacherli:2018:LSP

[AS18] Angelo Abächerli and Alain-Sol Sznitman. Level-set percolation for the Gaussian free field on a transient tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):173–201, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1519030824.

Assani:1987:QTE

[Ass87] I. Assani. Quelques théorèmes ergodiques dans les espaces L_{E}^{p}. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(2):209–224, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_2_209_0.

Andjel:1990:EPO

[ASS90] Enrique D. Andjel, Rinaldo B. Schinazi, and Roberto H. Schonmann. Edge processes of one dimensional stochastic growth models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):489–506, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1990__26_3_489_0.

Assani:1998:WPE

[Ass98] I. Assani. A weighted pointwise ergodic theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34 (1):139–150, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_1_139_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800216.

Assani:2000:MRT

[Ass00] I. Assani. Multiple return times theorems for weakly mixing systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(2):153–165, March 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_2_153_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001205.

Asselah:2010:STU

[Ass10] Amine Asselah. Shape transition under excess self-intersections for transient random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):250–278, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1267454116;   http://www.numdam.org/item?id=AIHPB_2010_
_46_1_250_0.

Asselah:2011:AUT

[Ass11] Amine Asselah. Annealed upper tails for the energy of a charged polymer. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):80–110, February 2011. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1294170231;   http://
www.numdam.org/item?id=AIHPB_2011__47_1_80_0.

Assiotis:2020:HPD

[Ass20] Theodoros Assiotis. Hua–Pickrell diffusions and Feller processes on the boundary of the graph of spectra. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1251–1283, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-2/HuaPickrell-diffusions-
and-Feller-processes-on-the-boundary-of-the/10.1214/
19-AIHP1001.full.

Astbury:1976:POS

[Ast76] K. A. Astbury. A potential operator and some ergodic properties of a positive L_{∞} contraction. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(2):151–162, April–June 1976. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1976__12_2_151_0.

Ayache:2011:MMB

[ASX11] Antoine Ayache, Narn-Rueih Shieh, and Yimin Xiao. Multiparameter multifractional Brownian motion: Local nondeterminism and joint continuity of the local times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1029–1054, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1317906500;   http://
www.numdam.org/item?id=AIHPB_2011__47_4_1029_0.

Angst:2020:PBR

[AT20] Jürgen Angst and Camille Tardif. On the Poisson boundary of the relativistic Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2792–2821, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/On-the-Poisson-boundary-
of-the-relativistic-Brownian-motion/10.1214/20-AIHP1059.
full.

Attal:1995:REE

[Att95] Stéphane Attal. Représentation des endomorphismes de l’espace de Wiener qui préservent les martingales. (French) [Representation of endomorphisms of the Wiener space that preserve martingales]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(3):467–484, ???? 1995. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1995_
_31_3_467_0.

Arizmendi:2020:SMF

[ATV20] Octavio Arizmendi, Pierre Tarrago, and Carlos Vargas. Subordination methods for free deconvolution. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2565–2594, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/Subordination-methods-
for-free-deconvolution/10.1214/20-AIHP1050.full.

Audibert:2004:AEE

[Aud04] Jean-Yves Audibert. Aggregated estimators and empirical complexity for least square regression. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(6):685–736, November/December 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_6_685_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000299.

Augeri:2018:LDT

[Aug18] Fanny Augeri. On the large deviations of traces of random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2239–2285, November 2018. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1539849798.

Austin:2015:ERM

[Aus15] Tim Austin. Exchangeable random measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3): 842–861, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1435759230.

Angel:2019:LLS

[AvdHH19] Omer Angel, Remco van der Hofstad, and Cecillia Holmgren. Limit laws for self-loops and multiple edges in the configuration model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1509–1530, August 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1569398877.

Aaronson:2018:DLP

[AW18] Jon Aaronson and Benjamin Weiss. Distributional limits of positive, ergodic stationary processes and infinite ergodic transformations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):879–906, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1524643233.

Aeckerle-Willems:2021:CSE

[AWS21] Cathrine Aeckerle-Willems and Claudia Strauch. Concentration of scalar ergodic diffusions and some statistical implications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):1857–1887, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/Concentration-of-scalar-
ergodic-diffusions-and-some-statistical-implications/
10.1214/20-AIHP1144.full.

Ayache:2008:JCL

[AWX08] Antoine Ayache, Dongsheng Wu, and Yimin Xiao. Joint continuity of the local times of fractional Brownian sheets. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4): 727–748, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1217964117;   http://www.numdam.org/
item?id=AIHPB_2008__44_4_727_0.

Atar:1997:ESN

[AZ97] Rami Atar and Ofer Zeitouni. Exponential stability for nonlinear filtering. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(6):697–725, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_6_697_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801100.

Arous:1999:IPC

[AZ99] G. Ben Arous and O. Zeitouni. Increasing propagation of chaos for mean field models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(1):85–102, January/February 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_1_85_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399800065.

Aaronson:2014:LTS

[AZ14] Jon Aaronson and Roland Zweimüller. Limit theory for some positive stationary processes with infinite mean. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1): 256–284, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1388545274.

Azais:1989:ATT

[Aza89] Jean-Marc Azaïs. Approximation des trajectoires et temps local des diffusions. (French) [Approximation of trajectories and local times of diffusions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2):175–194, ???? 1989. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1989_
_25_2_175_0.

Ben-Ari:2009:LDP

[BA09] Iddo Ben-Ari. Large deviations for partition functions of directed polymers in an IID field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):770–792, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1249391384;   http://www.numdam.org/item?id=AIHPB_2009_
_45_3_770_0.

Babillot:1988:TDR

[Bab88] M. Babillot. Théorie du renouvellement pour des chaînes semimarkoviennes transientes. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(4):507–569, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1988__24_4_507_0.

Babillot:1991:CAD

[Bab91] M. Babillot. Comportement asymptotique du mouvement brownien sur une variété homogène à courbure négative ou nulle. (French) [Asymptotic behavior of Brownian motion on a homogeneous variety with negative or zero curvature]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(1): 61–90, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_1_61_0.

Berenfeld:2021:SRP

[BAC21] Clément Berenfeld and Ery Arias-Castro. Some random paths with angle constraints. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):116–131, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/Some-random-paths-with-
angle-constraints/10.1214/20-AIHP1073.full.

Bagchi:1985:CCM

[Bag85] Sitadri Bagchi. On a. s. convergence of classes of multivalued asymptotic martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(4):313–321, ???? 1985. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1985_
_21_4_313_0.

Bailleul:2010:SAR

[Bai10] Ismaël Bailleul. A stochastic approach to relativistic diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3):760–795, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1281100398;   http://
www.numdam.org/item?id=AIHPB_2010__46_3_760_0.

Baldi:1974:EPF

[Bal74] Paolo Baldi. Sur l’équation de Poisson. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(4):423–434, October–December 1974. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1974__10_4_423_0.

Baldi:1976:TDG

[Bal76] Paolo Baldi. Sur la transience du groupe des déplacements de l’espace. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(3):285–290, July–September 1976. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1976__12_3_285_0.

Baldi:1981:CGL

[Bal81] P. Baldi. Caractérisation des groupes de Lie connexes récurrents. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(3):281–308, July–September 1981. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1981__17_3_281_0.

Balazs:2003:GFC

[Bal03] Márton Balázs. Growth fluctuations in a class of deposition models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(4):639–685, July/August 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_2_255_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000190. See correction [Bal04].

Balazs:2004:CPG

[Bal04] Márton Balázs. Correction to the paper “Growth fluctuations in a class of deposition models”: Annales de l’Institut Henri Poincaré Probabilités et statistiques 39 (2003) 639– 685. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2):255–258, March/April 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_4_639_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000633. See [Bal03].

Banos:2018:BEL

[Bañ18] David Baños. The Bismut-Elworthy-Li formula for meanfield stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):220–233, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1519030826.

Barra:1981:PRB

[Bar81] J. R. Barra. À propos d’un résultat de Brailovsky concernant une probabilité d’erreur en analyse discriminante. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(1):21–29, January– March 1981. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1981__17_1_21_0.

Barral:2000:DMP

[Bar00] Julien Barral. Differentiability of multiplicative processes related to branching random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(4):407–417, July 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_4_407_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001254.

Barret:2015:SAM

[Bar15] Florent Barret. Sharp asymptotics of metastable transition times for one dimensional SPDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):129–166, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1421244401.

Bass:1969:FSF

[Bas69] J. Bass. Fonctions stationnaires. Fonctions de corrélation. Application à la représentation spatio-temporelle de la turbulence. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(2): 135–193, April–June 1969. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1969__
5_2_135_0.

Bastero:1982:RFT

[Bas82] J. Bastero. A restricted form of the theorem of Maurey–Pisier for the cotype in p-Banach spaces. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(3):293– 304, July–September 1982. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1982__
18_3_293_0.

Bass:1989:FCF

[Bas89] J. Bass. Fonctions de corrélation des fonctions pseudoaléatoires. (French) [Correlation functions of pseudorandom functions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):503–515, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1989__25_4_503_0.

Basdevant:2008:ESE

[Bas08] Anne-Laure Basdevant. On the equivalence of some eternal additive coalescents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1020–1037, December 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1227287563;   http://www.numdam.org/item?id=AIHPB_2008_
_44_6_1020_0.

Basile:2014:KED

[Bas14] Giada Basile. From a kinetic equation to a diffusion under an anomalous scaling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1301–1322, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1413555501.

Batakis:2000:CPD

[Bat00] Athanassios Batakis. A continuity property of the dimension of the harmonic measure of Cantor sets under perturbations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(1):87–107, January 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_1_87_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001175.

Bauernschubert:2013:PTR

[Bau13] Elisabeth Bauernschubert. Perturbing transient random walk in a random environment with cookies of maximal strength.

Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):638–653, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1375967013.

Baxter:1996:DAF

[Bax96] Martin Baxter. Discounted additive functionals of Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(5):623–644, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1996__32_5_623_0.

Bayart:2015:CLT

[Bay15] Frédéric Bayart. Central limit theorems in linear dynamics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):1131–1158, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1435759243.

Bosq:1978:ECE

[BB78]  Denis Bosq and Jacques Bleuez.  Étude d’une classe

d’estimateurs non-paramétriques de la densité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4):479–498, October–December 1978. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1978__14_4_
479_0.

Barlow:1989:CBM

[BB89] Martin T. Barlow and Richard F. Bass. The construction of Brownian motion on the Sierpiński carpet. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(3): 225–257, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1989__25_3_225_0.

Bass:2001:SBL

[BB01] Richard F. Bass and Krzysztof Burdzy. The supremum of Brownian local times on Hölder curves. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(6):627–642, November/December 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_5_799_0;   http://

www.sciencedirect.com/science/article/pii/S0246020300010724.

See erratum [BB02].

Bass:2002:ESB

[BB02] Richard F. Bass and Krzysztof Burdzy. Erratum to “The supremum of Brownian local times on Hölder curves”. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(5):799–800, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_6_627_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302011214.

See [BB01].

Benjamini:2011:ITT

[BB11] Itai Benjamini and Nathanaël Berestycki. An integral test for the transience of a Brownian path with limited local time. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2):539–558, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887281;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_539_0.

Bansaye:2014:SPV

[BB14a] Vincent Bansaye and Christian Böinghoff. Small positive values for supercritical branching processes in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):770–805, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1403276998.

Baraud:2014:ECF

[BB14b] Yannick Baraud and Lucien Birgé. Estimating composite functions by model selection. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):285–314, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1388545275.

Basu:2017:LTL

[BB17a] Riddhipratim Basu and Nayantara Bhatnagar. Limit theorems for longest monotone subsequences in random Mallows permutations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1934–1951, November 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1511773732.

Borodin:2017:ILM

[BB17b] Alexei Borodin and Alexey Bufetov. An irreversible local Markov chain that preserves the six vertex model on a torus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):451–463, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1486544898.

Balazs:2018:PBM

[BB18] Márton Balázs and Ross Bowen. Product blocking measures and a particle system proof of the Jacobi triple product. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):514–528, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1519030837.

Bank:2021:SRT

[BB21] Peter Bank and David Besslich. On a stochastic representation theorem for Meyer-measurable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1336–1368, August 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-3/On-a-stochastic-representation-theorem-
for-Meyer-measurable-processes/10.1214/20-AIHP1113.
full.

Baldi:1978:TCL

[BBC78] Paolo Baldi, Philippe Bougerol, and Pierre Crepel. Théorème central limite local sur les extensions compactes de R^{d}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(1):99–111, January–March 1978. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1978__14_1_
99_0.

Bass:2005:URB

[BBC05] Richard F. Bass, Krzysztof Burdzy, and Zhen-Qing Chen. Uniqueness for reflecting Brownian motion in lip domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(2):197–235, March/April 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_2_197_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000615.

Barilari:2021:SPS

[BBCH21] Davide Barilari, Ugo Boscain, Daniele Cannarsa, and Karen Habermann. Stochastic processes on surfaces in threedimensional contact sub-Riemannian manifolds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57 (3):1388–1410, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-
3/Stochastic-processes-on-surfaces-in-three-dimensional-
contact-sub-Riemannian/10.1214/20-AIHP1124.full.

Brofferio:2012:IMR

[BBD12] Sara Brofferio, Dariusz Buraczewski, and Ewa Damek. On the invariant measure of the random difference equation X_{n} = A_{n}X_{n−1} + B_{n} in the critical case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):377–395, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1334148204.

Banerjee:2019:GVB

[BBD19] Sayan Banerjee, Krzysztof Burdzy, and Mauricio Duarte. Gravitation versus Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1531–1565, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1569398878.

Berger:2008:AHK

[BBHK08] N. Berger, M. Biskup, C. E. Hoffman, and G. Kozma. Anomalous heat-kernel decay for random walk among bounded random conductances. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2):374–392, April 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/

1207948225;   http://www.numdam.org/item?id=AIHPB_2008_
_44_2_374_0.

Berestycki:2014:ASF

[BBL14] Julien Berestycki, Nathanaël Berestycki, and Vlada Limic. Asymptotic sampling formulae for Λ-coalescents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3): 715–731, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1403276996.

Berestycki:2008:STB

[BBS08] Julien Berestycki, Nathanaël Berestycki, and Jason Schweinsberg. Small-time behavior of beta coalescents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2): 214–238, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1207948217;   http://www.numdam.org/
item?id=AIHPB_2008__44_2_214_0.

Berestycki:2015:CBB

[BBS15] Julien Berestycki, Nathanaël Berestycki, and Jason Schweinsberg. Critical branching Brownian motion with absorption: Particle configurations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1215–1250, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1445432041.

Barbour:2017:GRW

[BC17] A. D. Barbour and Andrea Collevecchio. General random walk in a random environment defined on Galton–Watson trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1657–1674, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1511773722.

Boedihardjo:2019:IBB

[BC19] Horatio Boedihardjo and Ilya Chevyrev. An isomorphism between branched and geometric rough paths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1131–1148, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820845.

Bates:2020:FLB

[BC20] Erik Bates and Sourav Chatterjee. Fluctuation lower bounds in planar random growth models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2406–2427, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/Fluctuation-lower-
bounds-in-planar-random-growth-models/10.1214/19-AIHP1043.
full.

Bertini:2002:SGG

[BCC02] Lorenzo Bertini, Nicoletta Cancrini, and Filippo Cesi. The spectral gap for a Glauber-type dynamics in a continuous gas. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(1):91–108, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_1_91_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010858.

Biau:2015:NIA

[BCG15] Gérard Biau, Frédéric Cérou, and Arnaud Guyader. New insights into Approximate Bayesian Computation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1): 376–403, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1421244410.

Basrak:2021:EES

[BCHJ21] Bojan Basrak, Yeonok Cho, Johannes Heiny, and Paul Jung. Extreme eigenvalue statistics of m -dependent heavytailed matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2100–2127, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/Extreme-eigenvalue-statistics-
of-m-dependent-heavy-tailed-matrices/10.1214/21-AIHP1152.
full.

Bretagnolle:1966:LSE

[BCK66] Jean Bretagnolle, Didier Dacunha Castelle, and Jean-Louis Krivine. Lois stables et espaces L^{p}. (French) [Stable laws and L^{p} spaces]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(3): 231–259, January–March 1966. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1966__
2_3_231_0.

Bertoin:2021:CSS

[BCK21] Jean Bertoin, Nicolas Curien, and Igor Kortchemski. On conditioning a self-similar growth-fragmentation by its intrinsic area. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):1136–1156, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/On-conditioning-a-self-similar-growth-
fragmentation-by-its-intrinsic/10.1214/20-AIHP1110.
full.

Binder:2015:RCC

[BCL15] I. Binder, L. Chayes, and H. K. Lei. On the rate of convergence for critical crossing probabilities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):672–715, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1428672687.

Bally:2019:TED

[BCP19] Vlad Bally, Lucia Caramellino, and Paolo Pigato. Tube estimates for diffusions under a local strong Hörmander condition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2320–2369, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/Tube-estimates-for-diffusions-
under-a-local-strong-H%c3%b6rmander-condition/10.1214/
18-AIHP950.full.

Bufetov:2021:CMP

[BCQ21] Alexander I. Bufetov, Fabio Deelan Cunden, and Yanqi Qiu. Conditional measures for Pfaffian point processes: Conditioning on a bounded domain. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):856–871, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-2/Conditional-measures-
for-Pfaffian-point-processes--Conditioning-on-a/10.
1214/20-AIHP1099.full.

Bass:2009:LDR

[BCR09] Richard Bass, Xia Chen, and Jay Rosen. Large deviations for Riesz potentials of additive processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3): 626–666, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1249391378;   http://www.numdam.org/
item?id=AIHPB_2009__45_3_626_0.

Borodin:2015:MFE

[BCR15] Alexei Borodin, Ivan Corwin, and Daniel Remenik. Multiplicative functionals on ensembles of non-intersecting paths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1): 28–58, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1421244398.

Beznea:2019:NAE

[BCR19] Lucian Beznea, Iulian Cîmpean, and Michael Röckner. A new approach to the existence of invariant measures for Markovian semigroups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):977–1000, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1557820838.

Beznea:2020:NEM

[BCR20] Lucian Beznea, Iulian Cîmpean, and Michael Röckner. A natural extension of Markov processes and applications to singular SDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2480–2506, November 2020.

CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/A-natural-extension-
of-Markov-processes-and-applications-to-singular/10.
1214/20-AIHP1047.full.

Bossy:2011:SRD

[BCT11] Mireille Bossy, Mamadou Cissé, and Denis Talay. Stochastic representations of derivatives of solutions of one-dimensional parabolic variational inequalities with Neumann boundary conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2):395–424, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1300887275;   http://
www.numdam.org/item?id=AIHPB_2011__47_2_395_0.

Benaim:2021:SAQ

[BCV21] Michel Benaïm, Nicolas Champagnat, and Denis Villemonais. Stochastic approximation of quasi-stationary distributions for diffusion processes in a bounded domain. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):726–739, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-2/Stochastic-approximation-
of-quasi-stationary-distributions-for-diffusion-processes-
in/10.1214/20-AIHP1093.full.

Bercu:1992:MCP

[BD92] B. Bercu and M. Duflo. Moindres carrés pondérés et poursuite. (French) [Weighted least squares and more]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(3): 403–430, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1992__28_3_403_0.

Brandiere:1996:ASC

[BD96a] Odile Brandière and Marie Duflo. Les algorithmes stochastiques contournent-ils les pièges?. (French) [Do stochastic algorithms bypass traps?]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(3):395–427, ???? 1996. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1996_
_32_3_395_0.

Bryc:1996:LDS

[BD96b] Włodzimierz Bryc and Amir Dembo. Large deviations and strong mixing. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(4):549–569, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1996__32_4_
549_0.

Bertoin:1997:SCR

[BD97] J. Bertoin and R. A. Doney. Spitzer’s condition for random walks and Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(2):167–178, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_2_167_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801203.

Borodin:2011:LDP

[BD11] Alexei Borodin and Maurice Duits. Limits of determinantal processes near a tacnode. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):243–258, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170237;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_243_0.

Bi:2016:TLG

[BD16] Hongwei Bi and Jean-François Delmas. Total length of the genealogical tree for quadratic stationary continuous-state branching processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1321–1350, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723522.

Bretagnolle:1967:CMA

[BDC67] Jean Bretagnolle and Didier Dacunha-Castelle. Sur une classe de marches aléatoires. (French) [On a class of random walks]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B.

Calcul des Probabilités et Statistique, 3(4):403–431, October– December 1967. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1967__3_4_403_0.

Bretagnolle:1968:TLD

[BDC68] Jean Bretagnolle and Didier Dacunha-Castelle. Théorèmes limites à distance finie pour les marches aléatoires. (French) [Finite-distance limit theorems for random walks]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(1):25–73, January–March 1968. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1968__4_1_25_0.

Bretagnolle:1969:DFL

[BDC69] Jean Bretagnolle and Didier Dacunha-Castelle. Le déterminisme des fonctions laplaciennes sur certains espaces de suites. (French) [Determinism of Laplace functions on some sequence spaces]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(1):1–12, January–March 1969. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1969__5_1_
1_0.

Benoist:2016:CIC

[BDCH16] Stéphane Benoist, Hugo Duminil-Copin, and Clément Hongler. Conformal invariance of crossing probabilities for the Ising model with free boundary conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1784–1798, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1479373248.

Benjamini:2020:IDL

[BDCKL20] Itai Benjamini, Hugo Duminil-Copin, Gady Kozma, and Cyrille Lucas. Internal diffusion-limited aggregation with uniform starting points. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):391–404, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Internal-diffusion-limited-
aggregation-with-uniform-starting-points/10.1214/19-
AIHP965.full.

Barnsley:1988:IMM

[BDEG88] M. F. Barnsley, S. G. Demko, J. H. Elton, and J. S. Geronimo. Invariant measures for Markov processes arising from iterated function systems with place-dependent probabilities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(3):367–394, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1989__25_4_589_0. See erratum

[BDEG89].

Barnsley:1989:EIM

[BDEG89] M. F. Barnsley, S. G. Demko, J. H. Elton, and J. S. Geronimo. Erratum: “Invariant measures for Markov processes arising from iterated function systems with place-dependent probabilities”. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):589–590, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//www.numdam.org/item?id=AIHPB_1988__24_3_367_0. See

[BDEG88].

Buraczewski:2019:LFC

[BDK19] Dariusz Buraczewski, Piotr Dyszewski, and Konrad Kolesko. Local fluctuations of critical Mandelbrot cascades. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1179–1202, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820848.

Budhiraja:2011:VRC

[BDM11] Amarjit Budhiraja, Paul Dupuis, and Vasileios Maroulas. Variational representations for continuous time processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3): 725–747, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834857;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_725_0.

Banos:2018:CMD

[BDMBP18] David R. Baños, Sindre Duedahl, Thilo Meyer-Brandis, and Frank Proske. Construction of Malliavin differentiable strong solutions of SDEs under an integrability condition on the drift without the Yamada–Watanabe principle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1464–1491, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1531296026.

Bishop:2020:PAS

[BDN20] Adrian N. Bishop, Pierre Del Moral, and Angèle Niclas. A perturbation analysis of stochastic matrix Riccati diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):884–916, May 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/A-perturbation-analysis-of-stochastic-
matrix-Riccati-diffusions/10.1214/19-AIHP987.full.

Brunel:1974:PEM

[BDNT74] A. Brunel, N. Dang-Ngoc, and J.-P. Thouvenot. Un problème d’existence de mesure invariante. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(2):211–227, April–June 1974. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1974__10_2_211_0.

Brassesco:1995:BFI

[BDP95] S. Brassesco, A. De Masi, and E. Presutti. Brownian fluctuations of the interface in the D = 1 Ginzburg–Landau equation with noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):81–118, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1995__31_1_81_0.

Bacro:1987:ECR

[BDS87] Jean-Noël Bacro, Paul Deheuvels, and Josef Steinebach. Exact convergence rates in Erdős–Rényi type theorems for renewal processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(2):195–207, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_2_195_0.

Borga:2021:ASP

[BDS21] Jacopo Borga, Enrica Duchi, and Erik Slivken. Almost square permutations are typically square. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):1834–1856, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/Almost-square-permutations-
are-typically-square/10.1214/20-AIHP1135.full.

Barbu:2011:KEA

[BDT11] Viorel Barbu, Giuseppe Da Prato, and Luciano Tubaro. Kolmogorov equation associated to the stochastic reflection problem on a smooth convex set of a Hilbert space II. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3): 699–724, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834856;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_699_0.

Bramson:2016:CLM

[BDZ16] Maury Bramson, Jian Ding, and Ofer Zeitouni. Convergence in law of the maximum of nonlattice branching random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1897–1924, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1479373253.

Bellow:1982:GFI

[BE82] A. Bellow and L. Egghe. Generalized Fatou inequalities. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(4):335–365, October–December 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1982__18_4_335_0.

Bougerol:1995:EPH

[BE95] Philippe Bougerol and Laure Elie. Existence of positive harmonic functions on groups and on covering manifolds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1): 59–80, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1995__31_1_59_0.

Becker:1974:DTF

[Bec74] R. Becker. Sur deux théorèmes fondamentaux du calcul des probabilités. (French) []. Annales de l’Institut Henri Poincaré.

Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(4):385–389, October–December 1974. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1974__10_4_385_0.

Becker:1989:SCF

[Bec89] R. Becker. Sur la séparabilité et la continuité des fonctions aléatoires. (French) [On separability and continuity of random functions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2):167–174, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1989__25_2_167_0.

Beffara:2003:CIS

[Bef03] Vincent Beffara. On conformally invariant subsets of the planar Brownian curve. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(5):793–821, September/October 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_5_793_0;   http://www.sciencedirect.com/science/
article/pii/S024602030300030X.

Bertin:2019:ADE

[BEK19] Karine Bertin, Salima El Kolei, and Nicolas Klutchnikoff. Adaptive density estimation on bounded domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):1916–1947, November 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-55/issue-4/Adaptive-density-estimation-on-bounded-
domains/10.1214/18-AIHP938.full.

Belinschi:2006:NRF

[Bel06] Serban Teodor Belinschi. A note on regularity for free convolutions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5):635–648, September/October 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_5_635_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000993.

Belton:2008:PSS

[Bel08] Alexander C. R. Belton. On the path structure of a semimartingale arising from monotone probability theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2): 258–279, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1207948219;   http://www.numdam.org/
item?id=AIHPB_2008__44_2_258_0.

Brossard:2009:MBM

[BÉL09] Jean Brossard, Michel Émery, and Christophe Leuridan. Maximal Brownian motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):876–886, August 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1249391390;   http://www.numdam.org/item?id=AIHPB_2009_
_45_3_876_0.

Benois:2003:EFL

[BEM03] O. Benois, R. Esposito, and R. Marra. Equilibrium fluctuations for lattice gases. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(5):743–777, September/October 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_5_743_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000165.

Benzecri:1967:LMA

[Ben67a] F. Benzécri. Linguistique mathématique: l’algèbre des constituants non-connexes. (French) [Mathematical linguistics: the algebra of non-related components]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(1):1–34, January–March 1967. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1967__3_1_1_0.

Benzecri:1967:CSS

[Ben67b] J. P. Benzécri. La catégorie des simplexes segmentés et ses applications au traitement des constituants non connexes en linguistique mathématique. (French) [The category of segmented simplices and its applications to the treatment of unconnected constituents in mathematical language ]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(3):227–275, July–September 1967. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1967__3_3_227_0.

Benzecri:1967:MGM

[Ben67c] J. P. Benzécri. Un modèle géométrique de milieu continu déformable: la théorie des deux métriques. (French) [A geometric model of a continuously-deformable medium: the theory of the two metrics]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(3):277– 321, July–September 1967. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1967__
3_3_277_0.

Benassi:1979:PGM

[Ben79] Albert Benassi. Processus gaussiens, markoviens d’ordre p, fortement markoviens d’ordre p et problème de Dirichlet stochastique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1):107–126, January–March 1979. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1979__15_1_107_0.

Benigni:2020:EDQ

[Ben20] L. Benigni. Eigenvectors distribution and quantum unique ergodicity for deformed Wigner matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2822–2867, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/Eigenvectors-distribution-
and-quantum-unique-ergodicity-for-deformed-Wigner-matrices/
10.1214/20-AIHP1060.full.

Bert:1968:CPL

[Ber68] M. C. Bert. Convergence d’un processus à liaisons complètes, et application à un processus d’apprentissage linéaire. (French) [Convergence of a completely-connected process and application to a linear learning process]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(1):1–24, January–March 1968. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1968__4_1_1_0.

Bernard:1970:QPT

[Ber70] P. Bernard. Quelques propriétés des trajectoires des fonctions aléatoires stables sur R^{k}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(2):131–151, April–June 1970. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1970__6_2_131_0.

Berman:1983:LNL

[Ber83] Simeon M. Berman. Local nondeterminism and local times of general stochastic processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(2):189–207, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1983__19_2_189_0.

Berman:1985:ABT

[Ber85a] Simeon M. Berman. An asymptotic bound for the tail of the distribution of the maximum of a Gaussian process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(1): 47–57, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1985__21_1_47_0.

Berman:1985:MGP

[Ber85b] Simeon M. Berman. The maximum of a Gaussian process with nonconstant variance. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(4):383–391, ???? 1985. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1985_
_21_4_383_0.

Bertoin:1988:ADC

[Ber88] Jean Bertoin. Une application du calcul du nombre de montées et de descentes aux fonctions de martingales locales continues. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):201–207, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1988__24_2_201_0.

Bertoin:1989:ATS

[Ber89] Jean Bertoin. Applications de la théorie spectrale des cordes vibrantes aux fonctionnelles additives principales d’un brownien réfléchi. (French) [Applications of spectral theory of vibrating strings to the main additive functionals of a reflected Brownian]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(3):307–323, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1989__25_3_307_0.

Bercu:1991:EMC

[Ber91a] B. Bercu. Sur l’estimateur des moindres carrés généralisé d’un modèle ARMAX. Application à l’identification des modèles ARMA. (French) [The generalized least squares estimator of an ARMAX model. Application to ARMA model identification]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(3):425–443, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1991__27_3_425_0.

Bertoin:1991:DTP

[Ber91b] Jean Bertoin. Sur la décomposition de la trajectoire d’un processus de Lévy spectralement positif en son infimum. (French) [On the decomposition of the trajectory of a spectrallypositive Lévy process in its infimum]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(4):537–547, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_4_537_0.

Bertoin:1995:LPC

[Ber95] Jean Bertoin. Lévy processes that can creep downwards never increase. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(2):379–391, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1995__31_2_379_0.

Bertoin:2002:SSF

[Ber02] Jean Bertoin. Self-similar fragmentations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(3):319–340, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_3_319_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300010736.

Bernardin:2012:HRL

[Ber12a] Cédric Bernardin. Homogenization results for a linear dynamics in random Glauber type environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):792–818, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1340714873.

Bertoin:2012:FT

[Ber12b] Jean Bertoin. Fires on trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):909–921, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1353098433.

Berlinkov:2013:RFI

[Ber13] Artemi Berlinkov. On random fractals with infinite branching: Definition, measurability, dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1080–1089, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1380718738.

Berestycki:2015:DPL

[Ber15a] Nathanaël Berestycki. Diffusion in planar Liouville quantum gravity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):947–964, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1435759236.

Bertoin:2015:CTL

[Ber15b] Jean Bertoin. The cut-tree of large recursive trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51 (2):478–488, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1428672677.

Bertoin:2020:NRL

[Ber20] Jean Bertoin. Noise reinforcement for Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2236–2252, August 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/Noise-reinforcement-for-L%c3%a9vy-
processes/10.1214/19-AIHP1037.full.

Blondel:2020:HLF

[BESS20] Oriane Blondel, Clément Erignoux, Makiko Sasada, and Marielle Simon. Hydrodynamic limit for a facilitated exclusion process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):667–714, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Hydrodynamic-limit-for-
a-facilitated-exclusion-process/10.1214/19-AIHP977.
full.

Baryshnikov:2008:MDS

[BESY08] Yu Baryshnikov, P. Eichelsbacher, T. Schreiber, and J. E. Yukich. Moderate deviations for some point measures in geometric probability. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3):422–446, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1211819419;   http://
www.numdam.org/item?id=AIHPB_2008__44_3_422_0.

Bettinelli:2015:SLR

[Bet15] Jérémie Bettinelli. Scaling limit of random planar quadrangulations with a boundary. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):432–477, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1428672676.

Bettinelli:2016:GBS

[Bet16] Jérémie Bettinelli. Geodesics in Brownian surfaces (Brownian maps). Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):612–646, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1462367887.

Berestycki:2013:LSB

[BEV13] N. Berestycki, A. M. Etheridge, and A. Véber. Large scale behaviour of the spatial Λ–Fleming–Viot process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2): 374–401, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1366117651.

Bewley:1971:EBN

[Bew71] Truman Bewley. Extension of the Birkhoff and von Neumann ergodic theorems to semigroup actions. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(4):283–291, October–December 1971. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1971__7_4_283_0.

Benassi:1988:HLA

[BF88] Albert Benassi and Jean-Pierre Fouque. Hydrodynamical limit for the asymmetric zero-range process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):189–200, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1988__24_2_189_0.

Bezandry:1992:PLC

[BF92] Paul Hubert Bézandry and Xavier Fernique. Sur la propriété de la limite centrale dans D[0, 1]. (French) [On the central limit property in D[0, 1]]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(1):31–46, ???? 1992. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1992__28_1_31_0.

Borodin:1995:FOR

[BF95] A. N. Borodin and M. I. Freidlin. Fast oscillating random perturbations of dynamical systems with conservation laws. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(3):485–525, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1995__31_3_485_0.

Bertini:2015:LDE

[BFG15] Lorenzo Bertini, Alessandra Faggionato, and Davide Gabrielli. Large deviations of the empirical flow for continuous time Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):867–900, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1435759232.

Berthet:2020:CLT

[BFK20] Philippe Berthet, Jean-Claude Fort, and Thierry Klein. A Central Limit Theorem for Wasserstein type distances between two distinct univariate distributions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):954–982, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/A-Central-Limit-Theorem-
for-Wasserstein-type-distances-between-two/10.1214/
19-AIHP990.full.

Betz:1985:PAC

[BG85] Cristina Betz and Henryk Gzyl. Probabilistic approach for comparing first eigenvalues. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(2):147–156, ???? 1985. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1985_
_21_2_147_0.

Bouhaik:1992:TLC

[BG92] Mostafa Bouhaik and Léonard Gallardo. Un théorème limite central dans un hypergroupe bidimensionnel. (French) [A central limit theorem in a two-dimensional hypergroup]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(1): 47–61, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1992__28_1_47_0.

Bezuidenhout:1999:CLT

[BG99a] Carol Bezuidenhout and Geoffrey Grimmett. A central limit theorem for random walks in random labyrinths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35 (5):631–683, September 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_5_631_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399001107.

Bodineau:1999:ASS

[BG99b] Thierry Bodineau and Alice Guionnet. About the stationary states of vortex systems. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 35(2):205–237, March/April 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_2_205_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399800119.

Benaych-Georges:2010:SRB

[BG10] Florent Benaych-Georges. On a surprising relation between the Marchenko–Pastur law, rectangular and square free convolutions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3):644–652, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1281100393;   http://
www.numdam.org/item?id=AIHPB_2010__46_3_644_0.

Butucea:2016:SDS

[BG16] Cristina Butucea and Ghislaine Gayraud. Sharp detection of smooth signals in a high-dimensional sparse matrix with indirect observations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1564–1591, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1479373240.

Belomestny:2020:NDE

[BG20] Denis Belomestny and Alexander Goldenshluger. Nonparametric density estimation from observations with multiplicative measurement errors. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):36–67, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Nonparametric-density-
estimation-from-observations-with-multiplicative-measurement-
errors/10.1214/18-AIHP954.full.

Benaych-Georges:2020:SRS

[BGBK20] Florent Benaych-Georges, Charles Bordenave, and Antti Knowles. Spectral radii of sparse random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2141–2161, August 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-

linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/Spectral-radii-of-sparse-random-matrices/
10.1214/19-AIHP1033.full.

Baraud:2014:ESG

[BGH14] Yannick Baraud, Christophe Giraud, and Sylvie Huet. Estimator selection in the Gaussian setting. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):1092–1119, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1403277009.

Basu:2019:ARW

[BGHR19] Riddhipratim Basu, Shirshendu Ganguly, Christopher Hoffman, and Jacob Richey. Activated random walk on a cycle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1258–1277, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1569398869.

Bernardin:2018:IPB

[BGJS18] Cédric Bernardin, Patrícia Gonçalves, Milton Jara, and Marielle Simon. Interpolation process between standard diffusion and fractional diffusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1731–1757, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1531296034.

Bernardin:2022:EFD

[BGJS22] C. Bernardin, P. Gonçalves, M. Jara, and S. Scotta. Equilibrium fluctuations for diffusive symmetric exclusion with long jumps and infinitely extended reservoirs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):303–342, February 2022. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-58/issue-1/Equilibrium-fluctuations-for-diffusive-
symmetric-exclusion-with-long-jumps-and/10.1214/21-
AIHP1176.full.

Bigot:2017:GPW

[BGKL17] Jérémie Bigot, Raúl Gouet, Thierry Klein, and Alfredo López. Geodesic PCA in the Wasserstein space by convex PCA. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):1–26, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1486544882.

Baudoin:2020:GBK

[BGM20] Fabrice Baudoin, Maria Gordina, and Phanuel Mariano. Gradient bounds for Kolmogorov type diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):612–636, February 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Gradient-bounds-for-Kolmogorov-type-
diffusions/10.1214/19-AIHP975.full.

Berger:2019:SBR

[BGN19] Noam Berger, Nina Gantert, and Jan Nagel. The speed of biased random walk among random conductances. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):862–881, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820834.

Benaych-Georges:2014:LDH

[BGP14] Florent Benaych-Georges and Sandrine Péché. Localization and delocalization for heavy tailed band matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (4):1385–1403, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555505.

Bahadoran:2014:EHA

[BGRS14] C. Bahadoran, H. Guiol, K. Ravishankar, and E. Saada. Euler hydrodynamics for attractive particle systems in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):403–424, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1395856133.

Bose:2010:LSD

[BGS10] Arup Bose, Sreela Gangopadhyay, and Arnab Sen. Limiting spectral distribution of XX’ matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3): 677–707, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1281100395;   http://www.numdam.org/
item?id=AIHPB_2010__46_3_677_0.

Bru:1980:EVAb

[BH80a] B. Bru and H. Heinich. Sur l’espérance des variables aléatoires à valeurs dans les espaces de Banach réticulés. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(3):197–210, July–September 1980. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1980__16_3_
197_0.

Bru:1980:EVAa

[BH80b] B. Bru and H. Heinich. Sur l’espérance des variables aléatoires vectorielles. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(3):177–196, July–September 1980. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1980__16_3_177_0.

Bru:1981:IPP

[BH81] B. Bru and H. Heinich. Isométries positives et propriétés ergodiques de quelques espaces de Banach. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(4):377–405, October–December 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_4_377_0.

Bru:1985:MAM

[BH85] Bernard Bru and Henri Heinich. Meilleures approximations et médianes conditionnelles. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(3):197–224, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1985__21_3_197_0.

Barlow:1997:TDE

[BH97] M. T. Barlow and B. M. Hambly. Transition density estimates for Brownian motion on scale irregular Sierpiński gaskets. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(5):531–557, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_5_531_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801045.

Baladi:2008:LLT

[BH08] Viviane Baladi and Aïcha Hachemi. A local limit theorem with speed of convergence for Euclidean algorithms and Diophantine costs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4):749–770, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1217964118;   http://
www.numdam.org/item?id=AIHPB_2008__44_4_749_0.

Benjamini:2020:RMC

[BH20]  Itai Benjamini and Jonathan Hermon.  Recurrence of

Markov chain traces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):734–759, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Recurrence-of-Markov-
chain-traces/10.1214/19-AIHP979.full.

Brunel:1993:SMP

[BHL93] Antoine Brunel, Shlomo Horowitz, and Michael Lin. On subinvariant measures for positive operators in L_{1}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(1): 105–117, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_1_105_0.

Ben-Hamou:2019:CMT

[BHLP19] Anna Ben-Hamou, Eyal Lubetzky, and Yuval Peres. Comparing mixing times on sparse random graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1116–1130, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1557820844.

Bhattacharya:2017:PPC

[BHR17] Ayan Bhattacharya, Rajat Subhra Hazra, and Parthanil Roy. Point process convergence for branching random walks with regularly varying steps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):802–818, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1491897746.

Bose:2012:PES

[BHS12] Arup Bose, Rajat Subhra Hazra, and Koushik Saha. Product of exponentials and spectral radius of random k-circulants. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):424–443, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1334148206.

Bauerschmidt:2021:GRW

[BHS21] Roland Bauerschmidt, Tyler Helmuth, and Andrew Swan. The geometry of random walk isomorphism theorems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):408–454, February 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-1/The-geometry-of-random-walk-isomorphism-
theorems/10.1214/20-AIHP1083.full.

Biane:1986:REP

[Bia86] Ph. Biane. Relations entre pont et excursion du mouvement brownien réel. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(1):1–7, ???? 1986. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1986__22_1_1_0.

Birge:1984:SID

[Bir84] Lucien Birge. Stabilité et instabilité du risque minimax pour des variables indépendantes équidistribuées. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(3):201–223, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1984__20_3_201_0.

Birge:2006:MST

[Bir06] Lucien Birgé. Model selection via testing: an alternative to (penalized) maximum likelihood estimators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(3): 273–325, May/June 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_3_273_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000841.

Bonnefont:2020:C

[BJ20] Michel Bonnefont and Nicolas Juillet. Couplings in L^{p} distance of two Brownian motions and their Lévy area. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):543–565, February 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Couplings-in-Lp-distance-of-two-Brownian-
motions-and-their/10.1214/19-AIHP972.full.

Bruneau:2010:IPR

[BJM10] Laurent Bruneau, Alain Joye, and Marco Merkli. Infinite products of random matrices and repeated interaction dynamics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):442–464, May 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1273584130;   http://
www.numdam.org/item?id=AIHPB_2010__46_2_442_0.

Bass:1993:ILT

[BK93] Richard F. Bass and Davar Khoshnevisan. Intersection local times and Tanaka formulas. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(3):419–451, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1993__29_3_419_0.

Bretagnolle:1995:ESV

[BK95] Jean Bretagnolle and Andrzej Klopotowski. Sur l’existence des suites de variables aléatoires s à s indépendantes échangeables ou stationnaires. (French) [On the existence of sequences of random variables s to independently exchangeable or stationary s ]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(2):325–350, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1995__31_2_325_0.

Bovier:2004:DGRa

[BK04a] Anton Bovier and Irina Kurkova. Derrida’s generalised random energy models 1: models with finitely many hierarchies. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(4):439–480, July/August 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_4_439_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000196.

Bovier:2004:DGRb

[BK04b] Anton Bovier and Irina Kurkova. Derrida’s generalized random energy models 2: models with continuous hierarchies. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(4):481–495, July/August 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_4_481_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000202.

Bufetov:2018:ARD

[BK18] Alexey Bufetov and Alisa Knizel. Asymptotics of random domino tilings of rectangular Aztec diamonds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1250–1290, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1531296019.

Bertin:2021:ARB

[BK21] Karine Bertin and Nicolas Klutchnikoff. Adaptive regression with Brownian path covariate. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1495–1520, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-

statistiques/volume-57/issue-3/Adaptive-regression-
with-Brownian-path-covariate/10.1214/20-AIHP1128.full.

Bass:2010:SJP

[BKK10] Richard F. Bass, Moritz Kassmann, and Takashi Kumagai. Symmetric jump processes: localization, heat kernels and convergence. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):59–71, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1267454108;   http://
www.numdam.org/item?id=AIHPB_2010__46_1_59_0.

Bonami:1971:PDA

[BKRR71] A. Bonami, N. Karoui, B. Roynette, and H. Reinhard. Processus de diffusion associé à un opérateur elliptique dégénéré. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(1): 31–80, January–March 1971. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_1_31_0.

Balazs:2012:MCF

[BKS12a] Márton Balázs, Júlia Komjáthy, and Timo Seppäläinen. Microscopic concavity and fluctuation bounds in a class of deposition processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):151–187, February 2012. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1327328018.

Burdzy:2012:SDJ

[BKS12b] K. Burdzy, T. Kulczycki, and R. L. Schilling. Stationary distributions for jump processes with memory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):609–630, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1340714865.

Benjamini:2010:ERW

[BKYY10] Itai Benjamini, Gady Kozma, Ariel Yadin, and Amir Yehudayoff. Entropy of random walk range. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1080–1092, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878339;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_1080_0.

Brodeau:1979:IPP

[BL79] F. Brodeau and A. Le Breton. Identification de paramètres pour un système excité par des bruits gaussien et poissonien. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1): 1–23, January–March 1979. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1979__
15_1_1_0.

Bleher:1995:VNL

[BL95] Pavel M. Bleher and Joel L. Lebowitz. Variance of number of lattice points in random narrow elliptic strip. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):27– 58, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1995__31_1_27_0.

Berthet:2002:SER

[BL02] Philippe Berthet and Mikhail Lifshits. Some exact rates in the functional law of the iterated logarithm. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):811–824, November/December 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_6_811_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011317.

Betz:2003:UGM

[BL03] Volker Betz and József Lőrinczi. Uniqueness of Gibbs measures relative to Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(5):877–889, September/October 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_5_877_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000219.

Bertoin:2005:SFA

[BL05] Jean Bertoin and Jean-François Le Gall. Stochastic flows associated to coalescent processes II: stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):307–333, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_307_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000130.

Beltran:2008:LGM

[BL08a] J. Beltrán and C. Landim. A lattice gas model for the incompressible Navier–Stokes equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5): 886–914, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1222261917;   http://www.numdam.org/
item?id=AIHPB_2008__44_5_886_0.

Berzin:2008:EMD

[BL08b] Corinne Berzin and José R. León. Estimation in models driven by fractional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2): 191–213, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1207948216;   http://www.numdam.org/
item?id=AIHPB_2008__44_2_191_0.

Brossard:2012:DOT

[BL12] Jean Brossard and Christophe Leuridan. Densité des orbites des trajectoires browniennes sous l’action de la transformation de Lévy. (French) [Density of orbits of Brownian trajectories under the action of the Lévy transformation]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48 (2):477–517, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1334148208.

Boissard:2014:MSC

[BL14] Emmanuel Boissard and Thibaut Le Gouic. On the mean speed of convergence of empirical and occupation measures in Wasserstein distance. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):539–563, May 2014. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1395856140.

Beltran:2015:TKD

[BL15] J. Beltrán and C. Landim. Tunneling of the Kawasaki dynamics at low temperatures in two dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1): 59–88, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1421244399.

Biskup:2019:ILS

[BL19] Marek Biskup and Oren Louidor. On intermediate level sets of two-dimensional discrete Gaussian free field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):1948–1987, November 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-55/issue-4/On-intermediate-level-sets-of-two-
dimensional-discrete-Gaussian-free/10.1214/18-AIHP939.
full.

Baik:2020:FEB

[BL20] Jinho Baik and Ji Oon Lee. Free energy of bipartite spherical Sherrington–Kirkpatrick model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2897–2934, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/Free-energy-of-bipartite-
spherical-SherringtonKirkpatrick-model/10.1214/20-AIHP1062.
full.

Blanchard:1973:PPM

[Bla73] F. Blanchard. Processus de points marqués et processus ramifiés. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(3):259– 275, July–September 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_3_259_0.

Blanchard:1975:CIF

[Bla75] François Blanchard. Conditions d’isomorphisme des flots spéciaux. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(2):173–185, April–June 1975. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1975__11_2_173_0.

Blanchard:1999:PME

[Bla99] Gilles Blanchard. The “progressive mixture” estimator for regression trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(6):793–820, November 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_6_793_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001156.

Blachere:2002:ALP

[Bla02] Sébastien Blachère. Agrégation limitée par diffusion interne sur Z^{d}. (French) [Internal-diffusion-limited aggregation on Z^{d}]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(4):613–648, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_4_613_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301011037.

Blachere:2003:CTR

[Bla03] Sébastien Blachère. Cut times for random walks on the discrete Heisenberg group. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(4):621–638, July/August 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_4_621_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000177.

Blitvic:2014:TPN

[Bli14] Natasha Blitvić. Two-parameter non-commutative Central Limit Theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1456–1473, November 2014. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1413555507.

Bitouze:1999:DKW

[BLM99] D. Bitouzé, B. Laurent, and P. Massart. A Dvoretzky– Kiefer–Wolfowitz type inequality for the Kaplan–Meier estimator. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(6):735–763, November 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_6_735_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001120.

Beliaev:2021:CST

[BLM21] Dmitry Beliaev, Terry J. Lyons, and Vlad Margarint. Continuity in κ in SLE_{κ} theory using a constructive method and Rough Path Theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):455–468, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/Continuity-in-%ce%ba-
in-SLE%ce%ba-theory-using-a-constructive-method/10.
1214/20-AIHP1084.full.

Benaim:2015:QPC

[BLMZ15] Michel Benaïm, Stéphané Le Borgne, Florent Malrieu, and Pierre-André Zitt. Qualitative properties of certain piecewise deterministic Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):1040–1075, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1435759239.

Basdevant:2011:SMD

[BLNR11] Anne-Laure Basdevant, Philippe Laurençot, James R. Norris, and Clément Rau. A stochastic min-driven coalescence process and its hydrodynamical limit. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2): 329–357, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887272;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_329_0.

Bertin:2011:ADD

[BLR11] K. Bertin, E. Le Pennec, and V. Rivoirard. Adaptive Dantzig density estimation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):43–74, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170229;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_43_0.

Bertin:2016:APE

[BLR16] Karine Bertin, Claire Lacour, and Vincent Rivoirard. Adaptive pointwise estimation of conditional density function. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):939–980, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1462367901.

Baik:2022:LOP

[BLS22] Jinho Baik, Zhipeng Liu, and Guilherme L. F. Silva. Limiting one-point distribution of periodic TASEP. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):248–302, February 2022. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-58/issue-1/Limiting-one-point-distribution-of-
periodic-TASEP/10.1214/21-AIHP1171.full.

Blanc-Lapierre:1969:CCC

[BLT69] André Blanc-Lapierre and Albert Tortrat. Compensation de corrélations dans le cas de fonctions aléatoires ou de suites de variables aléatoires de second ordre. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(2):101–112, April–June 1969. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1969__5_2_101_0.

Bretagnolle:1989:AMA

[BM89] Jean Bretagnolle and Jan Mielniczuk. On asymptotic minimaxity of the adaptative kernel estimate of a density function. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2):143–152, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1989__25_2_143_0.

Burdzy:1993:NPP

[BM93] Krzysztof Burdzy and Donald E. Marshall. Non-polar points for reflected Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(2):199–228, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1993__29_2_199_0.

Barthe:2000:SRI

[BM00] F. Barthe and B. Maurey. Some remarks on isoperimetry of Gaussian type. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(4):419–434, July 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_4_419_0;   http://www.sciencedirect.com/science/
article/pii/S024602030000131X.

Bakhtin:2008:CHM

[BM08] Yuri Bakhtin and Matilde Martínez. A characterization of harmonic measures on laminations by hyperbolic Riemann surfaces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1078–1089, December 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1227287565;   http://www.numdam.org/item?id=AIHPB_2008_
_44_6_1078_0.

Barral:2009:FMP

[BM09] Julien Barral and Benoît Mandelbrot. Fractional multiplicative processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1116–1129, November 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1257529895;   http://www.numdam.org/item?id=AIHPB_2009_
_45_4_1116_0.

Buraczewski:2016:PLD

[BM16] Dariusz Buraczewski and Sebastian Mentemeier. Precise large deviation results for products of random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3): 1474–1513, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1469723527.

Boldrighini:1994:IRWa

[BMP94a] Carlo Boldrighini, Robert A. Minlos, and Alessandro Pellegrinotti. Interacting random walk in a dynamical random environment. I. Decay of correlations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(4):519–558, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1994__30_4_559_0.

Boldrighini:1994:IRWb

[BMP94b] Carlo Boldrighini, Robert A. Minlos, and Alessandro Pellegrinotti. Interacting random walk in a dynamical random environment. II. Environment from the point of view of the particle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(4):559–605, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1994__30_4_519_0.

Butucea:2009:AGF

[BMP09] Cristina Butucea, Catherine Matias, and Christophe Pouet. Adaptive goodness-of-fit testing from indirect observations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2):352–372, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024673;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_352_0.

Bahadoran:2017:SBA

[BMRS17] C. Bahadoran, T. Mountford, K. Ravishankar, and E. Saada. Supercritical behavior of asymmetric zero-range process with sitewise disorder. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):766–801, May 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1491897745.

Beliaev:2021:RSW

[BMW21] D. Beliaev, S. Muirhead, and I. Wigman. Russo–Seymour– Welsh estimates for the Kostlan ensemble of random polynomials. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2189–2218, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/RussoSeymourWelsh-estimates-
for-the-Kostlan-ensemble-of-random-polynomials/10.1214/
20-AIHP1142.full.

Bernard:1990:RNW

[BN90] Pierre Bernard and David Nualart. Régularité C^{∞} des noyaux de Wiener d’une diffusion. (French) [Regularity of C^{∞} Wiener kernels of a diffusion]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(2):287–297, ???? 1990. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1990_
_26_2_287_0.

Bercovici:2018:EDD

[BNNS18] Hari Bercovici, Alexandru Nica, Michael Noyes, and Kamil Szpojankowski. Eta-diagonal distributions and infinite divisibility for R-diagonals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):907–937, May 2018. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1524643234.

Baudoin:2011:EFB

[BO11] Fabrice Baudoin and Neil O’Connell. Exponential functionals of Brownian motion and class-one Whittaker functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1096–1120, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1317906503;   http://
www.numdam.org/item?id=AIHPB_2011__47_4_1096_0.

Bodiou:1970:TBM

[Bod70] Georges Bodiou. Sur les treillis booléens métriques et le conditionnement général des probabilités. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(1):15–26, January–March 1970. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1970__6_1_15_0.

Bodiou:1971:CAF

[Bod71] G. Bodiou. Complément à l’article: ^{≪} Sur les treillis booléens métriques et le conditionnement général des probabilités ^{≫}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(1): 21–22, January–March 1971. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_1_21_0.

Bodineau:1997:IOD

[Bod97] Thierry Bodineau. Interface for one-dimensional random Kac potentials. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(5):559–590, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_5_559_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801057.

Boivin:1998:ETS

[Boi98] Daniel Boivin. Ergodic theorems for surfaces with minimal random weights. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(5):567–599, September/October 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_5_567_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800010.

Borell:1977:NGM

[Bor77] Christer Borell. A note on Gauss measures which agree on small balls. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(3):231– 238, July–September 1977. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1977__
13_3_231_0.

Borodin:1987:DRW

[Bor87] A. N. Borodin. On the distribution of random walk local time. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(1):63–89, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1987__23_1_63_0.

Borell:1996:NPC

[Bor96] Christer Borell. A note on parabolic convexity and heat conduction. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(3):387–393, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1996__32_3_387_0.

Borodin:2008:LFM

[Bor08] Alexei Borodin. Loop-free Markov chains as determinantal point processes. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 44(1):19–28, February 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1203969866;   http://www.numdam.org/item?id=AIHPB_2008_
_44_1_19_0.

Borda:2021:ERW

[Bor21] Bence Borda. Equidistribution of random walks on compact groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):54–72, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-1/Equidistribution-of-random-walks-
on-compact-groups/10.1214/20-AIHP1070.full.

Bosq:1975:DQI

[Bos75] D. Bosq. Une définition qualitative de l’indépendance et ses applications. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(3):225–252, July–September 1975. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1975__11_3_225_0.

Baudoin:2014:UBD

[BOT14] Fabrice Baudoin, Cheng Ouyang, and Samy Tindel. Upper bounds for the density of solutions to stochastic differential equations driven by fractional Brownian motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1): 111–135, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1388545268.

Bouleau:1980:RSB

[Bou80] N. Bouleau. Une remarque sur le schéma de Bernoulli et quelques extensions. (French) [A remark on Bernoulli’s scheme and some extensions]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(1):1–16, January–March 1980. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1980__16_1_1_0.

Bougerol:1983:ETL

[Bou83] Philippe Bougerol. Exemples de théorèmes locaux sur les groupes résolubles. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(4):369–391, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1983__19_4_369_0.

Bouleau:1984:FCV

[Bou84] N. Bouleau. Formules de changement de variables. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(2):133–145, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1984__20_2_133_0.

Bougerol:1988:CEL

[Bou88a] Philippe Bougerol. Comparaison des exposants de Lyapounov des processus markoviens multiplicatifs. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(4): 439–489, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1988__24_4_439_0.

Bouton:1988:AGA

[Bou88b] Catherine Bouton. Approximation gaussienne d’algorithmes stochastiques à dynamique markovienne. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(1): 131–155, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1988__24_1_131_0.

Bousch:2000:PAF

[Bou00] Thierry Bousch. Le poisson n’a pas d’arêtes. (French) [The fish has no bones]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(4):489–508, July 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_4_489_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001321.

Bousch:2007:RFR

[Bou07] Thierry Bousch. Sur les retardateurs. (French) [On retardants]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1):31–46, January/February 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_1_31_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000161.

Baudoin:2016:SER

[BOZ16] Fabrice Baudoin, Cheng Ouyang, and Xuejing Zhang. Smoothing effect of rough differential equations driven by fractional Brownian motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):412–428, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1452089274.

Bafico:1985:CGW

[BP85] R. Bafico and G. Pistone. G-convergence of generators and weak convergence of diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(1):1–13, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1985_
_21_1_1_0.

Benjamini:1992:RWT

[BP92] Itai Benjamini and Yuval Peres. Random walks on a tree and capacity in the interval. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(4):557–592, ???? 1992. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1992_
_28_4_557_0.

Bhat:1995:MDN

[BP95] B. V. Rajarama Bhat and K. R. Parthasarathy. Markov dilations of nonconservative dynamical semigroups and a quantum boundary theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(4):601–651, ???? 1995. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1995_
_31_4_601_0.

Bourgeat:2004:AEC

[BP04] Alain Bourgeat and Andrey Piatnitski. Approximations of effective coefficients in stochastic homogenization. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40 (2):153–165, March/April 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_2_153_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000657.

Barczy:2006:FTG

[BP06] Mátyás Barczy and Gyula Pap. Fourier transform of a Gaussian measure on the Heisenberg group. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5): 607–633, September/October 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_5_607_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000981.

Ba:2015:BPI

[BP15] Mamadou Ba and Etienne Pardoux. Branching processes with interaction and a generalized Ray–Knight theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51 (4):1290–1313, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1445432043.

Barlow:2012:CRW

[BPS12] Martin T. Barlow, Yuval Peres, and Perla Sousi. Collisions of random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):922–946, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1353098434.

Bouchard:2018:UAP

[BPTZ18] Bruno Bouchard, Dylan Possamaï, Xiaolu Tan, and Chao Zhou. A unified approach to a priori estimates for supersolutions of BSDEs in general filtrations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):154–172, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1519030823.

BenArous:2003:IDR

[BQR03] Gérard Ben Arous, Jeremy Quastel, and Alejandro F. Ramírez. Internal DLA in a random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39 (2):301–324, March/April 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_2_301_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302000031.

Bordenave:2020:SGS

[BQZ20] Charles Bordenave, Yanqi Qiu, and Yiwei Zhang. Spectral gap of sparse bistochastic matrices with exchangeable rows. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2971–2995, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Spectral-gap-of-sparse-
bistochastic-matrices-with-exchangeable-rows/10.1214/
20-AIHP1065.full.

Brunel:1974:QAP

[BR74] A. Brunel and D. Revuz. Quelques applications probabilistes de la quasi-compacité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(3):301–337, July–September 1974. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1974__10_3_301_0.

Brunel:1979:MHG

[BR79] A. Brunel and D. Revuz. Marches de Harris sur les groupes localement compacts V. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(3):205–234, July–September 1979. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_3_205_0.

Benaim:2003:SID

[BR03] Michel Benaïm and Olivier Raimond. Self-interacting diffusions II: convergence in law. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):1043–1055, November/December 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_6_1043_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020303000281.

Benaim:2008:ECF

[BR08] Michel Benaïm and Raphaël Rossignol. Exponential concentration for first passage percolation through modified Poincaré inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3):544–573, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1211819424;   http://
www.numdam.org/item?id=AIHPB_2008__44_3_544_0.

Berger:2015:RWD

[BR15] Noam Berger and Ron Rosenthal. Random walks on discrete point processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):727–755, May 2015. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1428672689.

Brancovan:1973:QPR

[Bra73] M. Brancovan. Quelques propriétés des résolvantes récurrentes au sens de Harris. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(1):1–18, January–March 1973. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1973__9_1_1_0.

Bravo:2009:FAT

[Bra09] Gerónimo Uribe Bravo. The falling apart of the tagged fragment and the asymptotic disintegration of the Brownian height fragmentation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1130–1149, November 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1257529896;   http://www.numdam.org/item?id=AIHPB_2009_
_45_4_1130_0.

Bretagnolle:1983:LLD

[Bre83] J. Bretagnolle. Lois limites du bootstrap de certaines fonctionnelles. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(3):281–296, ???? 1983. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1983_
_19_3_281_0.

Bretagnolle:1987:NLO

[Bre87] Jean Bretagnolle. Notes de lecture sur l’œuvre de Paul Lévy. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2):239–243, ???? 1987. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_S2_239_0.

Bretagnolle:2002:BJB

[Bre02] Jean Bretagnolle. Biographie de Jean Bretagnolle. (French) [Biography of Jean Bretagnolle]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):803, November/ December 2002. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S024602030201141X. En l’honneur

de J. Bretagnolle, D. Dacunha-Castelle, I. Ibragimov.

Bremont:2004:RWR

[Bré04] Julien Brémont. Random walks in random medium on Z and Lyapunov spectrum. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):309–336, May/June 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_3_309_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000712.

Bremont:2009:ODF

[Bré09] Julien Brémont. One-dimensional finite range random walk in random medium and invariant measure equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1): 70–103, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469972;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_70_0.

Bronner:1975:MPP

[Bro75] François Bronner. Méthodes probabilistes pour la détermination de résolvantes sous-markoviennes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(3):253–264, July–September 1975. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1975__11_3_253_0.

Brodeau:1983:IOP

[Bro83] F. Brodeau. Identifications optimales de paramètres pour un système linéaire excité par un bruit gaussien. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(2):123–152, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1983__19_2_123_0.

Brofferio:2003:HCR

[Bro03] Sara Brofferio. How a centred random walk on the affine group goes to infinity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(3):371–384, May/June 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_3_371_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000158.

Broman:2020:EPT

[Bro20] Erik I. Broman. The existence phase transition for scale invariant Poisson random fractal models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):715–733, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/The-existence-phase-
transition-for-scale-invariant-Poisson-random-fractal/
10.1214/19-AIHP978.full.

Bickel:2002:HMM

[BRR02] Peter J. Bickel, Ya’acov Ritov, and Tobias Rydén. Hidden Markov model likelihoods and their derivatives behave like i.i.d. ones. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):825–846, November/December 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_6_825_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011342.

Barbu:2017:DPR

[BRR17] Viorel Barbu, Michael Röckner, and Francesco Russo. Doubly probabilistic representation for the stochastic porous media type equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):2043–2073, November 2017. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1511773738.

Bardina:2003:OMF

[BRT03] Xavier Bardina, Carles Rovira, and Samy Tindel. Onsager– Machlup functional for stochastic evolution equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(1):69–93, January/February 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_1_69_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302000092.

Blomker:2007:PRS

[BRT07] D. Blömker, M. Romito, and R. Tribe. A probabilistic representation for the solutions to some non-linear PDEs using pruned branching trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(2):175–192, March/April 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_2_175_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000422.

Balazs:2014:MFP

[BRT14] Márton Balázs, Miklós Z. Rácz, and Bálint Tóth. Modeling flocks and prices: Jumping particles with an attractive interaction. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):425–454, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1395856134.

Bruter:1971:SDB

[Bru71] C. Bruter. Le système dynamique de Beboutoff. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(3):235–251, July– September 1971. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1971__7_3_235_0.

Brunel:1973:TEP

[Bru73] A. Brunel. Théorème ergodique ponctuel pour un semi-groupe commutatif finiment engendré de contractions de L^{1}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(4):327–343, October–December 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_4_327_0.

Brunhes:1976:PSO

[Bru76] Denis Brunhes. Prédiction scalaire ou prédiction vectorielle pour des processus faiblement stationnaires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(4):291–298, October–December 1976. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1976__
12_4_291_0.

Brugiere:1993:TLC

[Bru93] Pierre Brugière. Théorème de limite centrale pour un estimateur non paramétrique de la variance d’un processus de diffusion multidimensionnelle. (French) [Central Limit Theorem for a non-parametric estimator for the variance of a multidimensional diffusion process]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(3):357–389, ???? 1993. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1993_
_29_3_357_0.

Belius:2019:BEC

[BRZ19] David Belius, Jay Rosen, and Ofer Zeitouni. Barrier estimates for a critical Galton–Watson process and the cover time of the binary tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):127–154, February 2019. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1547802397.

Biane:2001:FDF

[BS01] Philippe Biane and Roland Speicher. Free diffusions, free entropy and free Fisher information. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(5):581–606, September/October 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_5_581_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300010748.

Beghdadi-Sakrani:2003:SRP

[BS03] S. Beghdadi-Sakrani. Some remarkable pure martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(2):287–299, March/April 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_2_287_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302000183.

Birkner:2010:AVQ

[BS10a] Matthias Birkner and Rongfeng Sun. Annealed vs quenched critical points for a random walk pinning model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2): 414–441, May 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1273584129;   http://www.numdam.org/
item?id=AIHPB_2010__46_2_414_0.

Bjorklund:2010:ASA

[BS10b] Michael Björklund and Daniel Schnellmann. Almost sure absolute continuity of Bernoulli convolutions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3): 888–893, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1281100403;   http://www.numdam.org/
item?id=AIHPB_2010__46_3_888_0.

Birkner:2011:DRR

[BS11] Matthias Birkner and Rongfeng Sun. Disorder relevance for the random walk pinning model in dimension 3. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1): 259–293, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170238;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_259_0.

Benjamini:2013:PHC

[BS13] Itai Benjamini and Alexandre Stauffer. Perturbing the hexagonal circle packing: A percolation perspective. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49 (4):1141–1157, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1380718741.

Basu:2017:CPC

[BS17] Deepan Basu and Artem Sapozhnikov. Crossing probabilities for critical Bernoulli percolation on slabs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1921–1933, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773731.

Bendikov:2004:SPD

[BSC04] A. Bendikov and L. Saloff-Coste. On the sample paths of diagonal Brownian motions on the infinite dimensional torus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2):227–254, March/April 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_2_227_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000487.

Betz:2020:ISA

[BST20] Volker Betz, Helge Schäfer, and Lorenzo Taggi. Interacting self-avoiding polygons. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1321–1335, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Interacting-self-avoiding-
polygons/10.1214/19-AIHP1003.full.

Bolthausen:2003:CPD

[BSZ03] Erwin Bolthausen, Alain-Sol Sznitman, and Ofer Zeitouni. Cut points and diffusive random walks in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(3):527–555, May/June 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_3_527_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000195.

Blanchard:2000:SPC

[BT00] F. Blanchard and P. Tisseur. Some properties of cellular automata with equicontinuity points. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(5):569–582, September 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2000__36_5_569_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300001412.

Butucea:2008:NES

[BT08] Cristina Butucea and Marie-Luce Taupin. New M-estimators in semi-parametric regression with errors in variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3):393–421, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1211819418;   http://www.numdam.org/
item?id=AIHPB_2008__44_3_393_0.

Berger:2013:HPM

[BT13] Quentin Berger and Fabio Lucio Toninelli. Hierarchical pinning model in correlated random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):781–816, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1372772644.

Baldi:2015:FCI

[BT15] P. Baldi and S. Trapani. Fourier coefficients of invariant random fields on homogeneous spaces of compact Lie groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):648–671, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1428672686.

Broman:2016:CPC

[BT16] Erik I. Broman and Johan Tykesson. Connectedness of Poisson cylinders in Euclidean space. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):102–126, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089262.

Benjamini:2017:HS

[BT17] Itai Benjamini and Vincent Tassion. Homogenization via sprinkling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):997–1005, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1491897754.

Baldasso:2018:HCC

[BT18a] Rangel Baldasso and Augusto Teixeira. How can a clairvoyant particle escape the exclusion process? Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2177–2202, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849796.

Belomestny:2018:LRD

[BT18b] Denis Belomestny and Mathias Trabs. Low-rank diffusion matrix estimation for high-dimensional time-changed Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1583–1621, August 2018. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1531296030.

Blondel:2018:KCL

[BT18c] O. Blondel and C. Toninelli. Kinetically constrained lattice gases: Tagged particle diffusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2335–2348, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849801.

Baldasso:2020:SIZ

[BT20] Rangel Baldasso and Augusto Teixeira. Spread of an infection on the zero range process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1898–1928, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Spread-of-an-infection-
on-the-zero-range-process/10.1214/19-AIHP1021.full.

Bromberg:2018:TCL

[BU18] Michael Bromberg and Corinna Ulcigrai. A temporal central limit theorem for real-valued cocycles over rotations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2304–2334, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1539849800.

Budhiraja:2003:ASE

[Bud03] A. Budhiraja. Asymptotic stability, ergodicity and other asymptotic properties of the nonlinear filter. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):

919–941, November/December 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_6_919_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000220.

Budzinski:2020:IGH

[Bud20] Thomas Budzinski. Infinite geodesics in hyperbolic random triangulations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1129–1161, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Infinite-geodesics-in-
hyperbolic-random-triangulations/10.1214/19-AIHP996.
full.

Buraczewski:2007:TSM

[Bur07] Dariusz Buraczewski. On tails of stationary measures on a class of solvable groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(4):417–440, July/August 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_4_417_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000719.

Backhausz:2017:SMF

[BV17] Ágnes Backhausz and Bálint Virág. Spectral measures of factor of i.i.d. processes on vertex-transitive graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53 (4):2260–2278, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1511773745.

Bolthausen:2018:LED

[BvdHK18] Erwin Bolthausen, Remco van der Hofstad, and Gady Kozma. Lace expansion for dummies. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):141–153, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1519030822.

Bozejko:2001:RTM

[BW01] Marek Bożejko and Janusz Wysoczański. Remarks on ttransformations of measures and convolutions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(6): 737–761, November/December 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_6_737_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010846.

Brofferio:2005:GKE

[BW05] Sara Brofferio and Wolfgang Woess. Green kernel estimates and the full Martin boundary for random walks on lamplighter groups and Diestel–Leader graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(6):1101–1123, November/December 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_6_773_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000609. See erratum [BW06].

Brofferio:2006:EGK

[BW06] Sara Brofferio and Wolfgang Woess. Erratum to: “Green kernel estimates and the full Martin boundary for random walks on lamplighter groups and Diestel–Leader graphs” [Ann. I. H. Poincaré — PR 41 (2005) 1101–1123]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6): 773–774, November/December 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_6_1101_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020306000562.

See [BW05].

Belton:2015:ACQ

[BW15] Alexander C. R. Belton and Stephen J. Wills. An algebraic construction of quantum flows with unbounded generators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):349–375, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1421244409.

Bakhtin:2019:TFF

[BW19a] Yuri Bakhtin and Wei Wu. Transversal fluctuations for a first passage percolation model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1042–1060, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1557820841.

Bryc:2019:LFD

[BW19b] Włodzimierz Bryc and Yizao Wang. Limit fluctuations for density of asymmetric simple exclusion processes with open boundaries. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2169–2194, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/Limit-fluctuations-for-
density-of-asymmetric-simple-exclusion-processes-with/
10.1214/18-AIHP945.full.

Biane:1987:VFP

[BY87] Ph. Biane and M. Yor. Variations sur une formule de Paul Lévy. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2):359–377, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_S2_359_0.

Bai:2008:CLT

[BY08] Zhidong Bai and Jian-Feng Yao. Central limit theorems for eigenvalues in a spiked population model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3): 447–474, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1211819420;   http://www.numdam.org/
item?id=AIHPB_2008__44_3_447_0.

Benjamini:2016:HMP

[BY16] Itai Benjamini and Ariel Yadin. Harmonic measure in the presence of a spectral gap. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1050–1060, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723510.

Berestycki:2019:CSA

[BY19] Nathanaël Berestycki and Ariel Yadin. Condensation of a selfattracting random walk. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 55(2):835–861, May 2019. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1557820833.

Buzzi:2012:MMI

[BZ12] J. Buzzi and L. Zambotti. Mean mutual information and symmetry breaking for finite random fields. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):343–367, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1334148202.

Cabane:1979:TLM

[Cab79] R. Cabane. Sur les transformations linéaires modulo un. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(2): 187–193, April–June 1979. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1979__
15_2_187_0.

Cairoli:1982:ESP

[Cai82] R. Cairoli. Enveloppe de Snell d’un processus à paramètre bidimensionnel. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 18(1):47–53, January–March 1982. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1982__18_1_47_0.

Calbrix:1981:MNF

[Cal81] J. Calbrix. Mesures non σ-finies: désintégration et quelques autres propriétés. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(1):75–95, January–March 1981. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1981__17_1_75_0.

Cuesta-Albertos:1998:ADB

[CAM98] J. A. Cuesta-Albertos and C. Matrán. The asymptotic distribution of the bootstrap sample mean of an infinitesimal array. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(1):23–48, ???? 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_1_23_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800125.

Carnal:1980:ICP

[Car80] Étienne Carnal. Indépendance conditionnelle permutable. (French) [Conditional permutable independence]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(1):39–47, January–March 1980. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1980__16_1_39_0.

Carnal:1981:DCP

[Car81] E. Carnal. Description de certains processus markoviens indexés par des sous-ensembles du cercle. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(2):229–248, April–June 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_2_229_0.

Carne:1985:BMS

[Car85] T. K. Carne. Brownian motion and stereographic projection. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(2):187–196, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1985__21_2_187_0.

Carinci:2013:RHL

[Car13] Gioia Carinci. Random hysteresis loops. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):307–339, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1366117649.

Carrapatoso:2015:QQK

[Car15] Kleber Carrapatoso. Quantitative and qualitative Kac’s chaos on the Boltzmann’s sphere. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):993–1039, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1435759238.

Caraceni:2016:SLR

[Car16] Alessandra Caraceni. The scaling limit of random outerplanar maps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1667–1686, November 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1479373244.

Castell:2004:MDD

[Cas04] Fabienne Castell. Moderate deviations for diffusions in a random Gaussian shear flow drift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):337–366, May/June 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2004__40_3_337_0;   http://www.sciencedirect.
com/science/article/pii/S0246020303000670.

Casal:2007:SEU

[Cas07] Alberto Rodríguez Casal. Set estimation under convexity type assumptions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6):763–774, November/December 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_6_763_0;   http://www.sciencedirect.com/science/
article/pii/S0246020307000027.

Castillo:2017:PTP

[Cas17] Ismaël Castillo. Pólya tree posterior distributions on densities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):2074–2102, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1511773739.

Cattiaux:1986:HHP

[Cat86] Patrick Cattiaux. Hypoellipticité et hypoellipticité partielle pour les diffusions avec une condition frontière. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(1):67–112, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1986__22_1_67_0.

Catoni:1991:SLD

[Cat91a] O. Catoni. Sharp large deviations estimates for simulated annealing algorithms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(3):291–383, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1991__27_3_
291_0.

Catoni:1991:ASL

[Cat91b] Olivier Catoni. Applications of sharp large deviations estimates to optimal cooling schedules. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(4):463–518, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_4_463_0.

Catoni:2003:LTE

[Cat03] Olivier Catoni. Laplace transform estimates and deviation inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(1):1–26, January/February 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_2003__39_1_
1_0;   http://www.sciencedirect.com/science/article/
pii/S0246020302000171.

Catoni:2012:CEM

[Cat12] Olivier Catoni. Challenging the empirical mean and empirical variance: A deviation study. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1148–1185, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1353098444.

Caubet:1965:SGG

[Cau65] Jean-Pierre Caubet. Semi-groupes généralisés de matrices positives. (French) [Generalized semigroups of positive matrices]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 1(3):239–310, January– March 1965. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1965__1_3_239_0.

Caubet:1969:PVC

[Cau69] Jean-Pierre Caubet. Sur les processus à valeurs dans le cube fondamental de Hilbert. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(3):245–254, July–September 1969. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1969__5_3_245_0.

Caravenna:2008:IPR

[CC08] Francesco Caravenna and Loïc Chaumont. Invariance principles for random walks conditioned to stay positive. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1): 170–190, February 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1203969873;   http://www.numdam.org/
item?id=AIHPB_2008__44_1_170_0.

Chen:2016:CDF

[CCC16] Yu-Ting Chen, Jihyeok Choi, and J. Theodore Cox. On the convergence of densities of finite voter models to the Wright–Fisher diffusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):286–322, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1452089270.

Carlen:2013:ACE

[CCEL13] Eric A. Carlen, Dario Cordero-Erausquin, and Elliott H. Lieb. Asymmetric covariance estimates of Brascamp–Lieb type and related inequalities for log-concave measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1): 1–12, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1359470123.

Csaki:2009:RWL

[CCFR09] Endre Csáki, Miklós Csörgő, Antónia Földes, and Pál Révész. Random walk local time approximated by a Brownian sheet combined with an independent Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45 (2):515–544, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024679;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_515_0.

Chen:2021:QIP

[CCKW21] Xin Chen, Zhen-Qing Chen, Takashi Kumagai, and Jian Wang. Quenched invariance principle for long range random walks in balanced random environments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2243–2267, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/Quenched-invariance-
principle-for-long-range-random-walks-in-balanced/10.
1214/21-AIHP1150.full.

Carmona:2003:SIR

[CCM03] Philippe Carmona, Laure Coutin, and Gérard Montseny. Stochastic integration with respect to fractional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(1):27–68, January/February 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_1_27_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011111.

Chazottes:2019:TSQ

[CCM19] J.-R. Chazottes, P. Collet, and S. Méléard. On time scales and quasi-stationary distributions for multitype birth-and-death processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2249–2294, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/On-time-scales-and-quasi-
stationary-distributions-for-multitype-birth/10.1214/
18-AIHP948.full.

Caravenna:2012:DTP

[CCP12] Francesco Caravenna, Philippe Carmona, and Nicolas Pétrélis. The discrete-time parabolic Anderson model with heavy-tailed potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1049–1080, November 2012. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1353098440.

Csaki:1992:FTI

[CCS92] Endre Csáki, Miklós Csörgő, and Qi-Man Shao. Fernique type inequalities and moduli of continuity for l^{2}-valued Ornstein– Uhlenbeck processes. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 28(4):479–517, ???? 1992. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1992_
_28_4_479_0.

Cabanal-Duvillard:2001:FLE

[CD01] Thierry Cabanal-Duvillard. Fluctuations de la loi empirique de grandes matrices aléatoires. (French) [Fluctuations in the empirical law of large random matrices]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(3): 373–402, May/June 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_3_373_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300010712.

Cotar:2012:DCU

[CD12] Codina Cotar and Jean-Dominique Deuschel. Decay of covariances, uniqueness of ergodic component and scaling limit for a class of ∇ϕ systems with non-convex potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3): 819–853, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1340714874.

Chiarini:2016:IPS

[CD16] Alberto Chiarini and Jean-Dominique Deuschel. Invariance principle for symmetric diffusions in a degenerate and unbounded stationary and ergodic random medium. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52 (4):1535–1563, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1479373239.

Cerqueti:2018:SIM

[CD18] Roy Cerqueti and Emilio De Santis. Stochastic Ising model with flipping sets of spins and fast decreasing temperature. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):757–789, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1524643228.

Cerrai:2019:LDT

[CD19a] Sandra Cerrai and Arnaud Debussche. Large deviations for the two-dimensional stochastic Navier–Stokes equation with vanishing noise correlation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):211–236, February 2019. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1547802400.

Cloez:2019:ISM

[CD19b] Bertrand Cloez and Claire Delplancke. Intertwinings and Stein’s magic factors for birth-death processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55 (1):341–377, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1547802403.

Cerou:2011:NTU

[CDG11] F. Cérou, P. Del Moral, and A. Guyader. A nonasymptotic theorem for unnormalized Feynman–Kac particle models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3): 629–649, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834852;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_629_0.

Cerou:2020:CLT

[CDGR20] Frédéric Cérou, Bernard Delyon, Arnaud Guyader, and Mathias Rousset. A central limit theorem for Fleming-Viot particle systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):637–666, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/A-central-limit-theorem-
for-FlemingViot-particle-systems/10.1214/19-AIHP976.
full.

Chaumont:2000:ULL

[CDH00] L. Chaumont, R. A. Doney, and Y. Hu. Upper and lower limits of doubly perturbed Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(2):219–249, March 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2000__36_2_219_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300001230.

Coupier:2020:APP

[CDL20] David Coupier, David Dereudre, and Simon Le Stum. Absence of percolation for Poisson outdegree-one graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1179–1202, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-56/issue-
2/Absence-of-percolation-for-Poisson-outdegree-one-
graphs/10.1214/19-AIHP998.full.

Callaway:2017:DSR

[CDLR17] Mark Callaway, Thai Son Doan, Jeroen S. W. Lamb, and Martin Rasmussen. The dichotomy spectrum for random dynamical systems and pitchfork bifurcations with additive noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1548–1574, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1511773718.

Carpentier:2021:TVD

[CDM21] Alexandra Carpentier, Céline Duval, and Ester Mariucci. Total variation distance for discretely observed Lévy processes: A Gaussian approximation of the small jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):901–939, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-
2/Total-variation-distance-for-discretely-observed-
L%c3%a9vy-processes--A/10.1214/20-AIHP1102.full.

Capitaine:2012:CLT

[CDMF12] M. Capitaine, C. Donati-Martin, and D. Féral. Central limit theorems for eigenvalues of deformations of Wigner matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):107–133, February 2012. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1327328016.

Caputo:2009:CED

[CDP09] Pietro Caputo, Paolo Dai Pra, and Gustavo Posta. Convex entropy decay via the Bochner–Bakry–Emery approach. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3): 734–753, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1249391382;   http://www.numdam.org/
item?id=AIHPB_2009__45_3_734_0.

Cohen:2000:NSA

[CE00] Serge Cohen and Anne Estrade. Non-symmetric approximations for manifold-valued semimartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(1): 45–70, January 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2000__36_1_45_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300001060.

Chen:2020:SSE

[CE20] Zhen-Qing Chen and János Engländer. Superdiffusions with super-exponential growth: Construction, mass and spread. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1809–1840, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/Superdiffusions-with-super-exponential-
growth-Construction-mass-and-spread/10.1214/19-AIHP1018.
full.

Cellarosi:2011:LCM

[Cel11] Francesco Cellarosi. Limiting curlicue measures for theta sums. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2):466–497, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1300887278;   http://
www.numdam.org/item?id=AIHPB_2011__47_2_466_0.

Cerf:1996:DMS

[Cer96] Raphaël Cerf. The dynamics of mutation-selection algorithms with large population sizes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(4):455–508, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1996__32_4_455_0.

Cerny:2004:CPA

[Čer04] Jiří Černý. Critical path analysis for continuum percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(6):661–675, November/December 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_6_661_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000470.

Cottrell:1987:EPA

[CF87] Marie Cottrell and Jean-Claude Fort. Étude d’un processus d’auto-organisation. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(1):1–20, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1987_
_23_1_1_0.

Chen:2009:UET

[CF09] Zhen-Qing Chen and Masatoshi Fukushima. On unique extension of time changed reflecting Brownian motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3): 864–875, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1249391389;   http://www.numdam.org/
item?id=AIHPB_2009__45_3_864_0.

Cerrai:2017:SND

[CF17] Sandra Cerrai and Mark Freidlin. SPDEs on narrow domains and on graphs: An asymptotic approach. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2): 865–899, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1491897749.

Chouk:2018:STS

[CF18] K. Chouk and P. K. Friz. Support theorem for a singular SPDE: The case of gPAM. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 54(1):202–219, February 2018. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1519030825.

Corwin:2012:USD

[CFP12] Ivan Corwin, Patrik L. Ferrari, and Sandrine Péché. Universality of slow decorrelation in KPZ growth. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1): 134–150, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1327328017.

Caputo:2013:IPM

[CFP13] P. Caputo, A. Faggionato, and T. Prescott. Invariance principle for Mott variable range hopping and other walks on point processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):654–697, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1372772640.

Courtade:2019:ESK

[CFP19] Thomas A. Courtade, Max Fathi, and Ashwin Pananjady. Existence of Stein kernels under a spectral gap, and discrepancy bounds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):777–790, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1557820831.

Csaki:1987:JDM

[CFS87] Endre Csáki, Antónia Földes, and Paavo Salminen. On the joint distribution of the maximum and its location for a linear diffusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(2):179–194, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_2_179_0.

Cohen:1968:DCL

[CG68] J. W. Cohen and I. Greenberg. Distribution of crossings of level K in a busy cycle of the M/G/1 queue. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(1):75–81, January–March 1968. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1968__4_1_75_0.

Chacon:1979:EBM

[CG79] R. V. Chacon and N. A. Ghoussoub. Embeddings in Brownian motion. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(3):287– 292, July–September 1979. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1979__
15_3_287_0.

Coquio:1992:CMR

[CG92] A. Coquio and J. B. Gravereaux. Calcul de Malliavin et régularité de la densité d’une probabilité invariante d’une chaîne de Markov. (French) [Malliavin calculus and density regularity of an invariant probability of a Markov chain]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(4): 431–478, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1992__28_4_431_0.

Castell:1996:ODE

[CG96] Fabienne Castell and Jessica Gaines. The ordinary differential equation approach to asymptotically efficient schemes for solution of stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(2): 231–250, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_2_231_0.

Chen:2004:FMD

[CG04] Xia Chen and Arnaud Guillin. The functional moderate deviations for Harris recurrent Markov chains and applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):89–124, January/February 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_1_89_0;   http://www.sciencedirect.com/science/
article/pii/S024602030300061X.

Cattiaux:2009:TET

[CG09] Patrick Cattiaux and Arnaud Guillin. Trends to equilibrium in total variation distance. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):117–145, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469974;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_117_0.

Couronne:2014:BSP

[CG14] Olivier Couronné and Lucas Gerin. A branching-selection process related to censored Galton–Walton processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1): 84–94, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1388545266.

Clement:2019:EFS

[CG19] Emmanuelle Clément and Arnaud Gloter. Estimating functions for SDE driven by stable Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1316–1348, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1569398871.

Ciech:2021:LPP

[CG21] Federico Ciech and Nicos Georgiou. Last passage percolation in an exponential environment with discontinuous rates. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2165–2188, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/Last-passage-percolation-
in-an-exponential-environment-with-discontinuous-rates/
10.1214/21-AIHP1172.full.

Caballero:2017:APM

[CGB17] M. Emilia Caballero, José Luis Pérez Garmendia, and Gerónimo Uribe Bravo. Affine processes on R_{+}^{m} × R^{n} and multiparameter time changes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1280–1304, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1500624042.

Comte:2010:NAE

[CGC10] F. Comte and V. Genon-Catalot. Nonparametric adaptive estimation for pure jump Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3): 595–617, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1281100391;   http://www.numdam.org/
item?id=AIHPB_2010__46_3_595_0.

Caravenna:2010:LSB

[CGG10] Francesco Caravenna, Giambattista Giacomin, and Massimiliano Gubinelli. Large scale behavior of semiflexible heteropolymers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):97–118, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1267454110;   http://
www.numdam.org/item?id=AIHPB_2010__46_1_97_0.

Comte:2011:AEC

[CGG11] F. Comte, S. Gaïffas, and A. Guilloux. Adaptive estimation of the conditional intensity of marker-dependent counting processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1171–1196, November 2011. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1317906507;   http://www.numdam.org/item?id=AIHPB_2011_
_47_4_1171_0.

Can:2022:AIM

[CGGvdH22] Van Hao Can, Cristian Giardinà, Claudio Giberti, and Remco van der Hofstad. Annealed Ising model on configuration models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):134–163, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Annealed-Ising-model-
on-configuration-models/10.1214/21-AIHP1156.full.

Collet:2020:PSM

[CGK20] Francesca Collet, Matthias Gorny, and Richard C. Kraaij. Path-space moderate deviations for a Curie-Weiss model of self-organized criticality. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):765–781, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/

annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Path-space-moderate-
deviations-for-a-CurieWeiss-model-of-self/10.1214/19-
AIHP981.full.

Campi:2021:PGM

[CGL21] Luciano Campi, Maddalena Ghio, and Giulia Livieri. Nplayer games and mean-field games with smooth dependence on past absorptions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):1901–1939, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/N-Player-games-and-mean-
field-games-with-smooth-dependence/10.1214/20-AIHP1138.
full.

Camia:2016:PIM

[CGN16] Federico Camia, Christophe Garban, and Charles M. Newman. Planar Ising magnetization field II. Properties of the critical and near-critical scaling limits. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):146–161, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089264.

Cohen:2006:SFF

[CGPP06] Serge Cohen, Xavier Guyon, Olivier Perrin, and Monique Pontier. Singularity functions for fractional processes: application to the fractional Brownian sheet. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(2):187–205, March/ April 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2006__42_2_187_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000580.

Castell:2013:LTO

[CGPP13] Fabienne Castell, Nadine Guillotin-Plantard, and Françoise Pène. Limit theorems for one and two-dimensional random walks in random scenery. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):506–528, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1366117656.

Cocozza:1979:EDE

[CGR79] C. Cocozza, A. Galves, and M. Roussignol. Étude de deux évolutions markoviennes de processus ponctuels sur R par des méthodes d’association. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(3):235–259, July–September 1979. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_3_235_0.

Cattiaux:2013:PIH

[CGZ13] Patrick Cattiaux, Arnaud Guillin, and Pierre André Zitt. Poincaré inequalities and hitting times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):95–118, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1359470127.

Chartrand:1967:PPG

[CH67] Gary Chartrand and Frank Harary. Planar permutation graphs. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(4):433–438, October– December 1967. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1967__3_4_433_0.

Csorgo:1990:DNW

[CH90] Miklós Csörgő and Lajos Horváth. On the distributions of L_{p} norms of weighted quantile processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):65–85, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1990__26_1_65_0.

Carmona:2006:USK

[CH06] Philippe Carmona and Yueyun Hu. Universality in Sherrington– Kirkpatrick’s spin glass model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(2):215–222, March/ April 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2006__42_2_215_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000634.

Coupier:2012:CPL

[CH12] David Coupier and Philippe Heinrich. Coexistence probability in the last passage percolation model is 6 − 8 log 2. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4): 973–988, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1353098436.

Carmona:2014:SCB

[CH14] Philippe Carmona and Yueyun Hu. The spread of a catalytic branching random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):327–351, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1395856130.

Cipriani:2017:TPG

[CH17] Alessandra Cipriani and Rajat Subhra Hazra. Thick points for Gaussian free fields with different cut-offs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1): 79–97, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1486544885.

Chen:2020:LDM

[CH20] Xinxin Chen and Hui He. Lower deviation and moderate deviation probabilities for maximum of a branching random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2507–2539, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Lower-deviation-and-
moderate-deviation-probabilities-for-maximum-of-a/10.
1214/20-AIHP1048.full.

Chatard:1970:APM

[Cha70] Jacqueline Chatard. Applications des propriétés de moyenne d’un groupe localement compact à la théorie ergodique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(4):307– 326, October–December 1970. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_1_81_0.

Chatard:1971:EAM

[Cha71] Jacqueline Chatard. Erratum à l’article de Mme J. Chatard. (French) [Erratum for the article by J. Chatard]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(1):81–82, January–March 1971. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1970__6_4_307_0.

Chauvin:1986:APB

[Cha86a] B. Chauvin. Arbres et processus de Bellman–Harris. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2):209–232, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1986__22_2_209_0.

Chauvin:1986:PBF

[Cha86b] B. Chauvin. Sur la propriété de branchement. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2):233–236, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1986__22_2_233_0.

Chan:1991:IQF

[Cha91] Terence Chan. Indefinite quadratic functionals of Gaussian processes and least-action paths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(2):239–271, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_2_239_0.

Chan:1994:OTC

[Cha94] Terence Chan. Occupation times of compact sets by planar Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):317–329, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1994__30_2_
317_0.

Chapon:2012:ADP

[Cha12] François Chapon. Affine Dunkl processes of type Ã_{1}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3): 854–870, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1340714875.

Chase:2021:NLB

[Cha21] Zachary Chase. New lower bounds for trace reconstruction. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):627–643, May 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/New-lower-bounds-for-trace-reconstruction/
10.1214/20-AIHP1089.full.

Chevalier:1976:EDS

[Che76] Jacques Chevalier. Estimation du support et du contour du support d’une loi de probabilité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(4):339–364, October–December 1976. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1976__12_4_339_0.

Chen:1987:CPD

[Che87] Louis H. Y. Chen. Characterization of probability distributions by Poincaré-type inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(1):91–110, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1987_
_23_1_91_0.

Chen:1997:APR

[Che97] Dayue Chen. Average properties of random walks on Galton– Watson trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(3):359–369, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_3_359_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397800969.

Cheliotis:2006:ODD

[Che06] Dimitrios Cheliotis. One-dimensional diffusion in an asymmetric random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6):715–726, November/December 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_6_715_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000057.

Chen:2008:LLE

[Che08] Xia Chen. Limit laws for the energy of a charged polymer. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4):638–672, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1217964114;   http://
www.numdam.org/item?id=AIHPB_2008__44_4_638_0.

Chen:2014:MES

[Che14] Wei-Kuo Chen. On the mixed even-spin Sherrington– Kirkpatrick model with ferromagnetic interaction. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1): 63–83, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1388545265.

Cheliotis:2015:MSB

[Che15] Dimitris Cheliotis. Metastable states in Brownian energy landscape. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):917–934, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1435759234.

Chen:2017:MAP

[Che17a] Xia Chen. Moment asymptotics for parabolic Anderson equation with fractional time-space noise: In Skorokhod regime. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):819–841, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1491897747.

Chen:2017:PUE

[Che17b] Xinxing Chen. Pointwise upper estimates for transition probabilities of continuous time random walks on graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1): 27–45, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1486544883.

Chen:2018:TPH

[Che18] Linan Chen. Thick points of high-dimensional Gaussian free fields. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1492–1526, August 2018. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1531296027.

Chen:2019:PAM

[Che19] Xia Chen. Parabolic Anderson model with rough or critical Gaussian noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):941–976, May 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1557820837.

Chen:2020:PAM

[Che20] Xia Chen. Parabolic Anderson model with a fractional Gaussian noise that is rough in time. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):792–825, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Parabolic-Anderson-model-
with-a-fractional-Gaussian-noise-that-is/10.1214/19-
AIHP983.full.

Chen:2015:EAT

[CHSX15] Xia Chen, Yaozhong Hu, Jian Song, and Fei Xing. Exponential asymptotics for time-space Hamiltonians. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1529–1561, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1445432052.

Chung:1984:LCB

[Chu84] Kai Lai Chung. The lifetime of conditional Brownian motion in the plane. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(4):349–351, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1984__20_4_349_0.

Csoka:2020:EE

[CHV20] Endre Csóka, Viktor Harangi, and Bálint Virág. Entropy and expansion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2428–2444, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Entropy-and-expansion/
10.1214/19-AIHP1044.full.

Chybiryakov:2008:SPR

[Chy08] Oleksandr Chybiryakov. Skew-product representations of multidimensional Dunkl Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4): 593–611, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1217964112;   http://www.numdam.org/
item?id=AIHPB_2008__44_4_593_0.

Caputo:2003:FVA

[CI03] Pietro Caputo and Dmitry Ioffe. Finite volume approximation of the effective diffusion matrix: the case of independent bond disorder. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(3):505–525, May/June 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_3_505_0;   http://www.sciencedirect.com/science/
article/pii/S024602030200016X.

Conus:2014:IMS

[CJKS14] Daniel Conus, Mathew Joseph, Davar Khoshnevisan, and Shang-Yuan Shiu. Initial measures for the stochastic heat equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):136–153, February 2014. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1388545269.

Cocozza:1977:EPM

[CK77] C. Cocozza and C. Kipnis. Existence de processus markoviens pour des systèmes infinis de particules. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(3):239–257, July–September 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_3_239_0.

Chen:2012:BMP

[CK12] Xia Chen and Alexey M. Kulik. Brownian motion and parabolic Anderson model in a renormalized Poisson potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3): 631–660, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1340714866.

Chen:2017:CPS

[CK17] Le Chen and Kunwoo Kim. On comparison principle and strict positivity of solutions to the nonlinear stochastic fractional heat equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):358–388, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1486544895.

Cerenzia:2020:HEA

[CK20] Mark Cerenzia and Jeffrey Kuan. Hard-edge asymptotics of the Jacobi growth process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2329–2355, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Hard-edge-asymptotics-
of-the-Jacobi-growth-process/10.1214/19-AIHP1040.full.

Cebron:2022:FBM

[CK22] Guillaume Cébron and Todd Kemp. Fluctuations of Brownian motions on GL_{N}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):524–547, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Fluctuations-of-Brownian-
motions-on-GLN/10.1214/21-AIHP1165.full.

Chen:2017:BTS

[CKK17] Le Chen, Davar Khoshnevisan, and Kunwoo Kim. A boundedness trichotomy for the stochastic heat equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53 (4):1991–2004, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1511773735.

Crisan:2014:CDE

[CKL14] Dan Crisan, Thomas G. Kurtz, and Yoonjung Lee. Conditional distributions, exchangeable particle systems, and stochastic partial differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):946–974, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1403277004.

Carlen:1987:UBS

[CKS87] E. A. Carlen, S. Kusuoka, and D. W. Stroock. Upper bounds for symmetric Markov transition functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2):245–287, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1987__23_S2_245_0.

Cairoli:1984:TFA

[CL84] R. Cairoli and M. Ledoux. Une topologie fine associée au produit de deux processus markoviens. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(4): 299–307, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1984__20_4_299_0.

Cattiaux:1994:MKI

[CL94] Patrick Cattiaux and Christian Léonard. Minimization of the Kullback information of diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(1):83– 132, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1995__31_4_705_0. See erratum [CL95].

Cattiaux:1995:ECM

[CL95] P. Cattiaux and C. Léonard. Erratum / correction to: “Minimization of the Kullback information of diffusion processes”. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(4):705–707, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1994__30_1_83_0. See [CL94].

Cuny:2009:PET

[CL09] Christophe Cuny and Michael Lin. Pointwise ergodic theorems with rate and application to the CLT for Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):710–733, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1249391381;   http://
www.numdam.org/item?id=AIHPB_2009__45_3_710_0.

Comte:2013:AAK

[CL13a] F. Comte and C. Lacour. Anisotropic adaptive kernel deconvolution. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):569–609, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1366117659.

Conze:2013:AEC

[CL13b] Jean-Pierre Conze and Michael Lin. Almost everywhere convergence of convolution powers on compact Abelian groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):550–568, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1366117658.

Curien:2017:SLP

[CL17] Nicolas Curien and Jean-François Le Gall. Scaling limits for the peeling process on random maps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):322–357, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544894.

Cass:2021:SRI

[CL21] Thomas Cass and Nengli Lim. Skorohod and rough integration for stochastic differential equations driven by Volterra processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):132–168, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/Skorohod-and-rough-integration-
for-stochastic-differential-equations-driven-by/10.
1214/20-AIHP1074.full.

Castelle:1998:SAB

[CLB98] Nathalie Castelle and Françoise Laurent-Bonvalot. Strong approximations of bivariate uniform empirical processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34 (4):425–480, July/August 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_4_425_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800241.

Cranston:1989:NTP

[ClJ89] M. Cranston and Y. le Jan. On the noncoalescence of a two point Brownian motion reflecting on a circle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2):99– 107, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1989__25_2_99_0.

Curien:2013:BCS

[CLM13] Nicolas Curien, Jean-François Le Gall, and Grégory Miermont. The Brownian cactus I. Scaling limits of discrete cactuses. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):340–373, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1366117650.

Chauvin:2014:LDM

[CLP14] Brigitte Chauvin, Quansheng Liu, and Nicolas Pouyanne. Limit distributions for multitype branching processes of m-ary search trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):628–654, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1395856143.

Conze:1970:CTA

[CM70] Jean-Pierre Conze and Jean-Claude Marcuard. Conjugaison topologique des automorphismes et des translations ergodiques de nilvariétés. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(2):153–157, April–June 1970. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1970__
6_2_153_0.

Caillot:1972:MBF

[CM72] Philippe Caillot and Françoise Martin. Le modèle bayésien. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(2): 183–210, April–June 1972. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1972__
8_2_183_0.

Cohen:2008:IPM

[CM08] Serge Cohen and Renaud Marty. Invariance principle, multifractional Gaussian processes and long-range dependence. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3):475–489, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1211819421;   http://www.numdam.org/
item?id=AIHPB_2008__44_3_475_0.

Chafai:2010:FPM

[CM10] Djalil Chafaï and Florent Malrieu. On fine properties of mixtures with respect to concentration of measure and Sobolev type inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):72–96, February 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1267454109;   http://www.numdam.org/item?id=AIHPB_2010_
_46_1_72_0.

Cammarota:2015:LBN

[CM15] Valentina Cammarota and Domenico Marinucci. On the limiting behaviour of needlets polyspectra. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):1159–1189, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1435759244.

Chaumont:2016:ABD

[CM16] Loïc Chaumont and Jacek Malecki. On the asymptotic behavior of the density of the supremum of Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3): 1178–1195, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1469723516.

Castillo:2021:SSP

[CM21] Ismaël Castillo and Romain Mismer. Spike and slab Pólya tree posterior densities: Adaptive inference. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1521–1548, August 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-3/Spike-and-slab-P%c3%b3lya-tree-posterior-
densities-Adaptive-inference/10.1214/20-AIHP1132.full.

Cancrini:2002:LSC

[CMR02] N. Cancrini, F. Martinelli, and C. Roberto. The logarithmic Sobolev constant of Kawasaki dynamics under a mixing condition revisited. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(4):385–436, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_4_385_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010962.

Cheridito:2005:SID

[CN05] Patrick Cheridito and David Nualart. Stochastic integral of divergence type with respect to fractional Brownian motion with Hurst parameter H ∈ (0, \frac{1}{2}). Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(6):1049–1081, November/December 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_6_1049_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020304000913.

Cosco:2021:GFD

[CN21] Clément Cosco and Shuta Nakajima. Gaussian fluctuations for the directed polymer partition function in dimension. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):872–889, May 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/Gaussian-fluctuations-for-the-directed-
polymer-partition-function-in-dimension/10.1214/20-
AIHP1100.full.

Catoni:2021:CPZ

[COBZ21] Olivier Catoni, Miquel Oliu-Barton, and Bruno Ziliotto. Constant payoff in zero-sum stochastic games. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):1888–1900, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/Constant-payoff-in-
zero-sum-stochastic-games/10.1214/20-AIHP1146.full.

Cohen:1968:EVD

[Coh68] J. W. Cohen. Extreme value distribution for the M/G/1 and the G/M/1 queueing systems. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(1):83–98, January–March 1968. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1968__4_1_83_0.

Cohen:1972:TSW

[Coh72] J. W. Cohen. On the tail of the stationary waiting time distribution and limit theorems for the M/G/1 queue. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(3):255–263, July–September 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1972__8_3_255_0.

Collombier:1978:AAA

[Col78] Dominique Collombier. Un algorithme d’aide à l’analyse des tables de contingence et des plans d’expériences incomplets. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(2): 215–231, April–June 1978. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1978__
14_2_215_0.

Collins:2004:MBT

[Col04] Benoît Collins. Martin boundary theory of some quantum random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):367–384, May/June 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_3_367_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000682.

Comets:1987:NLR

[Com87] Francis Comets. Nucleation for a long range magnetic model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(2):135–178, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1987__23_2_135_0.

Conolly:1965:MAD

[Con65] B. W. Conolly. Marche aléatoire dont la répartition de la longueur des étapes suit une loi exponentielle négative. (French) [Random walk whose stage-length distribution follows a negative exponential law]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(2):173–184, October–December 1965. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1965__2_2_173_0.

Conze:1972:ETA

[Con72a] Jean-Pierre Conze. Entropie des transformations affines et des flots sur les espaces homogènes compacts. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(1):67–81, January–March 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1972__8_1_67_0.

Conze:1972:ESD

[Con72b] Jean-Pierre Conze. Extensions de systèmes dynamiques par des endomorphismes de groupes compacts. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(1):33–66, January–March 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1972__8_1_33_0.

Concordet:1994:EDD

[Con94] D. Concordet. Estimation de la densité du recuit simulé. (French) [Estimation of the density of simulated annealing]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):265–302, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1994__30_2_265_0.

Conforti:2018:FBR

[Con18] Giovanni Conforti. Fluctuations of bridges, reciprocal characteristics and concentration of measure. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1432–1463, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1531296025.

Cook:2019:CLR

[Coo19] Nicholas Cook. The Circular Law for random regular digraphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2111–2167, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/The-Circular-Law-for-
random-regular-digraphs/10.1214/18-AIHP943.full.

Coquio:1993:FDE

[Coq93] Agnès Coquio. Forme de Dirichlet sur l’espace canonique de Poisson et applications aux équations différentielles stochastiques. (French) [Dirichlet shape on the Poisson canonical space and applications to stochastic differential equations]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(1): 1–36, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1993__29_1_1_0.

Coste:2021:SGS

[Cos21] Simon Coste. The spectral gap of sparse random digraphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):644–684, May 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/The-spectral-gap-of-sparse-random-
digraphs/10.1214/20-AIHP1090.full.

Courtade:2020:BPC

[Cou20] Thomas A. Courtade. Bounds on the Poincaré constant for convolution measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):566–579, February 2020.

CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Bounds-on-the-Poincar%
c3%a9-constant-for-convolution-measures/10.1214/19-
AIHP973.full.

Coston:2020:OSP

[COW20] Natalie Coston, Sean O’Rourke, and Philip Matchett Wood. Outliers in the spectrum for products of independent random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1284–1320, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/Outliers-in-the-spectrum-for-products-
of-independent-random-matrices/10.1214/19-AIHP1002.
full.

Charlot:1978:RSS

[CP78] François Charlot and Guy Pujolle. Recurrence in single server queues with impatient customers. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4):399–410, October–December 1978. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1978__14_4_399_0.

Cerf:2001:PCI

[CP01] Raphaël Cerf and Ágoston Pisztora. Phase coexistence in Ising, Potts and percolation models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(6):643–724, November/December 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_6_643_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010834.

Cox:2004:AVM

[CP04] J. Theodore Cox and Edwin A. Perkins. An application of the voter model-super-Brownian motion invariance principle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):25–32, January/February 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_

_40_1_25_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000463.

Coulibaly-Pasquier:2011:BMR

[CP11] Koléhè A. Coulibaly-Pasquier. Brownian motion with respect to time-changing Riemannian metrics, applications to Ricci flow. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2):515–538, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1300887280;   http://
www.numdam.org/item?id=AIHPB_2011__47_2_515_0.

Comets:2012:BRR

[CP12] Francis Comets and Serguei Popov. Ballistic regime for random walks in random environment with unbounded jumps and Knudsen billiards. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):721–744, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1340714870.

Chen:2017:SEQ

[CP17] Wei-Kuo Chen and Dmitry Panchenko. Some examples of quenched self-averaging in models with Gaussian disorder. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):243–258, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1486544891.

Courrege:1971:RLE

[CPP71] Ph. Courrège, J.-L. Philoche, and P. Priouret. Régression linéaire et estimation par la méthode des moindres carrés. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(4):253– 270, October–December 1971. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_4_253_0.

Cutting:2022:TUH

[CPV22] Christine Cutting, Davy Paindaveine, and Thomas Verdebout. Testing uniformity on high-dimensional spheres: The non-null behaviour of the Bingham test. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):567–602, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Testing-uniformity-on-
high-dimensional-spheres--The-non-null/10.1214/21-AIHP1168.
full.

Cosso:2017:BDC

[CPX17] Andrea Cosso, Huyên Pham, and Hao Xing. BSDEs with diffusion constraint and viscous Hamilton–Jacobi equations with unbounded data. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1528–1547, November 2017. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1511773717.

Carmona:2004:TLC

[CPY04] Philippe Carmona, Frédérique Petit, and Marc Yor. A trivariate law for certain processes related to perturbed Brownian motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(6):737–758, November/December 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_6_737_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000251.

Comets:2007:FFS

[CQR07] Francis Comets, Jeremy Quastel, and Alejandro F. Ramírez. Fluctuations of the front in a stochastic combustion model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(2):147–162, March/April 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_2_147_0;   http://
www.sciencedirect.com/science/article/pii/S024602030600046X.

Crepel:1978:TCL

[CR78] Pierre Crépel and Albert Raugi. Théorème central limite sur les groupes nilpotents. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(2):145–164, April–June 1978. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_2_145_0.

Cocozza:1979:UPN

[CR79] C. Cocozza and M. Roussignol. Unicité d’un processus de naissance et mort sur la droite réelle. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1):93–106, January–March 1979. CODEN AHPBAR. ISSN 0020-2347. URL http://

www.numdam.org/item?id=AIHPB_1979__15_1_93_0.

Cocozza:1980:TEP

[CR80] C. Cocozza and M. Roussignol. Théorèmes ergodiques pour un processus de naissance et mort sur la droite réelle. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(1):75–85, January– March 1980. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1980__16_1_75_0.

Cramer:1996:CBT

[CR96] M. Cramer and L. Rüschendorf. Convergence of a branching type recursion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(6):725–741, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1996__32_6_
725_0.

Cerrai:2005:LDI

[CR05a] Sandra Cerrai and Michael Röckner. Large deviations for invariant measures of stochastic reaction-diffusion systems with multiplicative noise and non-Lipschitz reaction term. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(1):69–105, January/February 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_1_69_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000500.

Chen:2005:EAI

[CR05b] Xia Chen and Jay Rosen. Exponential asymptotics for intersection local times of stable processes and random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5):901–928, September/October 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_

_41_5_901_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000925.

Chassagneux:2020:ORB

[CR20] Jean-François Chassagneux and Adrien Richou. Obliquely reflected backward stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2868–2896, November 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-4/Obliquely-reflected-backward-stochastic-
differential-equations/10.1214/20-AIHP1061.full.

Croydon:2008:CSR

[Cro08] David Croydon. Convergence of simple random walks on random discrete trees to Brownian motion on the continuum random tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):987–1019, December 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1227287562;   http://www.numdam.org/item?id=AIHPB_2008_
_44_6_987_0.

Croydon:2018:SLS

[Cro18] D. A. Croydon. Scaling limits of stochastic processes associated with resistance forms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):1939–1968, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1539849789.

Csaki:1998:FLI

[CRR98] Endre Csáki, Pál Révész, and Jay Rosen. Functional laws of the iterated logarithm for local times of recurrent random walks on Z^{2}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(4):545–563, July/August 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_4_545_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800277.

Chiang:2002:SPD

[CS02] Tzuu-Shuh Chiang and Shuenn-Jyi Sheu. Small perturbation of diffusions in inhomogeneous media. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(3):285–318, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_3_285_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301011013.

Collins:2007:NSI

[CŚ07] Benoît Collins and Piotr Śniady. New scaling of Itzykson– Zuber integrals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(2):139–146, March/April 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_2_139_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000574.

Candellero:2021:CCF

[CS21] Elisabetta Candellero and Alexandre Stauffer. Coexistence of competing first passage percolation on hyperbolic graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2128–2164, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/Coexistence-of-competing-
first-passage-percolation-on-hyperbolic-graphs/10.1214/
20-AIHP1134.full.

Coulhon:1990:POR

[CSC90] Thierry Coulhon and Laurent Saloff-Coste. Puissances d’un opérateur régularisant. (French) [Powers of a smoothing operator]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):419–436, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1990__26_3_419_0.

Comtet:2011:EDP

[CT11] Alain Comtet and Yves Tourigny. Excursions of diffusion processes and continued fractions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3):

850–874, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834861;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_850_0.

Cerf:2016:WST

[CT16] Raphaël Cerf and Marie Théret. Weak shape theorem in first passage percolation with infinite passage times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3): 1351–1381, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1469723523.

Candellero:2018:PIR

[CT18] Elisabetta Candellero and Augusto Teixeira. Percolation and isoperimetry on roughly transitive graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):1819–1847, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849785.

Cocozza-Thivent:1995:TCF

[CTR95] C. Cocozza-Thivent and M. Roussignol. Techniques de couplage en fiabilité. (French) [Coupling techniques in reliability]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):119–141, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1995__31_1_119_0.

Cerny:2011:GVC

[ČTW11] Jiří Černý, Augusto Teixeira, and David Windisch. Giant vacant component left by a random walk in a random dregular graph. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):929–968, November 2011. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1317906496;   http://www.numdam.org/item?id=AIHPB_2011_
_47_4_929_0.

Cuenca:2018:MPD

[Cue18] Cesar Cuenca. Markov processes on the duals to infinitedimensional classical Lie groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1359–1407, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1531296023.

Cuny:2003:TAV

[Cun03] Christophe Cuny. Un TCL avec vitesse pour la marche aléatoire gauche sur le groupe affine de R^{d}. (French) [A TCL with speed for the left random walk on the affine group of R^{d}]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(3):487–503, May/June 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_3_487_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000213.

Cuppens:1969:DFC

[Cup69] Roger Cuppens. Décomposition des fonctions caractéristiques indéfiniment divisibles de plusieurs variables à spectre de Poisson continu. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(2):123–133, April–June 1969. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1969__
5_2_123_0.

Coletti:2014:CBW

[CV14] Cristian Coletti and Glauco Valle. Convergence to the Brownian Web for a generalization of the drainage network model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):899–919, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1403277002.

Chen:2013:REP

[CW13] Bo Chen and Matthias Winkel. Restricted exchangeable partitions and embedding of associated hierarchies in continuum random trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):839–872, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1372772646.

Chen:2018:PNL

[CW18] Zhen-Qing Chen and Jie-Ming Wang. Perturbation by non-local operators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):606–639, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:

//projecteuclid.org/euclid.aihp/1524643224. See errata

[CW20].

Chen:2020:EPN

[CW20] Zhen-Qing Chen and Jie-Ming Wang. Errata for Perturbation by non-local operators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):760–763, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Errata-for-Perturbation-
by-non-local-operators/10.1214/20-AIHP1045.full. See

[CW18].

Cerrai:2021:IVF

[CX21] Sandra Cerrai and Guangyu Xi. Incompressible viscous fluids in R^{2} and SPDEs on graphs, in presence of fast advection and non smooth noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1636–1664, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Incompressible-viscous-
fluids-in-R2-and-SPDEs-on-graphs-in/10.1214/20-AIHP1118.
full.

Chen:2007:MCA

[CY07] Xinxing Chen and Jiangang Ying. The Markov chain asymptotics of random mapping graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(3):353–374, May/June 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2007__43_3_353_0;   http://www.sciencedirect.
com/science/article/pii/S0246020306000756.

Chen:2002:GFK

[CZ02] Zhen-Qing Chen and Tu-Sheng Zhang. Girsanov and Feynman–Kac type transformations for symmetric Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(4):475–505, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_4_475_0;   http://www.sciencedirect.com/science/
article/pii/S024602030101086X.

Dabrowski:2014:FSP

[Dab14] Yoann Dabrowski. A free stochastic partial differential equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1404–1455, November 2014. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1413555506.

Dahlqvist:2017:IFB

[Dah17] Antoine Dahlqvist. Integration formulas for Brownian motion on classical compact Lie groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1971–1990, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773734.

Dalalyan:2020:EWM

[Dal20] Arnak S. Dalalyan. Exponential weights in multivariate regression and a low-rankness favoring prior. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1465–1483, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-2/Exponential-weights-
in-multivariate-regression-and-a-low-rankness-favoring/
10.1214/19-AIHP1010.full.

Delode:1976:CMM

[DAP76] C. Delode, O. Arino, and J.-P. Penot. Champs mesurables et multisections. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(1):11–42, January–March 1976. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1976__12_1_11_0.

Darling:1985:CMM

[Dar85] R. W. R. Darling. Convergence of martingales on manifolds of negative curvature. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 21(2):157–175, ???? 1985. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1985_
_21_2_157_0.

Darling:1996:MNM

[Dar96] R. W. R. Darling. Martingales on noncompact manifolds: maximal inequalities and prescribed limits. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(4):431–454, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1996__32_4_431_0.

Dario:2021:QHD

[Dar21] Paul Dario. Quantitative homogenization of differential forms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):1157–1202, May 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/Quantitative-homogenization-of-differential-
forms/10.1214/20-AIHP1111.full.

Daures:1973:VMD

[Dau73] J.-P. Daurès. Version multivoque du théorème de Doob. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(2): 167–176, April–June 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_2_167_0.

Davydov:1974:CFP

[Dav74] Iu. A. Davydov. Sur une classe des fonctionnelles des processus stables et des marches aléatoires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(1):1–29, January–March 1974. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1974__10_1_1_0.

Davies:1975:MME

[Dav75] E. B. Davies. Markovian master equations. III. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(3):265–273, July–September 1975. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1975__11_3_265_0.

Davydov:2002:ECP

[Dav02] Youri Davydov. Enveloppes convexes des processus gaussiens. (French) [Convex hulls of Gaussian processes]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6): 847–861, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_847_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302011299.

Daviaud:2005:TPC

[Dav05] Olivier Daviaud. Thick points for the Cauchy process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5): 953–970, September/October 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_5_953_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000901.

delBarrio:1999:NCB

[dBMCA99] E. del Barrio, C. Matrán, and J. A. Cuesta-Albertos. Necessary conditions for the bootstrap of the mean of a triangular array. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(3):371–386, May/June 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_3_371_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399800168.

Dacunha-Castelle:1973:RIE

[DC73] D. Dacunha-Castelle. Remarques sur les isomorphismes entre espaces d’Orlicz. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(1):59–75, January–March 1973. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1973__9_1_59_0.

Dacunha-Castelle:2002:BDD

[DC02] Didier Dacunha-Castelle. Biographie de Didier Dacunha-Castelle. (French) [Biography of Didier Dacunha-Castelle]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):805–806, November/December 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020302011421. En l’honneur de J. Bre-

tagnolle, D. Dacunha-Castelle, I. Ibragimov.

Dacunha-Castelle:1990:MEP

[DCG90] Didier Dacunha-Castelle and Fabrice Gamboa. Maximum d’entropie et problème des moments. (French) [Maximum entropy and moment problem]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(4):567–596, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1990__26_4_567_0.

Duminil-Copin:2014:SSA

[DCKY14] Hugo Duminil-Copin, Gady Kozma, and Ariel Yadin. Supercritical self-avoiding walks are space-filling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2): 315–326, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1395856129.

Dacunha-Castelle:1974:TPP

[DCS74] Didier Dacunha-Castelle and Michel Schreiber. Techniques probabilistes pour l’étude de problèmes d’isomorphismes entre espaces de Banach. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(2):229–277, April–June 1974. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1974__10_2_229_0.

Dembo:2007:AID

[DD07] Amir Dembo and Jean-Dominique Deuschel. Aging for interacting diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(4):461–480, July/August 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_4_461_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000689.

Dembo:2010:MPR

[DD10] Amir Dembo and Jean-Dominique Deuschel. Markovian perturbation, response and fluctuation dissipation theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3): 822–852, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1281100400;   http://www.numdam.org/
item?id=AIHPB_2010__46_3_822_0.

Dembo:2013:PIP

[DDG13] Amir Dembo, Jian Ding, and Fuchang Gao. Persistence of iterated partial sums. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):873–884, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1372772647.

Dehling:1987:ASI

[DDP87] Herold Dehling, Manfred Denker, and Walter Philipp. The almost sure invariance principle for the empirical process of U-statistic structure. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(2):121–134, ???? 1987. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1987_
_23_2_121_0.

DaPrato:2004:ISC

[DDT04] Giuseppe Da Prato, Arnaud Debussche, and Luciano Tubaro. Irregular semi-convex gradient systems perturbed by noise and application to the stochastic Cahn–Hilliard equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40 (1):73–88, January/February 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_1_73_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000608.

deLaPena:1995:ECB

[de 95] Victor H. de La Peña. Erratum / correction to: “A bound on the moment generating function of a sum of dependent variables with an application to simple random sampling without replacement”. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(4):703–704, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//www.numdam.org/item?id=AIHPB_1994__30_2_197_0. See

[dlP94].

DeMeyer:1998:MVB

[De 98] Bernard De Meyer. The maximal variation of a bounded martingale and the central limit theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(1): 49–59, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_1_49_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800174.

DeMeyer:1999:RGB

[De 99] Bernard De Meyer. From repeated games to Brownian games. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(1):1–48, January/February 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1999__35_1_
1_0;   http://www.sciencedirect.com/science/article/
pii/S0246020399800041.

DeSantis:2002:GDS

[De 02] Emilio De Santis. Glauber dynamics of spin glasses at low and high temperature. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(5):681–710, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_5_681_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011068.

DeAngelis:2018:OSB

[De 18] Tiziano De Angelis. From optimal stopping boundaries to Rost’s reversed barriers and the Skorokhod embedding. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):1098–1133, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1524643241.

Diaconis:2002:DCG

[DE02] Persi Diaconis and Steven N. Evans. A different construction of Gaussian fields from Markov chains: Dirichlet covariances. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):863–878, November/December 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_

_38_6_863_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011238.

Dumitriu:2005:EHL

[DE05] Ioana Dumitriu and Alan Edelman. Eigenvalues of Hermite and Laguerre ensembles: large beta asymptotics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(6): 1083–1099, November/December 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//www.numdam.org/item?id=AIHPB_2005__41_6_1083_0;
http://www.sciencedirect.com/science/article/pii/S0246020305000646.

Drewitz:2016:TVS

[DE16] Alexander Drewitz and Dirk Erhard. Transience of the vacant set for near-critical random interlacements in high dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):84–101, February 2016. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1452089261.

Decreusefond:2005:SIR

[Dec05] L. Decreusefond. Stochastic integration with respect to Volterra processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(2):123–149, March/April 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_2_123_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000536.

Defosseux:2010:OMR

[Def10] Manon Defosseux. Orbit measures, random matrix theory and interlaced determinantal processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1): 209–249, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1267454115;   http://www.numdam.org/
item?id=AIHPB_2010__46_1_209_0.

Defosseux:2016:FCR

[Def16] Manon Defosseux. Fusion coefficients and random walks in alcoves. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1515–1534, November 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1479373238.

Degerine:1979:LML

[Deg79a] Serge Degerine. Lois de von Mises et lois liées. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1):63–77, January– March 1979. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1979__15_1_63_0.

Degerine:1979:TOP

[Deg79b] Serge Degerine. Tests optimaux sur les paramètres des lois de von Mises. (French) [Optimal tests on the parameters of the von Mises laws]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(4):375–392, October–December 1979. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_4_375_0.

Degerine:1982:PMC

[Deg82] Serge Degerine. Propriétés de moindres carrés de la matrice de covariance d’un vecteur aléatoire et applications statistiques. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(3):237–247, July–September 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1982__18_3_237_0.

Deheuvels:1974:VEE

[Deh74] Paul Deheuvels. Valeurs extrémales d’échantillons croissants d’une variable aléatoire réelle. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(1):89–114, January–March 1974. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1974__10_1_89_0.

Deheuvels:2000:CTF

[Deh00] Paul Deheuvels. Chung-type functional laws of the iterated logarithm for tail empirical processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(5):583–616, September 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2000__36_5_583_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300001436.

Delporte:1964:FAP

[Del64] Jean Delporte. Fonctions aléatoires presque sûrement continues sur un intervalle fermé. (French) [Random functions almost surely continuous over a closed interval]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 1(2):111–215, 1964. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1964__1_2_111_0.

Delasnerie:1977:FMM

[Del77a] Michel Delasnerie. Flot mélangeant et mesures de Palm. (French) [Mixing flow and Palm measures]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(4):357–369, October–December 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_4_357_0.

Delode:1977:CME

[Del77b] Claude Delode. Champs mesurables d’espaces sousliniens. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(2): 181–191, April–June 1977. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1977__
13_2_181_0.

Delesalle:1987:CDF

[Del87] Dominique Delesalle. Classes de Donsker et fonction d’entropie relatives aux mesures aléatoires. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(3): 531–559, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1987__23_3_531_0.

Delarue:2008:HTC

[Del08] F. Delarue. Hitting time of a corner for a reflected diffusion in the square. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5):946–961, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1222261919;   http://
www.numdam.org/item?id=AIHPB_2008__44_5_946_0.

Depaix:1965:EDC

[Dep65] Michel Depaix. Étude des distributions de certains processus de marche au hasard dans l’espace Z^{n} limité par des barrières. (French) [Study of the distributions of some random-walk process in the space Z^{n} bounded by barriers]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 1(4):311–439, April–June 1965. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1965__1_4_311_0.

Depauw:1999:FMC

[Dep99] Jérôme Depauw. Flux moyen d’un courant électrique dans un réseau aléatoire stationnaire de résistances. (French) [Average flow of an electric current in a stationary random resistor network]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(3):355–370, May/June 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_3_355_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399800156.

Derriennic:1973:FMP

[Der73] Yves Derriennic. Sur la frontière de Martin des processus de Markov à temps discret. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(3):233–258, July–September 1973. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1973__9_3_233_0.

Derriennic:1976:LFZ

[Der76] Yves Derriennic. Lois ^{≪} zéro ou deux ^{≫} pour les processus de Markov. Applications aux marches aléatoires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(2):111–129, April– June 1976. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1976__12_2_111_0.

Dermoune:1990:DEP

[Der90] A. Dermoune. Distributions sur l’espace de P. Lévy et calcul stochastique. (French) [Distributions on the space of P. Lévy and stochastic calculus]. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 26(1):101–119, ???? 1990. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1990_
_26_1_101_0.

Derrien:2000:ECC

[Der00] Jean-Marc Derrien. On the existence of cohomologous continuous cocycles for cocycles with values in some Lie groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(3):291–300, May 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2000__36_3_291_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300001242.

Devroye:1989:DKM

[Dev89] Luc Devroye. The double kernel method in density estimation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):533–580, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1989__25_4_533_0.

Devulder:2016:PSA

[Dev16] Alexis Devulder. Persistence of some additive functionals of Sinai’s walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1076–1105, August 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1469723512.

Deya:2020:NLF

[Dey20] Aurélien Deya. On a non-linear 2D fractional wave equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):477–501, February 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/On-a-non-linear-2D-fractional-wave-
equation/10.1214/19-AIHP969.full.

Diaconis:1987:DFS

[DF87a] Persi Diaconis and David Freedman. A dozen de Finettistyle results in search of a theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2):397–423, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1987__23_S2_397_0.

Dynkin:1987:SPR

[DF87b] E. B. Dynkin and P. J. Fitzsimmons. Stochastic processes on random domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2):379–396, ???? 1987. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1987_
_23_S2_379_0.

Delattre:2020:PCL

[DF20] Sylvain Delattre and Aurélie Fischer. On principal curves with a length constraint. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2108–2140, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/On-principal-curves-
with-a-length-constraint/10.1214/19-AIHP1030.full.

Dawson:2003:MCB

[DFM^{+}03] Donald A. Dawson, Klaus Fleischmann, Leonid Mytnik, Edwin A. Perkins, and Jie Xiong. Mutually catalytic branching in the plane: uniqueness. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(1):135–191, January/February 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2003__39_1_135_0;   http://www.sciencedirect.
com/science/article/pii/S0246020302000067.

Durmus:2016:SRC

[DFM16] Alain Durmus, Gersende Fort, and Éric Moulines. Subgeometric rates of convergence in Wasserstein distance for Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1799–1822, November 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1479373249.

DelMoral:2001:SIP

[DG01] Pierre Del Moral and Alice Guionnet. On the stability of interacting processes with applications to filtering and genetic algorithms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(2):155–194, March 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_2_155_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010645.

Dareiotis:2019:SED

[DG19] Konstantinos Dareiotis and Benjamin Gess. Supremum estimates for degenerate, quasilinear stochastic partial differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1765–1796, August 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1569398885.

Dawson:2008:RTT

[DGdH^{+}08] D. A. Dawson, A. Greven, F. den Hollander, Rongfeng Sun, and J. M. Swart. The renormalization transformation for two-type branching models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1038–1077, December 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1227287564;   http://www.numdam.org/
item?id=AIHPB_2008__44_6_1038_0.

Dareiotis:2021:PME

[DGG21] K. Dareiotis, M. Gerencsér, and B. Gess. Porous media equations with multiplicative space-time white noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2354–2371, November 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-4/Porous-media-equations-with-multiplicative-
spacetime-white-noise/10.1214/20-AIHP1139.full.

Dweik:2021:SOT

[DGKP21] Samer Dweik, Nassif Ghoussoub, Young-Heon Kim, and Aaron Zeff Palmer. Stochastic optimal transport with free end time. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):700–725, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/

volume-57/issue-2/Stochastic-optimal-transport-with-
free-end-time/10.1214/20-AIHP1092.full.

Douc:2008:BRT

[DGM08] Randal Douc, Arnaud Guillin, and Eric Moulines. Bounds on regeneration times and limit theorems for subgeometric Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2):239–257, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1207948218;   http://
www.numdam.org/item?id=AIHPB_2008__44_2_239_0.

Dedecker:2010:SAS

[DGM10] J. Dedecker, S. Gouëzel, and F. Merlevède. Some almost sure results for unbounded functions of intermittent maps and their associated Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3): 796–821, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1281100399;   http://www.numdam.org/
item?id=AIHPB_2010__46_3_796_0.

Durmus:2021:PDM

[DGM21] Alain Durmus, Arnaud Guillin, and Pierre Monmarché. Piecewise deterministic Markov processes and their invariant measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1442–1475, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Piecewise-deterministic-
Markov-processes-and-their-invariant-measures/10.1214/
20-AIHP1125.full.

Delepouve:2016:UBS

[DGR16] Thibault Delepouve, Razvan Gurau, and Vincent Rivasseau. Universality and Borel summability of arbitrary quartic tensor models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):821–848, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1462367895.

Deuschel:2018:QIP

[DGR18] Jean-Dominique Deuschel, Xiaoqin Guo, and Alejandro F. Ramírez. Quenched invariance principle for random walk in time-dependent balanced random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54 (1):363–384, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1519030832.

Djellout:2006:MDE

[DGW06] H. Djellout, A. Guillin, and L. Wu. Moderate deviations of empirical periodogram and non-linear functionals of moving average processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(4):393–416, July/August 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_4_393_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000877.

Dembo:2003:MDS

[DGZ03] A. Dembo, A. Guionnet, and O. Zeitouni. Moderate deviations for the spectral measure of certain random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):1013–1042, November/December 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2003__39_6_1013_0;   http://www.sciencedirect.
com/science/article/pii/S0246020303000244.

Dereudre:2021:SPT

[DH21] David Dereudre and Pierre Houdebert. Sharp phase transition for the continuum Widom–Rowlinson model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):387–407, February 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-1/Sharp-phase-transition-for-the-continuum-
WidomRowlinson-model/10.1214/20-AIHP1082.full.

denHollander:2014:SRW

[dHdS14] F. den Hollander and R. S. dos Santos. Scaling of a random walk on a supercritical contact process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1276–1300, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1413555500.

Damron:2020:LBF

[DHHX20] Michael Damron, Jack Hanson, Christian Houdré, and Chen Xu. Lower bounds for fluctuations in first-passage percolation for general distributions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1336–1357, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Lower-bounds-for-fluctuations-
in-first-passage-percolation-for-general/10.1214/19-
AIHP1004.full.

Dhariwal:2021:GMS

[DHJ^{+}21] Gaurav Dhariwal, Florian Huber, Ansgar Jüngel, Christian Kuehn, and Alexandra Neamtu. Global martingale solutions for quasilinear SPDEs via the boundedness-by-entropy method. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):577–602, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/Global-martingale-solutions-
for-quasilinear-SPDEs-via-the-boundedness-by/10.1214/
20-AIHP1088.full.

Dembo:2017:TGS

[DHMP17] Amir Dembo, Ruojun Huang, Ben Morris, and Yuval Peres. Transience in growing subgraphs via evolving sets. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3): 1164–1180, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1500624034.

Dhombres:1972:OML

[Dho72] Jean G. Dhombres. Sur les opérateurs multiplicativement liés dans les algèbres de dimension finie. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(4):333–363, October–December 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://

www.numdam.org/item?id=AIHPB_1972__8_4_333_0.

Delbaen:2011:USQ

[DHR11] Freddy Delbaen, Ying Hu, and Adrien Richou. On the uniqueness of solutions to quadratic BSDEs with convex generators and unbounded terminal conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2): 559–574, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887282;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_559_0.

Diaconis:2005:ABE

[Dia05] Persi Diaconis. Analysis of a Bose–Einstein Markov chain. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):409–418, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_3_409_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000142.

Dies:1976:ICF

[Die76] Jacques-Édouard Dies. Information et complexité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(4):365–390, October–December 1976. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1976__
12_4_365_0.

Dies:1978:ICF

[Die78] Jacques-Edouard Dies. Information et complexité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(1):113–118, January– March 1978. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1978__14_1_113_0.

Diebolt:1981:LDM

[Die81] Jean Diebolt. Sur la loi du maximum de certains processus gaussiens sur le tore. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(2):165–179, April–June 1981. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1981__17_2_165_0.

Dies:1982:QPL

[Die82] Jacques-Édouard Dies. Quelques propriétés des librairies d’Aven, Boguslavsky et Kogan. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(2): 115–148, April–June 1982. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1982__
18_2_115_0.

Dimitrov:2018:KAL

[Dim18] Evgeni Dimitrov. KPZ and Airy limits of Hall–Littlewood random plane partitions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):640–693, May 2018. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1524643225.

Dinwoodie:1992:MDT

[Din92] I. H. Dinwoodie. Mesures dominantes et théorème de Sanov. (French) [Dominant measures and theorem of Sanov]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(3): 365–373, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1992__28_3_365_0.

delaPena:2002:SBR

[dIS02] Victor H. de la Peña, Rustam Ibragimov, and Shaturgun Sharakhmetov. On sharp Burkholder–Rosenthal-type inequalities for infinite-degree U-statistics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):973–990, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_973_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302011275.

En l’honneur de J. Bretagnolle, D. Dacunha-Castelle, I. Ibragimov.

Dembo:2019:IAP

[DJO19] Amir Dembo, Milton Jara, and Stefano Olla. The infinite Atlas process: Convergence to equilibrium. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):607–619, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1557820825.

Dani:1979:EIM

[DK79] S. G. Dani and M. Keane. Ergodic invariant measures for actions of SL(2, Z). Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15 (1):79–84, January–March 1979. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1979__15_1_79_0.

Deuschel:2008:QIP

[DK08] Jean-Dominique Deuschel and Holger Kösters. The quenched invariance principle for random walks in random environments admitting a bounded cycle representation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3): 574–591, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1211819425;   http://www.numdam.org/
item?id=AIHPB_2008__44_3_574_0.

Dobler:2019:FMC

[DK19] Christian Döbler and Kai Krokowski. On the fourth moment condition for Rademacher chaos. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):61–97, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1547802395.

Drapeau:2016:DRM

[DKGT16] Samuel Drapeau, Michael Kupper, Emanuela Rosazza Gianin, and Ludovic Tangpi. Dual representation of minimal supersolutions of convex BSDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):868–887, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1462367897.

DelMoral:2016:PGS

[DKP16] P. Del Moral, R. Kohn, and F. Patras. On particle Gibbs samplers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1687–1733, November 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1479373245.

Doss:1978:EUC

[DL78] Halim Doss and Erik Lenglart. Sur l’existence, l’unicité et le comportement asymptotique des solutions d’équations différentielles stochastiques. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(2):189–214, April–June 1978. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1978__14_2_189_0.

Derriennic:1984:CAP

[DL84] Y. Derriennic and M. Lin. Sur le comportement asymptotique des puissances de convolution d’une probabilité. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(2):127–132, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1984__20_2_127_0.

Derrien:1996:TEP

[DL96] Jean-Marc Derrien and Emmanuel Lesigne. Un théorème ergodique polynômial ponctuel pour les endomorphismes exacts et les K-systèmes. (French) [A polynomial-time ergodic theorem for exact endomorphisms and K-systems]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(6): 765–778, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_6_765_0.

Prato:2015:MSR

[DL15] Giuseppe Da Prato and Alessandra Lunardi. Maximal Sobolev regularity in Neumann problems for gradient systems in infinite dimensional domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):1102–1123, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1435759241.

Dolgopyat:2021:GOR

[DLN21] Dmitry Dolgopyat, Marco Lenci, and Péter Nándori. Global observables for RW: Law of large numbers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):94–115, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-

et-statistiques/volume-57/issue-1/Global-observables-
for-RW-Law-of-large-numbers/10.1214/20-AIHP1072.full.

Doukhan:1986:MDQ

[DLP86] P. Doukhan, J. Leon, and F. Portal. Une mesure de la déviation quadratique d’estimateurs non paramétriques. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(1):37–66, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1986__22_1_37_0.

delaPena:1994:BMG

[dlP94] Victor H. de la Peña. A bound on the moment generating function of a sum of dependent variables with an application to simple random sampling without replacement. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2): 197–211, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1995__31_4_703_0. See erratum [de 95].

DeMasi:2007:TSH

[DLP07] Anna De Masi, Stephan Luckhaus, and Errico Presutti. Two scales hydrodynamic limit for a model of malignant tumor cells. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(3):257–297, May/June 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_3_257_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000410.

delaRue:1998:LIT

[dlR98] Thierry de la Rue. L’ergodicité induit un type spectral maximal équivalent à la mesure de Lebesgue. (French) [Ergodicity induces a maximum spectral type equivalent to the Lebesgue measure]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(2):249–263, ???? 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_2_249_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800320.

delaRue:2002:VDP

[dlR02] Thierry de la Rue. Vitesse de dispersion pour une classe de martingales. (French) [Velocity dispersion for a class of martingales]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(4):465–474, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_4_465_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010883.

delaRue:2003:TGP

[dlRd03] Thierry de la Rue and José de Sam Lazaro. Une transformation générique peut être insérée dans un flot. (French) [A generic transformation can be inserted in a flow]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(1): 121–134, January/February 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_1_121_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302000122.

Doukhan:2002:AWE

[DLS02] Paul Doukhan, Gabriel Lang, and Donatas Surgailis. Asymptotics of weighted empirical processes of linear fields with longrange dependence. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):879–896, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_6_879_0;   http://www.sciencedirect.
com/science/article/pii/S0246020302011391.

Du:2020:SHC

[DLZ20] Kai Du, Jiakun Liu, and Fu Zhang. Stochastic Hölder continuity of random fields governed by a system of stochastic PDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1230–1250, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/Stochastic-H%c3%b6lder-continuity-
of-random-fields-governed-by-a-system/10.1214/19-AIHP1000.
full.

Doney:2004:MPT

[DM04] R. A. Doney and R. A. Maller. Moments of passage times for Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):279–297, May/June 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_3_279_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000694.

Dalang:2009:IPH

[DM09] Robert C. Dalang and Carl Mueller. Intermittency properties in a hyperbolic Anderson problem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1150–1164, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1257529897;   http://www.numdam.org/
item?id=AIHPB_2009__45_4_1150_0.

Donati-Martin:2018:CEF

[DMGM18] Catherine Donati-Martin, Benjamin Groux, and Mylène Maïda. Convergence to equilibrium in the free Fokker–Planck equation with a double-well potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):1805–1818, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849784.

Dedecker:2014:QWI

[DMP14] Jérôme Dedecker, Florence Merlevède, and Magda Peligrad. A quenched weak invariance principle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):872–898, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1403277001.

Dedecker:2009:MDS

[DMPU09] Jérôme Dedecker, Florence Merlevède, Magda Peligrad, and Sergey Utev. Moderate deviations for stationary sequences of bounded random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2):453–476, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/

1241024676;   http://www.numdam.org/item?id=AIHPB_2009_
_45_2_453_0.

Doukhan:1994:FCL

[DMR94] Paul Doukhan, Pascal Massart, and Emmanuel Rio. The functional central limit theorem for strongly mixing processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(1):63–82, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1994__30_1_63_0.

Doukhan:1995:IPA

[DMR95] P. Doukhan, P. Massart, and E. Rio. Invariance principles for absolutely regular empirical processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(2):393–427, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1995__31_2_393_0.

Deheuvels:1993:SRI

[DMS93] Paul Deheuvels, David M. Mason, and Galen R. Shorack. Some results on the influence of extremes on the bootstrap. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(1):83–103, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_1_83_0.

Dembo:2005:LDM

[DMS05] Amir Dembo, Peter Mörters, and Scott Sheffield. Large deviations of Markov chains indexed by random trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(6):971–996, November/December 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_6_971_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000937.

Donati-Martin:1994:SSR

[DMSY94] Catherine Donati-Martin, Shiqi Song, and Marc Yor. On symmetric stable random variables and matrix transposition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(3):397–413, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1994__30_3_397_0.

Donati-Martin:1991:FTD

[DMY91] C. Donati-Martin and M. Yor. Fubini’s theorem for double Wiener integrals and the variance of the Brownian path. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(2):181–200, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1991__27_2_181_0.

Doney:2001:PNP

[DN01] R. A. Doney and Y. B. Nakhi. Perturbed and nonperturbed Brownian taboo processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(6):725–736, November/December 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_6_725_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010780.

Dang-Ngoc-Nghiem:1973:CDS

[DNN73] Dang-Ngoc-Nghiem. On the classification of dynamical systems. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(4):397–425, October–December 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_4_397_0.

Deya:2012:MTS

[DNT12] A. Deya, A. Neuenkirch, and S. Tindel. A Milstein-type scheme without Lévy area terms for SDEs driven by fractional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):518–550, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1334148209.

DeMasi:2020:QSL

[DO20] Anna De Masi and Stefano Olla. Quasi-static large deviations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):524–542, February 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/

volume-56/issue-1/Quasi-static-large-deviations/10.
1214/19-AIHP971.full.

Doss:1977:LEE

[Dos77] Halim Doss. Liens entre équations différentielles stochastiques et ordinaires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(2):99–125, April–June 1977. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1977__
13_2_99_0.

Doss:1980:QFA

[Dos80] Halim Doss. Quelques formules asymptotiques pour les petites perturbations de systèmes dynamiques. (French) [Some asymptotic formulas for little perturbations of dynamical systems]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(1):17–28, January– March 1980. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1980__16_1_17_0.

Doss:1991:NPG

[Dos91] H. Doss. Un nouveau principe de grandes déviations en théorie du filtrage non linéaire. (French) [A new large deviation principle in nonlinear filtering theory]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(3): 407–423, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1991__27_3_407_0.

Davis:2018:PSC

[DOS18] Mark Davis, Jan Obłój, and Pietro Siorpaes. Pathwise stochastic calculus with local times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):1–21, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1519030817.

Dozzi:1983:PMP

[Doz83] Markus Dozzi. Propriétés markoviennes de processus sur R^{2}. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(2):209–221, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1983__19_2_209_0.

Dauxois:1975:EAC

[DP75] J. Dauxois and A. Pousse. Une extension de l’analyse canonique. Quelques applications. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(4):355–379, October–December 1975. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1975__11_4_355_0.

DeMasi:1983:PES

[DP83] A. De Masi and E. Presutti. Probability estimates for symmetric simple exclusion random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):71–85, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1983__19_1_71_0.

Deshayes:1984:LAT

[DP84a] Jean Deshayes and Dominique Picard. Lois asymptotiques des tests et estimateurs de rupture dans un modèle statistique classique. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(4):309–327, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1984__20_4_
309_0.

Deshayes:1984:PIP

[DP84b] Jean Deshayes and Dominique Picard. Principe d’invariance sur le processus de vraisemblance. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(1):1–20, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1984__20_1_1_0.

Ding:2011:MTI

[DP11] Jian Ding and Yuval Peres. Mixing time for the Ising model: A uniform lower bound for all graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1020–1028, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1317906499;   http://www.numdam.org/
item?id=AIHPB_2011__47_4_1020_0.

Dembo:2000:TPB

[DPRZ00] Amir Dembo, Yuval Peres, Jay Rosen, and Ofer Zeitouni. Thin points for Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(6):749–774, November 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2000__36_6_749_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300001394.

DeMasi:1989:WAS

[DPS89] A. De Masi, E. Presutti, and E. Scacciatelli. The weakly asymmetric simple exclusion process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(1):1–38, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1989_
_25_1_1_0.

Dong:2016:SFP

[DPSZ16] Zhao Dong, Xuhui Peng, Yulin Song, and Xicheng Zhang. Strong Feller properties for degenerate SDEs with jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):888–897, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1462367898.

Duflo:1969:PAP

[DR69] M. Duflo and D. Revuz. Propriétés asymptotiques des probabilités de transition des processus de Markov récurrents. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(3):233– 244, July–September 1969. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1969__
5_3_233_0.

Duby:1982:ENP

[DR82] Camille Duby and Alain Rouault. Estimation non paramétrique de l’espérance et de la variance de la loi de reproduction d’un processus de ramification. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(2): 149–163, April–June 1982. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1982__
18_2_149_0.

Dedecker:2000:FCL

[DR00] Jérôme Dedecker and Emmanuel Rio. On the functional central limit theorem for stationary processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(1):1–34, January 2000. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_2000__36_1_1_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300001114.

Dedecker:2008:MCL

[DR08] Jérôme Dedecker and Emmanuel Rio. On mean central limit theorems for stationary sequences. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4): 693–726, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1217964116;   http://www.numdam.org/
item?id=AIHPB_2008__44_4_693_0.

deRaphelis:2017:SLM

[dR17a] Loïc de Raphélis. Scaling limit of multitype Galton–Watson trees with infinitely many types. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):200–225, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544889.

deRaynal:2017:SEU

[dR17b] P. E. Chaudru de Raynal. Strong existence and uniqueness for degenerate SDE with Hölder drift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):259–286, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544892.

Dereudre:2017:PDI

[DR17c] David Dereudre and Sylvie Roelly. Path-dependent infinitedimensional SDE with non-regular drift: An existence result. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):641–657, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1491897739.

Dimitrov:2022:GCP

[DR22] Evgeni Dimitrov and Mark Rychnovsky. GUE corners process in boundary-weighed six-vertex models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):188–219, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-58/issue-1/GUE-corners-process-
in-boundary-weighed-six-vertex-models/10.1214/21-AIHP1162.
full.

Drewitz:2014:LPP

[DRS14] Alexander Drewitz, Balázs Ráth, and Artëm Sapozhnikov. Local percolative properties of the vacant set of random interlacements with small intensity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1165–1197, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1413555496.

Dembo:2021:LLC

[DRZ21]  Amir Dembo, Jay Rosen, and Ofer Zeitouni.  Limit

law for the cover time of a random walk on a binary tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):830–855, May 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/Limit-law-for-the-cover-time-of-a-
random-walk/10.1214/20-AIHP1098.full.

Dunau:1984:QFC

[DS84] Jean-Louis Dunau and Henri Sénateur. Quelles fonctions changent toute loi uniforme en une loi uniforme?. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(3):247–250, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1984__20_3_247_0.

Deheuvels:1990:SPB

[DS90] Paul Deheuvels and Josef Steinebach. On the sample path behavior of the first passage time process of a Brownian motion with drift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):145–179, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1990__26_1_145_0.

Delbaen:1996:ACP

[DS96] F. Delbaen and W. Schachermayer. Attainable claims with p’th moments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(6):743–763, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1996__32_6_743_0.

Delbaen:1997:BSW

[DS97] Freddy Delbaen and Walter Schachermayer. The Banach space of workable contingent claims in arbitrage theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33 (1):113–144, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1997__33_1_113_0;   http://
www.sciencedirect.com/science/article/pii/S0246020397801185.

deSaporta:2003:RTS

[dS03] Benoîte de Saporta. Renewal theorem for a system of renewal equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(5):823–838, September/October 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_5_823_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000189.

Decreusefond:2006:ACR

[DS06] L. Decreusefond and N. Savy. Anticipative calculus with respect to filtered Poisson processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(3):343–372, May/June 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2006__42_3_343_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000865.

Danilenko:2007:MRO

[DS07] Alexandre I. Danilenko and Cesar E. Silva. Mixing rankone actions of locally compact Abelian groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43 (4):375–398, July/August 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_4_375_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000677.

Dereich:2013:CQA

[DSS13] Steffen Dereich, Michael Scheutzow, and Reik Schottstedt. Constructive quantization: Approximation by empirical measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1183–1203, November 2013. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1380718743.

Duflo:1990:LGN

[DST90] M. Duflo, R. Senoussi, and A. Touati. Sur la loi des grands nombres pour les martingales vectorielles et l’estimateur des moindres carrés d’un modèle de régression. (French) [On the law of large numbers for vector martingales and least-squares estimator of a regression model]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(4):549–566, ???? 1990. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1990_
_26_4_549_0.

Duflo:1991:PAP

[DST91] M. Duflo, R. Senoussi, and A. Touati. Propriétés asymptotiques presque sûres de l’estimateur des moindres carrés d’un modèle autorégressif vectoriel. (French) [Almost sure asymptotic properties of the least squares estimator of a vector autoregression model]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(1):1–25, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1991__27_1_1_0.

Durot:2003:DBE

[DT03] Cécile Durot and Anne-Sophie Tocquet. On the distance between the empirical process and its concave majorant in a monotone regression framework. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(2):217–240, March/ April 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2003__39_2_217_0;   http://www.sciencedirect.
com/science/article/pii/S0246020302000134.

deTiliere:2007:SLI

[dT07] Béatrice de Tilière. Scaling limit of isoradial dimer models and the case of triangular quadri-tilings. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6): 729–750, November/December 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_6_729_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000859.

deTiliere:2016:CIM

[dT16] Béatrice de Tilière. Critical Ising model and spanning trees partition functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1382–1405, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1469723524.

Dembin:2020:SMC

[DT20] Barbara Dembin and Marie Théret. Size of a minimal cutset in supercritical first passage percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1419–1439, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-2/Size-of-a-minimal-
cutset-in-supercritical-first-passage-percolation/10.
1214/19-AIHP1008.full.

Das:2021:FMS

[DT21] Sayan Das and Li-Cheng Tsai. Fractional moments of the stochastic heat equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):778–799, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-2/Fractional-moments-of-
the-stochastic-heat-equation/10.1214/20-AIHP1095.full.

Dellacherie:1998:NPA

[DTMS98] C. Dellacherie, D. Taïbi, S. Martinez, and J. San Martin. Noyaux potentiels associés à une filtration. (French) [Potential kernels associated with filtration]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(6):707–725, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1998__34_6_707_0;   http://www.sciencedirect.
com/science/article/pii/S0246020399800016.

Dubuc:1975:MSP

[Dub75] Serge Dubuc. Méthodes stochastiques pour la détermination de polynômes de Bernstein. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(3):203–223, July–September 1975. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1975__11_3_203_0.

Dubedat:2004:RPB

[Dub04] Julien Dubédat. Reflected planar Brownian motions, intertwining relations and crossing probabilities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(5): 539–552, September/October 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_5_539_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000275.

Duchamps:2020:TWT

[Duc20] Jean-Jil Duchamps. Trees within trees II: Nested fragmentations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1203–1229, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/Trees-within-trees-II-Nested-fragmentations/
10.1214/19-AIHP999.full.

Duheille:1997:PPE

[Duh97] Frédérique Duheille. Une preuve probabiliste élémentaire d’un résultat de P. Baird et J. C. Wood. (French) [An elementary probabilistic proof of a result of P. Baird and J. C. Wood]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(2):283–291, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_2_283_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801252.

Duquesne:2003:PDR

[Duq03] Thomas Duquesne. Path decompositions for real Levy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(2):339–370, March/April 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_2_339_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000043.

Davydov:1998:RCM

[DV98] Yu. Davydov and A. M. Vershik. Réarrangements convexes des marches aléatoires. (French) [Convex rearrangements of random walks]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(1):73–95, ???? 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_1_73_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800198.

Delmas:2005:NLN

[DV05] Jean-François Delmas and Pascal Vogt. Non-linear Neumann’s condition for the heat equation: a probabilistic representation using catalytic super-Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5): 817–849, September/October 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_5_817_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000464.

Durieu:2008:CBC

[DV08] Olivier Durieu and Dalibor Volný. Comparison between criteria leading to the weak invariance principle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2): 324–340, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1207948222;   http://www.numdam.org/
item?id=AIHPB_2008__44_2_324_0.

Dumaz:2013:RTE

[DV13] Laure Dumaz and Bálint Virág. The right tail exponent of the Tracy–Widom β distribution. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):915–933, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1380718731.

Deijfen:2021:SM

[DV21] Maria Deijfen and Timo Vilkas. The Schelling model on Z. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):800–814, May 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/The-Schelling-model-on-Z/10.1214/
20-AIHP1096.full.

Deijfen:2013:SFP

[DvdHH13] Maria Deijfen, Remco van der Hofstad, and Gerard Hooghiemstra. Scale-free percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):817–838, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1372772645.

Dhara:2020:HTC

[DvdHvLS20] Souvik Dhara, Remco van der Hofstad, Johan S. H. van Leeuwaarden, and Sanchayan Sen. Heavy-tailed configuration models at criticality. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1515–1558, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Heavy-tailed-configuration-
models-at-criticality/10.1214/19-AIHP980.full.

Delgado-Vences:2020:CLT

[DVNZ20] Francisco Delgado-Vences, David Nualart, and Guangqu Zheng. A central limit theorem for the stochastic wave equation with fractional noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):3020–3042, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/A-Central-Limit-Theorem-
for-the-stochastic-wave-equation-with/10.1214/20-AIHP1069.
full.

Dawson:2000:SPD

[DVW00] D. A. Dawson, J. Vaillancourt, and H. Wang. Stochastic partial differential equations for a class of interacting measurevalued diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(2):167–180, March 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_2_167_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001217.

Dieker:2008:DTK

[DW08] A. B. Dieker and J. Warren. Determinantal transition kernels for some interacting particles on the line. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1162–1172, December 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1227287569;   http://www.numdam.org/
item?id=AIHPB_2008__44_6_1162_0.

Djellout:2011:LNE

[DW11] Hacene Djellout and Liming Wu. Lipschitzian norm estimate of one-dimensional Poisson equations and applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2):450–465, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887277;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_450_0.

Denisov:2015:ETI

[DW15] Denis Denisov and Vitali Wachtel. Exit times for integrated random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):167–193, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1421244402.

Duquesne:2017:DLT

[DW17] Thomas Duquesne and Minmin Wang. Decomposition of Lévy trees along their diameter. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):539–593, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1491897736.

Dalalyan:2011:SOA

[DY11] Arnak Dalalyan and Nakahiro Yoshida. Second-order asymptotic expansion for a non-synchronous covariation estimator. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3):748–789, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1308834858;   http:/
/www.numdam.org/item?id=AIHPB_2011__47_3_748_0.

deZelicourt:1981:MMP

[dZ81] Catherine de Zelicourt. Une méthode de martingales pour la convergence d’une suite de processus de sauts markoviens vers une diffusion associée à une condition frontière. Application aux systèmes de files d’attente. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(4):351–376, October–December 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_4_351_0.

Doney:2005:PSE

[DZ05] R. A. Doney and T. Zhang. Perturbed Skorohod equations and perturbed reflected diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(1): 107–121, January/February 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_1_107_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000548.

Edgar:1979:US

[Edg79] G. A. Edgar. Uniform semiamarts. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(3):197–203, July–September 1979. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_3_197_0.

Erhard:2014:PAM

[EdHM14] D. Erhard, F. den Hollander, and G. Maillard. The parabolic Anderson model in a dynamic random environment: Basic properties of the quenched Lyapunov exponent. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (4):1231–1275, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555499.

ElKaroui:2005:NLR

[EF05] Nicole El Karoui and Hans Föllmer. A non-linear Riesz respresentation in probabilistic potential theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):269–283, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_3_269_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000154.

Ellis:2016:BWT

[EF16] Tom Ellis and Ohad N. Feldheim. The Brownian web is a two-dimensional black noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):162–172, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089265.

Erhard:2020:NEF

[EFG^{+}20] D. Erhard, T. Franco, P. Gonçalves, A. Neumann, and M. Tavares. Non-equilibrium fluctuations for the SSEP with a slow bond. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1099–1128, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Non-equilibrium-fluctuations-
for-the-SSEP-with-a-slow-bond/10.1214/19-AIHP995.full.

Erlihson:2008:LSG

[EG08] Michael M. Erlihson and Boris L. Granovsky. Limit shapes of Gibbs distributions on the set of integer partitions: The expansive case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5):915–945, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1222261918;   http://
www.numdam.org/item?id=AIHPB_2008__44_5_915_0.

Egghe:1980:SNC

[Egg80] L. Egghe. Some new Chacon–Edgar-type inequalities for stochastic processes, and characterizations of Vitali-conditions. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(4):327–337, October–December 1980. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1980__
16_4_327_0.

Eisenbaum:2003:FIR

[EH03] Nathalie Eisenbaum and Yueyun Hu. A family of integral representations for the Brownian variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):1083–1096, November/December 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_6_1083_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020303000293.

Erhard:2019:DRS

[EH19] Dirk Erhard and Martin Hairer. Discretisation of regularity structures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2209–2248, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/Discretisation-of-regularity-
structures/10.1214/18-AIHP947.full.

Eaton:2007:PSA

[EHJ07] Morris L. Eaton, James P. Hobert, and Galin L. Jones. On perturbations of strongly admissible prior distributions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(5):633–653, September/October 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_5_633_0;   http://www.sciencedirect.com/science/
article/pii/S0246020307000064.

Englander:2010:SLL

[EHK10] János Engländer, Simon C. Harris, and Andreas E. Kyprianou. Strong Law of Large Numbers for branching diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1): 279–298, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1267454117;   http://www.numdam.org/
item?id=AIHPB_2010__46_1_279_0.

Ehrhard:1986:EEP

[Ehr86] Antoine Ehrhard. Éléments extrémaux pour les inégalités de Brunn–Minkowski gaussiennes. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2): 149–168, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1986__22_2_149_0.

Eichelsbacher:2001:MDF

[Eic01] Peter Eichelsbacher. Moderate deviations for functional U-processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(2):245–273, March 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_2_245_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010633.

ElAbdalaoui:2000:CGO

[El 00] El Houcein El Abdalaoui. A class of generalized Ornstein transformations with the weak mixing property. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36 (6):775–786, November 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2000__36_6_775_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300010566.

Eldan:2014:EPH

[Eld14] Ronen Eldan. Extremal points of high-dimensional random walks and mixing times of a Brownian motion on the sphere. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):95–110, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1388545267.

Eldan:2016:SES

[Eld16] Ronen Eldan. Skorokhod embeddings via stochastic flows on the space of Gaussian measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1259–1280, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723520.

Enriquez:2010:ALL

[ELS10] N. Enriquez, C. Lucas, and F. Simenhaus. The arcsine law as the limit of the internal DLA cluster generated by Sinai’s walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):991–1000, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1288878334;   http://
www.numdam.org/item?id=AIHPB_2010__46_4_991_0.

Eldan:2020:SLS

[ELS20] Ronen Eldan, Joseph Lehec, and Yair Shenfeld. Stability of the logarithmic Sobolev inequality via the Föllmer process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2253–2269, August 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/Stability-of-the-logarithmic-Sobolev-
inequality-via-the-F%c3%b6llmer-process/10.1214/19-
AIHP1038.full.

Eweda:1983:QMA

[EM83] E. Eweda and O. Macchi. Quadratic mean and almostsure convergence of unbounded stochastic approximation algorithms with correlated observations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(3):235–255, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1983__19_3_235_0.

Eichelsbacher:2015:RCC

[EM15] Peter Eichelsbacher and Bastian Martschink. On rates of convergence in the Curie–Weiss–Potts model with an external field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):252–282, February 2015. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1421244405.

Emery:1985:DFB

[Eme85] M. Emery. Une définition faible de BMO. (French) [A weak definition of BMO]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(1):59–71, ???? 1985. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1985__21_1_59_0.

Emery:2005:CAB

[Éme05] M. Émery. On certain almost Brownian filtrations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41 (3):285–305, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_3_285_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000166.

Emilion:1981:PAP

[Émi81a] R. Émilion. Processus additifs positifs dans L_{∞}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(2):185–189, April– June 1981. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1981__17_2_185_0.

Emilion:1981:TEL

[Émi81b] R. Émilion. Théorème ergodique local dans L_{p} (1 < p < ∞). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(2): 181–184, April–June 1981. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1981__
17_2_181_0.

Emilion:1985:CZS

[Émi85] R. Émilion. Continuity at zero of semi-groups on L_{1} and differentiation of additive processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(4):305–312, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1985__21_4_305_0.

Emilion:1986:ASL

[Émi86] R. Émilion. Additive and superadditive local theorems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(1): 19–36, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1986__22_1_19_0.

Eskenazis:2020:CBL

[EN20] Alexandros Eskenazis and Evita Nestoridi. Cutoff for the Bernoulli–Laplace urn model with o(n) swaps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2621–2639, November 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-4/Cutoff-for-the-BernoulliLaplace-urn-
model-with-on-swaps/10.1214/20-AIHP1052.full.

Englander:2008:QLL

[Eng08] János Engländer. Quenched law of large numbers for branching Brownian motion in a random medium. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3): 490–518, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1211819422;   http://www.numdam.org/
item?id=AIHPB_2008__44_3_490_0.

Englander:2009:LLN

[Eng09] János Engländer. Law of large numbers for superdiffusions: The non-ergodic case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):1–6, February 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1234469969;   http://www.numdam.org/item?id=AIHPB_2009_
_45_1_1_0.

Enriquez:1996:EBB

[Enr96] Nathanaël Enriquez. Effet d’un bruit blanc sur l’oscillateur harmonique de dimension d. (French) [Effect of white noise on the d-dimensional harmonic oscillator]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(5):601–622, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1996__32_5_601_0.

Enriquez:2007:IPA

[Enr07] Nathanaël Enriquez. An invariance principle for Azéma martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6):717–727, November/December 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_6_717_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000835.

Evans:2022:ELM

[EO22] Steven N. Evans and Mehdi Ouaki. Excursions away from the Lipschitz minorant of a Lévy process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):424–454, February 2022. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-58/issue-1/Excursions-away-from-the-Lipschitz-
minorant-of-a-L%c3%a9vy-process/10.1214/21-AIHP1155.
full.

Evans:1998:CMC

[EP98] Steven N. Evans and Jim Pitman. Construction of Markovian coalescents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(3):339–383, May/June 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_3_339_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800150.

Englander:2006:CSP

[EP06] János Engländer and Ross G. Pinsky. The compact support property for measure-valued processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5): 535–552, September/October 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_5_535_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000932.

ElAbdalaoui:2006:NCO

[EPP06] E. H. El Abdalaoui, F. Parreau, and A. A. Prikhod’ko. A new class of Ornstein transformations with singular spectrum. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6):671–681, November/December 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_6_671_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000033.

Evans:2017:SBM

[EPT17] Steven Evans, Jim Pitman, and Wenpin Tang. The spans in Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1108–1135, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1500624032.

Ehrhardt:2013:PTO

[ER13] Torsten Ehrhardt and Brian Rider. Perturbed Toeplitz operators and radial determinantal processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):934–960, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1380718732.

Erbar:2010:HEM

[Erb10] Matthias Erbar. The heat equation on manifolds as a gradient flow in the Wasserstein space. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1): 1–23, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1267454105;   http://www.numdam.org/
item?id=AIHPB_2010__46_1_1_0.

Erbar:2014:GFE

[Erb14] Matthias Erbar. Gradient flows of the entropy for jump processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):920–945, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1403277003.

ElKaroui:1971:PMR

[ERR71] Nicole El Karoui, Hervé Reinhard, and Bernard Roynette. Processus de Markov: retournement des trajectoires à un temps d’entrée dans un presque-borélien. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(2):131–144, April–June 1971. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1971__7_2_131_0.

Eichelsbacher:2015:MDS

[ERS15] P. Eichelsbacher, M. Raič, and T. Schreiber. Moderate deviations for stabilizing functionals in geometric probability.

Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):89–128, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1421244400.

Englander:2016:WEV

[ERS16] János Engländer, Yan-Xia Ren, and Renming Song. Weak extinction versus global exponential growth of total mass for superdiffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):448–482, February 2016. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1452089276.

Ellis:1979:CSS

[ES79] Martin H. Ellis and J. Michael Steele. Catching small sets under flows. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1): 33–40, January–March 1979. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1979__
15_1_33_0.

Eichelsbacher:2002:LDE

[ES02]  Peter Eichelsbacher and Uwe Schmock.  Large devia-

tions of U-empirical measures in strong topologies and applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(5):779–797, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_5_779_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011160.

Erdos:2017:URM

[ES17] László Erdős and Kevin Schnelli. Universality for random matrix flows with time-dependent density. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1606–1656, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773721.

Estrade:1992:ESI

[Est92] Anne Estrade. Exponentielle stochastique et intégrale multiplicative discontinues. (French) [Stochastic exponential and multiplicative discontinuous integral]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(1):107–129, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1992__28_1_107_0.

Erdos:2012:LRF

[ESYY12] László Erdős, Benjamin Schlein, Horng-Tzer Yau, and Jun Yin. The local relaxation flow approach to universality of the local statistics for random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):1–46, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1327328013.

Elliott:1990:TRN

[ET90] Robert J. Elliott and Allanus H. Tsoi. Time reversal of non-Markov point processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(2):357–373, ???? 1990. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1990_
_26_2_357_0.

ElMachkouri:2003:CET

[EV03] Mohamed El Machkouri and Dalibor Volný. Contre-exemple dans le théorème central limite fonctionnel pour les champs aléatoires réels. (French) [Counterexample of the central limit theorem for real functional random fields]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39 (2):325–337, March/April 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_2_325_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302000110.

Evans:1987:PTF

[Eva87] Steven N. Evans. Potential theory for a family of several Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(3):499–530, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_3_499_0.

Evans:1991:TMV

[Eva91] Steven N. Evans. Trapping a measure-valued Markov branching process conditioned on non-extinction. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(2):215–220, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_2_215_0.

Evans:1994:MRS

[Eva94] Steven N. Evans. Multiplicities of a random sausage. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(3): 501–518, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1994__30_3_501_0.

Evans:1997:CML

[Eva97] Steven N. Evans. Coalescing Markov labelled partitions and a continuous sites genetics model with infinitely many types. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(3):339–358, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1997__33_3_339_0;   http://
www.sciencedirect.com/science/article/pii/S0246020397800957.

Englander:2006:LLN

[EW06] János Engländer and Anita Winter. Law of large numbers for a class of superdiffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(2):171–185, March/April 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_2_171_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000579.

Eberle:2019:SCM

[EZ19] Andreas Eberle and Raphael Zimmer. Sticky couplings of multidimensional diffusions with different drifts. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2370–2394, November 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-55/issue-4/Sticky-couplings-of-multidimensional-
diffusions-with-different-drifts/10.1214/18-AIHP951.
full.

Fehrman:2019:ETS

[Feh19] Benjamin Fehrman. On the exit time and stochastic homogenization of isotropic diffusions in large domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):720–755, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820829.

Ferenczi:1984:SLR

[Fer84a] S. Ferenczi. Systèmes localement de rang un. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(1):35–51, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1984__20_1_35_0.

Fernique:1984:CMG

[Fer84b] Xavier Fernique. Comparaison de mesures gaussiennes et de mesures produit. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(2):165–175, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1984__20_2_165_0.

Ferenczi:1985:SRG

[Fer85] S. Ferenczi. Systèmes de rang un gauche. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(2):177–186, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1985__21_2_177_0.

Fernique:1990:RCF

[Fer90] X. Fernique. Sur la régularité de certaines fonctions aléatoires d’Ornstein–Uhlenbeck. (French) [On the regularity of some Ornstein–Uhlenbeck random functions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):399–417, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_3_399_0.

Ferre:2015:PFO

[Fer15] D. Ferré. Parametric first-order Edgeworth expansion for Markov additive functionals. Application to M-estimations.

Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):781–808, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1428672691.

Ferreira:2018:EEI

[FF18] Helena Ferreira and Marta Ferreira. Estimating the extremal index through local dependence. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):587–605, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1524643223.

Favaro:2019:BNA

[FFJ19] Stefano Favaro, Shui Feng, and Paul A. Jenkins. Bayesian nonparametric analysis of Kingman’s coalescent. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1087–1115, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820843.

Ferrari:2000:ASE

[FFV00] P. A. Ferrari, L. R. G. Fontes, and M. E. Vares. The asymmetric simple exclusion model with multiple shocks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(2):109–126, March 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2000__36_2_109_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300001187.

Freitas:2017:EVL

[FFV17] Ana Cristina Moreira Freitas, Jorge Milhazes Freitas, and Sandro Vaienti. Extreme value laws for non stationary processes generated by sequential and random dynamical systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1341–1370, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1500624044.

Ferrari:2005:TDP

[FFW05] P. A. Ferrari, L. R. G. Fontes, and Xian-Yuan Wu. Two-dimensional Poisson trees converge to the Brownian web. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5):851–858, September/October 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_5_851_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000883.

Foguel:1979:OMB

[FG79] Shaul R. Foguel and Nassif A. Ghoussoub. Ornstein–Metivier– Brunel theorem revisited. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(3):293–301, July–September 1979. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_3_293_0.

Fitzsimmons:1992:LTV

[FG92] P. J. Fitzsimmons and R. K. Getoor. Limit theorems and variation properties for fractional derivatives of the local time of a stable process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2):311–333, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1992__28_2_
311_0.

Faggionato:2012:RFL

[FG12] Alessandra Faggionato and Davide Gabrielli. A representation formula for large deviations rate functionals of invariant measures on the one dimensional torus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):212–234, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1327328020.

Friedlander:2016:ILS

[FGHV16] Susan Friedlander, Nathan Glatt-Holtz, and Vlad Vicol. Inviscid limits for a stochastically forced shell model of turbulent flow. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1217–1247, August 2016. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1469723518.

Fontbona:2017:VAS

[FGJ17] Joaquin Fontbona, Nathael Gozlan, and Jean-François Jabir. A variational approach to some transport inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1719–1746, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1511773724.

Ferrari:1995:EWT

[FGL95] P. A. Ferrari, A. Galves, and T. M. Liggett. Exponential waiting time for filling a large interval in the symmetric simple exclusion process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):155–175, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1995__31_1_155_0.

Ferrari:2009:CPR

[FGM09] Pablo A. Ferrari, Patricia Gonçalves, and James B. Martin. Collision probabilities in the rarefaction fan of asymmetric exclusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1048–1064, November 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1257529891;   http://www.numdam.org/item?id=AIHPB_2009_
_45_4_1048_0.

Franco:2013:HBS

[FGN13] Tertuliano Franco, Patrícia Gonçalves, and Adriana Neumann. Hydrodynamical behavior of symmetric exclusion with slow bonds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):402–427, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1366117652.

Ferrari:2019:LLS

[FGN19] P. L. Ferrari, P. Ghosal, and P. Nejjar. Limit law of a second class particle in TASEP with non-random initial condition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1203–1225, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1569398867.

Flandoli:2009:RSC

[FGR09] Franco Flandoli, Massimiliano Gubinelli, and Francesco Russo. On the regularity of stochastic currents, fractional Brownian motion and applications to a turbulence model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2):

545–576, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024680;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_545_0.

Faggionato:2018:VMV

[FGS18] Alessandra Faggionato, Nina Gantert, and Michele Salvi. The velocity of 1 d Mott variable-range hopping with external field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1165–1203, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1531296017.

Faggionato:2019:ERL

[FGS19] Alessandra Faggionato, Nina Gantert, and Michele Salvi. Einstein relation and linear response in one-dimensional Mott variable-range hopping. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1477–1508, August 2019. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1569398876.

Fattler:2016:CAS

[FGV16] Torben Fattler, Martin Grothaus, and Robert Voßhall. Construction and analysis of a sticky reflected distorted Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):735–762, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1462367892.

Ferre:2012:LTS

[FHL12] Déborah Ferré, Loïc Hervé, and James Ledoux. Limit theorems for stationary Markov processes with L^{2}-spectral gap. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):396–423, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1334148205.

Flegel:2019:HTR

[FHS19] Franziska Flegel, Martin Heida, and Martin Slowik. Homogenization theory for the random conductance model with degenerate ergodic weights and unbounded-range jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1226–1257, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1569398868.

Funaki:2021:APR

[FHSX21] Tadahisa Funaki, Masato Hoshino, Sunder Sethuraman, and Bin Xie. Asymptotics of PDE in random environment by paracontrolled calculus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1702–1735, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Asymptotics-of-PDE-in-
random-environment-by-paracontrolled-calculus/10.1214/
20-AIHP1129.full.

Frangos:1988:SIS

[FI88] Nikos E. Frangos and Peter Imkeller. Some inequalities for strong martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(3):395–402, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1988__24_3_
395_0.

Follmer:1993:ACG

[FI93] Hans Föllmer and Peter Imkeller. Anticipation cancelled by a Girsanov transformation: a paradox on Wiener space. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(4):569–586, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_4_569_0.

Fontes:2006:CNM

[FINR06] L. R. G. Fontes, M. Isopi, C. M. Newman, and K. Ravishankar. Coarsening, nucleation, and the marked Brownian web. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1):37–60, January/February 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_1_37_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000592.

Fischler:1973:QTL

[Fis73] Roger Fischler. Quelques théorèmes limites du calcul des probabilités dont la valeur limite dépend d’une variable aléatoire. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(4):345– 349, October–December 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_4_345_0.

Fischler:1976:CFA

[Fis76] Roger Fischler. Convergence faible avec indices aléatoires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(4): 391–399, October–December 1976. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1976__12_4_391_0.

Fischer:2001:CML

[Fis01] Torsten Fischer. Coupled map lattices with asynchronous updatings. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(4):421–479, July/August 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_4_421_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010809.

Fitzsimmons:1998:QSR

[Fit98] P. J. Fitzsimmons. The quasi-sure ratio ergodic theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(3):385–405, May/June 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_3_385_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800162.

Friesen:2021:EDS

[FJR21] Martin Friesen, Peng Jin, and Barbara Rüdiger. Existence of densities for stochastic differential equations driven by Lévy processes with anisotropic jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):250–271, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-

statistiques/volume-57/issue-1/Existence-of-densities-
for-stochastic-differential-equations-driven-by-L%c3%
a9vy/10.1214/20-AIHP1077.full.

Fortet:1969:RPM

[FK69] R. Fortet and M. Kambouzia. Les répartitions ponctuelles markoviennes; estimations et tests concernant leurs lois de probabilité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(3):195–212, July–September 1969. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1969__5_3_195_0.

Fortet:1975:EAR

[FK75] R. Fortet and M. Kambouzia. Ensembles aléatoires, répartitions ponctuelles aléatoires, problèmes de recouvrement. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(4):299–316, October–December 1975. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1975__
11_4_299_0.

Fan:1993:RIR

[FK93] Ai Hua Fan and Jean-Pierre Kahane. Rareté des intervalles recouvrant un point dans un recouvrement aléatoire. (French) [Scarcity of intervals covering a point in a random collection]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(3):453–466, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1993__29_3_453_0.

Fleischmann:1994:LDP

[FK94] Klaus Fleischmann and Ingemar Kaj. Large deviation probabilities for some rescaled superprocesses. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(4):607–645, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1994__30_4_607_0.

Ferrari:1995:SCP

[FK95] P. A. Ferrari and C. Kipnis. Second class particles in the rarefaction fan. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):143–154, ???? 1995. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1995__31_1_143_0.

Foondun:2010:GMS

[FK10] Mohammud Foondun and Davar Khoshnevisan. On the global maximum of the solution to a stochastic heat equation with compact-support initial data. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):895–907, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878328;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_895_0.

Fournier:2016:TMI

[FL16] Nicolas Fournier and Eva Löcherbach. On a toy model of interacting neurons. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1844–1876, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1479373251.

Frikha:2020:WUD

[FL20] Noufel Frikha and Libo Li. Weak uniqueness and density estimates for SDEs with coefficients depending on some path-functionals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1002–1040, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Weak-uniqueness-and-
density-estimates-for-SDEs-with-coefficients-depending/
10.1214/19-AIHP992.full.

Fernando:2021:EEW

[FL21] Kasun Fernando and Carlangelo Liverani. Edgeworth expansions for weakly dependent random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):469–505, February 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-1/Edgeworth-expansions-for-weakly-dependent-
random-variables/10.1214/20-AIHP1085.full.

Flandoli:2002:PDS

[Fla02] Franco Flandoli. On a probabilistic description of small scale structures in 3d fluids. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(2):207–228, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_2_207_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010925.

Fleury:2010:BPC

[Fle10] B. Fleury. Between Paouris concentration inequality and variance conjecture. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):299–312, May 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1273584125;   http://www.numdam.org/item?id=AIHPB_2010_
_46_2_299_0.

Fliess:1975:SRP

[Fli75] Michel Fliess. Séries rationnelles positives et processus stochastiques. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11 (1):1–21, January–March 1975. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1975__
11_1_1_0.

Flipo:1981:CDS

[Fli81] Daniel Flipo. Comparaison des disciplines de service des files d’attente G/G/1. (French) [Comparison of queue service disciplines G/G/1]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(2): 191–212, April–June 1981. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1981__
17_2_191_0.

Franchi:1984:TIP

[FlJ84] Jacques Franchi and Yves le Jan. Sur les trajectoires intrinsèques des processus de Markov et le théorème de Shih. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(2):103–126, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1984__20_2_103_0.

Florens:1982:EBI

[Flo82] J. P. Florens. Expériences bayésiennes invariantes. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(4):305–317, October–December 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1982__18_4_305_0.

Fouks:2006:EAM

[FLP06] Jean-Denis Fouks, Emmanuel Lesigne, and Marc Peigné. Étude asymptotique d’une marche aléatoire centrifuge. (French) [Asymptotic study of a centrifugal random walk]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42 (2):147–170, March/April 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_2_147_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000476.

Fromont:2011:ATH

[FLRB11] M. Fromont, B. Laurent, and P. Reynaud-Bouret. Adaptive tests of homogeneity for a Poisson process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1): 176–213, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170235;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_176_0.

Friz:2006:LAU

[FLS06] P. Friz, T. Lyons, and D. Stroock. Lévy’s area under conditioning. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1):89–101, January/February 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_1_89_0;   http://www.sciencedirect.com/science/
article/pii/S024602030500066X.

Ferrari:2004:PTS

[FLT04] P. A. Ferrari, C. Landim, and H. Thorisson. Poisson trees, succession lines and coalescing random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2): 141–152, March/April 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_2_141_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000669.

Ferrari:2009:MHA

[FM09] Pablo A. Ferrari and James B. Martin. Multiclass Hammersley– Aldous–Diaconis process and multiclass-customer queues. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):250–265, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1234469981;   http:/
/www.numdam.org/item?id=AIHPB_2009__45_1_250_0.

Foucart:2019:CSB

[FM19] Clément Foucart and Chunhua Ma. Continuous-state branching processes, extremal processes and super-individuals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1061–1086, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820842.

Fayolle:1992:RWQ

[FMM92] G. Fayolle, V. A. Malyshev, and M. V. Menshikov. Random walks in a quarter plane with zero drifts. I. Ergodicity and null recurrence. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2):179–194, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1992__28_2_179_0.

Figalli:2014:GAC

[FMP14] A. Figalli, F. Maggi, and A. Pratelli. A geometric approach to correlation inequalities in the plane. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):1–14, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1388545262.

Fontes:2000:LVD

[FNS00] Luiz Renato G. Fontes, Eduardo Jordão Neves, and Vladas Sidoravicius. Limit velocity for a driven particle in a random medium with mass aggregation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(6):787–805, November 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2000__36_6_787_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300010591.

Foguel:1975:ICN

[Fog75] S. R. Foguel. Iterates of a convolution on a non abelian group. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(2):199–202, April– June 1975. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1975__11_2_199_0.

Fortet:1968:RPA

[For68] R. Fortet. Sur les répartitions ponctuelles aléatoires, en particulier de Poisson. (French) [On random point distributions, in particular Poisson distributions]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(2):99–142, April–June 1968. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1968__4_2_99_0.

Fortet:1973:ENR

[For73] R. Fortet. Espaces à noyau reproduisant et lois de probabilités des fonctions aléatoires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(1):41–58, January–March 1973. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1973__9_1_41_0.

Fortet:1978:CDQ

[For78] R. Fortet. Condition de Doeblin et quasi-compacité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4):379–390, October–December 1978. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1978__
14_4_379_0.

Fouque:1984:CLP

[Fou84] Jean-Pierre Fouque. La convergence en loi pour les processus à valeurs dans un espace nucléaire. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(3): 225–245, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1984__20_3_225_0.

Fourati:1995:RMP

[Fou95] S. Fourati. Représentation des mesures par des fonctionnelles additives entre deux temps. (French) [Representation of measures by functionals additive between two times]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(3): 527–544, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1995__31_3_527_0.

Fougeres:2000:HIG

[Fou00] Pierre Fougères. Hypercontractivité et isopérimétrie gaussienne. Applications aux systèmes de spins. (French) [Hypercontractivity and Gaussian isoperimetry. Applications to spin systems]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(5):647–689, September 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_5_647_0;   http://www.sciencedirect.com/science/
article/pii/S024602030000145X.

Fournier:2001:SPS

[Fou01] Nicolas Fournier. Strict positivity of the solution to a 2dimensional spatially homogeneous Boltzmann equation without cutoff. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(4):481–502, July/August 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_4_481_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010670.

Fourati:2005:VLF

[Fou05] Sonia Fourati. Vervaat et Lévy. (French) [Vervaat and Lévy]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):461–478, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_3_461_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000178.

Fournier:2013:PUS

[Fou13] Nicolas Fournier. On pathwise uniqueness for stochastic differential equations driven by stable Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1): 138–159, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1359470129.

Frick:2008:RPU

[FP08] Sarah Bailey Frick and Karl Petersen. Random permutations and unique fully supported ergodicity for the Euler adic transformation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5):876–885, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1222261916;   http://
www.numdam.org/item?id=AIHPB_2008__44_5_876_0.

Fontbona:2017:RCE

[FP17] Joaquin Fontbona and Fabien Panloup. Rate of convergence to equilibrium of fractional driven stochastic differential equations with some multiplicative noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):503–538, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1491897735.

Feldheim:2018:RCD

[FP18a] Ohad Noy Feldheim and Ron Peled. Rigidity of 3-colorings of the discrete torus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):952–994, May 2018. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1524643236.

Fribergh:2018:BRW

[FP18b] Alexander Fribergh and Serguei Popov. Biased random walks on the interlacement set. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1341–1358, August 2018. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1531296022.

Ferrari:1988:NEF

[FPV88] P. A. Ferrari, E. Presutti, and M. E. Vares. Non equilibrium fluctuations for a zero range process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):237–268, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1988__24_2_237_0.

Furstenberg:1995:PFD

[FPW95] Hillel Furstenberg, Yuval Peres, and Benjamin Weiss. Perfect filtering and double disjointness. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(3):453–465, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1995__31_3_453_0.

Fournier:2003:LTA

[FR03] Nicolas Fournier and Bernard Roynette. On long time almost sure asymptotics of renormalized branching diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):979–991, November/December 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_6_979_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000311.

Friz:2014:CRF

[FR14] Peter Friz and Sebastian Riedel. Convergence rates for the full Gaussian rough paths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):154–194, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1388545270.

Franchi:1991:TRA

[Fra91] J. Franchi. Théorème des résidus asymptotique pour le mouvement brownien sur une surface riemannienne compacte. (French) [Asymptotic theorem of residues for Brownian motion on a compact Riemannian surface]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(4):445–462, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_4_445_0.

Friant:1967:LFC

[Fri67] Jean Friant. Les langages ^{≪} Context-Sensitive ^{≫}. (French) [“Context-sensitive” languages]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(1):35–120, January–March 1967. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1967__3_1_35_0.

Floreani:2022:OPD

[FRS22] Simone Floreani, Frank Redig, and Federico Sau. Orthogonal polynomial duality of boundary driven particle systems and non-equilibrium correlations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):220–247, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Orthogonal-polynomial-
duality-of-boundary-driven-particle-systems-and-non/
10.1214/21-AIHP1163.full.

Fitzsimmons:1989:CEM

[FS89] P. J. Fitzsimmons and Thomas S. Salisbury. Capacity and energy for multiparameter Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(3):325–350, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1989__25_3_325_0.

Fan:2003:ILS

[FS03] Aihua Fan and Dominique Schneider. Sur une inégalité de Littlewood–Salem. (French) [On an inequality of Littlewood– Salem]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(2):193–216, March/April 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_2_193_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000146.

Fleischmann:2004:SSM

[FS04] Klaus Fleischmann and Anja Sturm. A super-stable motion with infinite mean branching. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(5):513–537, September/October 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_5_513_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000287.

Faggionato:2017:RWQ

[FS17] Alessandra Faggionato and Vittoria Silvestri. Random walks on quasi one dimensional lattices: Large deviations and fluctuation theorems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):46–78, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1486544884.

Fremond:1971:CGL

[FSP71] C. Frémond and M. Sueur-Pontier. Caractérisation des groupes localement compacts de type (T) ayant la propriété de point fixe. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(4):293– 298, October–December 1971. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_4_293_0.

Flacke:1979:BSS

[FT79] Werner Flacke and Norbert Therstappen. Bayesian sufficient statistics and invariance. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(4):303–314, October–December 1979. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_4_303_0.

Fukushima:2005:PPP

[FT05] Masatoshi Fukushima and Hiroshi Tanaka. Poisson point processes attached to symmetric diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):419–459, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_3_419_0;   http://www.sciencedirect.
com/science/article/pii/S024602030500018X.

Fisher:2012:DAS

[FT12] Albert M. Fisher and Marina Talet. Dynamical attraction to stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):551–578, May 2012. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1334148210.

Fournier:2020:TNN

[FTV20] Nicolas Fournier, Etienne Tanré, and Romain Veltz. On a toy network of neurons interacting through their dendrites. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1041–1071, May 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/On-a-toy-network-of-neurons-interacting-
through-their-dendrites/10.1214/19-AIHP993.full.

Funaki:1999:FBP

[Fun99] T. Funaki. Free boundary problem from stochastic lattice gas model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(5):573–603, September 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_5_573_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001077.

Furman:1997:MET

[Fur97] Alex Furman. On the multiplicative ergodic theorem for uniquely ergodic systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(6):797–815, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_6_797_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801136.

Friz:2005:ABR

[FV05] Peter Friz and Nicolas Victoir. Approximations of the Brownian rough path with applications to stochastic analysis. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):703–724, July/August 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_4_703_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000603.

Friz:2007:LDP

[FV07] Peter Friz and Nicolas Victoir. Large deviation principle for enhanced Gaussian processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6):775–785, November/December 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_6_775_0;   http://
www.sciencedirect.com/science/article/pii/S024602030700009X.

Friz:2010:DED

[FV10] Peter Friz and Nicolas Victoir. Differential equations driven by Gaussian signals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):369–413, May 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1273584128;   http://www.numdam.org/item?id=AIHPB_2010_
_46_2_369_0.

Fontes:2006:RES

[FVY06] L. R. G. Fontes, M. Vachkovskaia, and A. Yambartsev. Repulsion of an evolving surface on walls with random heights. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(2):207–213, March/April 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_2_207_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000622.

Fan:2004:CSR

[FW04] Ai-Hua Fan and Jun Wu. On the covering by small random intervals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):125–131, January/February 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_1_125_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000566.

Fleischmann:2007:LDP

[FW07] Klaus Fleischmann and Vitali Wachtel. Lower deviation probabilities for supercritical Galton–Watson processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43 (2):233–255, March/April 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_2_233_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000434.

Fleischmann:2009:LTA

[FW09] Klaus Fleischmann and Vitali Wachtel. On the left tail asymptotics for the limit law of supercritical Galton–Watson processes in the Böttcher case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):201–225, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469978;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_201_0.

Follmer:2000:WBM

[FWY00] Hans Föllmer, Ching-Tang Wu, and Marc Yor. On weak Brownian motions of arbitrary order. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(4):447–487, July 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_4_447_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001333.

Fournier:2019:EBS

[FX19] Nicolas Fournier and Liping Xu. On the equivalence between some jumping SDEs with rough coefficients and some non-local PDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1163–1178, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1557820847.

Florens-Zmirou:1988:SDT

[FZ88] Danielle Florens-Zmirou. Statistiques de diffusions et temps local. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(1):99–130, ???? 1988. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1988__24_1_99_0.

Galambos:1970:DAF

[Gal70] J. Galambos. Distribution of arithmetical functions. A survey. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(4):281–305, October– December 1970. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1970__6_4_281_0.

Galtchouk:1976:RME

[Gal76] L. I. Galtchouk. Représentation des martingales engendrées par un processus à accroissements indépendants (cas des martingales de carré intégrable). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(3):199–211, July–September 1976. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1976__12_3_199_0.

Gallardo:1996:CMD

[Gal96] Léonard Gallardo. Chaînes de Markov à dérive stable et loi des grands nombres sur les hypergroupes. (French) [Stabledrift Markov chains and law of large numbers on hypergroups]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(6):701–723, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1996__32_6_701_0.

Gangolli:1967:PDK

[Gan67] Ramesh Gangolli. Positive definite kernels on homogeneous spaces and certain stochastic processes related to Lévy’s Brownian motion of several parameters. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(2):121–226, April–June 1967. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1967__3_2_121_0.

Gantert:2002:STA

[Gan02] Nina Gantert. Subexponential tail asymptotics for a random walk with randomly placed one-way nodes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(1):1– 16, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_2002__38_1_1_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301010901.

Ganesan:2013:SGC

[Gan13] Ghurumuruhan Ganesan. Size of the giant component in a random geometric graph. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1130–1140, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1380718740.

Garel:1978:TDV

[Gar78] B. Garel. Tests de détection de valeurs aberrantes multidimensionnelles. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):303–314, July–September 1978. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_3_303_0.

Garel:1979:DPV

[Gar79] B. Garel. Détection de plusieurs valeurs aberrantes dans un échantillon gaussien multidimensionnel. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1):51–62, January–March 1979. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1979__15_1_51_0.

Garet:1999:GFA

[Gar99] Olivier Garet. Gibbsian fields associated to exponentially decreasing quadratic potentials. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(3):387–415, May/June 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1999__35_3_387_0;   http://www.sciencedirect.
com/science/article/pii/S024602039980017X.

Gassiat:1990:ESP

[Gas90] Élisabeth Gassiat. Estimation semi-paramétrique d’un modèle autorégressif stationnaire multiindice non nécessairement causal. (French) [Semi-parametric estimation of a stationary autoregressive multiindex model that is not necessarily causal]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):181–205, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1990__26_1_181_0.

Gassiat:2002:LRI

[Gas02] Elisabeth Gassiat. Likelihood ratio inequalities with applications to various mixtures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):897–906, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_897_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302011251.

Gassiat:2021:NUR

[Gas21] Paul Gassiat. Non-uniqueness for reflected rough differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1369–1387, August 2021.

CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Non-uniqueness-for-reflected-
rough-differential-equations/10.1214/20-AIHP1121.full.

Gaunt:2020:SMF

[Gau20] Robert E. Gaunt. Stein’s method for functions of multivariate normal random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1484–1513, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Steins-method-for-functions-
of-multivariate-normal-random-variables/10.1214/19-
AIHP1011.full.

Gaveau:1975:DFM

[Gav75] B. Gaveau. Désintégration des fonctionnelles de Markov. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(4): 317–324, October–December 1975. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1975__11_4_317_0.

Gu:2016:WCA

[GB16] Yu Gu and Guillaume Bal. Weak convergence approach for parabolic equations with large, highly oscillatory, random potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):261–285, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1452089269.

Gillois:1969:CGP

[GBB69a] M. Gillois, J. Bouffette, and Ar. Bouffette. Covariance génotypique a priori dans les populations homogames. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(1):87–99, January– March 1969. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1969__5_1_87_0.

Gillois:1969:EPE

[GBB69b] M. Gillois, J. Bouffette, and Ar. Bouffette. Étude des populations d’effectif limité homogames, phénotypiques et panmictiques. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5 (1):69–86, January–March 1969. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1969__5_1_69_0.

Gauglhofer:1973:ESP

[GBR73] Margrit Gauglhofer and A. T. Bharucha-Reid. ϵ-entropy of sets of probability distribution functions and their Fourier–Stieltjes transforms. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(2): 113–144, April–June 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_2_113_0.

Goldman:2003:SFP

[GC03] André Goldman and Pierre Calka. On the spectral function of the Poisson–Voronoi cells. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):1057–1082, November/December 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_6_1057_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020303000256.

Genon-Catalot:1993:EDC

[GCJ93] Valentine Genon-Catalot and Jean Jacod. On the estimation of the diffusion coefficient for multi-dimensional diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(1):119–151, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_1_119_0.

Georgin:1978:CCM

[Geo78] J.-P. Georgin. Contrôle des chaînes de Markov sur des espaces arbitraires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):255–277, July–September 1978. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_3_255_0.

Georgii:1997:OC

[Geo97] Hans-Otto Georgii. Orbit coupling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(2):253–268, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1997__33_2_253_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397801239.

Gzyl:1979:OTS

[GG79] Cristina Gzyl and Henryk Gzyl. Occupation time sets of supports of continuous additive functionals. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1):41–50, January–March 1979. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1979__15_1_41_0.

Gine:2001:CKD

[GG01] Evarist Giné and Armelle Guillou. On consistency of kernel density estimators for randomly censored data: rates holding uniformly over adaptive intervals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(4): 503–522, July/August 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_4_503_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010810.

Gine:2002:RSU

[GG02] Evarist Giné and Armelle Guillou. Rates of strong uniform consistency for multivariate kernel density estimators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):907–921, November/December 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_6_907_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011287.

Gloter:2008:LPH

[GG08] Arnaud Gloter and Emmanuel Gobet. LAMN property for hidden processes: The case of integrated diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1): 104–128, February 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1203969870;   http://www.numdam.org/
item?id=AIHPB_2008__44_1_104_0.

Gao:2015:SLS

[GGL15] Zhiqiang Gao, Yves Guivarc’h, and Émile Le Page. Stable laws and spectral gap properties for affine random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1): 319–348, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1421244408.

Goegebeur:2015:UAP

[GGS15] Yuri Goegebeur, Armelle Guillou, and Gilles Stupfler. Uniform asymptotic properties of a nonparametric regression estimator of conditional tails. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):1190–1213, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1435759245.

Geman:1973:RPM

[GH73] Donald Geman and Joseph Horowitz. Remarks on Palm measures. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(3):215– 232, July–September 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_3_215_0.

Guivarch:1988:TLP

[GH88] Y. Guivarc’h and J. Hardy. Théorèmes limites pour une classe de chaînes de Markov et applications aux difféomorphismes d’Anosov. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(1):73–98, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1988__24_1_
73_0.

Goldschmidt:2010:BNE

[GH10] Christina Goldschmidt and Bénédicte Haas. Behavior near the extinction time in self-similar fragmentations I: the stable case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):338–368, May 2010. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1273584127;   http://
www.numdam.org/item?id=AIHPB_2010__46_2_338_0.

Grimmett:2012:GLP

[GH12] G. R. Grimmett and A. E. Holroyd. Geometry of Lipschitz percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):309–326, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1334148200.

Ghoussoub:1975:PHA

[Gho75] N. Ghoussoub. Processus de Harris abstraits. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(4):381–395, October–December 1975. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1975__
11_4_381_0.

Ghoussoub:1977:BLV

[Gho77] N. Ghoussoub. Banach lattices valued amarts. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(2):159–169, April–June 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_2_159_0.

Gradinaru:2001:SLD

[GHR01] Mihai Gradinaru, Samuel Herrmann, and Bernard Roynette. A singular large deviations phenomenon. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(5):555–580, September/October 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_5_555_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010755.

Gantert:2011:ASP

[GHS11] Nina Gantert, Yueyun Hu, and Zhan Shi. Asymptotics for the survival probability in a killed branching random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1): 111–129, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170232;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_111_0.

Gelenbe:1980:QSW

[GI80] E. Gelenbe and R. Iasnogorodski. A queue with server of walking type (autonomous service). Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(1):63–73, January–March 1980. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1980__16_1_63_0.

Greenberg:2005:IPP

[GI05] Lev Greenberg and Dmitry Ioffe. On an invariance principle for phase separation lines. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5):871–885, September/October 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_5_871_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000610.

Gillois:1965:RIGa

[Gil65a] M. Gillois. Relation d’identité en génétique I. — Postulats et axiomes mendéliens. (French) [Genetic identity relationship I.

— Assumptions and Mendelian axioms]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(1):1–38, July–September 1965. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1965__2_1_1_0.

Gillois:1965:RIGb

[Gil65b] M. Gillois. Relation d’identité en génétique II. — corrélation génétique dans le cas de dominance. (French) [ Genetic identity relationship II. Genetic correlation in the case of dominance]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(1):39–94, July– September 1965. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1965__2_1_1_0.

Gillois:1966:LPE

[Gil66a] M. Gillois. L’homogamie dans une population d’effectif limité. (French) [Homogamy in a limited-effect population]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(4):299–347, April–June 1966. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1966__2_4_299_0.

Gillois:1966:NVC

[Gil66b] M. Gillois. Note sur la variance et la covariance génotypiques entre apparentés. (French) [Note on the variance and genotypic covariance between related parties]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(4):349–352, April–June 1966. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1966__2_4_349_0.

Gillois:1966:RDG

[Gil66c] Michel Gillois. La relation de dépendance en génétique. (French) [The relation of dependence in genetics]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(3):261–278, January–March 1966. CODEN AHPBAR. ISSN 0020-2347. URL http://

www.numdam.org/item?id=AIHPB_1966__2_3_261_0.

Gine:1980:DPA

[Gin80] Evarist Giné. Domains of partial attraction in several dimensions. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(2):87– 100, April–June 1980. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1981__17_1_

143_0. See erratum [Gin81].

Gine:1981:EDP

[Gin81] Evarist Giné. Erratum to “Domains of [partial] attraction in several dimensions”. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17 (1):143–145, January–March 1981. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1980__16_2_87_0. See [Gin80].

Gine:1983:LDS

[Gin83] Evarist Giné. Large deviations in spaces of stable type. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(3):267–279, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1983__19_3_267_0.

Giroux:1970:LES

[Gir70] Gaston Giroux. Sur les lois d’entrée et de sortie d’une chaîne de Markov. (French) []. Annales de l’Institut Henri Poincaré.

Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(4):345–362, October–December 1970. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1970__6_4_345_0.

Giraud:2002:RPB

[Gir02] Christophe Giraud. On regular points in Burgers turbulence with stable noise initial data. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(2):229–251, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_2_229_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301010998.

Graczyk:2011:WHF

[GJ11] Piotr Graczyk and Tomasz Jakubowski. On Wiener–Hopf factors for stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):9–19, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1294170227;   http://www.numdam.org/item?id=AIHPB_2011_
_47_1_9_0.

Gheissari:2019:SGS

[GJ19] Reza Gheissari and Aukosh Jagannath. On the spectral gap of spherical spin glass dynamics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):756–776, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1557820830.

Guigues:2020:HTE

[GJN20] Vincent Guigues, Anatoli Juditsky, and Arkadi Nemirovski. Hypothesis testing via Euclidean separation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1929–1957, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-3/Hypothesis-testing-
via-Euclidean-separation/10.1214/19-AIHP1022.full.

Goldenshluger:2008:CPEb

[GJTZ08a] A. Goldenshluger, A. Juditsky, A. Tsybakov, and A. Zeevi. Change-point estimation from indirect observations. 2. Adaptation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5):819–836, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1222261914;   http://
www.numdam.org/item?id=AIHPB_2008__44_5_819_0.

Goldenshluger:2008:CPEa

[GJTZ08b] A. Goldenshluger, A. Juditsky, A. B. Tsybakov, and A. Zeevi. Change-point estimation from indirect observations. 1. Minimax complexity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5):787–818, October 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1222261913;   http://www.numdam.org/item?id=AIHPB_2008_
_44_5_787_0.

Goodman:1993:GCF

[GK93] V. Goodman and J. Kuelbs. Gaussian chaos and functional laws of the iterated logarithm for Itô–Wiener integrals. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(4):485–512, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_4_485_0.

Gruninger:2009:PCP

[GK09] Gabriela Grüninger and Wolfgang König. Potential confinement property of the parabolic Anderson model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3): 840–863, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1249391388;   http://www.numdam.org/
item?id=AIHPB_2009__45_3_840_0.

Grubel:2017:EEP

[GK17] Rudolf Grübel and Zakhar Kabluchko. Edgeworth expansions for profiles of lattice branching random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53 (4):2103–2134, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1511773740.

Gadat:2020:PRT

[GKMMR20] Sébastien Gadat, Jonas Kahn, Clément Marteau, and Cathy Maugis-Rabusseau. Parameter recovery in two-component contamination mixtures: The L^{2} strategy. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1391–1418, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-2/Parameter-recovery-
in-two-component-contamination-mixtures--The-L2/10.
1214/19-AIHP1007.full.

Gantert:2007:ADR

[GKS07] Nina Gantert, Wolfgang König, and Zhan Shi. Annealed deviations of random walk in random scenery. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1): 47–76, January/February 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_1_47_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000185.

Geiger:2003:LTS

[GKV03] Jochen Geiger, Götz Kersting, and Vladimir A. Vatutin. Limit theorems for subcritical branching processes in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(4):593–620, July/August 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_4_593_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000201.

Guerre:1980:QPM

[GL80] S. Guerre and J. T. Lapresté. Quelques propriétés des modèles étalés sur les espaces de Banach. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(4):339–347, October–December 1980. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1980__16_4_339_0.

Guyon:1989:CLH

[GL89] Xavier Guyon and José Leon. Convergence en loi des Hvariations d’un processus gaussien stationnaire sur R. (French)

[Convergence in the law of H-variations of a Gaussian stationary process on R]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(3):265–282, ???? 1989. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1989_
_25_3_265_0.

Guivarch:2016:SGP

[GL16] Y. Guivarc’h and É. Le Page. Spectral gap properties for linear random walks and Pareto’s asymptotics for affine stochastic recursions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):503–574, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1462367884.

Gassiat:2020:EDD

[GL20] Paul Gassiat and Cyril Labbé. Existence of densities for the dynamic. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):326–373, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Existence-of-densities-
for-the-dynamic-Phi4_3-model/10.1214/19-AIHP963.full.

Glasner:1976:CDM

[Gla76] Shmuel Glasner. On Choquet–Deny measures. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(1):1–10, January–March 1976. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1976__12_1_1_0.

Grama:2018:LTA

[GLL18] Ion Grama, Ronan Lauvergnat, and Émile Le Page. Limit theorems for affine Markov walks conditioned to stay positive. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):529–568, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1519030838.

Grasmair:2018:VMN

[GLM18] Markus Grasmair, Housen Li, and Axel Munk. Variational multiscale nonparametric regression: Smooth functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):1058–1097, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1524643240.

Glover:1991:MF

[Glo91] Joseph Glover. Markov functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(2):221–238, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_2_221_0.

Gheissari:2020:ESM

[GLP20] Reza Gheissari, Eyal Lubetzky, and Yuval Peres. Exponentially slow mixing in the mean-field Swendsen–Wang dynamics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):68–86, February 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Exponentially-slow-mixing-in-the-
mean-field-SwendsenWang-dynamics/10.1214/18-AIHP955.
full.

Guyon:1973:PEA

[GLS73a] X. Guyon, E. Lesquoy, and F. Schaeffer. Plans d’expérience adaptés à un champ d’hypothèses a priori. Construction. Précision globale. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(4):379–396, October–December 1973. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1973__9_4_379_0.

Guyon:1973:RPE

[GLS73b] X. Guyon, E. Lesquoy, and F. Schaeffer. Recherche de plans d’expérience adaptés à des hypothèses a priori dans certains cas particuliers. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(4):369–377, October–December 1973. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1973__9_4_369_0.

Gantert:2005:VMA

[GLS05] Nina Gantert, Matthias Löwe, and Jeffrey E. Steif. The voter model with anti-voter bonds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):767–780, July/August 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_4_767_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000640.

Goncalves:2009:HLP

[GLT09] P. Gonçalves, C. Landim, and C. Toninelli. Hydrodynamic limit for a particle system with degenerate rates. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4): 887–909, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1257529884;   http://www.numdam.org/
item?id=AIHPB_2009__45_4_887_0.

Giacomin:2011:DRM

[GLT11] Giambattista Giacomin, Hubert Lacoin, and Fabio Lucio Toninelli. Disorder relevance at marginality and critical point shift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):148–175, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1294170234;   http://
www.numdam.org/item?id=AIHPB_2011__47_1_148_0.

Goldstein:2017:OBH

[GLTZ17] Sheldon Goldstein, Joel L. Lebowitz, Roderich Tumulka, and Nino Zanghî. Any orthonormal basis in high dimension is uniformly distributed over the sphere. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):701–717, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1491897742.

Griffin:2013:SLT

[GM13] Philip S. Griffin and Ross A. Maller. Small and large time stability of the time taken for a Lévy process to cross curved boundaries. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):208–235, February 2013. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1359470132.

Giunti:2018:QHD

[GM18a] Arianna Giunti and Jean-Christophe Mourrat. Quantitative homogenization of degenerate random environments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):22– 50, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1519030818.

Gouere:2018:CPH

[GM18b] Jean-Baptiste Gouéré and Régine Marchand. Continuum percolation in high dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):1778–1804, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1539849783.

Gwynne:2019:CFB

[GM19] Ewain Gwynne and Jason Miller. Convergence of the free Boltzmann quadrangulation with simple boundary to the Brownian disk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):551–589, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1547802409.

Glazman:2020:SAW

[GM20] Alexander Glazman and Ioan Manolescu. Self-avoiding walk on Z^{2} with Yang–Baxter weights: Universality of critical fugacity and 2-point function. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2281–2300, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Self-avoiding-walk-on-
mathbbZ2-with-YangBaxter-weights--Universality/10.1214/
19-AIHP1024.full.

Gwynne:2021:CCL

[GM21] Ewain Gwynne and Jason Miller. Conformal covariance of the Liouville quantum gravity metric for γ ∈ (0, 2). Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):1016–1031, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-

poincare-probabilites-et-statistiques/volume-57/issue-
2/Conformal-covariance-of-the-Liouville-quantum-gravity-
metric-for-%ce%b302/10.1214/20-AIHP1105.full.

Garcia:1978:SCF

[GMP78] M. Garcia, P. Maillard, and Y. Peltraut. Sur la structure de certaines familles de tribus. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):335–341, July–September 1978. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_3_335_0.

Gassiat:2002:PFP

[GMP02] Elisabeth Gassiat, Pascal Massart, and Dominique Picard. Préface. (French) [Preface]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):801, November/ December 2002. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0246020302011378.

Gwynne:2019:SLC

[GMS19] Ewain Gwynne, Cheng Mao, and Xin Sun. Scaling limits for the critical Fortuin–Kasteleyn model on a random planar map I: Cone times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):1–60, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1547802394.

Gorny:2018:GCP

[GMSS18] Matthias Gorny, Édouard Maurel-Segala, and Arvind Singh. The geometry of a critical percolation cluster on the UIPT. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2203–2238, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1539849797.

Giuliani:2017:HFI

[GMT17] Alessandro Giuliani, Vieri Mastropietro, and Fabio Lucio Toninelli. Height fluctuations in interacting dimers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1): 98–168, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1486544886.

Griffin:1993:CBM

[GMV93] Philip S. Griffin, Terry R. McConnell, and Gregory Verchota. Conditioned Brownian motion in simply connected planar domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(2):229–249, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_2_229_0.

Grama:2002:FHC

[GN02] Ion Grama and Michael Nussbaum. A functional Hungarian construction for sums of independent random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):923–957, November/December 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_6_923_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011445.

Gradinaru:2009:MTS

[GN09] Mihai Gradinaru and Ivan Nourdin. Milstein’s type schemes for fractional SDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1085–1098, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1257529893;   http://www.numdam.org/item?id=AIHPB_2009_
_45_4_1085_0.

Gradinaru:2005:OIG

[GNRV05] Mihai Gradinaru, Ivan Nourdin, Francesco Russo, and Pierre Vallois. m-order integrals and generalized Itô’s formula; the case of a fractional Brownian motion with any Hurst index. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):781–806, July/August 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_4_781_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000652.

Gach:2013:SAD

[GNS13] Florian Gach, Richard Nickl, and Vladimir Spokoiny. Spatially adaptive density estimation by localised Haar projections. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):900–914, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1372772649.

Grorud:1994:HVA

[GNSS94] Axel Grorud, David Nualart, and Marta Sanz-Solé. Hilbertvalued anticipating stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(1): 133–161, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1994__30_1_133_0.

Gradinaru:2013:EAB

[GO13] Mihai Gradinaru and Yoann Offret. Existence and asymptotic behaviour of some time-inhomogeneous diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1): 182–207, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1359470131.

Gobet:2002:LPE

[Gob02] Emmanuel Gobet. LAN property for ergodic diffusions with discrete observations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(5):711–737, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_5_711_0;   http://www.sciencedirect.com/science/
article/pii/S024602030201107X.

Goldman:1989:TBP

[Gol89] André Goldman. Techniques biharmoniques pour l’étude du mouvement brownien de P. Lévy à trois paramètres. (French) [Biharmonic techniques for the study of Brownian movement of P. Lévy with three parameters]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):351–381, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1989__25_4_351_0.

Goldstein:1995:AFR

[Gol95] Sheldon Goldstein. Antisymmetric functionals of reversible Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):177–190, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1995__31_1_
177_0.

Gold:2018:ISB

[Gol18] Julian Gold. Isoperimetry in supercritical bond percolation in dimensions three and higher. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2092–2158, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849794.

Gorny:2017:DCW

[Gor17] Matthias Gorny. A dynamical Curie–Weiss model of SOC: The Gaussian case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):658–678, May 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1491897740.

Gouezel:2005:BET

[Gou05] Sébastien Gouëzel. Berry–Esseen theorem and local limit theorem for non uniformly expanding maps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(6):997–1024, November/December 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_6_997_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000871.

Gourcy:2007:LDP

[Gou07] Mathieu Gourcy. Large deviation principle of occupation measure for stochastic Burgers equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(4): 441–459, July/August 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_4_441_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000744.

Grunewald:2009:TSA

[GOVW09] Natalie Grunewald, Felix Otto, Cédric Villani, and Maria G. Westdickenberg. A two-scale approach to logarithmic Sobolev inequalities and the hydrodynamic limit. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2): 302–351, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024672;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_302_0.

Gozlan:2010:PID

[Goz10] Nathael Gozlan. Poincaré inequalities and dimension free concentration of measure. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3):708–739, August 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1281100396;   http://www.numdam.org/item?id=AIHPB_2010_
_46_3_708_0.

Glorennec:1974:TRP

[GP74a] P. Y. Glorennec and J. Pellaumail. Théorème de Riesz pour des processus réels. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(3):355–367, July–September 1974. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1974__10_3_355_0.

Gravereaux:1974:FIP

[GP74b] J. B. Gravereaux and J. Pellaumail. Formule de Itô pour des processus non continus à valeurs dans des espaces de Banach. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(4): 399–422, October–December 1974. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1974__10_4_399_0.

Gaudron:2001:ECL

[GP01] Guillaume Gaudron and Etienne Pardoux. EDSR, convergence en loi et homogénéisation d’EDP paraboliques semilinéaires. (French) [EDSR, convergence in distribution and parabolic semi-linear EDP homogenization]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(1):1– 42, January 2001. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_2001__37_1_1_0;   http://www.sciencedirect.
com/science/article/pii/S024602030001058X.

Guedon:2007:CMS

[GP07] O. Guédon and G. Paouris. Concentration of mass on the Schatten classes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1):87–99, January/February 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_1_87_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000203.

Gantert:2011:MDB

[GP11] Nina Gantert and Jonathon Peterson. Maximal displacement for bridges of random walks in a random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3): 663–678, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834854;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_663_0.

Gadat:2014:LTB

[GP14] Sébastien Gadat and Fabien Panloup. Long time behaviour and stationary regime of memory gradient diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (2):564–601, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1395856141.

Ganguly:2020:RTD

[GP20] Shirshendu Ganguly and Soumik Pal. The random transposition dynamics on random regular graphs and the Gaussian free field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2935–2970, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/The-random-transposition-
dynamics-on-random-regular-graphs-and-the/10.1214/20-
AIHP1063.full.

Guillotin-Plantard:2016:QLD

[GPdS16] Nadine Guillotin-Plantard and Renato Soares dos Santos. The quenched limiting distributions of a charged-polymer model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):703–725, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1462367890.

Guillotin-Plantard:2010:LTR

[GPP10] Nadine Guillotin-Plantard and Clémentine Prieur. Limit theorem for random walk in weakly dependent random scenery. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1178–1194, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1288878342;   http://
www.numdam.org/item?id=AIHPB_2010__46_4_1178_0.

Gantert:2010:IVR

[GPS10] Nina Gantert, Yuval Peres, and Zhan Shi. The infinite valley for a recurrent random walk in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):525–536, May 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1273584133;   http://www.numdam.org/
item?id=AIHPB_2010__46_2_525_0.

Godinho:2015:PCS

[GQ15] David Godinho and Cristobal Quiñinao. Propagation of chaos for a subcritical Keller–Segel model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):965–992, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1435759237.

Geng:2016:ITS

[GQ16] X. Geng and Z. Qian. On an inversion theorem for Stratonovich’s signatures of multidimensional diffusion paths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):429–447, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1452089275.

Genest:2008:VPB

[GR08] Christian Genest and Bruno Rémillard. Validity of the parametric bootstrap for goodness-of-fit testing in semiparametric models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1096–1127, December 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1227287567;   http://www.numdam.org/item?id=AIHPB_2008_
_44_6_1096_0.

Gaudilliere:2011:SFS

[GR11] A. Gaudillière and J. Reygner. Sampling the Fermi statistics and other conditional product measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3): 790–812, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834859;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_790_0.

Gravey:1984:TAN

[Gra84] A. Gravey. Temps d’attente et nombre de clients dans une file nd/d/1. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(1):53–73, ???? 1984. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1984__20_1_53_0.

Graham:1985:BPC

[Gra85] Carl Graham. Boundary processes: the calculus of processes diffusing on the boundary. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(1):73–102, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1985__21_1_
73_0.

Graham:1988:MPS

[Gra88a] Carl Graham. The martingale problem with sticky reflection conditions, and a system of particles interacting at the boundary. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(1):45–72, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1988__24_1_45_0.

Gravereaux:1988:CMP

[Gra88b] J. B. Gravereaux. Calcul de Malliavin et probabilité invariante d’une chaîne de Markov. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):159–188, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1988__24_2_159_0.

Graham:1992:NDJ

[Gra92] Carl Graham. Nonlinear diffusion with jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(3):

393–402, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1992__28_3_393_0.

Grabiner:1999:BMW

[Gra99] David J. Grabiner. Brownian motion in a Weyl chamber, non-colliding particles, and random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35 (2):177–204, March/April 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_2_177_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399800107.

Gretete:2008:SDC

[Gre08] Driss Gretete. Stabilité du comportement des marches aléatoires sur un groupe localement compact. (French) [Stability behavior of random walks on a locally compact group]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1):129–142, February 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1203969871;   http:/
/www.numdam.org/item?id=AIHPB_2008__44_1_129_0.

Griffeath:1977:ETC

[Gri77] David Griffeath. An ergodic theorem for a class of spin systems. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(2):141–157, April– June 1977. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1977__13_2_141_0.

Grigorescu:2004:IDC

[Gri04] Ilie Grigorescu. An infinite dimensional central limit theorem for correlated martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2):167–196, March/ April 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2004__40_2_167_0;   http://www.sciencedirect.
com/science/article/pii/S0246020303000451.

Gozlan:2018:CCW

[GRS^{+}18] Nathael Gozlan, Cyril Roberto, Paul-Marie Samson, Yan Shu, and Prasad Tetali. Characterization of a class of weak transport-entropy inequalities on the line. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1667–1693, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1531296032.

Grunig:1966:PCR

[Gru66a] Roland Grunig. Probabilités conditionnelles régulières sur des tribus de type non dénombrable. (French) [Regular conditional probabilities on tribes of uncountable types]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(3):227–229, January–March 1966. CODEN AHPBAR. ISSN 0020-2347. URL http://

www.numdam.org/item?id=AIHPB_1966__2_3_227_0.

Grunig:1966:PPT

[Gru66b] Roland Grunig. Probabilités parfaites. Théorème de Kolmogorov. (French) [Perfect probabilities. Kolmogorov’s theorem]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(3):221– 225, January–March 1966. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1966__
2_3_221_0.

Gradinaru:1999:ATI

[GRVY99] Mihai Gradinaru, Bernard Roynette, Pierre Vallois, and Marc Yor. Abel transform and integrals of Bessel local times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(4):531–572, July 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_4_531_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001053.

Getoor:1987:EFB

[GS87] R. K. Getoor and J. Steffens. The energy functional, balayage, and capacity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2):321–357, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_S2_321_0.

Gamet:1997:TEM

[GS97] Catherine Gamet and Dominique Schneider. Théorèmes ergodiques multidimensionnels et suites aléatoires universellement représentatives en moyenne. (French) [Multidimensional ergodic theorems and universally representative random sequences in average]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(2):269–282, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_2_269_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801240.

Gobron:2010:CAH

[GS10a] Thierry Gobron and Ellen Saada. Couplings, attractiveness and hydrodynamics for conservative particle systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1132–1177, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1288878341;   http://
www.numdam.org/item?id=AIHPB_2010__46_4_1132_0.

Goia:2010:ONT

[GS10b] Aldo Goia and Ernesto Salinelli. Optimal nonlinear transformations of random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3):653–676, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1281100394;   http://www.numdam.org/item?id=AIHPB_2010_
_46_3_653_0.

Gartner:2015:SLL

[GS15a] Jürgen Gärtner and Adrian Schnitzler. Stable limit laws for the parabolic Anderson model between quenched and annealed behaviour. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):194–206, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1421244403.

Gorin:2015:LMT

[GS15b] Vadim Gorin and Mykhaylo Shkolnikov. Limits of multilevel TASEP and similar processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):18–27, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1421244397.

Gobet:2018:ODS

[GS18] Emmanuel Gobet and Uladzislau Stazhynski. Optimal discretization of stochastic integrals driven by general Brownian semimartingale. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1556–1582, August 2018. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1531296029.

Goswami:2005:DCQ

[GSS05] Debashish Goswami, Lingaraj Sahu, and Kalyan B. Sinha. Dilation of a class of quantum dynamical semigroups with unbounded generators on UHF algebras. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):505–522, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_3_505_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000191.

Gualtierotti:1977:MGC

[Gua77] A. F. Gualtierotti. Mixtures of Gaussian cylinder set measures and abstract Wiener spaces as models for detection. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(4):333–356, October–December 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_4_333_0.

Guenais:1999:SEA

[Gue99] Mélanie Guenais. Spectres de M-extensions aléatoires. (French) [Spectra of random M-extensions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(2):239–259, March/ April 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1999__35_2_239_0;   http://www.sciencedirect.
com/science/article/pii/S0246020399800120.

Gugushvili:2012:NID

[Gug12] Shota Gugushvili. Nonparametric inference for discretely sampled Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):282–307, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1327328023.

Guilbart:1979:PSE

[Gui79] C. Guilbart. Produits scalaires sur l’espace des mesures. (French) [Scalar products on measure space]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(4):333–354, October–December 1979. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1979__15_4_333_0.

Guivarch:1990:ENL

[Gui90] Yves Guivarc’h. Sur une extension de la notion de loi semistable. (French) [On an extension of the concept of the semistable law]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(2):261–285, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1990__26_2_261_0.

Guiol:1997:RPP

[Gui97] H. Guiol. Un résultat pour le processus d’exclusion à longue portée. (French) [A result for the process of longrange exclusion]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(4):387–405, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_4_387_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397800982.

Guillotin:2000:ADR

[Gui00] N. Guillotin. Asymptotics of a dynamic random walk in a random scenery: I. Law of large numbers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(2):127–151, March 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2000__36_2_127_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300001199.

Guionnet:2002:LDU

[Gui02] Alice Guionnet. Large deviations upper bounds and central limit theorems for non-commutative functionals of Gaussian large random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(3):341–384, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_

_38_3_341_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010937.

Guo:2014:LVR

[Guo14] Xiaoqin Guo. On the limiting velocity of random walks in mixing random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):375–402, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1395856132.

Gurau:2014:URT

[Gur14] Razvan Gurau. Universality for random tensors. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (4):1474–1525, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555508.

Gustavsson:2005:GFE

[Gus05] Jonas Gustavsson. Gaussian fluctuations of eigenvalues in the GUE. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(2):151–178, March/April 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_2_151_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000573.

Gill:1995:IEB

[GvdLW95] Richard D. Gill, Mark J. van der Laan, and Jon A. Wellner. Inefficient estimators of the bivariate survival function for three models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(3):545–597, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1995__31_3_545_0.

Gugushvili:2020:BLI

[GvdVY20] Shota Gugushvili, Aad van der Vaart, and Dong Yan. Bayesian linear inverse problems in regularity scales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2081–2107, August 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/

volume-56/issue-3/Bayesian-linear-inverse-problems-
in-regularity-scales/10.1214/19-AIHP1029.full.

Goldstein:2018:SMP

[GW18] Larry Goldstein and Nathakhun Wiroonsri. Stein’s method for positively associated random variables with applications to the Ising and voter models, bond percolation, and contact process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):385–421, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1519030833.

Gantert:1998:LMD

[GZ98] N. Gantert and O. Zeitouni. Large and moderate deviations for the local time of a recurrent Markov chain on Z^{2}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(5):687–704, September/October 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_5_687_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800046.

Garcia-Zelada:2019:LDP

[GZ19] David García-Zelada. A large deviation principle for empirical measures on Polish spaces: Application to singular Gibbs measures on manifolds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1377–1401, August 2019. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1569398873.

Haas:2004:RFD

[Haa04] Bénédicte Haas. Regularity of formation of dust in self-similar fragmentations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(4):411–438, July/August 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_4_411_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000736.

Hachem:1982:QTE

[Hac82] Bachar Hachem. Quelques théorèmes ergodiques sur-additifs dans L^{p} (1 ≤ p < ∞). (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(3):201– 222, July–September 1982. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1982__
18_3_201_0.

Haezendonck:1972:PBF

[Hae72] Jean Haezendonck. Un processus de branchement fellerien discernable. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(3):265–306, July–September 1972. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1972__8_3_265_0.

Hafouta:2020:NMD

[Haf20] Yeor Hafouta. Nonconventional moderate deviations theorems and exponential concentration inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):428–448, February 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Nonconventional-moderate-deviations-
theorems-and-exponential-concentration-inequalities/
10.1214/19-AIHP967.full.

Haggstrom:2003:FPM

[Häg03] Olle Häggström. Is the fuzzy Potts model Gibbsian? Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(5): 891–917, September/October 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_5_891_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000268.

Haiman:1981:VES

[Hai81] George Haiman. Valeurs extrémales de suites stationnaires de variables aléatoires m-dépendantes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(3):309–330, July–September 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_3_309_0.

Haiman:1987:EES

[Hai87] George Haiman. Étude des extrêmes d’une suite stationnaire m-dépendante avec une application relative aux accroissements du processus de Wiener. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(3):425–457, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1987__23_3_425_0.

Hamadene:1996:EDS

[Ham96] S. Hamadene. Équations différentielles stochastiques rétrogrades: le cas localement lipschitzien. (French) [Backward-stochastic differential equations: the locally Lipschitz case]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(5): 645–659, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_5_645_0.

Hanen:1971:PPS

[Han71] A. Hanen. Processus ponctuels stationnaires et flots spéciaux. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(1): 23–30, January–March 1971. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_1_23_0.

Hansel:1972:TRM

[Han72] G. Hansel. Théorème de relèvement et mesures bivalentes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(4):395– 401, October–December 1972. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1972__
8_4_395_0.

Hanen:1973:MAS

[Han73] A. Hanen. Mesures aléatoires stationnaires et mesure de Palm. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(3):311– 325, July–September 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_3_311_0.

Handa:2004:RDM

[Han04] Kenji Handa. Reversible distributions of multi-allelic Gillespie– Sato diffusion models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(5):569–597, September/October 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2004__40_5_569_0;   http://www.sciencedirect.
com/science/article/pii/S0246020304000226.

Hanin:2017:PZC

[Han17] Boris Hanin. Pairing of zeros and critical points for random polynomials. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1498–1511, August 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1500624049.

Harge:2001:VLP

[Har01] Gilles Hargé. Valeurs limites pour les éléments des chaos. (French) [Limit values for the elements of chaos]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(5): 523–554, September/October 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_5_523_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010822.

Harge:2005:CEC

[Har05] Gilles Hargé. Characterization of equality in the correlation inequality for convex functions, the U-conjecture. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41 (4):753–765, July/August 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_4_753_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000639.

Hardy:2015:ACP

[Har15] Adrien Hardy. Average characteristic polynomials of determinantal point processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):283–303, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1421244406.

Hassenforder:1990:SAP

[Has90] C. Hassenforder. Seules les affinités préservent le type de la loi gamma à paramètre entier. (French) [Only affinities preserve the type of the gamma distribution with integer parameter]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(4):541–548, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1990__26_4_541_0.

Hazod:1993:PCL

[Haz93] W. Hazod. Probabilities on contractible locally compact groups: the existence of universal distributions in the sense of W. Doeblin. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(3):339–355, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_3_339_0.

Heinich:1971:IPM

[Hei71] H. Heinich. Intégration pour une mesure à valeurs dans un groupe topologique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(3):177–192, July–September 1971. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1971__7_3_177_0.

Heinich:1974:ICE

[Hei74a] H. Heinich. Intégration dans certains espaces de Riesz à distance concave. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(2):185–200, April–June 1974. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1974__10_2_185_0.

Heinich:1974:MVE

[Hei74b] H. Heinich. Mesures à valeurs dans un espace de Banach complètement réticulé. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(3):339–344, July–September 1974. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1974__10_3_339_0.

Heinich:1978:MAP

[Hei78] H. Heinich. Martingales asymptotiques pour l’ordre. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):315–333, July–September 1978. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1978__14_3_
315_0.

Heinich:1983:MVE

[Hei83] Henri Heinich. Mesures vectorielles dans les espaces réticulés. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(2):153–174, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1983__19_2_153_0.

Heinich:1990:MVF

[Hei90] Henri Heinich. Médianes vectorielles. (French) [Vector medians]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):375–385, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1990__26_3_375_0.

Helffer:1999:RDC

[Hel99] Bernard Helffer. Remarks on decay of correlations and Witten Laplacians III. Application to logarithmic Sobolev inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(4):483–508, July 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_4_483_0;   http://www.sciencedirect.com/science/
article/pii/S024602039900103X.

Hennion:1982:TCC

[Hen82] Hubert Hennion. Transience de certaines chaînes semimarkoviennes. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(3):277–291, July– September 1982. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1982__18_3_
277_0.

Hennion:1991:DDP

[Hen91] Hubert Hennion. Dérivabilité du plus grand exposant caractéristique des produits de matrices aléatoires indépendantes à coefficients positifs. (French) [of the largest characteristic exponent differentiability of independent products of random matrices with positive coefficients]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(1):27–59, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1991_
_27_1_27_0.

Hering:1977:MMC

[Her77] H. Hering. Minimal moment conditions in the limit theory for general Markov branching processes. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(4):299–319, October–December 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_4_299_0.

Herve:1994:EOQ

[Her94] Loïc Hervé. Étude d’opérateurs quasi-compacts positifs. Applications aux opérateurs de transfert. (French) [Study of positive quasi-compact operators. Applications to transfer operators]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(3):437–466, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1994__30_3_437_0.

Herve:2005:TLP

[Her05] Loïc Hervé. Théorème local pour chaînes de Markov de probabilité de transition quasi-compacte. Applications aux chaînes V - géométriquement ergodiques et aux modèles itératifs. (French) [Local theorem for Markov chains of quasi-compact transition probability. Applications to V -geometrically ergodic chains and to iterative models]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(2):179–196, March/April 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_2_179_0;   http://www.sciencedirect.com/science/
article/pii/S024602030400055X.

Herve:2008:QCM

[Her08a] Loïc Hervé. Quasi-compactness and mean ergodicity for Markov kernels acting on weighted supremum normed spaces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1090–1095, December 2008. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1227287566;   http://
www.numdam.org/item?id=AIHPB_2008__44_6_1090_0.

Herve:2008:VCT

[Her08b] Loïc Hervé. Vitesse de convergence dans le théorème limite central pour des chaînes de Markov fortement ergodiques. (French) [Speed of convergence in the central limit theorem for strongly-ergodic Markov chains]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2): 280–292, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1207948220;   http://www.numdam.org/
item?id=AIHPB_2008__44_2_280_0.

Hermon:2018:SUM

[Her18] Jonathan Hermon. On sensitivity of uniform mixing times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):234–248, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1519030827.

Hermon:2019:RNB

[Her19] Jonathan Hermon. Reversibility of the non-backtracking random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2295–2319, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/Reversibility-of-the-
non-backtracking-random-walk/10.1214/18-AIHP949.full.

Heuchenne:1979:LCM

[Heu79] C. Heuchenne. L’analyse de covariance mise en question. (French) [Analysis of covariance questioned]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(4):315–332, October–December 1979. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1979__15_4_315_0.

Heurteaux:1998:EDI

[Heu98] Yanick Heurteaux. Estimations de la dimension inférieure et de la dimension supérieure des mesures. (French) [Estimates of the lower dimension and the greater dimension of measures]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(3):309–338, May/June 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_3_309_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800149.

Heyer:1968:LFN

[Hey68] Herbert Heyer. L’analyse de Fourier non-commutative et applications à la théorie des probabilités. (French) [Fourier noncommutative analysis and applications to probability theory]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(2):143–164, April–June 1968. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1968__4_2_143_0.

Heydenreich:2011:LRS

[Hey11] Markus Heydenreich. Long-range self-avoiding walk converges to α-stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):20–42, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1294170228;   http://www.numdam.org/item?id=AIHPB_2011_
_47_1_20_0.

Hering:1981:CBD

[HH81] H. Hering and F. M. Hoppe. Critical branching diffusions: proper normalization and conditioned limit. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(3):251–274, July–September 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_3_251_0.

Harris:2009:BBM

[HH09] J. W. Harris and S. C. Harris. Branching Brownian motion with an inhomogeneous breeding potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3): 793–801, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1249391385;   http://www.numdam.org/
item?id=AIHPB_2009__45_3_793_0.

Harris:2006:FPA

[HHK06] J. W. Harris, S. C. Harris, and A. E. Kyprianou. Further probabilistic analysis of the Fisher–Kolmogorov–Petrovskii– Piscounov equation: one sided travelling-waves. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1): 125–145, January/February 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_1_125_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000695.

Hudde:2021:SGI

[HHM21] Anselm Hudde, Martin Hutzenthaler, and Sara Mazzonetto. A stochastic Gronwall inequality and applications to moments, strong completeness, strong local Lipschitz continuity, and perturbations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):603–626, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-2/A-stochastic-Gronwall-
inequality-and-applications-to-moments-strong-completeness/
10.1214/20-AIHP1064.full.

Hilico:1973:PPM

[Hil73] C. Hilico. Processus ponctuels marqués stationnaires. Application à l’interaction sélective de deux processus ponctuels stationnaires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(2):177–192, April–June 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_2_177_0.

Hillion:1977:QIR

[Hil77] Alain Hillion. Quelques inégalités sur les régions de confiance. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(4): 371–384, October–December 1977. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1977__13_4_371_0.

Hilico:1978:MAS

[Hil78] C. Hilico. Mesures aléatoires stationnaires sur un espace produit. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(1): 25–31, January–March 1978. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1978__
14_1_25_0.

Hirsch:2016:APL

[Hir16] Christian Hirsch. On the absence of percolation in a linesegment based lilypond model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):127–145, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089263.

Hayashi:2011:ISC

[HJY11] Takaki Hayashi, Jean Jacod, and Nakahiro Yoshida. Irregular sampling and central limit theorems for power variations: the continuous case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1197–1218, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1317906508;   http://www.numdam.org/
item?id=AIHPB_2011__47_4_1197_0.

Hamana:2002:LDR

[HK02] Yuji Hamana and Harry Kesten. Large deviations for the range of an integer valued random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(1):17–58, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_1_17_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010871.

Husseini:2009:JPH

[HK09] Ryad Husseini and Moritz Kassmann. Jump processes, L-harmonic functions, continuity estimates and the Feller property. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1099–1115, November 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1257529894;   http://www.numdam.org/item?id=AIHPB_2009_
_45_4_1099_0.

Hairer:2015:GVN

[HK15] Martin Hairer and David Kelly. Geometric versus nongeometric rough paths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):207–251, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1421244404.

Harris:2010:SLL

[HKK10] S. C. Harris, R. Knobloch, and A. E. Kyprianou. Strong law of large numbers for fragmentation processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1): 119–134, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1267454111;   http://www.numdam.org/
item?id=AIHPB_2010__46_1_119_0.

Heyne:2014:MSB

[HKM14] Gregor Heyne, Michael Kupper, and Christoph Mainberger. Minimal supersolutions of BSDEs with lower semicontinuous generators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):524–538, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1395856139.

Hairer:2016:AHC

[HKPG16] Martin Hairer, Leonid Koralov, and Zsolt Pajor-Gyulai. From averaging to homogenization in cellular flows — an exact description of the transition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1592–1613, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1479373241.

Hopfner:2005:REI

[HL05] Reinhard Höpfner and Eva Löcherbach. Remarks on ergodicity and invariant occupation measure in branching diffusions with immigration. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(6):1025–1047, November/December 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_6_1025_0;   http://www.sciencedirect.
com/science/article/pii/S0246020304000858.

Huang:2020:SSS

[HL20] Jiaoyang Huang and Benjamin Landon. Spectral statistics of sparse Erdős–Rényi graph Laplacians. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):120–154, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-1/Spectral-statistics-
of-sparse-Erd%c5%91sR%c3%a9nyi-graph-Laplacians/10.
1214/19-AIHP957.full.

Hu:2022:ADP

[HL22] Yaozhong Hu and Khoa Lê. Asymptotics of the density of parabolic Anderson random fields. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):105–133, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Asymptotics-of-the-density-
of-parabolic-Anderson-random-fields/10.1214/21-AIHP1148.
full.

Holden:2022:MCB

[HLLS22] Nina Holden, Gregory F. Lawler, Xinyi Li, and Xin Sun. Minkowski content of Brownian cut points. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):455–488, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-58/issue-1/Minkowski-content-
of-Brownian-cut-points/10.1214/21-AIHP1159.full.

Hachem:2006:EDE

[HLN06] W. Hachem, P. Loubaton, and J. Najim. The empirical distribution of the eigenvalues of a Gram matrix with a given variance profile. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6):649–670, November/December 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_6_649_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000021.

Huang:2017:LTA

[HLN17] Jingyu Huang, Khoa Lê, and David Nualart. Large time asymptotics for the parabolic Anderson model driven by spatially correlated noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1305–1340, August 2017. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1500624043.

Hachem:2013:BFB

[HLNV13] Walid Hachem, Philippe Loubaton, Jamal Najim, and Pascal Vallet. On bilinear forms based on the resolvent of large random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):36–63, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1359470125.

Henry-Labordere:2019:BDR

[HLOT^{+}19] Pierre Henry-Labordère, Nadia Oudjane, Xiaolu Tan, Nizar Touzi, and Xavier Warin. Branching diffusion representation of semilinear PDEs and Monte Carlo approximation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):184–210, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1547802399.

Holden:2022:NPP

[HLS22] Nina Holden, Xinyi Li, and Xin Sun. Natural parametrization of percolation interface and pivotal points. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):7– 25, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-58/issue-1/Natural-parametrization-
of-percolation-interface-and-pivotal-points/10.1214/
21-AIHP1160.full.

Hopfner:2016:ESH

[HLT16] R. Höpfner, E. Löcherbach, and M. Thieullen. Ergodicity for a stochastic Hodgkin–Huxley model driven by Ornstein– Uhlenbeck type input. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):483–501, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089277.

Hu:2021:RAB

[HLT21] Yiping Hu, Russell Lyons, and Pengfei Tang. A reverse Aldous–Broder algorithm. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):890–900, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-2/A-reverse-AldousBroder-
algorithm/10.1214/20-AIHP1101.full.

Huang:2016:PAS

[HM16] Lorick Huang and Stéphane Menozzi. A parametrix approach for some degenerate stable driven SDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1925–1975, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1479373254.

Hairer:2018:SFP

[HM18] M. Hairer and J. Mattingly. The strong Feller property for singular stochastic PDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1314–1340, August 2018. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1531296021.

Haiman:1998:RBR

[HMNP98] George Haiman, Nelly Mayeur, Valéry Nevzorov, and Madan L. Puri. Records and 2-block records of 1-dependent stationary sequences under local dependence. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(4): 481–503, July/August 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_4_481_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800253.

Holroyd:2020:SMH

[HMP20a] Alexander E. Holroyd, James B. Martin, and Yuval Peres. Stable matchings in high dimensions via the Poisson-weighted infinite tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):826–846, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/Stable-matchings-in-high-dimensions-
via-the-Poisson-weighted-infinite/10.1214/19-AIHP984.
full.

Honore:2020:NAG

[HMP20b] I. Honoré, S. Menozzi, and G. Pagès. Non-asymptotic Gaussian estimates for the recursive approximation of the invariant distribution of a diffusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1559–1605, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Non-asymptotic-Gaussian-
estimates-for-the-recursive-approximation-of-the/10.
1214/19-AIHP985.full.

Hoffmann:2012:AWE

[HMSH12] M. Hoffmann, A. Munk, and J. Schmidt-Hieber. Adaptive wavelet estimation of the diffusion coefficient under additive error measurements. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1186–1216, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1353098445.

Holden:2019:HRC

[HNPS19] Nina Holden, Serban Nacu, Yuval Peres, and Thomas S. Salisbury. How round are the complementary components of planar Brownian motion? Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):882–908, May 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1557820835.

Hobson:1994:AAT

[Hob94] David Hobson. Asymptotics for an arcsin type result. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):235–243, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1994__30_2_235_0.

Hocquet:2019:FTS

[Hoc19] Antoine Hocquet. Finite-time singularity of the stochastic harmonic map flow. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1011–1041, May 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1557820840.

Hoffmann:2001:EDC

[Hof01] Marc Hoffmann. On estimating the diffusion coefficient: parametric versus nonparametric. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(3):339–372, May/June 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2001__37_3_339_0;   http://www.sciencedirect.
com/science/article/pii/S0246020300010700.

Hofmanova:2015:BGK

[Hof15] Martina Hofmanová. A Bhatnagar–Gross–Krook approximation to stochastic scalar conservation laws. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1500–1528, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1445432051.

Hognas:1974:MAD

[Hög74a] Göran Högnäs. Marches aléatoires sur un demi-groupe compact. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(1): 115–154, January–March 1974. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1974__
10_1_115_0.

Hognas:1974:RMA

[Hög74b] Göran Högnäs. Remarques sur les marches aléatoires dans un demi-groupe avec un idéal compact ayant une probabilité positive. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10 (3):345–354, July–September 1974. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1974__10_3_345_0.

Holmes:2012:EAT

[Hol12] Mark Holmes. Excited against the tide: A random walk with competing drifts. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):745–773, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1340714871.

Holroyd:2017:ODC

[Hol17] Alexander E. Holroyd. One-dependent coloring by finitary factors. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):753–765, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1491897744.

Horowitz:1974:SGM

[Hor74] S. Horowitz. Semi-groups of Markov operators. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(1):155–166, January–March 1974. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1974__10_1_155_0.

Hoshino:2018:GWP

[Hos18] Masato Hoshino. Global well-posedness of complex Ginzburg– Landau equation with a space-time white noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54 (4):1969–2001, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1539849790.

Hiai:2000:LDT

[HP00] Fumio Hiai and Dénes Petz. A large deviation theorem for the empirical eigenvalue distribution of random unitary matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(1):71–85, January 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_1_71_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001163.

Hairer:2011:EHS

[HP11] M. Hairer and N. S. Pillai. Ergodicity of hypoelliptic SDEs driven by fractional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2): 601–628, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887284;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_601_0.

Hermon:2017:PAT

[HP17] Jonathan Hermon and Yuval Peres. The power of averaging at two consecutive time steps: Proof of a mixing conjecture by Aldous and Fill. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):2030–2042, November 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1511773737.

Hughes:2019:BZS

[HP19] Thomas Hughes and Edwin Perkins. On the boundary of the zero set of super-Brownian motion and its local time. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2395–2422, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/On-the-boundary-of-the-
zero-set-of-super-Brownian/10.1214/18-AIHP952.full.

Holden:2021:CWC

[HP21] Nina Holden and Ellen Powell. Conformal welding for critical Liouville quantum gravity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1229–1254, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Conformal-welding-for-
critical-Liouville-quantum-gravity/10.1214/20-AIHP1116.
full.

Holroyd:2009:PM

[HPPS09] Alexander E. Holroyd, Robin Pemantle, Yuval Peres, and Oded Schramm. Poisson matching. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):266–287, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469982;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_266_0.

Haggstrom:1997:DP

[HPS97] Olle Häggström, Yuval Peres, and Jeffrey E. Steif. Dynamical percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(4):497–528, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_4_497_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801033.

Hall:2002:RBM

[HPT02] Peter Hall, Byeong U. Park, and Berwin A. Turlach. Rollingball method for estimating the boundary of the support of a point-process intensity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):959–971, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_959_0;   http://
www.sciencedirect.com/science/article/pii/S0246020302011329.

Hu:2017:GEP

[HQZ17] Ying Hu, Zhongmin Qian, and Zichen Zhang. Gradient estimates for porous medium and fast diffusion equations by martingale method. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1793–1820, November 2017. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1511773726.

Harris:2012:UPB

[HR12] Simon C. Harris and Matthew I. Roberts. The unscaled paths of branching Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):579–608, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1334148211.

Harris:2017:MFL

[HR17] Simon C. Harris and Matthew I. Roberts. The many-to-few lemma and multiple spines. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):226–242, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544890.

Hu:2021:MFL

[HRSS21] Kaitong Hu, Zhenjie Ren, David Siska, and Lukasz Szpruch. Mean-field Langevin dynamics and energy landscape of neural networks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2043–2065, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/Mean-field-Langevin-
dynamics-and-energy-landscape-of-neural-networks/10.
1214/20-AIHP1140.full.

Huang:2018:LQG

[HRV18] Yichao Huang, Rémi Rhodes, and Vincent Vargas. Liouville quantum gravity on the unit disk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1694–1730, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1531296033.

Hirsch:1996:IBS

[HS96] Francis Hirsch and Shiqi Song. Inequalities for Bochner’s subordinates of two-parameter symmetric Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(5):589–600, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_5_589_0.

Hu:1999:SEL

[HS99] Yueyun Hu and Zhan Shi. Shortest excursion lengths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(1): 103–120, January/February 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_1_103_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399800077.

Hasebe:2011:MC

[HS11] Takahiro Hasebe and Hayato Saigo. The monotone cumulants. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1160–1170, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1317906506;   http://
www.numdam.org/item?id=AIHPB_2011__47_4_1160_0.

Hasebe:2017:UFL

[HS17] Takahiro Hasebe and Noriyoshi Sakuma. Unimodality for free Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):916–936, May 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1491897751.

Hammersley:2021:MVS

[HSS21] William R. P. Hammersley, David Siska, and Lukasz Szpruch. McKean–Vlasov SDEs under measure dependent Lyapunov conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):1032–1057, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/McKeanVlasov-SDEs-under-measure-dependent-
Lyapunov-conditions/10.1214/20-AIHP1106.full.

Hasebe:2020:SPF

[HSW20] Takahiro Hasebe, Thomas Simon, and Min Wang. Some properties of the free stable distributions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):296–325, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Some-properties-of-the-
free-stable-distributions/10.1214/19-AIHP962.full.

Horridge:2004:SDK

[HT04] Paul Horridge and Roger Tribe. On stationary distributions for the KPP equation with branching noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(6):759–770, November/December 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_6_759_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000561.

Hu:2002:PSP

[Hu02] Yaozhong Hu. Probability structure preserving and absolute continuity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(4):557–580, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_4_557_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301011049.

Hiai:2009:LST

[HU09] Fumio Hiai and Yoshimichi Ueda. A log-Sobolev type inequality for free entropy of two projections. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1): 239–249, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469980;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_239_0.

Hu:2016:HBM

[Hu16] Yueyun Hu. How big is the minimum of a branching random walk? Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):233–260, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1452089268.

Huang:2019:RWG

[Hua19] Ruojun Huang. On random walk on growing graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1149–1162, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820846.

Hudson:2005:ICQ

[Hud05]  R. L. Hudson.  Itô calculus and quantisation of Lie

bialgebras. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):375–390, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_375_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000208.

Huesmann:2016:OTB

[Hue16] Martin Huesmann. Optimal transport between random measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):196–232, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1452089267.

Hutchcroft:2019:SPP

[Hut19] Tom Hutchcroft. Statistical physics on a product of trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):1001–1010, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820839.

Hutchcroft:2020:ICT

[Hut20] Tom Hutchcroft. Indistinguishability of collections of trees in the uniform spanning forest. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):917–927, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Indistinguishability-
of-collections-of-trees-in-the-uniform-spanning-forest/
10.1214/19-AIHP988.full.

Holcomb:2017:OAP

[HV17] Diane Holcomb and Benedek Valkó. Overcrowding asymptotics for the Sine_{β} process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1181–1195, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1500624035.

Hu:2022:SHE

[HW22] Yaozhong Hu and Xiong Wang. Stochastic heat equation with general rough noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):379–423, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Stochastic-heat-equation-
with-general-rough-noise/10.1214/21-AIHP1161.full.

Ibragimov:2002:IIB

[Ibr02] Ildar Ibragimov. Ildar Ibragimov’s biography. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6): 807–810, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0246020302011433.

En l’honneur de J. Bretagnolle, D. Dacunha-Castelle, I. Ibragimov.

Iksanov:2020:FBM

[IKM20] Alexander Iksanov, Konrad Kolesko, and Matthias Meiners. Fluctuations of Biggins’ martingales at complex parameters. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2445–2479, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Fluctuations-of-Biggins-
martingales-at-complex-parameters/10.1214/20-AIHP1046.
full.

Ichiba:2018:SIE

[IKPY18] Tomoyuki Ichiba, Ioannis Karatzas, Vilmos Prokaj, and Minghan Yan. Stochastic integral equations for Walsh semimartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):726–756, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1524643227.

Istas:1997:QVE

[IL97] Jacques Istas and Gabriel Lang. Quadratic variations and estimation of the local Hölder index of a Gaussian process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(4):407–436, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_4_407_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397800994.

Inglot:2006:AON

[IL06] Tadeusz Inglot and Teresa Ledwina. Asymptotic optimality of new adaptive test in regression model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5): 579–590, September/October 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_5_579_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000968.

Imkeller:1984:SAL

[Imk84] Peter Imkeller. Stochastic analysis and local times for (N, d)- Wiener process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(1):75–101, ???? 1984. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1984__20_1_75_0.

Iftimie:2008:HSR

[IPP08] Bogdan Iftimie, Étienne Pardoux, and Andrey Piatnitski. Homogenization of a singular random one-dimensional PDE. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(3):519–543, June 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1211819423;   http://www.numdam.org/
item?id=AIHPB_2008__44_3_519_0.

Ignatiouk-Robert:1998:LDR

[IR98] Irina Ignatiouk-Robert. Large deviations for a random walk in dynamical random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(5):601–636, September/October 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_5_601_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800022.

Ishikawa:1997:LDE

[Ish97] Yasushi Ishikawa. Large deviation estimate of transition densities for jump processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(2):179–222, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_2_179_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801215.

Istas:1992:WCG

[Ist92] Jacques Istas. Wavelet coefficients of a Gaussian process and applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(4):537–556, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1992__28_4_537_0.

Sato:2004:MLE

[iSW04] Ken iti Sato and Toshiro Watanabe. Moments of last exit times for Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2):207–225, March/April 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_2_207_0;   http://www.sciencedirect.com/science/
article/pii/S024602030300044X.

Sato:2005:LET

[iSW05] Ken iti Sato and Toshiro Watanabe. Last exit times for transient semistable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5):929–951, September/October 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_5_929_0;   http://
www.sciencedirect.com/science/article/pii/S024602030400086X.

Iyer:2020:TVI

[IY20] Srikanth K. Iyer and D. Yogeshwaran. Thresholds for vanishing of ‘isolated’ faces in random Cech and Vietoris– Rips complexes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1869–1897, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Thresholds-for-vanishing-
of-Isolated-faces-in-random-%c4%8cech-and/10.1214/19-
AIHP1020.full.

Jacod:1971:TRC

[Jac71a] J. Jacod. Théorème de renouvellement et classification pour les chaînes semi-markoviennes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(2):83–129, April–June 1971. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1974__10_2_201_0. See corrections [Jac74].

Jacod:1971:FMS

[Jac71b] Jean Jacod. Fonctionnelles multiplicatives et sous-processus des P -processus. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(4):299–325, October–December 1971. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1971__7_4_299_0.

Jacod:1971:GIP

[Jac71c] Jean Jacod. Générateurs infinitésimaux des processus à accroissements semi-markoviens. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(3):219–233, July–September 1971. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1971__7_3_219_0.

Jacod:1973:SGM

[Jac73] Jean Jacod. Semi-groupes et mesures invariantes pour les processus semi-markoviens à espace d’état quelconque. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(1):77–112, January– March 1973. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1973__9_1_77_0.

Jacod:1974:CCA

[Jac74] Jean Jacod. Corrections et compléments à l’article: ^{≪} Théorème de renouvellement et classification pour les chaînes semi-markoviennes ^{≫}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(2):201–209, April–June 1974. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1971__7_2_83_0. See [Jac71a].

Jacob:1979:CUD

[Jac79] P. Jacob. Convergence uniforme à distance finie de mesures signées. (French) [Uniform finite-distance convergence for signed measures]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(4):355–373, October–December 1979. CODEN AHPBAR.

ISSN 0020-2347. URL http://www.numdam.org/item?id=
AIHPB_1979__15_4_355_0.

Jacob:1982:RCI

[Jac82] Pierre Jacob. Représentations convergentes des intégrales stochastiques. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(3):249–275, July– September 1982. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1982__18_3_
249_0.

Jacod:1990:PVP

[Jac90] Jean Jacod. Sur le processus de vraisemblance partielle. (French) [On the process of partial likelihood]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(2): 299–329, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1990__26_2_299_0.

Jacod:1998:RCL

[Jac98] Jean Jacod. Rates of convergence to the local time of a diffusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(4):505–544, July/August 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_4_505_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800265.

Jacob:2010:EIB

[Jac10] Emmanuel Jacob. Excursions of the integral of the Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3):869–887, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1281100402;   http://
www.numdam.org/item?id=AIHPB_2010__46_3_869_0.

Jackson:2018:LBM

[Jac18] Henry Jackson. Liouville Brownian motion and thick points of the Gaussian free field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):249–279, February 2018. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1519030828.

Jaffuel:2012:CBS

[Jaf12] Bruno Jaffuel. The critical barrier for the survival of branching random walk with absorption. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):989–1009, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1353098437.

Jaibi:1980:EFA

[Jai80] M. R. Jaibi. Évolution d’une file d’attente avec priorité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(3): 211–223, July–September 1980. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1980__16_3_211_0.

Jakubowski:1986:ST

[Jak86] Adam Jakubowski. On the Skorokhod topology. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(3): 263–285, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1986__22_3_263_0.

Janssen:1970:GDC

[Jan70] Jacques Janssen. Sur une généralisation du concept de promenade aléatoire sur la droite réelle. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(3):249–269, July–September 1970. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1970__6_3_249_0.

Janssen:2018:JER

[JD18] Anja Janßen and Holger Drees. Joint exceedances of random products. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):437–465, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1519030835.

Janvresse:2004:PAT

[JdlR04] Élise Janvresse and Thierry de la Rue. The Pascal adic transformation is loosely Bernoulli. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2):133–139, March/ April 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2004__40_2_133_0;   http://www.sciencedirect.
com/science/article/pii/S0246020303000621.

Janvresse:2012:ZKE

[JdlR12] Élise Janvresse and Thierry de la Rue. Zero Krengel entropy does not kill Poisson entropy. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):368–376, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1334148203.

Jeulin:1993:PGV

[Jeu93] T. Jeulin. Processus gaussiens à variation finie. (French) [Gaussian processes with finite variation]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(1):153–160, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1993__29_1_153_0.

Jiao:2012:ZBT

[Jia12] Ying Jiao. Zero bias transformation and asymptotic expansions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):258–281, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1327328022.

Jin:2014:UDR

[Jin14] Xiong Jin. A uniform dimension result for two-dimensional fractional multiplicative processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):512–523, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1395856138.

Jirina:1970:SPS

[Jiř70] Miloslav Jiřina. A simplified proof of the Sevastyanov theorem on branching processes. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(1):1–7, January–March 1970. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1970__
6_1_1_0.

Jirina:1978:BR

[Jir78] Miloslav Jirina. A biased roulette. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(1):1–23, January–March 1978. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_1_1_0.

Johnson:2018:SOF

[JJ18] Tobias Johnson and Matthew Junge. Stochastic orders and the frog model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):1013–1030, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1524643238.

Jarvenpaa:2014:HDA

[JJK^{+}14] Esa Järvenpää, Maarit Järvenpää, Henna Koivusalo, Bing Li, and Ville Suomala. Hausdorff dimension of affine random covering sets in torus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1371–1384, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1413555504.

Jacod:1982:TLC

[JKM82] J. Jacod, A. Klopotowski, and J. Mémin. Théorème de la limite centrale et convergence fonctionnelle vers un processus à accroissements indépendants: la méthode des martingales. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 18(1): 1–45, January–March 1982. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1982__
18_1_1_0.

Jacod:2005:AEM

[JKMP05] Jean Jacod, Thomas G. Kurtz, Sylvie Méléard, and Philip Protter. The approximate Euler method for Lévy driven stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):523–558, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_523_0;   http://www.sciencedirect.com/science/
article/pii/S024602030500021X.

Jara:2006:NCL

[JL06] M. D. Jara and C. Landim. Nonequilibrium central limit theorem for a tagged particle in symmetric simple exclusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5):567–577, September/October 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_5_567_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000956.

Jara:2008:QNE

[JL08] M. D. Jara and C. Landim. Quenched non-equilibrium central limit theorem for a tagged particle in the exclusion process with bond disorder. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2):341–361, April 2008. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1207948223;   http://www.numdam.org/item?id=AIHPB_2008_
_44_2_341_0.

Jara:2013:NFT

[JLS13] Milton Jara, Claudio Landim, and Sunder Sethuraman. Nonequilibrium fluctuations for a tagged particle in onedimensional sublinear zero-range processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3): 611–637, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1372772638.

Jara:2021:DVB

[JLT21] M. Jara, C. Landim, and K. Tsunoda. Derivation of viscous Burgers equations from weakly asymmetric exclusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):169–194, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/Derivation-of-viscous-
Burgers-equations-from-weakly-asymmetric-exclusion-
processes/10.1214/20-AIHP1075.full.

Johnstone:1993:AME

[JM93] Iain M. Johnstone and K. Brenda Macgibbon. Asymptotically minimax estimation of a constrained Poisson vector via polydisc transforms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(2):289–319, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_2_289_0.

Jourdain:1998:PCF

[JM98] B. Jourdain and S. Méléard. Propagation of chaos and fluctuations for a moderate model with smooth initial data. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(6):727–766, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_6_727_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399800028.

Jaramillo:2019:FLT

[JN19] Arturo Jaramillo and David Nualart. Functional limit theorem for the self-intersection local time of the fractional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):480–527, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1547802407.

Johansson:2012:UCH

[Joh12] Kurt Johansson. Universality for certain Hermitian Wigner matrices under weak moment conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):47–79, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1327328014.

Johnson:2017:DLS

[Joh17] Oliver Johnson. A discrete log-Sobolev inequality under a Bakry–Émery type condition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1952–1970, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773733.

Jara:2017:HLS

[JP17] Milton Jara and Jonathon Peterson. Hydrodynamic limit for a system of independent, sub-ballistic random walks in a common random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1747–1792, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1511773725.

Janvresse:2010:ASG

[JRdlR10] Élise Janvresse, Benoît Rittaud, and Thierry de la Rue. Almostsure growth rate of generalized random Fibonacci sequences. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):135–158, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1267454112;   http:/
/www.numdam.org/item?id=AIHPB_2010__46_1_135_0.

Jacod:1996:JMP

[JS96] J. Jacod and A. V. Skorokhod. Jumping Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(1):11–67, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1996__32_1_11_0.

Jonckheere:2020:LLN

[JS20] Matthieu Jonckheere and Santiago Saglietti. On laws of large numbers in L^{2} for supercritical branching Markov processes beyond λ-positivity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):265–295, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/On-laws-of-large-numbers-
in-L2-for-supercritical-branching/10.1214/19-AIHP961.
full.

Jaccottet:1976:MAF

[JT76] Monique Jaccottet and Richard Tomassone. Méthodes d’analyse factorielle en théorie des plans d’expérience. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(3):233–256, July–September 1976. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1976__12_3_
233_0.

Juillet:2016:SSP

[Jui16] Nicolas Juillet. Stability of the shadow projection and the left-curtain coupling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1823–1843, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1479373250.

Jumarie:1980:NMI

[Jum80] Guy Jumarie. Sur une nouvelle méthode d’introduction des facteurs de subjectivité dans les problèmes d’estimation statistique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(4):349–369, October–December 1980. CODEN AHPBAR.

ISSN 0020-2347. URL http://www.numdam.org/item?id=
AIHPB_1980__16_4_349_0.

Johnson:2007:SRC

[JV07] Oliver Johnson and Christophe Vignat. Some results concerning maximum Rényi entropy distributions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(3): 339–351, May/June 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_3_339_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000665.

Jin:2005:CTM

[JYZ05] Hanqing Jin, Jia-An Yan, and Xun Yu Zhou. Continuoustime mean–risk portfolio selection. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):559–580, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_3_559_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000221.

Kahane:1987:MAD

[Kah87] Jean-Pierre Kahane. Multiplications aléatoires et dimensions de Hausdorff. (French) [Random multiplications and Hausdorff dimensions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2):289–296, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1987__23_S2_289_0.

Kalinin:2020:MIE

[Kal20] Alexander Kalinin. Markovian integral equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):155–174, February 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Markovian-integral-equations/10.1214/
19-AIHP958.full.

Kannan:1972:OVSa

[Kan72a] D. Kannan. An operator-valued stochastic integral, II. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(1):9–32, January–March 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1972__8_3_217_0.

Kannan:1972:OVSb

[Kan72b] D. Kannan. An operator-valued stochastic integral, III. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(3):217–228, July–September 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1972__8_1_9_0.

Kasprzak:2020:FAS

[Kas20] Mikołaj J. Kasprzak. Functional approximations via Stein’s method of exchangeable pairs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2540–2564, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/Functional-approximations-
via-Steins-method-of-exchangeable-pairs/10.1214/20-
AIHP1049.full.

Kazamaki:1971:SPM

[Kaz71] Norihiko Kazamaki. Some properties of martingale integrals. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(1):9–19, January– March 1971. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1971__7_1_9_0.

Kannan:1971:PMS

[KBR71] D. Kannan and A. T. Bharucha-Reid. Probability measures on H_{p} spaces. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(3):205– 217, July–September 1971. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_3_205_0.

Keller:2017:SPD

[Kel17] Gerhard Keller. A simpler proof for the dimension of the graph of the classical Weierstrass function. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):169–181, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544887.

Kenyon:1997:LSL

[Ken97] Richard Kenyon. Local statistics of lattice dimers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33 (5):591–618, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1997__33_5_591_0;   http://
www.sciencedirect.com/science/article/pii/S0246020397801069.

Kesten:1986:SBR

[Kes86] Harry Kesten. Subdiffusive behavior of random walk on a random cluster. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(4):425–487, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1986__22_4_425_0.

Kaimanovich:1998:PFH

[KF98] Vadim A. Kaimanovich and Albert Fisher. A Poisson formula for harmonic projections. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(2):209–216, ???? 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_2_209_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800307.

Kohatsu-Higa:2012:MCM

[KHT12] Arturo Kohatsu-Higa and Akihiro Tanaka. A Malliavin Calculus method to study densities of additive functionals of SDE’s with irregular drifts. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):871–883, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1340714876.

Kifer:2014:NLT

[Kif14] Yuri Kifer. Nonconventional limit theorems in averaging. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):236–255, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1388545273.

Kigami:2013:TNC

[Kig13] Jun Kigami. Transitions on a noncompact Cantor set and random walks on its defining tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1090–1129, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1380718739.

Kipnis:1974:MSG

[Kip74] Claude Kipnis. Majoration des semi-groupes de contractions de L_{1} et applications. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(4):369–384, October–December 1974. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1974__10_4_369_0.

Kipnis:1987:FTO

[Kip87] C. Kipnis. Fluctuations des temps d’occupation d’un site dans l’exclusion simple symétrique. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(1):21–35, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1987__23_1_21_0.

Kannan:1975:CGM

[KK75] D. Kannan and Pl. Kannappan. On a characterization of Gaussian measures in a Hilbert space. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(4):397–404, October–December 1975. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1975__11_4_397_0.

Komorowski:2003:LLN

[KK03] Tomasz Komorowski and Grzegorz Krupa. The law of large numbers for ballistic, multi-dimensional random walks on random lattices with correlated sites. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(2):263–285, March/ April 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2003__39_2_263_0;   http://www.sciencedirect.
com/science/article/pii/S024602030200002X.

Knobloch:2014:SHF

[KK14] Robert Knobloch and Andreas E. Kyprianou. Survival of homogeneous fragmentation processes with killing. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2): 476–491, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1395856136.

Kieburg:2019:PRM

[KK19] Mario Kieburg and Holger Kösters. Products of random matrices from polynomial ensembles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):98–126, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1547802396.

Kallianpur:1985:ASF

[KKK85] G. Kallianpur, D. Kannan, and R. L. Karandikar. Analytic and sequential Feynman integrals on abstract Wiener and Hilbert spaces, and a Cameron–Martin formula. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(4): 323–361, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1985__21_4_323_0.

Kassel:2015:RTC

[KKW15] Adrien Kassel, Richard Kenyon, and Wei Wu. Random two-component spanning forests. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1457–1464, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1445432048.

Kipnis:1995:GDP

[KL95] C. Kipnis and C. Léonard. Grandes déviations pour un système hydrodynamique asymétrique de particules indépendantes.

(French) [Large deviations for asymmetric hydrodynamic system of independent particles]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):223–248, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1995__31_1_223_0.

Khoshnevisan:1996:CLI

[KL96] Davar Khoshnevisan and Thomas M. Lewis. Chung’s law of the iterated logarithm for iterated Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(3): 349–359, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_3_349_0.

Kwiatkowski:2002:FRT

[KL02] Jan Kwiatkowski and Yves Lacroix. Finite rank transformation and weak closure theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(2):155–192, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_2_155_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301010913.

Knopfel:2005:FSI

[KL05a] Holger Knöpfel and Matthias Löwe. Fluctuations in a p-spin interaction model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):807–815, July/August 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_4_807_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000664.

Kondratiev:2005:GDC

[KL05b] Yuri Kondratiev and Eugene Lytvynov. Glauber dynamics of continuous particle systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):685–702, July/August 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_4_685_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000597.

Kyprianou:2010:RLP

[KL10] A. E. Kyprianou and R. L. Loeffen. Refracted Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):24–44, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1267454106;   http://
www.numdam.org/item?id=AIHPB_2010__46_1_24_0.

Koltchinskii:2016:ACB

[KL16] Vladimir Koltchinskii and Karim Lounici. Asymptotics and concentration bounds for bilinear forms of spectral projectors of sample covariance. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1976–2013, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1479373255.

Kouritzin:2020:LTL

[KL20] Michael A. Kouritzin and Khoa Lê. Long-time limits and occupation times for stable Fleming–Viot processes with decaying sampling rates. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2595–2620, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Long-time-limits-and-
occupation-times-for-stable-FlemingViot-processes/10.
1214/20-AIHP1051.full.

Kliem:2009:DSD

[Kli09] Sandra Kliem. Degenerate stochastic differential equations for catalytic branching networks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):943–980, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1257529887;   http://www.numdam.org/
item?id=AIHPB_2009__45_4_943_0.

Komjathy:2021:PTH

[KLL21] Júlia Komjáthy, John Lapinskas, and Johannes Lengler. Penalising transmission to hubs in scale-free spatial random graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):1968–2016, November 2021.

CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/Penalising-transmission-
to-hubs-in-scale-free-spatial-random-graphs/10.1214/
21-AIHP1149.full.

Kyprianou:2012:SSB

[KLMSR12] A. E. Kyprianou, R.-L. Liu, A. Murillo-Salas, and Y.-X. Ren. Supercritical super-Brownian motion with a general branching mechanism and travelling waves. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):661–687, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1340714867.

Kipnis:1995:MPS

[KLO95] C. Kipnis, C. Landim, and S. Olla. Macroscopic properties of a stationary non-equilibrium distribution for anon-gradient interacting particle system. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):191–221, ???? 1995. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1995_
_31_1_191_0.

Kabluchko:2021:FPF

[KLS21] Zakhar Kabluchko, Matthias Löwe, and Kristina Schubert. Fluctuations for the partition function of Ising models on Erdős–Rényi random graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2017–2042, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/Fluctuations-for-
the-partition-function-of-Ising-models-on-Erd%c3%b6sR%
c3%a9nyi/10.1214/20-AIHP1137.full.

Kesten:1999:SOL

[KM99] Harry Kesten and R. A. Maller. Stability and other limit laws for exit times of random walks from a strip or a halfplane. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(6):685–734, November 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_

_35_6_685_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001089.

Koranyi:2002:AME

[KM02] Adam Korányi and K. Brenda MacGibbon. Asymptotically minimax estimation of order-constrained parameters and eigenfunctions of the Laplacian on the ball. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(2):193–206, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_2_193_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301010986.

Kosygina:2011:LLT

[KM11] Elena Kosygina and Thomas Mountford. Limit laws of transient excited random walks on integers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2): 575–600, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887283;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_575_0.

Kusuoka:2014:SPT

[KM14] Seiichiro Kusuoka and Carlo Marinelli. On smoothing properties of transition semigroups associated to a class of SDEs with jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1347–1370, November 2014. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1413555503.

Konig:2017:MFI

[KM17] Wolfgang König and Chiranjib Mukherjee. Mean-field interaction of Brownian occupation measures, I: Uniform tube property of the Coulomb functional. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):2214–2228, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773743.

Klenke:2020:IRS

[KM20] Achim Klenke and Leonid Mytnik. Infinite rate symbiotic branching on the real line: The tired frogs model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):847–883, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-56/issue-
2/Infinite-rate-symbiotic-branching-on-the-real-line-
-The/10.1214/19-AIHP986.full.

Kutoyants:1992:EPP

[KMB92] Yu. A. Kutoyants, T. Mourid, and D. Bosq. Estimation paramétrique d’un processus de diffusion avec retards. (French) [Parametric estimation of a diffusion process with delays]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(1):95–106, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1992__28_1_95_0.

Konakov:2010:EPL

[KMM10] Valentin Konakov, Stéphane Menozzi, and Stanislav Molchanov. Explicit parametrix and local limit theorems for some degenerate diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):908–923, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1288878329;   http://www.numdam.org/item?id=AIHPB_2010_
_46_4_908_0.

Komjathy:2014:UMT

[KMP14] Júlia Komjáthy, Jason Miller, and Yuval Peres. Uniform mixing time for random walk on lamplighter graphs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (4):1140–1160, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555494.

Kwapien:2006:TRC

[KMR06] Stanisław Kwapień, Michael B. Marcus, and Jan Rosiński. Two results on continuity and boundedness of stochastic convolutions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5):553–566, September/October 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_5_553_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000944.

Kondah:1997:VCV

[KMS97] Abdelaziz Kondah, Véronique Maume, and Bernard Schmitt. Vitesse de convergence vers l’état d’équilibre pour des dynamiques markoviennes non höldériennes. (French) [Speed of convergence to a steady state for non-Hölderian Markov dynamics]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(6):675–695, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_6_675_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801094.

Keller:2014:GII

[KMS14] Nathan Keller, Elchanan Mossel, and Arnab Sen. Geometric influences II: Correlation inequalities and noise sensitivity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1121–1139, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555493.

Kobylanski:2013:LDP

[Kob13] Magdalena Kobylanski. Large deviations principle by viscosity solutions: The case of diffusions with oblique Lipschitz reflections. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):160–181, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1359470130.

Koldobsky:1992:GLR

[Kol92] Alexander L. Koldobsky. Generalized Lévy representation of norms and isometric embeddings into L_{p}-spaces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(3): 335–353, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1992__28_3_335_0.

Koltchinskii:2003:BMD

[Kol03] Vladimir Koltchinskii. Bounds on margin distributions in learning problems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):943–978, November/December 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_

_39_6_943_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000232.

Koltchinskii:2009:SPE

[Kol09] Vladimir Koltchinskii. Sparsity in penalized empirical risk minimization. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):7–57, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1234469970;   http://
www.numdam.org/item?id=AIHPB_2009__45_1_7_0.

Komorowski:1992:APS

[Kom92] T. Komorowski. Asymptotic periodicity of some stochastically perturbed dynamical systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2):165–178, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1992__28_2_165_0.

Komoriya:1998:HLA

[Kom98] Keishi Komoriya. Hydrodynamic limit for asymmetric mean zero exclusion processes with speed change. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(6): 767–797, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_6_767_0;   http://
www.sciencedirect.com/science/article/pii/S024602039980003X.

Kopf:1982:NNT

[Kop82] Christoph Kopf. Negative nonsingular transformations. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 18(1):81–102, January– March 1982. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1982__18_1_81_0.

Kortchemski:2015:LTC

[Kor15] Igor Kortchemski. Limit theorems for conditioned non-generic Galton–Watson trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):489–511, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1428672678.

Kozma:2016:ERW

[KOS16] Gady Kozma, Tal Orenshtein, and Igor Shinkar. Excited random walk with periodic cookies. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1023–1049, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723509.

Kuntz:2019:DLR

[KOS19] Juan Kuntz, Michela Ottobre, and Andrew M. Stuart. Diffusion limit for the random walk Metropolis algorithm out of stationarity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1599–1648, August 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1569398880.

Kozma:2011:TOT

[Koz11] Gady Kozma. The triangle and the open triangle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1): 75–79, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170230;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_75_0.

Kozlov:1996:DED

[KP96] Serguei M. Kozlov and Andrei L. Piatnitski. Degeneration of effective diffusion in the presence of periodic potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(5):571–587, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_5_571_0.

Kyprianou:2011:CTT

[KP11] A. E. Kyprianou and P. Patie. A Ciesielski–Taylor type identity for positive self-similar Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3): 917–928, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834863;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_917_0.

Konev:2012:ERN

[KP12] Victor Konev and Serguei Pergamenshchikov. Efficient robust nonparametric estimation in a semimartingale regression model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1217–1244, November 2012. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1353098446.

Kosygina:2017:ERW

[KP17] Elena Kosygina and Jonathon Peterson. Excited random walks with Markovian cookie stacks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1458–1497, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1500624048.

Kulik:2019:NGL

[KP19] Alexei Kulik and Ilya Pavlyukevich. Non-Gaussian limit theorem for non-linear Langevin equations driven by Lévy noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1278–1315, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1569398870.

Kurtz:1995:SSD

[KPP95] Thomas G. Kurtz, Étienne Pardoux, and Philip Protter. Stratonovich stochastic differential equations driven by general semimartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(2):351–377, ???? 1995. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1995_
_31_2_351_0.

Kyprianou:2018:UCF

[KPR18] A. E. Kyprianou, S. W. Pagett, and T. Rogers. Universality in a class of fragmentation-coalescence processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):1134–1151, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1524643242.

Karatzas:2016:SBP

[KPS16] Ioannis Karatzas, Soumik Pal, and Mykhaylo Shkolnikov. Systems of Brownian particles with asymmetric collisions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1): 323–354, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1452089271.

Kabluchko:2020:STL

[KPT20] Zakhar Kabluchko, Joscha Prochno, and Christoph Thäle. Sanov-type large deviations in Schatten classes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):928–953, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-56/issue-
2/Sanov-type-large-deviations-in-Schatten-classes/10.
1214/19-AIHP989.full.

Karatzas:2016:PSS

[KR16] Ioannis Karatzas and Johannes Ruf. Pathwise solvability of stochastic integral equations with generalized drift and non-smooth dispersion functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):915–938, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1462367900.

Krengel:1976:RRA

[Kre76] Ulrich Krengel. On Rudolph’s representation of aperiodic flows. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(4):319– 338, October–December 1976. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1976__
12_4_319_0.

Karoui:1970:PTP

[KRR70] N. Karoui, H. Reinhard, and B. Roynette. Processus tués de processus de Hunt conservatifs et prolongements de processus standards. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(3):201–236, July–September 1970. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1970__6_3_201_0.

Khoshnevisan:2004:ELT

[KRS04] Davar Khoshnevisan, Pál Révész, and Zhan Shi. On the explosion of the local times along lines of Brownian sheet. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):1–24, January/February 2004. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_2004__40_1_
1_0;   http://www.sciencedirect.com/science/article/
pii/S0246020303000578.

Kyprianou:2018:DFS

[KRS18] Andreas E. Kyprianou, Victor Rivero, and Bati Sengül. Deep factorisation of the stable process II: Potentials and applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):343–362, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1519030831.

Kyprianou:2021:SLP

[KRS21] Andreas E. Kyprianou, Victor Rivero, and Weerapat Satitkanitkul. Stable Lévy processes in a cone. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2066–2099, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/Stable-L%c3%a9vy-
processes-in-a-cone/10.1214/20-AIHP1143.full.

Krokowski:2016:BEB

[KRT16] Kai Krokowski, Anselm Reichenbachs, and Christoph Thäle. Berry–Esseen bounds and multivariate limit theorems for functionals of Rademacher sequences. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):763–803, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1462367893.

Kious:2021:ORR

[KSS21] Daniel Kious, Bruno Schapira, and Arvind Singh. Once reinforced random walk on Z × Γ. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2219–2242, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/Once-reinforced-random-
walk-on-Z%ce%93/10.1214/21-AIHP1151.full.

Katok:1997:SET

[KT97] Anatole Katok and Jean-Paul Thouvenot. Slow entropy type invariants and smooth realization of commuting measurepreserving transformations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(3):323–338, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1997__33_3_323_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397800945.

Kuksin:2013:WNS

[Kuk13] Sergei B. Kuksin. Weakly nonlinear stochastic CGL equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1033–1056, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1380718736.

Kunita:1986:CSF

[Kun86] Hiroshi Kunita. Convergence of stochastic flows with jumps and Lévy processes in diffeomorphisms group. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(3): 287–321, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1986__22_3_287_0.

Kurtzmann:2010:OMS

[Kur10] Aline Kurtzmann. The ODE method for some self-interacting diffusions on R ^{d}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3):618–643, August 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1281100392;   http://www.numdam.org/item?id=AIHPB_2010_
_46_3_618_0.

Kosloff:2022:LLT

[KV22] Zemer Kosloff and Dalibor Volny. Local limit theorem in deterministic systems. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 58(1):548–566, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Local-limit-theorem-
in-deterministic-systems/10.1214/21-AIHP1169.full.

Kenyon:2020:GFD

[KW20] Richard W. Kenyon and David B. Wilson. The Green’s function on the double cover of the grid and application to the uniform spanning tree trunk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1841–1868, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/The-Greens-function-
on-the-double-cover-of-the-grid/10.1214/19-AIHP1019.
full.

Kyprianou:2004:TWS

[Kyp04] A. E. Kyprianou. Travelling wave solutions to the K-P-P equation: alternatives to Simon Harris’ probabilistic analysis. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):53–72, January/February 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_1_53_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000554.

Kovchegov:2017:HSS

[KZ17] Yevgeniy Kovchegov and Ilya Zaliapin. Horton self-similarity of Kingman’s coalescent tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1069–1107, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1500624031.

Koltchinskii:2021:EES

[KZ21] Vladimir Koltchinskii and Mayya Zhilova. Efficient estimation of smooth functionals in Gaussian shift models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):351–386, February 2021. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-1/Efficient-estimation-of-smooth-functionals-
in-Gaussian-shift-models/10.1214/20-AIHP1081.full.

Labbe:2014:GFC

[Lab14] Cyril Labbé. Genealogy of flows of continuous-state branching processes via flows of partitions and the Eve property. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3): 732–769, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1403276997.

Lacaze:1973:EMF

[Lac73] M. B. Lacaze. Estimation de moyennes et fonctions de répartition de suites d’échantillonnage. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(2):145–165, April–June 1973. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1973__9_2_145_0.

Lachal:1991:PIP

[Lac91] A. Lachal. Sur le premier instant de passage de l’intégrale du mouvement brownien. (French) [On the first passage time of the integral of Brownian motion]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(3):385–405, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_3_385_0.

Lacey:1997:RTT

[Lac97a] Michael T. Lacey. The return time theorem fails on infinite measure-preserving systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(4):491–495, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1997__33_4_491_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397801021.

Lachal:1997:TPS

[Lac97b] Aimé Lachal. Les temps de passage successifs de l’intégrale du mouvement brownien. (French) [The passage of successive times in the integral of Brownian motion]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(1):1– 36, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1997__33_1_1_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397801148.

Lacaux:2004:RHM

[Lac04] Céline Lacaux. Real harmonizable multifractional Lévy motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):259–277, May/June 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_3_259_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000645.

Lacour:2007:AET

[Lac07] Claire Lacour. Adaptive estimation of the transition density of a Markov chain. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(5):571–597, September/October 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_5_571_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000720.

Lacoin:2012:SBMa

[Lac12a] Hubert Lacoin. Superdiffusivity for Brownian motion in a Poissonian potential with long range correlation I: Lower bound on the volume exponent. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1010–1028, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1353098438.

Lacoin:2012:SBMb

[Lac12b] Hubert Lacoin. Superdiffusivity for Brownian motion in a Poissonian potential with long range correlation II: Upper bound on the volume exponent. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1029–1048, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1353098439.

Lacoin:2017:SEP

[Lac17] Hubert Lacoin. The simple exclusion process on the circle has a diffusive cutoff window. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1402–1437, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1500624046.

Lafon:1972:GCF

[Laf72] Monique Lafon. Géométries combinatoires finies. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(4):307–317, October– December 1972. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1972__8_4_307_0.

Lalley:1998:PFG

[Lal98] Steven P. Lalley. Percolation on Fuchsian groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34 (2):151–177, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_2_151_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800228.

Lambert:2000:CAL

[Lam00] A. Lambert. Completely asymmetric Lévy processes confined in a finite interval. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(2):251–274, March 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_2_251_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001266.

Lamarre:2020:CRT

[Lam20] Pierre Yves Gaudreau Lamarre. On the convergence of random tridiagonal matrices to stochastic semigroups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2686–2731, November 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-4/On-the-convergence-of-random-tridiagonal-
matrices-to-stochastic-semigroups/10.1214/20-AIHP1055.
full.

Lambert:2021:PSG

[Lam21] Gaultier Lambert. Poisson statistics for Gibbs measures at high temperature. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):326–350, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/Poisson-statistics-for-
Gibbs-measures-at-high-temperature/10.1214/20-AIHP1080.
full.

Lanery:1982:SBT

[Lan82] E. Lanery. Solutions bayésiennes en théorie de la décision statistique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 18 (1):55–79, January–March 1982. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1982__18_1_55_0.

Landim:1991:HLA

[Lan91] C. Landim. Hydrodynamical limit for asymmetric attractive particle systems on Z^{d}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(4):559–581, ???? 1991. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1991_
_27_4_559_0.

Lapeyre:1986:EGD

[Lap86] H. Lapeyre. Estimations de grandes déviations pour des systèmes où apparaissent un bruit gaussien et un bruit non gaussien. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(1):9–17, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1986__22_1_
9_0.

Larisse:1972:MAHa

[Lar72a] Jean Larisse. Marches au hasard sur les demi-groupes discrets, I. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(2): 107–125, April–June 1972. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1972__
8_2_127_0.

Larisse:1972:MAHb

[Lar72b] Jean Larisse. Marches au hasard sur les demi-groupes discrets, II. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(2): 127–173, April–June 1972. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1972__
8_3_229_0.

Larisse:1972:MAHc

[Lar72c] Jean Larisse. Marches au hasard sur les demi-groupes discrets, III. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(3):229– 240, July–September 1972. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1972__
8_2_127_0.

Laroche:1993:ICF

[Lar93] Étienne Laroche. Inégalités de corrélation sur {−1, 1}^{n} et dans R^{n}. (French) [Correlation inequalities {−1, 1}^{n} and R^{n}]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(4):531–567, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1993__29_4_531_0.

Luedtke:2020:PGP

[LC20] Alex Luedtke and Antoine Chambaz. Performance guarantees for policy learning. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2162–2188, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Performance-guarantees-
for-policy-learning/10.1214/19-AIHP1034.full.

Lazaro-Cami:2009:SRS

[LCO09] Joan-Andreu Lázaro-Camí and Juan-Pablo Ortega. Superposition rules and stochastic Lie–Scheffers systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4): 910–931, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1257529885;   http://www.numdam.org/
item?id=AIHPB_2009__45_4_910_0.

LeCam:1985:AFM

[Le 85] Lucien Le Cam. Sur l’approximation de familles de mesures par des familles gaussiennes. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(3):225–287, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1985__21_3_225_0.

LePage:1989:RDP

[Le 89] Émile Le Page. Régularité du plus grand exposant caractéristique des produits de matrices aléatoires indépendantes et applications. (French) [Regularity of the greatest characteristic exponent of products of independent random matrices and applications]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2):109–142, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1989__25_2_109_0.

LeJan:1991:SPL

[Le 91] Y. Le Jan. Superprocesses and projective limits of branching Markov process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(1):91–106, ???? 1991. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1991__27_1_91_0.

LeBorgne:2002:PIP

[Le 02] Stéphane Le Borgne. Principes d’invariance pour les flots diagonaux sur SL(d, R)/SL(d, Z). (French) [Invariance principles for diagonal waves on SL(d, R)/SL(d, Z)]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(4):581–612, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_4_581_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301011050.

LeGall:2019:BDB

[Le 19] Jean-François Le Gall. Brownian disks and the Brownian snake. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):237–313, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1547802401.

Leandre:1985:RPS

[Léa85] R. Léandre. Régularité de processus de sauts dégénérés. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(2):125–146, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1985__21_2_125_0.

Leandre:1988:RPS

[Léa88] Rémi Léandre. Régularité de processus de sauts dégénérés. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):209–236, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1988__24_2_209_0.

Ledrappier:1970:PRC

[Led70] François Ledrappier. Des produits de Riesz comme mesures spectrales. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(4):335–344, October–December 1970. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1970__6_4_335_0.

Ledrappier:1978:GSF

[Led78] F. Ledrappier. En général, un semi-flot spécial est exact. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4): 465–478, October–December 1978. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1978__14_4_465_0.

Ledoux:1981:CMF

[Led81] Michel Ledoux. Classe L log L et martingales fortes à paramètre bidimensionnel. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(3):275–280, July–September 1981. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1981__17_3_275_0.

Ledoux:1983:TLC

[Led83] M. Ledoux. Théorème limite central dans les espaces l_{p}(B)(1 ≤ p < ∞). (French) []. Annales de l’Institut Henri Poincaré.

Probabilités et Statistiques, 19(4):393–411, ???? 1983. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1983_
_19_4_393_0.

Ledoux:1992:DMS

[Led92] Michel Ledoux. Sur les déviations modérées des sommes de variables aléatoires vectorielles indépendantes de même loi. (French) [On moderate deviations of sums of independent random vector variables of the same law]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2):267–280, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1992__28_2_267_0.

Ledoux:2009:RFM

[Led09] M. Ledoux. A recursion formula for the moments of the Gaussian orthogonal ensemble. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):754–769, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1249391383;   http://www.numdam.org/item?id=AIHPB_2009_
_45_3_754_0.

Leguesdron:1989:MAS

[Leg89] J. P. Leguesdron. Marche aléatoire sur le semi-groupe des contractions de R^{d}. Cas de la marche aléatoire sur R_{+} avec choc élastique en zéro. (French) [Random walk on the semi-group of contractions of R^{d}. Case of the random walks on R_{+} with elastic shock in zero]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):483–502, ???? 1989. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1989_
_25_4_483_0.

Lehec:2013:RFE

[Leh13] Joseph Lehec. Representation formula for the entropy and functional inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):885–899, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1372772648.

Lemaire:1973:NDEa

[Lem73a] Jean Lemaire. Une nouvelle définition de l’équitabilité pour les jeux de hasard. Application au paradoxe de Saint-Pétersbourg. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(2): 193–204, April–June 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_2_205_0.

Lemaire:1973:NDEb

[Lem73b] Jean Lemaire. Une nouvelle définition de l’équitabilité pour les jeux de hasard. Application au paradoxe de Saint-Pétersbourg. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(2): 205–214, April–June 1973. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1973__
9_2_193_0.

Lemanczyk:1988:TE

[Lem88] Mariusz Lemańczyk. Toeplitz Z_{2}-extensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(1):1– 43, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1988__24_1_1_0.

Lemler:2016:OIL

[Lem16] Sarah Lemler. Oracle inequalities for the Lasso in the highdimensional Aalen multiplicative intensity model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2): 981–1008, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1462367902.

Lenglart:1977:RDE

[Len77] E. Lenglart. Relation de domination entre deux processus. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(2): 171–179, April–June 1977. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1977__
13_2_171_0.

Lenz:2004:ENU

[Len04] Daniel Lenz. Existence of non-uniform cocycles on uniquely ergodic systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(2):197–206, March/April 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_2_197_0;   http://www.sciencedirect.com/science/
article/pii/S024602030300058X.

Leonard:1986:LGN

[Léo86] Christian Léonard. Une loi des grands nombres pour des systèmes de diffusions avec interaction et à coefficients non bornés. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2):237–262, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1986__22_2_
237_0.

Leonard:1995:LDL

[Léo95] Christian Léonard. Large deviations for long range interacting particle systems with jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(2):289–323, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1995__31_2_289_0.

Lerasle:2012:OMS

[Ler12] Matthieu Lerasle. Optimal model selection in density estimation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):884–908, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1340714877.

Lesigne:1987:TEP

[Les87] Emmanuel Lesigne. Théorèmes ergodiques ponctuels pour des mesures diagonales. Cas des systèmes distaux. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(4):593–612, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1987__23_4_593_0.

Letac:1972:GSP

[Let72] Gérard Letac. Groupe de Stam d’une probabilité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(2):175–181, April–June 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1972__8_2_175_0.

Li:2019:MDP

[Li19] Songzi Li. Matrix Dirichlet processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):909–940, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1557820836.

Liang:2007:BRG

[Lia07] Zongxia Liang. Besov regularity for the generalized local time of the indefinite Skorohod integral. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1): 77–86, January/February 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_1_77_0;   http://
www.sciencedirect.com/science/article/pii/S0246020306000197.

Liao:2016:TST

[Lia16] Ming Liao. A three-series theorem on Lie groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52 (2):726–734, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1462367891.

Lichnerowicz:1969:LDF

[Lic69] Marc Lichnerowicz. Sur la loi de demande en fonction des prix. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(3):255– 273, July–September 1969. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1969__
5_3_255_0.

Lichnerowicz:1970:MEE

[Lic70]  Marc Lichnerowicz.  Un modèle d’échange économique

(Économie et thermodynamique). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(2):159–200, April–June 1970. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1970__6_2_159_0.

Lifshits:1994:TPG

[Lif94] M. A. Lifshits. Tail probabilities of Gaussian suprema and Laplace transform. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):163–179, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1994__30_2_
163_0.

Limic:2012:GRE

[Lim12] Vlada Limic. Genealogies of regular exchangeable coalescents with applications to sampling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):706–720, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1340714869.

Lin:1975:QCU

[Lin75] Michael Lin. Quasi-compactness and uniform ergodicity of Markov operators. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(4):345–354, October–December 1975. CODEN AHPBAR.

ISSN 0020-2347. URL http://www.numdam.org/item?id=
AIHPB_1975__11_4_345_0.

Lin:1976:SRL

[Lin76] Michael Lin. Strong ratio limit theorems for mixing Markov operators. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(2):181– 191, April–June 1976. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1976__12_2_
181_0.

Lin:1977:EPO

[Lin77] Michael Lin. Ergodic properties of an operator obtained from a continuous representation. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(4):321–331, October–December 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1977__13_4_321_0.

Lin:2017:TBH

[Lin17] Shen Lin. Typical behavior of the harmonic measure in critical Galton–Watson trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):718–752, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1491897743.

Liu:1996:SPB

[Liu96] Quansheng Liu. On the survival probability of a branching process in a random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(1):1–10, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1996_
_32_1_1_0.

Liu:2001:LDB

[Liu01] Quansheng Liu. Local dimensions of the branching measure on a Galton–Watson tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(2):195–222, March 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_2_195_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010657.

Liu:2018:HFS

[Liu18] Zhipeng Liu. Height fluctuations of stationary TASEP on a ring in relaxation time scale. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):1031–1057, May 2018. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1524643239.

leJan:1987:ESP

[lJ87] Y. le Jan. Équilibre statistique pour les produits de difféomorphismes aléatoires indépendants. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(1): 111–120, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1987__23_1_111_0.

Loukianova:2008:UDE

[LL08] D. Loukianova and O. Loukianov. Uniform deterministic equivalent of additive functionals and non-parametric drift estimation for one-dimensional recurrent diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4): 771–786, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1217964119;   http://www.numdam.org/
item?id=AIHPB_2008__44_4_771_0.

Lewko:2015:EAS

[LL15] Allison Lewko and Mark Lewko. An exact asymptotic for the square variation of partial sum processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1597–1619, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1445432054.

Locherbach:2011:PBE

[LLL11] Eva Löcherbach, Dasha Loukianova, and Oleg Loukianov. Polynomial bounds in the ergodic theorem for one-dimensional diffusions and integrability of hitting times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2): 425–449, May 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1300887276;   http://www.numdam.org/
item?id=AIHPB_2011__47_2_425_0.

Locherbach:2014:SCH

[LLL14] Eva Löcherbach, Oleg Loukianov, and Dasha Loukianova. Spectral condition, hitting times and Nash inequality. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (4):1213–1230, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555498.

Lachal:2006:MCR

[LLO06] A. Lachal, S. Leorato, and E. Orsingher. Minimal cyclic random motion in R^{n} and hyper-Bessel functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6):753–772, November/December 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_6_753_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000173.

Loukianov:2011:SGE

[LLS11] Oleg Loukianov, Dasha Loukianova, and Shiqi Song. Spectral gaps and exponential integrability of hitting times for linear diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3):679–698, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1308834855;   http://
www.numdam.org/item?id=AIHPB_2011__47_3_679_0.

Lepeltier:1976:PME

[LM76] J.-P. Lepeltier and B. Marchal. Problème des martingales et équations différentielles stochastiques associées à un opérateur intégro-différentiel. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(1):43–103, January–March 1976. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1976__12_1_43_0.

Lepeltier:1977:EPO

[LM77] J.-P. Lepeltier and B. Marchal. Sur l’existence de politiques optimales dans le contrôle intégro-différentiel. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(1):45–97, January–March 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_1_45_0.

Linde:1983:CSS

[LM83] Werner Linde and Peter Mathé. Conditional symmetries of stable measures on R^{n}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):57–69, ???? 1983. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1983__19_1_57_0.

Landim:1997:HLM

[LM97] C. Landim and M. Mourragui. Hydrodynamic limit of mean zero asymmetric zero range processes in infinite volume. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(1):65–82, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_1_65_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801161.

Liu:2009:SGC

[LM09] Wei Liu and Yutao Ma. Spectral gap and convex concentration inequalities for birth–death processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1): 58–69, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469971;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_58_0.

Li:2011:SSS

[LM11] Zenghu Li and Leonid Mytnik. Strong solutions for stochastic differential equations with jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1055–1067, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1317906501;   http://www.numdam.org/
item?id=AIHPB_2011__47_4_1055_0.

Lecue:2013:OER

[LM13] Guillaume Lecué and Shahar Mendelson. On the optimality of the empirical risk minimization procedure for the convex aggregation problem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):288–306, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1359470136.

Locherbach:2022:MSI

[LM22] Eva Löcherbach and Pierre Monmarché. Metastability for systems of interacting neurons. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):343–378, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Metastability-for-systems-
of-interacting-neurons/10.1214/21-AIHP1164.full.

Liverani:2003:LYM

[LMD03] Carlangelo Liverani and Véronique Maume-Deschamps. Lasota– Yorke maps with holes: conditionally invariant probability measures and invariant probability measures on the survivor set. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(3):385–412, May/June 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_3_385_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000055.

Laurent:2018:MTC

[LMMR18] Béatrice Laurent, Clément Marteau, and Cathy Maugis-Rabusseau. Multidimensional two-component Gaussian mixtures detection. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):842–865, May 2018. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1524643231.

Liang:2021:EES

[LMW21] Mingjie Liang, Mateusz B. Majka, and Jian Wang. Exponential ergodicity for SDEs and McKean–Vlasov processes with Lévy noise. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1665–1701, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Exponential-ergodicity-
for-SDEs-and-McKeanVlasov-processes-with-L%c3%a9vy-
noise/10.1214/20-AIHP1123.full.

Lee:2009:SPS

[LMX09] Kijung Lee, Carl Mueller, and Jie Xiong. Some properties of superprocesses under a stochastic flow. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2): 477–490, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024677;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_477_0.

Lepingle:1989:DSD

[LNS89] Dominique Lépingle, David Nualart, and Marta Sanz. Dérivation stochastique de diffusions réfléchies. (French) [Stochastic derivation of reflected diffusions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(3):283–305, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1989__25_3_283_0.

Locherbach:2002:LLS

[Löc02] Eva Löcherbach. LAN and LAMN for systems of interacting diffusions with branching and immigration. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38 (1):59–90, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_1_59_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010779.

Long:1969:CFC

[Lon69] Ton That Long. Sur le calcul fonctionnel d’une contraction complètement non unitaire. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(2):113–121, April–June 1969. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1969__5_3_213_0.

Long:1981:EMR

[Lon81] Jui Lin Long. Sur l’espace H^{p} de martingales régulières (0 < p ≤ 1). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17 (1):123–142, January–March 1981. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1981__17_1_123_0.

Lootgieter:1977:AAC

[Loo77a] J. C. Lootgieter. La σ-algèbre asymptotique d’une chaîne de Galton–Watson. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(3):193–230, July–September 1977. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1977__13_3_193_0.

Lootgieter:1977:PRC

[Loo77b] J.-C. Lootgieter. Problèmes de récurrence concernant des mouvements aléatoires de particules sur Z avec destruction. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(2): 127–139, April–June 1977. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1977__
13_2_127_0.

Lootgieter:1977:RSK

[Loo77c] J. C. Lootgieter. Sur la répartition des suites de Kakutani (I). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(4): 385–410, October–December 1977. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1978__14_3_279_0.

Lootgieter:1978:RSK

[Loo78] J.-C. Lootgieter. Sur la répartition des suites de Kakutani (II). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14 (3):279–302, July–September 1978. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1977__13_4_385_0.

Lootgieter:1981:CDF

[Loo81] J. C. Lootgieter. Sur une conjecture de L. E. Dubins. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(1):97–122, January–March 1981. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1981__17_1_
97_0.

Laurant:1969:ISM

[LOR69] F. Laurant, M. Oheix, and J.-P. Raoult. Introduction à la statistique mathématique. V. Tests d’hypothèses. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(4):385–414, October– December 1969. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1969__5_4_357_0.

LotiViaud:1985:PBD

[Lot85] Daniel Pierre Loti Viaud. Processus de branchements dépendant de la densité, markovien en temps continu. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(3):289–303, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1985__21_3_289_0.

LotiViaud:1991:GDP

[Lot91] Daniel Pierre Loti Viaud. Grandes déviations pour une famille de processus de Galton–Watson dépendant de l’effectif de la population. (French) [Large deviations for a family of Galton– Watson processes depending on the size of the population]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(2):141–179, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1991__27_2_141_0.

Louchard:1968:MBV

[Lou68] G. Louchard. Mouvement brownien et valeurs propres du laplacien. (French) [Brownian motion and eigenvalues of the Laplacian]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(4):331–342, October–December 1968. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1968__
4_4_331_0.

Louhichi:2000:WCE

[Lou00] Sana Louhichi. Weak convergence for empirical processes of associated sequences. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(5):547–567, September 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_5_547_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001400.

Loulakis:2005:MER

[Lou05] Michail Loulakis. Mobility and Einstein relation for a tagged particle in asymmetric mean zero random walk with simple exclusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(2):237–254, March/April 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_2_237_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000834.

Lamberton:1990:ARF

[LP90] Damien Lamberton and Gilles Pagès. Sur l’approximation des réduites. (French) [On the approximation of reductions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(2):331–355, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1990__26_2_331_0.

LePage:1997:LLT

[LP97] Émile Le Page and Marc Peigné. A local limit theorem on the semi-direct product of R^{∗+} and R^{d}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(2):223–252, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1997__33_2_223_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397801227.

Lucon:2017:LTD

[LP17] E. Luçon and C. Poquet. Long time dynamics and disorderinduced traveling waves in the stochastic Kuramoto model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1196–1240, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1500624040.

Lemaire:2015:IMD

[LPP15] Vincent Lemaire, Gilles Pagès, and Fabien Panloup. Invariant measure of duplicated diffusions and application to Richardson–Romberg extrapolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1562–1596, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1445432053.

Lanconelli:2020:LDC

[LPP20] Alberto Lanconelli, Stefano Pagliarani, and Andrea Pascucci. Local densities for a class of degenerate diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1440–1464, May 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/Local-densities-for-a-class-of-degenerate-
diffusions/10.1214/19-AIHP1009.full.

Littaye-Petit:1969:ISM

[LPPV69] M. Littaye-Petit, J.-L. Piednoir, and B. Van Cutsem. Introduction à la statistique mathématique. II. Exhaustivité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B.

Calcul des Probabilités et Statistique, 5(4):289–322, October– December 1969. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1969__5_4_323_0.

Lyons:2003:MCI

[LPS03] Russell Lyons, Yuval Peres, and Oded Schramm. Markov chain intersections and the loop-erased walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(5): 779–791, September/October 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_5_779_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000335.

Lyons:2020:IGU

[LPS20] Russell Lyons, Yuval Peres, and Xin Sun. Induced graphs of uniform spanning forests. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2732–2744, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Induced-graphs-of-uniform-
spanning-forests/10.1214/20-AIHP1056.full.

Lacey:1994:RET

[LPWR94] Michael Lacey, Karl Petersen, Mate Wierdl, and Dan Rudolph. Random ergodic theorems with universally representative sequences. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(3):353–395, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1994__30_3_353_0.

Landim:2002:SGL

[LPY02] C. Landim, G. Panizo, and H. T. Yau. Spectral gap and logarithmic Sobolev inequality for unbounded conservative spin systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(5):739–777, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_5_739_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011081.

Laredo:1983:GDD

[LR83] Catherine Laredo and Alain Rouault. Grandes déviations, dynamique de populations et phénomènes malthusiens. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(4):323–350, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1983__19_4_323_0.

Letac:1986:FAC

[LR86] G. Letac and Q. I. Rahman. A factorisation of the Askey’s characteristic function (1 − ∥t∥_{2n+1})_{+}^{n+1}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2):169–174, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1986__22_2_169_0.

Lorang:1993:TSP

[LR93] Gérard Lorang and Bernard Roynette. Un théorème de Schilder pour des fonctionnelles browniennes non régulières. (French) [A theorem of Schilder for irregular Brownian functionals]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(4):513–530, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_4_513_0.

Lorang:1996:EFL

[LR96] G. Lorang and B. Roynette. Étude d’une fonctionnelle liée au pont de Bessel. (French) [Study of a functional related to the Bessel bridge]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(1):107–133, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1996__32_1_
107_0.

Jan:2014:TEB

[LR14] Yves Le Jan and Olivier Raimond. Three examples of Brownian flows on R. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1323–1346, November 2014. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1413555502.

Locherbach:2017:EMJ

[LR17] Eva Löcherbach and Victor Rabiet. Ergodicity for multidimensional jump diffusions with position dependent jump rate. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1136–1163, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1500624033.

Lipshutz:2019:PDR

[LR19] David Lipshutz and Kavita Ramanan. Pathwise differentiability of reflected diffusions in convex polyhedral domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1439–1476, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1569398875.

Larsson:2020:CLS

[LR20] Martin Larsson and Johannes Ruf. Convergence of local supermartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2774–2791, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Convergence-of-local-
supermartingales/10.1214/20-AIHP1058.full.

Lakner:2019:RPR

[LRZ19] Peter Lakner, Josh Reed, and Bert Zwart. On the roughness of the paths of RBM in a wedge. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1566–1598, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1569398879.

Ladelli:1991:PAM

[LS91] L. Ladelli and H. Sadi. Propriété asymptotique d’un modèle statistique pour une chaîne de Markov à valeurs dans R^{d}. (French) [Asymptotic property of a statistical model for a Markov chain with values in R^{d}]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(4):519–535, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1991__27_4_519_0.

Lifshits:2005:SDF

[LS05a] Mikhail Lifshits and Thomas Simon. Small deviations for fractional stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):725–752, July/August 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_4_725_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000627.

Lindsay:2005:QSC

[LS05b] J. Martin Lindsay and Adam G. Skalski. Quantum stochastic convolution cocycles I. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):581–604, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_581_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000105.

Liggett:2006:SDC

[LS06] Thomas M. Liggett and Jeffrey E. Steif. Stochastic domination: the contact process, Ising models and FKG measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(2):223–243, March/April 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_2_223_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000658.

Lambert:2014:WCR

[LS14] Amaury Lambert and Florian Simatos. The weak convergence of regenerative processes using some excursion path decompositions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):492–511, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1395856137.

Lalley:2016:MDC

[LS16] Steven P. Lalley and Yuan Shao. Maximal displacement of critical branching symmetric stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3): 1161–1177, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1469723515.

Lamarre:2019:ESB

[LS19a] Pierre Yves Gaudreau Lamarre and Mykhaylo Shkolnikov. Edge of spiked beta ensembles, stochastic Airy semigroups and reflected Brownian motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1402–1438, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1569398874.

Landim:2019:MOD

[LS19b] C. Landim and I. Seo. Metastability of one-dimensional, non-reversible diffusions with periodic boundary conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):1850–1889, November 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-55/issue-4/Metastability-of-one-dimensional-
non-reversible-diffusions-with-periodic-boundary/10.
1214/18-AIHP936.full.

Li:2020:SPD

[LS20] Liping Li and Wenjie Sun. On stiff problems via Dirichlet forms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2051–2080, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/On-stiff-problems-via-Dirichlet-forms/
10.1214/19-AIHP1028.full.

Lawler:2002:VBI

[LSW02] Gregory F. Lawler, Oded Schramm, and Wendelin Werner. Values of Brownian intersection exponents III: Two-sided exponents. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(1):109–123, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_1_109_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010895.

Li:2021:CLT

[LSX21] Yiting Li, Kevin Schnelli, and Yuanyuan Xu. Central limit theorem for mesoscopic eigenvalue statistics of deformed Wigner matrices and sample covariance matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):506–546, February 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-1/Central-limit-theorem-for-mesoscopic-
eigenvalue-statistics-of-deformed-Wigner/10.1214/20-
AIHP1086.full.

Lucas:2006:HBM

[LT06] Alain Lucas and Emmanuel Thilly. Hausdorff–Besicovitch measure of fractal functional limit laws induced by Wiener process in Hölder norms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(3):373–392, May/June 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_3_373_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000889.

Landim:2018:HDL

[LT18] C. Landim and K. Tsunoda. Hydrostatics and dynamical large deviations for a reaction-diffusion model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):51– 74, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1519030819.

Liu:2020:DRP

[LT20] Yanghui Liu and Samy Tindel. Discrete rough paths and limit theorems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1730–1774, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Discrete-rough-paths-
and-limit-theorems/10.1214/19-AIHP1015.full.

Liddle:2021:SLF

[LT21] George Liddle and Amanda Turner. Scaling limits and fluctuations for random growth under capacity rescaling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):980–1015, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-
2/Scaling-limits-and-fluctuations-for-random-growth-
under-capacity-rescaling/10.1214/20-AIHP1104.full.

Last:2018:TRM

[LTT18] Günter Last, Wenpin Tang, and Hermann Thorisson. Transporting random measures on the line and embedding excursions into Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2286–2303, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1539849799.

Luschgy:1978:EPP

[Lus78] Harald Luschgy. Sur l’existence d’une plus petite sous-tribu exhaustive par paire. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4):391–398, October–December 1978. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1978__14_4_391_0.

Luh:2020:SRM

[LV20] Kyle Luh and Van Vu. Sparse random matrices have simple spectrum. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2307–2328, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Sparse-random-matrices-
have-simple-spectrum/10.1214/19-AIHP1032.full.

Lin:1996:CPS

[LW96] Michael Lin and Rainer Wittmann. Convolution powers of spread-out probabilities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(5):661–667, ???? 1996. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1996_
_32_5_661_0.

LeGall:2006:CBT

[LW06] Jean-François Le Gall and Mathilde Weill. Conditioned Brownian trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(4):455–489, July/August 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_4_455_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000907.

Li:2020:ERM

[LW20a] Housen Li and Frank Werner. Empirical risk minimization as parameter choice rule for general linear regularization methods. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):405–427, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Empirical-risk-minimization-
as-parameter-choice-rule-for-general-linear/10.1214/
19-AIHP966.full.

Liechty:2020:AFF

[LW20b] Karl Liechty and Dong Wang. Asymptotics of free fermions in a quadratic well at finite temperature and the Moshe–Neuberger-Shapiro random matrix model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1072–1098, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/Asymptotics-of-free-
fermions-in-a-quadratic-well-at-finite/10.1214/19-AIHP994.
full.

Liu:2016:BSE

[LWZ16] Dang-Zheng Liu, Dong Wang, and Lun Zhang. Bulk and soft-edge universality for singular values of products of Ginibre random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1734–1762, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1479373246.

Li:2019:RDE

[LY19] Liping Li and Jiangang Ying. Regular Dirichlet extensions of one-dimensional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):1815–1849, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-55/issue-4/Regular-Dirichlet-
extensions-of-one-dimensional-Brownian-motion/10.1214/
18-AIHP935.full.

Liu:2015:SSP

[LZ15] Huili Liu and Xiaowen Zhou. Some support properties for a class of Λ–Fleming–Viot processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):1076–1101, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1435759240.

Lygkonis:2022:EWF

[LZ22] Dimitris Lygkonis and Nikos Zygouras. Edwards–Wilkinson fluctuations for the directed polymer in the full. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58 (1):65–104, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-58/issue-
1/EdwardsWilkinson-fluctuations-for-the-directed-polymer-
in-the-full-L2/10.1214/21-AIHP1173.full.

Macchi:1978:RAE

[Mac78] Odile Macchi. Résolution adaptative de l’équation de Wiener– Hopf. Cas d’un canal de données affecté de gigue. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):355–377, July–September 1978. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1978__14_3_
355_0.

Mackevicius:1987:SSS

[Mac87] Vigirdas Mackevičius. S^{p} stability of solutions of symmetric stochastic differential equations with discontinuous driving semimartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(4):575–592, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1987__23_4_
575_0.

Madaule:2015:MLC

[Mad15] Thomas Madaule. Maximum of a log-correlated Gaussian field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1369–1431, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1445432046.

Maigret:1978:TLC

[Mai78] Nelly Maigret. Théorème de limite centrale fonctionnel pour une chaîne de Markov récurrente au sens de Harris et positive. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4): 425–440, October–December 1978. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1978__14_4_425_0.

Maillard:2013:NAI

[Mai13] Pascal Maillard. The number of absorbed individuals in branching Brownian motion with a barrier. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):428–455, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1366117653.

Majka:2019:TIN

[Maj19] Mateusz B. Majka. Transportation inequalities for nonglobally dissipative SDEs with jumps via Malliavin calculus and coupling. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2019–2057, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/Transportation-inequalities-
for-non-globally-dissipative-SDEs-with-jumps-via/10.
1214/18-AIHP941.full.

Malecot:1965:ECF

[Mal65] G. Malécot. Évolution continue des fréquences d’un gène mendélien (dans le cas de migration homogène entre groupes d’effectif fini constant). (French) [Continuous evolution of frequencies of a Mendelian gene (in the case of homogeneous migration between effective finite constant groups)]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(2):137–150, October–December 1965. CODEN AHPBAR. ISSN 0020-2347. URL http://

www.numdam.org/item?id=AIHPB_1965__2_2_137_0.

Malric:1990:FBB

[Mal90] Marc Malric. Filtrations browniennes et balayage. (French) [Brownian filtration and scanning]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(4):507–539, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_4_507_0.

Malric:1996:PEF

[Mal96] Marc Malric. Propriétés d’échange et fins d’ensembles optionnels. (French) [Exchange properties and ends of optional sets]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(3):291–297, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1996__32_3_291_0.

Malyarenko:2011:IRF

[Mal11] Anatoliy Malyarenko. Invariant random fields in vector bundles and application to cosmology. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1068–1095, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1317906502;   http://www.numdam.org/
item?id=AIHPB_2011__47_4_1068_0.

Martias:1988:SPV

[Mar88] Claude Martias. Sur les supports des processus à valeurs dans des espaces nucléaires. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(3):345–365, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1988__24_3_345_0.

Martini:2000:MLO

[Mar00] Claude Martini. On the marginal laws of one-dimensional stochastic integrals with uniformly elliptic integrand. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(1):35–43, January 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2000__36_1_35_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300001138.

Marchal:2015:CSS

[Mar15] P. Marchal. A class of special subordinators with nested ranges. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):533–544, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1428672680.

Marzouk:2016:SBF

[Mar16] Cyril Marzouk. On the sizes of burnt and fireproof components for fires on a large Cayley tree. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1):355–375, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1452089272.

Marzouk:2020:SLD

[Mar20] Cyril Marzouk. Scaling limits of discrete snakes with stable branching. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):502–523, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/Scaling-limits-of-discrete-
snakes-with-stable-branching/10.1214/19-AIHP970.full.

Marx:2021:IDR

[Mar21] Victor Marx. Infinite-dimensional regularization of McKean– Vlasov equation with a Wasserstein diffusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4): 2315–2353, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-

4/Infinite-dimensional-regularization-of-McKeanVlasov-
equation-with-a-Wasserstein-diffusion/10.1214/20-AIHP1136.
full.

Massart:1986:RCC

[Mas86] Pascal Massart. Rates of convergence in the central limit theorem for empirical processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(4):381–423, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1986__22_4_381_0.

Mason:1988:SIT

[Mas88] David M. Mason. A strong invariance theorem for the tail empirical process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(4):491–506, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1988__24_4_
491_0.

Masse:1993:PIP

[Mas93] Bruno Massé. Principes d’invariance pour la probabilité d’un dilaté de l’enveloppe convexe d’un échantillon. (French) [Invariance principles for the probability of an expansion of the convex hull of a sample]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(1):37–55, ???? 1993. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1993__29_1_37_0.

Massart:2002:TLY

[Mas02] Pascal Massart. Tusnady’s lemma, 24 years later. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):991–1007, November/December 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_6_991_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302011305.

Mathieu:1993:INP

[Mat93] Pierre Mathieu. Inégalités en norme L_{p} pour le produit des suprema de plusieurs martingales arrêtées à des temps aléatoires. Inégalités avec poids. (French) [Inequalities in the L_{p} norm for the product of suprema of several martingales stopped in random time. Inequalities with weight]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(4): 467–484, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_4_467_0.

Mathieu:1997:HTS

[Mat97] P. Mathieu. Hitting times and spectral gap inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(4):437–465, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_4_564_0;   http://
www.sciencedirect.com/science/article/pii/S0246020397801008.

See erratum [Mat98b].

Mathieu:1998:Ea

[Mat98a] Pierre Mathieu. Erratum. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(4):ii, July/August 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0246020398800289.

Mathieu:1998:EPM

[Mat98b] Pierre Mathieu. Erratum: P. Mathieu, “Hitting times and spectral gap inequalities”, Annales de l’Institut Henri Poincaré (B) Probability and Statistics, Volume 33, Issue 4, 1997, Pages 437–465. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(4):564, ???? 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_4_437_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800289. See [Mat97].

Matus:1998:CDM

[Mat98c] František Matús. Combining m-dependence with Markovness. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(4):407–423, July/August 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_4_407_0;   http://www.sciencedirect.com/science/
article/pii/S024602039880023X.

Matus:2003:CPP

[Mat03] František Matúš. Conditional probabilities and permutahedron. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(4):687–701, July/August 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_4_687_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000207.

Matus:2012:CIL

[Mat12] František Matúš. On conditional independence and logconvexity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):1137–1147, November 2012. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1353098443.

Mayer:1968:PMN

[May68] Claude Mayer. Processus de Markov non stationnaires et espace-temps. (French) [Non-stationary Markov processes and space–time]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(3):165– 177, July–September 1968. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1968__
4_3_165_0.

McDonald:1975:RTM

[McD75] David McDonald. Renewal theorem and Markov chains. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(2):187–197, April– June 1975. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1975__11_2_187_0.

McGill:1989:FBM

[McG89] Paul McGill. Factorising Brownian motion at two boundaries; an example. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):517–531, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1989__25_4_517_0.

Marquez-Carreras:2006:ABM

[MCRT06] David Márquez-Carreras, Carles Rovira, and Samy Tindel. Asymptotic behavior of the magnetization for the perceptron model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(3):327–342, May/June 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_3_327_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000853.

Mejane:2004:UBV

[Mej04] Olivier Mejane. Upper bound of a volume exponent for directed polymers in a random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):299–308, May/June 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2004__40_3_299_0;   http://www.sciencedirect.
com/science/article/pii/S0246020303000724.

Melard:1978:PDS

[Mél78] Guy Mélard. Propriétés du spectre évolutif d’un processus non stationnaire. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(4):411–424, October–December 1978. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_4_411_0.

Mela:2006:CNA

[Mél06] Xavier Méla. A class of nonstationary adic transformations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1):103–123, January/February 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_1_103_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000488.

Menard:1978:RUS

[Mén78] E. Ménard. Remarque sur l’uniforme sommabilité des suites de variables aléatoires à valeurs vectorielles normées. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(1):85–98, January– March 1978. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1978__14_1_85_0.

Menard:2010:TUI

[Mén10] Laurent Ménard. The two uniform infinite quadrangulations of the plane have the same law. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):190–208, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1267454114;   http://www.numdam.org/
item?id=AIHPB_2010__46_1_190_0.

Merzbach:1983:CCO

[Mer83] Ely Merzbach. Chemins croissants optionnels et théorème de section. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(3):223–234, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1983__19_3_
223_0.

Merle:2006:LBL

[Mer06] Mathieu Merle. Local behaviour of local times of super-Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(4):491–520, July/August 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_4_491_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000919.

Metivier:1972:TLQ

[Mét72] Michel Métivier. Théorèmes limite quotient pour chaînes de Markov récurrentes au sens de Harris. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(2):93–105, April–June 1972. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1972__8_2_93_0.

Metivier:1984:CFP

[Mét84] Michel Métivier. Convergence faible et principe d’invariance pour des martingales à valeurs dans des espaces de Sobolev. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(4):329–348, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1984__20_4_329_0.

Meyre:1991:EAD

[Mey91] Thierry Meyre. Étude asymptotique du temps passé par le mouvement brownien dans un cône. (French) [Asymptotic study of the time spent by Brownian motion in a cone]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(1): 107–124, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1991__27_1_107_0.

Muller-Gronbach:2020:PEM

[MGY20] Thomas Müller-Gronbach and Larisa Yaroslavtseva. On the performance of the Euler–Maruyama scheme for SDEs with discontinuous drift coefficient. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1162–1178, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/On-the-performance-of-
the-EulerMaruyama-scheme-for-SDEs-with/10.1214/19-AIHP997.
full.

Mikkelsen:2018:DFP

[MH18] Frederik Riis Mikkelsen and Niels Richard Hansen. Degrees of freedom for piecewise Lipschitz estimators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):819–841, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1524643230.

Miclo:1992:CAE

[Mic92a] Laurent Miclo. Comportement asymptotique de l’énergie libre spécifique. Application à l’ergodicité et au recuit simulé en dimension infinie. (French) [Asymptotic behavior of specific free energy. Application to ergodicity and simulated annealing in infinite dimension]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2):195–234, ???? 1992. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1992_
_28_2_195_0.

Miclo:1992:RSE

[Mic92b] Laurent Miclo. Recuit simulé sur R^{n}. Étude de l’évolution de l’énergie libre. (French) [Simulated annealing on R^{n}. Study of the evolution of free energy]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2):235–266, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1992__28_2_235_0.

Miclo:1999:MSE

[Mic99] Laurent Miclo. Une majoration sous-exponentielle pour la convergence de l’entropie des chaînes de Markov à trou spectral. (French) [A sub-exponential increase for the convergence of the entropy of Markov chains to a spectral hole]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(3):261–311, May/June 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_3_261_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399800132.

Miclo:2009:CAP

[Mic09] Laurent Miclo. Une condition asymptotique pour le calcul de constantes de Sobolev logarithmiques sur la droite. (French) [A condition for the asymptotic calculation of Sobolev logarithmic constants on the right]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1): 146–156, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469975;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_146_0.

Miclo:2017:SST

[Mic17] Laurent Miclo. Strong stationary times for one-dimensional diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):957–996, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1491897753.

Miermont:2008:IPS

[Mie08] Grégory Miermont. Invariance principles for spatial multitype Galton–Watson trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1128–1161, December 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1227287568;   http://www.numdam.org/item?id=AIHPB_2008_
_44_6_1128_0.

Mikami:1988:AEI

[Mik88] Toshio Mikami. Asymptotic expansions of the invariant density of a Markov process with a small parameter. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(3): 403–424, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1988__24_3_403_0.

Millet:1983:CRT

[Mil83] Annie Millet. On convergence and regularity of two-parameter (Δ1) submartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):25–42, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1983__19_1_
25_0.

Mimica:2014:HFS

[Mim14] Ante Mimica. On harmonic functions of symmetric Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):214–235, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1388545272.

Misturini:2016:EAM

[Mis16] Ricardo Misturini. Evolution of the ABC model among the segregated configurations in the zero-temperature limit. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):669–702, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1462367889.

Mkrtchyan:2014:ESW

[Mkr14] Sevak Mkrtchyan. Entropy of Schur–Weyl measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50 (2):678–713, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1395856145.

Macchi:1971:TIS

[MM71] César Macchi and Odile Macchi. Un théorème d’itération stochastique multidimensionnelle. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(3):193–204, July–September 1971. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1971__7_3_193_0.

Maillard:2009:LDV

[MM09] G. Maillard and T. Mountford. Large deviations for voter model occupation times in two dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2): 577–588, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024681;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_577_0.

Maillard:2013:EBS

[MM13] G. Maillard and T. S. Mountford. Ergodic behaviour of “signed voter models”. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):13–35, February 2013. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1359470124.

Mairesse:2014:PCA

[MM14] Jean Mairesse and Irène Marcovici. Probabilistic cellular automata and random fields with i.i.d. directions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2): 455–475, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1395856135.

Mallein:2018:BMR

[MM18] Bastien Mallein and Piotr Miło’s. Brownian motion and random walk above quenched random wall. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):1877–1916, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849787.

Mignotte:1983:SF

[MN83] M. Mignotte and J. L. Nicolas. Statistiques sur F_{q}[X]. (French) [Statistics on F_{q}[X]]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(2):113–121, ???? 1983. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1983_
_19_2_113_0.

Martirosyan:2018:LLD

[MN18] D. Martirosyan and V. Nersesyan. Local large deviations principle for occupation measures of the stochastic damped nonlinear wave equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2002–2041, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1539849791.

Miasojedow:2021:AML

[MNR21] Błazej Miasojedow, Wojciech Niemiro, and Wojciech Rejchel. Asymptotics of maximum likelihood estimators based on Markov chain Monte Carlo methods. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):815–829, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-2/Asymptotics-of-maximum-
likelihood-estimators-based-on-Markov-chain-Monte/10.
1214/20-AIHP1097.full.

Moulin-Ollagnier:1983:TEP

[MO83] Jean Moulin-Ollagnier. Théorème ergodique presque sousadditif et convergence en moyenne de l’information. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(3):257–266, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1983__19_3_257_0.

Mourragui:2007:LDM

[MO07] Mustapha Mourragui and Enza Orlandi. Large deviations from a macroscopic scaling limit for particle systems with Kac interaction and random potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6):677–715, November/December 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_6_677_0;   http://www.sciencedirect.com/science/
article/pii/S0246020307000088.

Mokkadem:1990:PMP

[Mok90] Abdelkader Mokkadem. Propriétés de mélange des processus autorégressifs polynomiaux. (French) [Mixing properties of polynomial autoregressive processes]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(2):219–260, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_2_219_0.

Mataramvura:2004:DDF

[MØP04] Sure Mataramvura, Bernt Øksendal, and Frank Proske. The Donsker delta function of a Lévy process with application to chaos expansion of local time. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(5):553–567, September/October 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_5_553_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000263.

Morien:1999:DSB

[Mor99] Pierre-Luc Morien. On the density for the solution of a Burgers-type SPDE. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(4):459–482, July 1999. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_4_459_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001028.

Morters:2001:ADS

[Mör01] Peter Mörters. The average density of super-Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(1):71–100, January 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_1_71_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010608.

Morters:2011:APA

[MOS11] Peter Mörters, Marcel Ortgiese, and Nadia Sidorova. Ageing in the parabolic Anderson model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):969–1000, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1317906497;   http://www.numdam.org/
item?id=AIHPB_2011__47_4_969_0.

Mourragui:1996:CHE

[Mou96] Mustapha Mourragui. Comportement hydrodynamique et entropie relative des processus de sauts, de naissances et de morts. (French) [Hydrodynamic behavior and relative entropy of process jumps, of births, and of deaths]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(3):361–385, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1996__32_3_361_0.

Mourrat:2011:SLR

[Mou11a] Jean-Christophe Mourrat. Scaling limit of the random walk among random traps on Z^{d}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3):813–849, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1308834860;   http://www.numdam.org/item?id=AIHPB_2011_
_47_3_813_0.

Mourrat:2011:VDF

[Mou11b] Jean-Christophe Mourrat. Variance decay for functionals of the environment viewed by the particle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1): 294–327, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170239;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_294_0.

Metcalfe:2009:IPC

[MOW09] Anthony P. Metcalfe, Neil O’Connell, and Jon Warren. Interlaced processes on the circle. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1165–1184, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1257529898;   http://www.numdam.org/
item?id=AIHPB_2009__45_4_1165_0.

Marcus:1984:SRC

[MP84] M. B. Marcus and G. Pisier. Some results on the continuity of stable processes and the domain of attraction of continuous stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(2):177–199, ???? 1984. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1984__20_2_177_0.

Mueller:2000:ETP

[MP00] C. Mueller and E. Perkins. Extinction for two parabolic stochastic PDE’s on the lattice. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(3):301–338, May 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_3_301_0;   http://www.sciencedirect.com/science/
article/pii/S024602030000128X.

Martin:2019:PDB

[MP19] Jörg Martin and Nicolas Perkowski. Paracontrolled distributions on Bravais lattices and weak universality of the 2d parabolic Anderson model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2058–2110, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-55/issue-4/Paracontrolled-distributions-
on-Bravais-lattices-and-weak-universality-of-the/10.
1214/18-AIHP942.full.

Martin:1973:ICM

[MPL73] F. Martin, J.-L. Petit, and M. Littaye. Indépendance conditionnelle dans le modèle statistique bayésien. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(1):19–40, January–March 1973. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1973__9_1_19_0.

Martin:1971:CEF

[MPPL71] Françoise Martin, Jean-Luc Petit, and Monique Petit-Littaye. Comparaison des expériences. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(2):145–176, April–June 1971. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1971__7_2_145_0.

Mastrolia:2017:MDB

[MPR17] Thibaut Mastrolia, Dylan Possamaï, and Anthony Réveillac. On the Malliavin differentiability of BSDEs. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1): 464–492, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1486544899.

Martin-Reyes:1986:WES

[MR86] F. J. Martin-Reyes. Weights for ergodic square functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(3):333–345, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1986__22_3_333_0.

Milhaud:1989:EED

[MR89] X. Milhaud and A. Raugi. Étude de l’estimateur du maximum de vraisemblance dans le cas d’un processus autorégressif: convergence, normalité asymptotique, vitesse de convergence. (French) [Study of the estimator of the maximum likelihood in the case of an autoregressive process: convergence, asymptotic normality, speed of convergence]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):383–428, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1989__25_4_383_0.

Matula:1990:IPN

[MR90] Przemysław Matuła and Zdzisław Rychlik. The invariance principle for nonstationary sequences of associated random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):387–397, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1990__26_3_387_0.

Marcus:1994:LIL

[MR94a] Michael B. Marcus and Jay Rosen. Laws of the iterated logarithm for the local times of recurrent random walks on Z^{2} and of Lévy processes and random walks in the domain of attraction of Cauchy random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(3):467–499, ???? 1994. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1994_
_30_3_467_0.

Mindlin:1994:CAS

[MR94b] D. S. Mindlin and B. A. Rubshtein. Convolutional attractors of stationary sequences of random measures on compact groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):213–233, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1994__30_2_213_0.

Marcus:1997:LIL

[MR97] Michael B. Marcus and Jay Rosen. Laws of the iterated logarithm for intersections of random walks on Z^{4}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(1):37–63, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1997__33_1_37_0;   http://
www.sciencedirect.com/science/article/pii/S024602039780115X.

Mallein:2021:BEG

[MR21] Bastien Mallein and Sanjay Ramassamy. Barak–Erdős graphs and the infinite-bin model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):1940–1967, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-4/BarakErd%c5%91s-graphs-
and-the-infinite-bin-model/10.1214/20-AIHP1141.full.

Mukherjee:2022:TPI

[MR22] Rajarshi Mukherjee and Gourab Ray. On testing for parameters in Ising models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):164–187, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/On-testing-for-parameters-
in-Ising-models/10.1214/21-AIHP1157.full.

Merkl:2021:RIV

[MRT21] Franz Merkl, Silke W. W. Rolles, and Pierre Tarrès. Random interlacements for vertex-reinforced jump processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):1058–1080, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-
2/Random-interlacements-for-vertex-reinforced-jump-
processes/10.1214/20-AIHP1107.full.

Ma:2016:LDN

[MRTZ16] Jin Ma, Zhenjie Ren, Nizar Touzi, and Jianfeng Zhang. Large deviations for non-Markovian diffusions and a path-dependent eikonal equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1196–1216, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1469723517.

Maillard:2016:LHK

[MRVZ16] P. Maillard, R. Rhodes, V. Vargas, and O. Zeitouni. Liouville heat kernel: Regularity and bounds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1281–1320, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723521.

Marinucci:2020:AES

[MRW20] Domenico Marinucci, Maurizia Rossi, and Igor Wigman. The asymptotic equivalence of the sample trispectrum and the nodal length for random spherical harmonics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):374–390, February 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/The-asymptotic-equivalence-of-the-
sample-trispectrum-and-the-nodal/10.1214/19-AIHP964.
full.

Ma:1998:AAD

[MRZ98] Zhi-Ming Ma, Michael Röckner, and Tu-Sheng Zhang. Approximation of arbitrary Dirichlet processes by Markov chains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(1):1–22, ???? 1998. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1998__34_1_
1_0;   http://www.sciencedirect.com/science/article/
pii/S0246020398800113.

Mandrekar:1970:SID

[MS70] V. Mandrekar and H. Salehi. Subordination of infinitedimensional stationary stochastic processes. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(2):115–130, April–June 1970. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1970__6_2_115_0.

Marcuard:1983:EIA

[MS83a] J.-C. Marcuard and B. Schmitt. Entropie et itinéraires des applications unimodales de l’intervalle. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(4): 351–367, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1983__19_4_351_0.

Millet:1983:DPI

[MS83b] Annie Millet and Louis Sucheston. Demiconvergence of processes indexed by two indices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(2):175–187, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1983__19_2_175_0.

Mohammed:1996:LEL

[MS96] Salah-Eldin A. Mohammed and Michael K. R. Scheutzow. Lyapunov exponents of linear stochastic functional differential equations driven by semimartingales. Part I: the multiplicative ergodic theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(1):69–105, ???? 1996. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1996__32_1_69_0.

Manita:2004:NTS

[MS04] A. Manita and F. Simonot. A note on the two-sided regulated random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):43–52, January/February 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_1_43_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000591.

Marteau:2017:MGF

[MS17] Clément Marteau and Theofanis Sapatinas. Minimax goodnessof-fit testing in ill-posed inverse problems with partially unknown operators. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1675–1718, November 2017. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1511773723.

Miller:2019:LQG

[MS19] Jason Miller and Scott Sheffield. Liouville quantum gravity spheres as matings of finite-diameter trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1712–1750, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1569398883.

Mukherjee:2021:TDC

[MS21] Rajarshi Mukherjee and Subhabrata Sen. Testing degree corrections in stochastic block models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1583–1635, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Testing-degree-corrections-
in-stochastic-block-models/10.1214/20-AIHP1120.full.

Murugan:2019:HKE

[MSC19] Mathav Murugan and Laurent Saloff-Coste. Heat kernel estimates for anomalous heavy-tailed random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):697–719, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820828.

Millet:2006:LDR

[MSS06] Annie Millet and Marta Sanz-Solé. Large deviations for rough paths of the fractional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42 (2):245–271, March/April 2006. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2006__42_2_245_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000683.

Martinelli:2020:DSK

[MST20] F. Martinelli, A. Shapira, and C. Toninelli. Diffusive scaling of the Kob–Andersen model in Z^{d}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2189–2210, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Diffusive-scaling-of-
the-KobAndersen-model-in-mathbbZd/10.1214/19-AIHP1035.
full.

Mehrdad:2015:SBW

[MSZ15] Behzad Mehrdad, Sanchayan Sen, and Lingjiong Zhu. The speed of a biased walk on a Galton–Watson tree without leaves is monotonic with respect to progeny distributions for high values of bias. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):304–318, February 2015. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1421244407.

Mountford:2008:ARB

[MT08] Thomas Mountford and Pierre Tarrès. An asymptotic result for Brownian polymers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1):29–46, February 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1203969867;   http://www.numdam.org/item?id=AIHPB_2008_
_44_1_29_0.

Melbourne:2020:RTM

[MT20] Ian Melbourne and Dalia Terhesiu. Renewal theorems and mixing for non Markov flows with infinite measure. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):449–476, February 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Renewal-theorems-and-mixing-for-non-
Markov-flows-with-infinite/10.1214/19-AIHP968.full.

Masamune:2011:CPS

[MU11] Jun Masamune and Toshihiro Uemura. Conservation property of symmetric jump processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3):650–662, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1308834853;   http://www.numdam.org/item?id=AIHPB_2011_
_47_3_650_0.

Mustapha:2006:BIP

[Mus06]  Sami Mustapha.  Bornes inférieures pour les marches

aléatoires sur les groupes p-adiques moyennables. (French) [Lower bounds for random walks on p-adic averageable groups]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1):81–88, January/February 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_1_81_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000671.

Martin:1969:ISM

[MV69] Françoise Martin and Daniel Vaguelsy. Introduction à la statistique mathématique. IV. Propriétés asymptotiques du modèle statistique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(4):357–384, October–December 1969. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1969__5_4_385_0.

Misiurewicz:1991:KSS

[MV91] M. Misiurewicz and E. Visinescu. Kneading sequences of skew tent maps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(1):125–140, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1991__27_1_125_0.

Marcuard:1992:MPS

[MV92] J. C. Marcuard and E. Visinescu. Monotonicity properties of some skew tent maps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(1):1–29, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1992__28_1_
1_0.

Mytnik:2007:SIL

[MV07] L. Mytnik and J. Villa. Self-intersection local time of (α, d, β)- superprocess. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(4):481–507, July/August 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_4_481_0;   http://www.sciencedirect.com/science/
article/pii/S0246020307000106.

Muirhead:2020:SPT

[MV20] Stephen Muirhead and Hugo Vanneuville. The sharp phase transition for level set percolation of smooth planar Gaussian fields. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):1358–1390, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-2/The-sharp-phase-transition-
for-level-set-percolation-of-smooth/10.1214/19-AIHP1006.
full.

Merkl:2002:IVA

[MW02] Franz Merkl and Mario V. Wüthrich. Infinite volume asymptotics of the ground state energy in a scaled Poissonian potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(3):253–284, ???? 2002. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2002_
_38_3_253_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301011001.

Morvai:2005:FEE

[MW05] Gusztáv Morvai and Benjamin Weiss. Forward estimation for ergodic time series. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5):859–870, September/October 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_5_859_0;   http://www.sciencedirect.
com/science/article/pii/S0246020304000846.

Morvai:2007:EMF

[MW07] Gusztáv Morvai and Benjamin Weiss. On estimating the memory for finitarily Markovian processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1): 15–30, January/February 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_1_15_0;   http://
www.sciencedirect.com/science/article/pii/S024602030600015X.

Morvai:2011:TSP

[MW11] Gusztáv Morvai and Benjamin Weiss. Testing stationary processes for independence. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1219–1225, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1317906509;   http://www.numdam.org/
item?id=AIHPB_2011__47_4_1219_0.

Mayer-Wolf:2004:LTO

[MWRZ04] Eddy Mayer-Wolf, Alexander Roitershtein, and Ofer Zeitouni. Limit theorems for one-dimensional transient random walks in Markov environments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(5):635–659, September/October 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2004__40_5_635_0;   http://www.sciencedirect.
com/science/article/pii/S0246020304000305.

Meyer:1984:TCL

[MZ84] P. A. Meyer and W. A. Zheng. Tightness criteria for laws of semimartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(4):353–372, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1984__20_4_
353_0.

Melbourne:2015:WCS

[MZ15] Ian Melbourne and Roland Zweimüller. Weak convergence to stable Lévy processes for nonuniformly hyperbolic dynamical systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):545–556, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1428672681.

Maillard:2016:SBB

[MZ16] Pascal Maillard and Ofer Zeitouni. Slowdown in branching Brownian motion with inhomogeneous variance. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3): 1144–1160, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1469723514.

Nagahata:2005:RDC

[Nag05] Yukio Nagahata. Regularity of the diffusion coefficient matrix for the lattice gas with energy. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(1):45–67, January/February 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_1_45_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000585.

Nakao:1986:WCS

[Nak86] Shintaro Nakao. On weak convergence of sequences of continuous local martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(3):371–380, ???? 1986. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1986_
_22_3_371_0.

Nakashima:2015:BRW

[Nak15] Makoto Nakashima. Branching random walks in random environment and super-Brownian motion in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(4):1251–1289, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1445432042.

Nakajima:2020:DSF

[Nak20] Shuta Nakajima. Divergence of shape fluctuation for general distributions in first-passage percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):782–791, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-2/Divergence-of-shape-
fluctuation-for-general-distributions-in-first-passage/
10.1214/19-AIHP982.full.

Nelson:2020:BMA

[NdS20] Peter Nelson and Renato Soares dos Santos. Brownian motion in attenuated or renormalized inverse-square Poisson potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):1–35, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Brownian-motion-in-attenuated-or-
renormalized-inverse-square-Poisson-potential/10.1214/
18-AIHP953.full.

Neveu:1969:HMF

[Nev69] J. Neveu. Harmonicité des moments d’une fonction aléatoire. (French) [Harmonicity of moments of a random function]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(1):13–29, January– March 1969. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1969__5_1_13_0.

Neveu:1972:CPS

[Nev72a] J. Neveu. Convergence presque sûre de martingales multivoques. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8 (1):1–7, January–March 1972. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1972__
8_1_1_0.

Neveu:1972:ICM

[Nev72b] J. Neveu. Sur l’irréductibilité des chaînes de Markov. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(3):249–254, July– September 1972. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1972__8_3_249_0.

Neveu:1976:ECP

[Nev76] J. Neveu. Sur l’espérance conditionnelle par rapport à un mouvement brownien. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(2):105–109, April–June 1976. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1976__12_2_105_0.

Neveu:1983:CDD

[Nev83] J. Neveu. Courte démonstration du théorème ergodique suradditif. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):87–90, ???? 1983. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1983__19_1_87_0.

Neveu:1986:APG

[Nev86] J. Neveu. Arbres et processus de Galton–Watson. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2):199–207, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1986__22_2_199_0.

Nghiem:1970:CFE

[Ngh70] Dang Ngoc Nghiem. Convergence forte des espérances conditionnelles et des projecteurs d’un espace de Hilbert. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(1):9–13, January– March 1970. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1970__6_1_9_0.

Nicaise:2003:ADT

[Nic03] Florent Nicaise. Anticipative direct transformations on the Poisson space. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(4):557–592, July/August 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_4_557_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000079.

Naik-Nimbalkar:1981:BPB

[NN81] Uttara Naik-Nimbalkar. Bochner property in Banach spaces. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(1):1–19, January– March 1981. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1981__17_1_1_0.

Nourdin:2016:FIF

[NN16] Ivan Nourdin and David Nualart. Fisher information and the fourth moment theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):849–867, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1462367896.

Nestoridi:2020:MTD

[NN20] Evita Nestoridi and Oanh Nguyen. On the mixing time of the Diaconis–Gangolli random walk on contingency tables over Z/qZ. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(2):983–1001, May 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-2/On-the-mixing-time-of-the-DiaconisGangolli-
random-walk-on/10.1214/19-AIHP991.full.

Neuenkirch:2009:TAE

[NNRT09] A. Neuenkirch, I. Nourdin, A. Rößler, and S. Tindel. Trees and asymptotic expansions for fractional stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):157–174, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1234469976;   http://
www.numdam.org/item?id=AIHPB_2009__45_1_157_0.

Nourdin:2010:CNC

[NNT10] Ivan Nourdin, David Nualart, and Ciprian A. Tudor. Central and non-central limit theorems for weighted power variations of fractional Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1055–1079, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878338;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_1055_0.

Nappo:1988:LLC

[NO88] G. Nappo and E. Orlandi. Limit laws for a coagulation model of interacting random particles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(3):319–344, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1988__24_3_319_0.

Nourdin:2010:MNA

[NPR10] Ivan Nourdin, Giovanni Peccati, and Anthony Réveillac. Multivariate normal approximation using Stein’s method and Malliavin calculus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):45–58, February 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1267454107;   http://www.numdam.org/item?id=AIHPB_2010_
_46_1_45_0.

Naor:2003:PSM

[NR03] Assaf Naor and Dan Romik. Projecting the surface measure of the sphere of ℓ_{p}^{n}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(2):241–261, March/April 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_2_241_0;   http://www.sciencedirect.com/science/
article/pii/S0246020302000080.

Nguyen:2017:NIB

[NR17] Gia Bao Nguyen and Daniel Remenik. Non-intersecting Brownian bridges and the Laguerre orthogonal ensemble. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53 (4):2005–2029, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1511773736.

Naor:2020:CMC

[NRR20] Assaf Naor, Shravas Rao, and Oded Regev. Concentration of Markov chains with bounded moments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):2270–2280, August 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/Concentration-of-Markov-chains-with-
bounded-moments/10.1214/19-AIHP1039.full.

Newman:2010:MPB

[NRS10] C. M. Newman, K. Ravishankar, and E. Schertzer. Marking (1, 2) points of the Brownian web and applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):537–574, May 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1273584134;   http://www.numdam.org/
item?id=AIHPB_2010__46_2_537_0.

Nualart:1981:CTT

[NS81] D. Nualart and M. Sanz. Changing time for two-parameter strong martingales. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17 (2):147–163, April–June 1981. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1981__
17_2_147_0.

Newman:1995:RWS

[NS95] C. M. Newman and D. L. Stein. Random walk in a strongly inhomogeneous environment and invasion percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1): 249–261, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1995__31_1_249_0.

Noack:2022:CID

[NS22] Christian Noack and Philippe Sosoe. Concentration for integrable directed polymer models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):34–64, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-58/issue-1/Concentration-for-integrable-
directed-polymer-models/10.1214/21-AIHP1154.full.

Nualart:1984:FIP

[Nua84] David Nualart. Une formule d’Itô pour les martingales continues à deux indices et quelques applications. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(3):251–275, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1984__20_3_251_0.

Nummelin:1978:URL

[Num78] E. Nummelin. Uniform and ratio limit theorems for Markov renewal and semi-regenerative processes on a general state space. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(2):119–143, April– June 1978. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1978__14_2_119_0.

Newman:2018:GFC

[NW18] Charles M. Newman and Wei Wu. Gaussian fluctuations for the classical XY model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):1759–1777, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1539849782.

Nikeghbali:2013:GWP

[NZ13] Ashkan Nikeghbali and Dirk Zeindler. The generalized weighted probability measure on the symmetric group and the asymptotic behavior of the cycles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):961–981, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1380718733.

Nualart:2021:TVE

[NZ21] David Nualart and Hongjuan Zhou. Total variation estimates in the Breuer–Major theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):740–777, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-2/Total-variation-estimates-
in-the-BreuerMajor-theorem/10.1214/20-AIHP1094.full.

Oz:2017:CSB

[ÖÇE17] Mehmet Öz, Mine Çaglar, and János Engländer. Conditional speed of branching Brownian motion, skeleton decomposition and application to random obstacles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2): 842–864, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1491897748.

Odasso:2006:ESC

[Oda06] Cyril Odasso. Ergodicity for the stochastic complex Ginzburg– Landau equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(4):417–454, July/August 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_4_417_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000890.

Oz:2019:OSS

[ÖE19] Mehmet Öz and János Engländer. Optimal survival strategy for branching Brownian motion in a Poissonian trap field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):1890–1915, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-55/issue-4/Optimal-survival-strategy-for-branching-
Brownian-motion-in-a-Poissonian/10.1214/18-AIHP937.
full.

Ogawa:1974:PMI

[Oga74] Shigeyoshi Ogawa. Processus de Markov en interaction et système semi-linéaire d’équations d’évolution. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(2):279–299, April– June 1974. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1974__10_2_279_0.

Okazaki:1985:ESD

[Oka85] Yoshiaki Okazaki. Equivalent-singular dichotomy for quasiinvariant ergodic measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(4):393–400, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1985__21_4_393_0.

Ocone:1989:GIV

[OP89] Daniel Ocone and Etienne Pardoux. A generalized Itô–Ventzell formula. Application to a class of anticipating stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(1):39–71, ???? 1989. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1989__25_1_39_0.

Ortgiese:2018:SLA

[OR18] Marcel Ortgiese and Matthew I. Roberts. Scaling limit and ageing for branching random walk in Pareto environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1291–1313, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1531296020.

Ost:2020:SST

[ORB20] G. Ost and P. Reynaud-Bouret. Sparse space-time models: Concentration inequalities and Lasso. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2377–2405, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/Sparse-spacetime-
models-Concentration-inequalities-and-Lasso/10.1214/
19-AIHP1042.full.

Orsingher:1982:RFV

[Ors82] Enzo Orsingher. Randomly forced vibrations of a string. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(4):367–394, October–December 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1982__18_4_367_0.

Osada:1998:IPM

[Osa98] Hirofumi Osada. An invariance principle for Markov processes and Brownian particles with singular interaction. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(2):217–248, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1998__34_2_217_0;   http://
www.sciencedirect.com/science/article/pii/S0246020398800319.

Ogura:2002:CLT

[OTT02] Y. Ogura, M. Tomisaki, and M. Tsuchiya. Convergence of local type Dirichlet forms to a non-local type one. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38 (4):507–556, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_4_507_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010950.

Ouchti:2005:RCC

[Ouc05] Lahcen Ouchti. On the rate of convergence in the central limit theorem for martingale difference sequences. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41 (1):35–43, January/February 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_1_35_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000524.

Ouvrard:1973:MLT

[Ouv73] Jean-Yves Ouvrard. Martingales locales et théorème de Girsanov dans les espaces de Hilbert réels séparables. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(4):351–368, October– December 1973. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1973__9_4_351_0.

Ondrejat:2020:TRS

[OV20] Martin Ondreját and Mark Veraar. On temporal regularity of stochastic convolutions in. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1792–1808, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/On-temporal-regularity-
of-stochastic-convolutions-in-2-smooth-Banach/10.1214/
19-AIHP1017.full.

Overbeck:1994:MBS

[Ove94] Ludger Overbeck. Martin boundaries of some branching processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):181–195, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1994__30_2_181_0.

ORourke:2017:SNH

[OW17] Sean O’Rourke and Philip Matchett Wood. Spectra of nearly Hermitian random matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3):1241–1279, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1500624041.

Paccaut:2000:SRT

[Pac00] Frédéric Paccaut. Statistics of return times for weighted maps of the interval. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(3):339–366, May 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_3_339_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001278.

Pain:2018:NCG

[Pai18] Michel Pain. The near-critical Gibbs measure of the branching random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1622–1666, August 2018. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1531296031.

Panchenko:2009:CMS

[Pan09] Dmitry Panchenko. Cavity method in the spherical SK model. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1020–1047, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1257529890;   http://
www.numdam.org/item?id=AIHPB_2009__45_4_1020_0.

Pan:2014:CBT

[Pan14] Guangming Pan. Comparison between two types of large sample covariance matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(2):655–677, May 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1395856144.

Parthasarathy:2005:EQS

[Par05] K. R. Parthasarathy. Extremal quantum states in coupled systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):257–268, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_257_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000233.

Payen:1967:FAD

[Pay67] R. Payen. Fonctions aléatoires du second ordre à valeurs dans un espace de Hilbert. (French) [Second-order random functions of values in a Hilbert space]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 3(4):323–396, October–December 1967. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1967__3_4_323_0.

Pazman:1977:PEP

[Páz77] Andrej Pázman. Plans d’expérience pour les estimations de fonctionnelles non-linéaires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(3):259–267, July–September 1977. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1977__13_3_259_0.

Penda:2014:DIM

[PD14] S. Valère Bitseki Penda and Hacène Djellout. Deviation inequalities and moderate deviations for estimators of parameters in bifurcating autoregressive models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):806–844, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1403276999.

Peccati:2001:RRT

[Pec01] Giovanni Peccati. A representation result for time-space Brownian chaos. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(5):607–625, September/October 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_5_607_0;   http://www.sciencedirect.com/science/
article/pii/S0246020301010767.

Peche:2012:UBS

[Péc12] Sandrine Péché. Universality in the bulk of the spectrum for complex sample covariance matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1): 80–106, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1327328015.

Peigne:1992:MMS

[Pei92] Marc Peigné. Marches de Markov sur le semi-groupe des contractions de R^{d}. Cas de la marche aléatoire à pas markoviens sur (R^{+})^{d} avec chocs élastiques sur les axes. (French) [Markov walks on the semi-group of contractions of R^{d}. Case of random walk with Markov steps on (R^{+})^{d} with elastic collisions on the axes]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(1):63–94, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1992__28_1_63_0.

Pellaumail:1972:AER

[Pel72] Jean Pellaumail. Application de l’existence d’un relèvement à un théorème sur la désintégration des mesures. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(3):211–215, July– September 1972. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1972__8_3_211_0.

Pellaumail:1979:FDP

[Pel79] J. Pellaumail. Formule du produit et décomposition de réseaux de files d’attente. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(3):261–286, July–September 1979. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_3_261_0.

Pellaumail:1990:GAC

[Pel90] J. Pellaumail. Graphes et algorithme de calcul de probabilités stationnaires d’un processus markovien discret. (French) [Graphs and calculation algorithms of stationary probabilities of a discrete Markov process]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):121–143, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_1_121_0.

Pellegrini:2010:MCA

[Pel10] Clément Pellegrini. Markov chains approximation of jump– diffusion stochastic master equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):924–948, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878330;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_924_0.

Pemantle:1997:PBM

[Pem97] Robin Pemantle. The probability that Brownian motion almost contains a line. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(2):147–165, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_2_147_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397801197.

Pene:2009:PLP

[Pèn09] Françoise Pène. Planar Lorentz process in a random scenery. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):818–839, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1249391387;   http:/
/www.numdam.org/item?id=AIHPB_2009__45_3_818_0.

Pene:2019:MDI

[Pèn19] Françoise Pène. Mixing and decorrelation in infinite measure: The case of the periodic Sinai billiard. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):378–411, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1547802404.

Perkins:1989:HMC

[Per89] Edwin Perkins. The Hausdorff measure of the closed support of super-Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2):205–224, ???? 1989. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1989_
_25_2_205_0.

Peres:1992:DAC

[Per92] Yuval Peres. Domains of analytic continuation for the top Lyapunov exponent. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(1):131–148, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1992__28_1_
131_0.

Peres:2000:PNP

[Per00] Yuval Peres. Percolation on nonamenable products at the uniqueness threshold. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(3):395–406, May 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_3_395_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001308.

Petit:1970:EAI

[Pet70] Jean-Luc Petit. Exhaustivité, ancillarité et invariance. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(4):327–334, October–December 1970. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1970__
6_4_327_0.

Petit:1979:TEC

[Pet79] B. Petit. Théorie ergodique: classification de certaines transformations réelles. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1):25–32, January–March 1979. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_1_25_0.

Petit:1996:TDS

[Pet96] Frédérique Petit. Théorème du support pour les diffusions réfléchies de type Ventcell. (French) [Theorem of support for Ventcell-type reflected diffusions]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(2):135–210, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1996__32_2_135_0.

Peterson:2009:QLT

[Pet09a] Jonathon Peterson. Quenched limits for transient, ballistic, sub-Gaussian one-dimensional random walk in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):685–709, August 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1249391380;   http://
www.numdam.org/item?id=AIHPB_2009__45_3_685_0.

Petrelis:2009:CSI

[Pet09b] Nicolas Petrelis. Copolymer at selective interfaces and pinning potentials: weak coupling limits. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(1):175–200, February 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1234469977;   http://www.numdam.org/
item?id=AIHPB_2009__45_1_175_0.

Philoche:1971:PDT

[Phi71] J.-L. Philoche. À propos du théorème de Gauss–Markov. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(4):271– 281, October–December 1971. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1971__
7_4_271_0.

Phuc:1981:UDI

[Phu81] Ho Dang Phuc. Universal distribution for infinitely divisible distributions on Frèchet space. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(2):219–227, April–June 1981. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1981__17_2_219_0.

Piau:1995:MBC

[Pia95] Didier Piau. Martingales browniennes et conjecture de Sakai. (French) [Brownian martingales and a conjecture of Sakai]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(3):429–452, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1995__31_3_429_0.

Picard:1964:EAE

[Pic64] Ph. Picard. Étude analytique de l’équation de diffusion des gènes et de certaines de ses généralisations. (French) [Analytical study of the diffusion equation of genes and some of its generalizations]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 1 (1):23–109, 1964. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1964__1_1_23_0.

Picard:1965:MSL

[Pic65] Ph. Picard. Sur les modèles stochastiques logistiques en démographie. (French) [On logistic stochastic models in demography]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(2):151–172, October–December 1965. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1965__
2_2_151_0.

Picard:1994:BMV

[Pic94] Jean Picard. Barycentres et martingales sur une variété. (French) [Centroids and martingales on a variety]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(4): 647–702, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1994__30_4_647_0.

Picard:1996:FDE

[Pic96] Jean Picard. Formules de dualité sur l’espace de Poisson. (French) [Duality formulas in Poisson space]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(4): 509–548, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_4_509_0.

Picard:2005:SCM

[Pic05] Jean Picard. Stochastic calculus and martingales on trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(4):631–683, July/August 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_4_631_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000512.

Pierre:2009:IDS

[Pie09] Patie Pierre. Infinite divisibility of solutions to some selfsimilar integro-differential equations and exponential functionals of Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):667–684, August 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1249391379;   http://www.numdam.org/item?id=AIHPB_2009_
_45_3_667_0.

Pigato:2018:TED

[Pig18] Paolo Pigato. Tube estimates for diffusion processes under a weak Hörmander condition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):299–342, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1519030830.

Pinsky:1985:NEH

[Pin85] Mark A. Pinsky. On non-Euclidean harmonic measure. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(1): 39–46, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1985__21_1_39_0.

Pinsky:1990:LCD

[Pin90] Ross G. Pinsky. The lifetimes of conditioned diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):87–99, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1990__26_1_87_0.

Pinsky:2010:TRS

[Pin10] Ross G. Pinsky. Transience/recurrence and the speed of a one-dimensional random walk in a “have your cookie and eat it” environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):949–964, November 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1288878331;   http://www.numdam.org/item?id=AIHPB_2010_
_46_4_949_0.

Pinelis:2014:BHI

[Pin14a] Iosif Pinelis. On the Bennett–Hoeffding inequality. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1): 15–27, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1388545263.

Pinsky:2014:TRS

[Pin14b] Ross G. Pinsky. Transience, recurrence and speed of diffusions with a non-Markovian two-phase “use it or lose it” drift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1198–1212, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555497.

Pisier:1984:RCV

[Pis84] Gilles Pisier. Remarques sur les classes de Vapnik–Cervonenkis. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(4):287–298, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1984__20_4_287_0.

Poisat:2013:QAC

[Poi13] Julien Poisat. On quenched and annealed critical curves of random pinning model with finite range correlations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49 (2):456–482, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1366117654.

Pollard:2002:MIB

[Pol02] David Pollard. Maximal inequalities via bracketing with adaptive truncation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):1039–1052, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_6_1039_0;   http://www.sciencedirect.
com/science/article/pii/S024602030201138X.

Pouyanne:2008:AAP

[Pou08] Nicolas Pouyanne. An algebraic approach to Pólya processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2):293–323, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1207948221;   http://
www.numdam.org/item?id=AIHPB_2008__44_2_293_0.

Povel:1998:ODA

[Pov98] Tobias Povel. The one dimensional annealed δ-Lyapounov exponent. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(1):61–72, ???? 1998. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_1_61_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800186.

Pedersen:1998:CEA

[PP98] J. L. Pedersen and G. Peškir. Computing the expectation of the Azéma–Yor stopping times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(2):265–276, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1998__34_2_265_0;   http://www.sciencedirect.
com/science/article/pii/S0246020398800332.

Pham-Phu-Hien:1975:MAD

[PPH75] Pham-Phu-Hien. Mesure asymptotique définie par une fonction à valeurs dans R^{n} ou dans un espace vectoriel topologique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(1): 23–107, January–March 1975. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1975__
11_1_23_0.

Pardoux:1997:PIS

[PPR97] Etienne Pardoux, Frédéric Pradeilles, and Zusheng Rao. Probabilistic interpretation of a system of semi-linear parabolic partial differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(4):467–490, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_4_467_0;   http://www.sciencedirect.com/science/
article/pii/S024602039780101X.

Pardo:2018:EMZ

[PPR18] J. C. Pardo, J. L. Pérez, and V. M. Rivero. The excursion measure away from zero for spectrally negative Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(1):75–99, February 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1519030820.

Pratelli:1992:CVL

[Pra92] L. Pratelli. Sur la convergence vers la loi de Poisson. (French) [Convergence to the Poisson law]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2):151–164, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1992__28_2_151_0.

Pratelli:2007:EBM

[Pra07] Aldo Pratelli. On the equality between Monge’s infimum and Kantorovich’s minimum in optimal mass transportation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1):1–13, January/February 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_2007__43_1_
1_0;   http://www.sciencedirect.com/science/article/
pii/S0246020306000148.

Priouret:1968:PMV

[Pri68] Pierre Priouret. Processus de Markov sur une variété à bord compacte. (French) [Markov processes on a variety of compact boundary]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(3):193– 253, July–September 1968. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1968__
4_3_193_0.

Privault:2012:IPM

[Pri12] Nicolas Privault. Invariance of Poisson measures under random transformations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(4):947–972, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1353098435.

Priola:2018:DTU

[Pri18] Enrico Priola. Davie’s type uniqueness for a class of SDEs with jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):694–725, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1524643226.

Protter:1986:SMP

[Pro86] Philip Protter. Semimartingales and measure preserving flows. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2):127–147, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1986__22_2_127_0.

Prokaj:2011:HCD

[PRS11] Vilmos Prokaj, Miklós Rásonyi, and Walter Schachermayer. Hiding a constant drift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(2):498–514, May 2011. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/

1300887279;   http://www.numdam.org/item?id=AIHPB_2011_
_47_2_498_0.

Pizzo:2013:FRD

[PRS13] Alessandro Pizzo, David Renfrew, and Alexander Soshnikov. On finite rank deformations of Wigner matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1): 64–94, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1359470126.

Pruss:1997:CBT

[Pru97a] Alexander R. Pruss. Comparisons between tail probabilities of sums of independent symmetric random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(5):651–671, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1997__33_5_651_0;   http://
www.sciencedirect.com/science/article/pii/S0246020397801082.

Pruss:1997:ODR

[Pru97b] Alexander R. Pruss. One-dimensional random walks, decreasing rearrangements and discrete Steiner symmetrization. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(1):83–112, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1997__33_1_83_0;   http://
www.sciencedirect.com/science/article/pii/S0246020397801173.

Pinsky:1992:SRE

[PS92] Ross Pinsky and Michael Scheutzow. Some remarks and examples concerning the transience and recurrence of random diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(4):519–536, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1992__28_4_519_0.

Pene:2009:QRT

[PS09] Françoise Pène and Benoît Saussol. Quantitative recurrence in two-dimensional extended processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1065–1084, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1257529892;   http://www.numdam.org/
item?id=AIHPB_2009__45_4_1065_0.

Peterson:2013:WQL

[PS13] Jonathon Peterson and Gennady Samorodnitsky. Weak quenched limiting distributions for transient one-dimensional random walk in a random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):722–752, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1372772642.

Peled:2014:OCH

[PS14] Ron Peled and Wojciech Samotij. Odd cutsets and the hardcore model on Z^{d}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):975–998, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1403277005.

Pillai:2019:MTC

[PS19] Natesh S. Pillai and Aaron Smith. Mixing times for a constrained Ising process on the two-dimensional torus at low density. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1649–1678, August 2019. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1569398881.

Peres:2009:DSI

[PSS09] Yuval Peres, Oded Schramm, and Jeffrey E. Steif. Dynamical sensitivity of the infinite cluster in critical percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2):491–514, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024678;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_491_0.

Peres:2016:MDT

[PSS16] Yuval Peres, Bruno Schapira, and Perla Sousi. Martingale defocusing and transience of a self-interacting random walk. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1009–1022, August 2016. CODEN AHPBAR.

ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1469723508.

Podolskij:2017:SDI

[PSV17] Mark Podolskij, Christian Schmidt, and Mathias Vetter. On Uand V -statistics for discontinuous Itô semimartingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(3): 1007–1050, August 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1500624029.

Perkins:1998:MSS

[PT98] Edwin A. Perkins and S. James Taylor. The multifractal structure of super-Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(1):97–138, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_1_97_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800204.

Picard:2002:EPD

[PT02] Dominique Picard and Karine Tribouley. Evolutionary Pareto distributions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):1023–1037, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_6_1023_0;   http://www.sciencedirect.
com/science/article/pii/S024602030201124X.

Pipiras:2008:IPC

[PT08] Vladas Pipiras and Murad S. Taqqu. Identification of periodic and cyclic fractional stable motions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4): 612–637, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1217964113;   http://www.numdam.org/
item?id=AIHPB_2008__44_4_612_0.

Pene:2021:PPT

[PT21] Françoise Pène and Damien Thomine. Probabilistic potential theory and induction of dynamical systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1736–1767, August 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-3/Probabilistic-potential-theory-and-
induction-of-dynamical-systems/10.1214/20-AIHP1122.
full.

Pellaumail:1979:IRS

[PW79] J. Pellaumail and A. Weron. Integrals related to stationary processes and cylindrical martingales. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(2):127–146, April–June 1979. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_2_127_0.

Pardoux:1994:SRD

[PW94] E. Pardoux and R. J. Williams. Symmetric reflected diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(1):13–62, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1994__30_1_13_0.

Perera:2002:IVS

[PW02] Gonzalo Perera and Mario Wschebor. Inference on the variance and smoothing of the paths of diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):1009–1022, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_1009_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020302011263.

Powell:2017:LLG

[PW17] Ellen Powell and Hao Wu. Level lines of the Gaussian free field with general boundary data. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):2229–2259, November 2017. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1511773744.

Perry:2020:SLS

[PWB20] Amelia Perry, Alexander S. Wein, and Afonso S. Bandeira. Statistical limits of spiked tensor models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):230–264, February 2020. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Statistical-limits-of-spiked-tensor-
models/10.1214/19-AIHP960.full.

Peng:2005:SSR

[PX05] Shige Peng and Mingyu Xu. The smallest g-supermartingale and reflected BSDE with single and double L^{2} obstacles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):605–630, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_3_605_0;   http://
www.sciencedirect.com/science/article/pii/S0246020305000245.

Privault:2004:MBR

[PZ04] Nicolas Privault and Jean-Claude Zambrini. Markovian bridges and reversible diffusion processes with jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(5): 599–633, September/October 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_5_599_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000214.

Pain:2021:TTO

[PZ21] Michel Pain and Olivier Zindy. Two-temperatures overlap distribution for the 2D discrete Gaussian free field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):685–699, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-
2/Two-temperatures-overlap-distribution-for-the-2D-
discrete-Gaussian-free/10.1214/20-AIHP1091.full.

Qian:2019:CBL

[Qia19] Wei Qian. Conditioning a Brownian loop-soup cluster on a portion of its boundary. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):314–340, February 2019. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1547802402.

Quastel:2015:TED

[QR15] Jeremy Quastel and Daniel Remenik. Tails of the endpoint distribution of directed polymers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(1):1–17, February 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1421244396.

R:1983:ABQ

[R.83] José Rafael Leon R. Asymptotic behaviour of the quadratic measure of deviation of multivariate density estimates. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(3):297–309, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1983__19_3_297_0.

Rabeherimanana:1994:PGD

[Rab94] T. J. Rabeherimanana. Principe de grandes déviations en théorie du filtrage non linéaire et algèbres de Lie nilpotentes. (French) [Large-deviation principle in nonlinear filtering theory and nilpotent Lie algebras]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(3):331–352, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1994__30_3_331_0.

Raimond:1999:FBI

[Rai99] Olivier Raimond. Flots browniens isotropes sur la sphère. (French) [Isotropic Brownian flows on the sphere]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35 (3):313–354, May/June 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_3_313_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399800144.

Raoult:1970:PAL

[Rao70] Jean-Pierre Raoult. Propriétés asymptotiques locales des tests. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(1): 61–113, January–March 1970. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1970__
6_1_61_0.

Rao:1972:PSS

[Rao72] M. M. Rao. Prediction sequences in smooth Banach spaces. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(4):319–332, October– December 1972. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1972__8_4_319_0.

Rassoul-Agha:2009:ASF

[RAS09] Firas Rassoul-Agha and Timo Seppäläinen. Almost sure functional central limit theorem for ballistic random walk in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2):373–420, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1241024674;   http://
www.numdam.org/item?id=AIHPB_2009__45_2_373_0.

Raschel:2011:GFK

[Ras11a] Kilian Raschel. Green functions for killed random walks in the Weyl chamber of Sp(4). Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1001–1019, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1317906498;   http://www.numdam.org/
item?id=AIHPB_2011__47_4_1001_0.

Rassoul-Agha:2011:PLQ

[RAS11b] Firas Rassoul-Agha and Timo Seppäläinen. Process-level quenched large deviations for random walk in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):214–242, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1294170236;   http://
www.numdam.org/item?id=AIHPB_2011__47_1_214_0.

Raugi:1983:DTC

[Rau83] Albert Raugi. Une démonstration d’un théorème de Choquet– Deny par les martingales. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):101–109, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1983__19_1_101_0.

Raugi:1992:TSO

[Rau92] Albert Raugi. Théorie spectrale d’un opérateur de transition sur un espace métrique compact. (French) [Spectral theory of a transition operator on a compact metric space]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(2): 281–309, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1992__28_2_281_0.

Raugi:2004:GCD

[Rau04] Albert Raugi. A general Choquet–Deny theorem for nilpotent groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(6):677–683, November/December 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_6_677_0;   http://www.sciencedirect.com/science/
article/pii/S024602030400024X.

Raugi:2009:PED

[Rau09] Albert Raugi. A probabilistic ergodic decomposition result. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):932–942, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1257529886;   http://
www.numdam.org/item?id=AIHPB_2009__45_4_932_0.

Raynaud:1968:GAF

[Ray68] Hervé Raynaud. Sur les graphes aléatoires. (French) [On random graphs]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(4):255– 329, October–December 1968. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1968__
4_4_255_0.

Ray:2015:LUM

[Ray15] Gourab Ray. Large unicellular maps in high genus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51 (4):1432–1456, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1445432047.

Rafferty:2018:MCH

[RCG18] Thomas Rafferty, Paul Chleboun, and Stefan Grosskinsky. Monotonicity and condensation in homogeneous stochastic particle systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):790–818, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1524643229.

Raad:2020:SMF

[RDL20] Mads Bonde Raad, Susanne Ditlevsen, and Eva Löcherbach. Stability and mean-field limits of age dependent Hawkes processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1958–1990, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-3/Stability-and-mean-field-limits-of-
age-dependent-Hawkes-processes/10.1214/19-AIHP1023.
full.

Rousseau-Egele:1979:LFG

[RE79] J. Rousseau-Egele. La loi forte des grands nombres pour les processus harmonisables. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(2):175–186, April–June 1979. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_2_175_0.

Rebolledo:2005:DQM

[Reb05]  Rolando Rebolledo.  Decoherence of quantum Markov

semigroups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):349–373, ???? 2005. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2005_
_41_3_349_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000257.

Recoules:1986:AMB

[Rec86] Raymond Recoules. Approximation d’un mouvement brownien et d’un pont brownien par un processus stationnaire. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(1):113–125, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1986__22_1_113_0.

Regnier:1970:TEI

[Reg70] A. Regnier. Théorèmes ergodiques individuels purement topologiques. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(3):271–280, July–September 1970. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1970__6_3_271_0.

Reinhard:1982:IDT

[Rei82] J. M. Reinhard. Identités du type Baxter–Spitzer pour une classe de promenades aléatoires semi-markoviennes. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(4):319–333, October–December 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1982__18_4_319_0.

Revuz:1983:LDT

[Rev83] D. Revuz. Lois du tout ou rien et comportement asymptotique pour les probabilités de transition des processus de Markov. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):9–24, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1983__19_1_9_0.

Rezakhanlou:2002:CLT

[Rez02] Fraydoun Rezakhanlou. A central limit theorem for the asymmetric simple exclusion process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(4):437–464, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_4_437_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301011025.

Renyi:1964:ESP

[RF64] Alfred Rényi and Maurice Fréchet. Sur les espaces simples des probabilités conditionnelles. (French) [On simple probability conditional spaces]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 1(1):3–21, 1964. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1964__1_1_3_0.

Rhee:1984:DNG

[Rhe84] Wansoo T. Rhee. On the distribution of the norm for a Gaussian measure. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(3):277–286, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1984__20_3_277_0.

Rhodes:2008:HST

[Rho08]  Rémi Rhodes.  On homogenization of space–time de-

pendent and degenerate random flows II. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(4): 673–692, August 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1217964115;   http://www.numdam.org/
item?id=AIHPB_2008__44_4_673_0.

Rhodes:2009:HLS

[Rho09] Rémi Rhodes. Homogenization of locally stationary diffusions with possibly degenerate diffusion matrix. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4): 981–1001, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1257529888;   http://www.numdam.org/
item?id=AIHPB_2009__45_4_981_0.

Richard:2013:LPC

[Ric13] Mathieu Richard. Lévy processes conditioned on having a large height process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):982–1013, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1380718734.

Rio:1993:CIS

[Rio93] Emmanuel Rio. Covariance inequalities for strongly mixing processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(4):587–597, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_4_587_0.

Rio:2002:IBP

[Rio02] Emmanuel Rio. Une inégalité de Bennett pour les maxima de processus empiriques. (French) [An inequality of Bennett for the maxima of empirical processes]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):1053–1057, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_1053_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020302011226.

Rio:2009:UBM

[Rio09] Emmanuel Rio. Upper bounds for minimal distances in the central limit theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(3):802–817, August 2009. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1249391386;   http://www.numdam.org/item?id=AIHPB_2009_
_45_3_802_0.

Rivero:2007:SCR

[Riv07] Víctor Rivero. Sinaï’s condition for real valued Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(3):299–319, May/June 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_3_299_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000446.

Rivero:2012:TAE

[Riv12] Víctor Rivero. Tail asymptotics for exponential functionals of Lévy processes: The convolution equivalent case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48 (4):1081–1102, November 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1353098441.

Rizzolo:2015:SLM

[Riz15] Douglas Rizzolo. Scaling limits of Markov branching trees and Galton–Watson trees conditioned on the number of vertices with out-degree in a given set. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):512–532, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1428672679.

Rajput:1981:LZP

[RLT81] Balram S. Rajput, Donald Louie, and A. Tortrat. Une loi de zéro-un pour une classe de mesures sur les groupes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(4):331–336, October–December 1981. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1981__
17_4_331_0.

Raimond:2021:SVR

[RN21] Olivier Raimond and Tuan-Minh Nguyen. Strongly vertexreinforced jump process on a complete graph. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57 (3):1549–1568, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/journals/annales-de-linstitut-henri-
poincare-probabilites-et-statistiques/volume-57/issue-
3/Strongly-vertex-reinforced-jump-process-on-a-complete-
graph/10.1214/20-AIHP1115.full.

Robert:1987:PVM

[Rob87] Philippe Robert. Sur un processus de vie et de mort de particules sur [0,1]. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(2):225–235, ???? 1987. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1987_
_23_2_225_0.

Rogers:1984:NIR

[Rog84] L. C. G. Rogers. A new identity for real Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(1): 21–34, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1984__20_1_21_0.

Rollin:2013:SMH

[Röl13] Adrian Röllin. Stein’s method in high dimensions with applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):529–549, May 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1366117657.

Romier:1969:ISMa

[Rom69a] G. Romier. Introduction à la statistique mathématique. I. Modèle d’expérimentation statistique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(4):275–288, October–December 1969. CODEN AHPBAR. ISSN 0020-2347. URL http://

www.numdam.org/item?id=AIHPB_1969__5_4_289_0.

Romier:1969:ISMb

[Rom69b] G. Romier. Introduction à la statistique mathématique. III. Décision statistique. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(4):323–355, October–December 1969. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1969__5_4_357_0.

Rosenthal:1986:WPP

[Ros86] Alain Rosenthal. Weak Pinsker property and Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(3):347–369, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1986__22_3_347_0.

Rosenthal:1989:SEM

[Ros89] A. Rosenthal. On strictly ergodic models for commuting ergodic transformations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(1):73–92, ???? 1989. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_1989__25_1_73_0.

Rosen:1996:JCR

[Ros96] Jay S. Rosen. Joint continuity of renormalized intersection local times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(6):671–700, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1996__32_6_671_0.

Rosen:1999:JCD

[Ros99] Jay Rosen. Joint continuity and a Doob–Meyer type decomposition for renormalized intersection local times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35 (2):143–176, March/April 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1999__35_2_143_0;   http://
www.sciencedirect.com/science/article/pii/S0246020399800090.

Rosen:2001:DPI

[Ros01] Jay Rosen. Dirichlet processes and an intrinsic characterization of renormalized intersection local times. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37 (4):403–420, July/August 2001. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2001__37_4_403_0;   http://
www.sciencedirect.com/science/article/pii/S0246020301010792.

Rosen:2010:CDR

[Ros10] Jay S. Rosen. Continuous differentiability of renormalized intersection local times in R^{1}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1025–1041, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878336;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_1025_0.

Roth:1990:REC

[Rot90] Jean-Pierre Roth. Reformulation et extension de certains théorèmes ergodiques. (French) [Reformulation and extension of some ergodic theorems]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):437–450, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1990__26_3_437_0.

Roudier:1972:CMC

[Rou72] J. Roudier. Chaîne de Markov μ-continue à l’infini. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(3):241–248, July– September 1972. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1972__8_3_241_0.

Rouault:1978:LZS

[Rou78] Alain Rouault. Lois de Zipf et sources markoviennes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(2):169–188, April– June 1978. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1979__15_2_195_0. See

[Rou79].

Rouault:1979:EAR

[Rou79] A. Rouault. Errata de l’article A. Rouault. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(2):195–196, April–June 1979. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1978__14_2_169_0. See [Rou78].

Roussignol:1980:PSI

[Rou80] M. Roussignol. Un processus de saut sur R a une infinité de particules. (French) [A jump process on R has an infinity of particles]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(2): 101–108, April–June 1980. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1980__
16_2_101_0.

Roussignol:1986:PSA

[Rou86] M. Roussignol. Processus de saut avec interaction selon les plus proches particules. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(2):175–198, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1986__22_2_175_0.

Rouault:1987:EMP

[Rou87a] Alain Rouault. Espérances et majorations pour un processus de branchement spatial markovien. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(3): 459–497, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1987__23_3_459_0.

Rouault:1987:PPP

[Rou87b] Alain Rouault. Probabilités de présence dans un processus de branchement spatial markovien. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(1):37– 61, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1987__23_1_37_0.

Roux:1997:AME

[Rou97] Daniel Roux. Analyse multi-échelle d’un processus gaussien markovien au voisinage d’une singularité. (French) [Multiscale analysis of a Markov Gaussian process in the vicinity of a singularity]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(3):295–322, ???? 1997. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1997_
_33_3_295_0;   http://www.sciencedirect.com/science/
article/pii/S0246020397800933.

Rousseau:2021:LCS

[Rou21] Jérôme Rousseau. Longest common substring for random subshifts of finite type. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1768–1785, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Longest-common-substring-
for-random-subshifts-of-finite-type/10.1214/20-AIHP1130.
full.

Roynette:1973:PPD

[Roy73] B. Roynette. Points polaires d’une diffusion. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 9(3):285–309, July– September 1973. CODEN AHPBAR. ISSN 0020-2347. URL

http://www.numdam.org/item?id=AIHPB_1973__9_3_285_0.

Roynette:1974:TCL

[Roy74] Bernard Roynette. Théorème central-limite pour le groupe des déplacements de R^{d}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(4):391–398, October–December 1974. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1974__10_4_391_0.

Royer:1980:CEP

[Roy80] Gilles Royer. Croissance exponentielle de produits markoviens de matrices aléatoires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(1):49–62, January–March 1980. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1980__16_1_49_0.

Reinhard:1970:RDP

[RR70a] Hervé Reinhard and Bernard Roynette. Recollement de deux processus de Markov et fonctionnelles additives. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(1):27–40, January– March 1970. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1970__6_1_27_0.

Reinhard:1970:RPM

[RR70b] Hervé Reinhard and Bernard Roynette. Retournement des processus de Markov à un temps fixe. Remarque sur l’hypothèse de dualité. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6 (1):41–60, January–March 1970. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1970__6_1_41_0.

Rodzik:1994:ASC

[RR94] Beata Rodzik and Zdzislaw Rychlik. An almost sure central limit theorem for independent random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(1):1– 11, ???? 1994. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1994__30_1_1_0.

Ramanan:2018:IBD

[RS18] Kavita Ramanan and Mykhaylo Shkolnikov. Intertwinings of beta-Dyson Brownian motions of different dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):1152–1163, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1524643243.

Ray:2019:AND

[RSH19] Kolyan Ray and Johannes Schmidt-Hieber. Asymptotic nonequivalence of density estimation and Gaussian white noise for small densities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):2195–2208, November 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-55/issue-4/Asymptotic-nonequivalence-
of-density-estimation-and-Gaussian-white-noise-for/
10.1214/18-AIHP946.full.

Roy:2016:RDF

[RSS16] Rahul Roy, Kumarjit Saha, and Anish Sarkar. Random directed forest and the Brownian web. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1106–1143, August 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1469723513.

Rockner:2021:SCO

[RSX21] Michael Röckner, Xiaobin Sun, and Yingchao Xie. Strong convergence order for slow-fast McKean–Vlasov stochastic differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):547–576, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/Strong-convergence-order-
for-slowfast-McKeanVlasov-stochastic-differential-equations/
10.1214/20-AIHP1087.full.

Rezakhanlou:1997:BGL

[RT97] Fraydoun Rezakhanlou and James E. Tarver. Boltzmann-grad limit for a particle system in continuum. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(6):753–796, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1997__33_6_753_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397801124.

Rossignol:2010:LLD

[RT10] Raphaël Rossignol and Marie Théret. Lower large deviations and laws of large numbers for maximal flows through a box in first passage percolation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1093–1131, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878340;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_1093_0.

Rolla:2018:NFB

[RT18] L. T. Rolla and L. Tournier. Non-fixation for biased Activated Random Walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):938–951, May 2018. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1524643235.

Ruckebusch:1980:TGR

[Ruc80] Guy Ruckebusch. Théorie géométrique de la représentation markovienne. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(3):225–297, July–September 1980. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1980__16_3_225_0.

Rhodes:2013:OTM

[RV13] Rémi Rhodes and Vincent Vargas. Optimal transportation for multifractal random measures and applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1): 119–137, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1359470128.

Rivera:2019:QIN

[RV19] Alejandro Rivera and Hugo Vanneuville. Quasi-independence for nodal lines. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1679–1711, August 2019. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL https://projecteuclid.org/euclid.aihp/
1569398882.

Radke:2020:TSC

[RV20] Jordan Radke and Beatrice-Helen Vritsiou. On the thin-shell conjecture for the Schatten classes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):87–119, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/On-the-thin-shell-conjecture-
for-the-Schatten-classes/10.1214/18-AIHP956.full.

Roynette:2009:BPR

[RVY09] B. Roynette, P. Vallois, and M. Yor. Brownian penalisations related to excursion lengths, VII. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2): 421–452, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024675;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_421_0.

Randrianantoanina:2017:NBR

[RW17] Narcisse Randrianantoanina and Lian Wu. Noncommutative Burkholder/Rosenthal inequalities associated with convex functions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):1575–1605, November 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1511773719.

Rachev:1989:SMM

[RY89] S. T. Rachev and J. E. Yukich. Smoothing metrics for measures on groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):429–441, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1989__25_4_429_0.

Rajeev:1995:LTA

[RY95] B. Rajeev and M. Yor. Local times and almost sure convergence of semi-martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(4):653–667, ???? 1995. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1995_
_31_4_653_0.

Roelly:1993:CMG

[RZ93] Sylvie Roelly and Hans Zessin. Une caractérisation des mesures de Gibbs sur C(0, 1)Z^{d} par le calcul des variations stochastiques. (French) [A characterization of Gibbs measures on C(0, 1)Z^{d} by calculating stochastic variations]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(3): 327–338, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1993__29_3_327_0.

Rezaei:2017:HMN

[RZ17] Mohammad A. Rezaei and Dapeng Zhan. Higher moments of the natural parameterization for SLE curves. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1): 182–199, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1486544888.

Robles:2020:RPL

[RZ20] Nicolas Robles and Dirk Zeindler. Random permutations with logarithmic cycle weights. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1991–2016, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Random-permutations-
with-logarithmic-cycle-weights/10.1214/19-AIHP1025.
full.

Saada:1988:MIP

[Saa88] Ellen Saada. Mesures invariantes pour les systèmes à une infinité de particules linéaires à valeurs dans [0, ∞[^{s}. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(4):427–437, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1988__24_4_427_0.

Saada:1990:PZR

[Saa90] Ellen Saada. Processus de zéro-rangé avec particule marquée. (French) [Zero-ranked process with tagged particle]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1): 5–17, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_1_5_0.

Saada:1995:MDV

[Saa95] Ellen Saada. Un modèle du votant en milieu aléatoire. (French) [A model of voting in random media]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):263–271, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1995__31_1_263_0.

Sabot:2001:IDS

[Sab01] Christophe Sabot. Integrated density of states of self-similar Sturm–Liouville operators and holomorphic dynamics in higher dimension. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(3):275–311, May/June 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_3_275_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010682.

Samson:2007:ICD

[Sam07] Paul-Marie Samson. Infimum-convolution description of concentration properties of product probability measures, with applications. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(3):321–338, May/June 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_3_321_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000690.

Sart:2014:ETD

[Sar14] Mathieu Sart. Estimation of the transition density of a Markov chain. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):1028–1068, August 2014. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1403277007.

Sarantsev:2017:ISC

[Sar17] Andrey Sarantsev. Infinite systems of competing Brownian particles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(4):2279–2315, November 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1511773746.

Sarantsev:2018:WCO

[Sar18] Andrey Sarantsev. Weak convergence of obliquely reflected diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1408–1431, August 2018. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

https://projecteuclid.org/euclid.aihp/1531296024.

Sart:2021:EDH

[Sar21] Mathieu Sart. Estimating a density, a hazard rate, and a transition intensity via the ρ-estimation method. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):195–249, February 2021. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-1/Estimating-a-density-a-hazard-rate-
and-a-transition-intensity/10.1214/20-AIHP1076.full.

Stasinski:2021:DMB

[SBM21] Roman Stasiński, Julien Berestycki, and Bastien Mallein. Derivative martingale of the branching Brownian motion in dimension. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1786–1810, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-3/Derivative-martingale-of-the-branching-
Brownian-motion-in-dimension-d1/10.1214/20-AIHP1131.
full.

Schmitt:1972:TEP

[Sch72a] Bernard Schmitt. Théorème ergodique ponctuel pour les suites uniformes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(4):387–394, October–December 1972. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1972__8_4_387_0.

Schreiber:1972:QRC

[Sch72b] Michel Schreiber. Quelques remarques sur les caractérisations des espaces L^{p}, 0 ≤ p < 1. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(1):83–92, January–March 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1972__8_1_83_0.

Schachermayer:1983:CBS

[Sch83] Walter Schachermayer. The class of Banach spaces, which do not have c_{0} as a spreading model, is not L^{2}-hereditary. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):

1–8, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1983__19_1_1_0.

Schwartz:1985:MBT

[Sch85] Laurent Schwartz. Le mouvement brownien sur R^{N}, en tant que semi-martingale dans S_{N}. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(1):15–25, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1985__21_1_15_0.

Schwartz:1989:EDS

[Sch89] Laurent Schwartz. Équation différentielle stochastique (EDS) sur R^{N} et sur R^{N} ∪ {∞} = S_{N}. (French) [Stochastic differential equation (SDE) on R^{N} and R^{N∪{∞}=S_{N}}]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(3): 259–263, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1989__25_3_259_0.

Schneider:2000:PTM

[Sch00] Dominique Schneider. Polynômes trigonométriques et marches aléatoires multidimensionnelles: application à la théorie ergodique. (French) [Trigonometric polynomials and multidimensional random walks: application to ergodic theory]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(5):617–646, September 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2000__36_5_617_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300001382.

Schmitz:2006:DRE

[Sch06] Tom Schmitz. Diffusions in random environment and ballistic behavior. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(6):683–714, November/December 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_6_683_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000045.

Schapira:2009:RWB

[Sch09] Bruno Schapira. Random walk on a building of type Ã_{r} and Brownian motion of the Weyl chamber. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(2): 289–301, May 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1241024671;   http://www.numdam.org/
item?id=AIHPB_2009__45_2_289_0.

Senoussi:1990:PIM

[Sen90a] R. Senoussi. Problème d’identification dans le modèle de Cox. (French) [Identification problem in the Cox model]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1): 45–64, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_1_45_0.

Senoussi:1990:SAP

[Sen90b] R. Senoussi. Statistique asymptotique presque-sûre de modèles statistiques convexes. (French) [Almost-sure asymptotic statistics of statistically-convex models]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):19–44, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1990_
_26_1_19_0.

Sengul:2017:SLC

[Sen17] Bati Sengül. Scaling limits of coalescent processes near time zero. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):616–640, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1491897738.

Seppalainen:2000:SLL

[Sep00] Timo Seppäläinen. Strong law of large numbers for the interface in ballistic deposition. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(6):691–736, November 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_6_691_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001370.

Sepulveda:2019:TLS

[Sep19] Avelio Sepúlveda. On thin local sets of the Gaussian free field. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(3):1797–1813, August 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1569398886.

Seppalainen:2020:EUC

[Sep20] Timo Seppäläinen. Existence, uniqueness and coalescence of directed planar geodesics: Proof via the increment-stationary growth process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(3):1775–1791, August 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-3/Existence-uniqueness-
and-coalescence-of-directed-planar-geodesics--Proof/
10.1214/19-AIHP1016.full.

Sethuraman:2001:EMC

[Set01] Sunder Sethuraman. On extremal measures for conservative particle systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(2):139–154, March 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_2_139_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010621.

Sethuraman:2007:DTP

[Set07] Sunder Sethuraman. On diffusivity of a tagged particle in asymmetric zero-range dynamics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(2):215–232, March/ April 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2007__43_2_215_0;   http://www.sciencedirect.
com/science/article/pii/S0246020306000458.

Souchet:2002:EYW

[SG02] Sandie Souchet and Xavier Guyon. Estimation de Yule–Walker d’un CAR(p) observé à temps discret. (French) [Yule-Walker estimate of a CAR(p) observed at discrete time]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):

1093–1100, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//www.numdam.org/item?id=AIHPB_2002__38_6_1093_0;
http://www.sciencedirect.com/science/article/pii/S0246020302011354.

En l’honneur de J. Bretagnolle, D. Dacunha-Castelle, I. Ibragimov.

Shen:2003:BDR

[She03] Lian Shen. On ballistic diffusions in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(5):839–876, September/October 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_3_385_0;   http://www.sciencedirect.com/science/
article/pii/S024602030300027X. See addendum [She04].

Shen:2004:AAB

[She04] Lian Shen. Addendum to the article “On ballistic diffusions in random environment”: Annales de l’Institut Henri Poincaré Probabilités et statistiques 39 (5) (2003) 839–876. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(3):385–386, May/June 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_5_839_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000111. See [She03].

Shen:2018:LPS

[She18] Yi Shen. Location of the path supremum for self-similar processes with stationary increments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2349–2360, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849802.

Shiraishi:2010:HKR

[Shi10] Daisuke Shiraishi. Heat kernel for random walk trace on Z^{3} and Z^{4}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1001–1024, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1288878335;   http://
www.numdam.org/item?id=AIHPB_2010__46_4_1001_0.

Shiraishi:2019:HDS

[Shi19] Daisuke Shiraishi. Hausdorff dimension of the scaling limit of loop-erased random walk in three dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):791–834, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820832.

Shi:2020:GFM

[Shi20] Quan Shi. A growth-fragmentation model related to Ornstein– Uhlenbeck type processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):580–611, February 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-1/A-growth-fragmentation-
model-related-to-OrnsteinUhlenbeck-type-processes/10.
1214/19-AIHP974.full.

Siebert:1984:SCS

[Sie84] Eberhard Siebert. Semistable convolution semigroups on measurable and topological groups. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(2):147–164, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1984__20_2_147_0.

Simenhaus:2007:ADR

[Sim07] François Simenhaus. Asymptotic direction for random walks in random environments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6):751–761, November/December 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_6_751_0;   http://
www.sciencedirect.com/science/article/pii/S0246020307000052.

Singh:2007:LBD

[Sin07] Arvind Singh. Limiting behavior of a diffusion in an asymptotically stable environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(1):101–138, January/February 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=

AIHPB_2007__43_1_101_0;   http://www.sciencedirect.
com/science/article/pii/S0246020306000215.

Shea:2014:EDP

[SJ14] Brian P. Shea and Galin L. Jones. Evaluating default priors with a generalization of Eaton’s Markov chain. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3): 1069–1091, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1403277008.

Saada:1995:PAN

[SL95] E. Saada and C. Landim. Préface au numéro spécial en l’honneur de Claude Kipnis. (French) [Preface to the special issue in honor of Claude Kipnis]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):3–12, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1995_
_31_1_3_0.

Slade:2022:SCP

[Sla22] Gordon Slade. A simple convergence proof for the lace expansion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):26–33, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-58/issue-1/A-simple-convergence-proof-for-the-
lace-expansion/10.1214/21-AIHP1166.full.

Slominski:1993:EUS

[Sło93] Leszek Słomínski. On existence, uniqueness and stability of solutions of multidimensional SDE’s with reflecting boundary conditions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(2):163–198, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_2_163_0.

Szpirglas:1979:MGF

[SM79] J. Szpirglas and G. Mazziotto. Modèle général de filtrage non linéaire et équations différentielles stochastiques associées. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(2):

147–173, April–June 1979. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1979__
15_2_147_0.

Smolenski:1983:LLM

[Smo83] W. Smolenski. Linear Lusin-measurable functionals in case of a continuous cylinder measure. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(4):311–321, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1983__19_4_311_0.

Smythe:1975:EPM

[Smy75] R. T. Smythe. Ergodic properties of marked point processes in R^{r}. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(2):109– 125, April–June 1975. CODEN AHPBAR. ISSN 0020-2347.

URL http://www.numdam.org/item?id=AIHPB_1975__11_2_
109_0.

Sato:1975:SGR

[SO75] Hiroshi Sato and Yoshiaki Okazaki. Separabilities of a Gaussian Radon measure. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(3):287–298, July–September 1975. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1975__11_3_287_0.

Sodin:2008:IIB

[Sod08] Sasha Sodin. An isoperimetric inequality on the ℓ_{p} balls. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(2):362–373, April 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1207948224;   http://www.numdam.org/
item?id=AIHPB_2008__44_2_362_0.

Sohier:2013:SLH

[Soh13] Julien Sohier. The scaling limits of a heavy tailed Markov renewal process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(2):483–505, May 2013. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1366117655.

Song:2013:CPS

[Son13] Yongsheng Song. Characterizations of processes with stationary and independent increments under G-expectation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1): 252–269, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1359470134.

Saint-Pierre:1975:DMN

[SP75] Jean Saint-Pierre. Désintégration d’une mesure non bornée. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(3): 275–286, July–September 1975. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1975__11_3_275_0.

Saint-Pierre:1976:MBV

[SP76] Jean Saint-Pierre. Multiapplications boréliennes à valeurs convexes de R^{n}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(2):163–180, April–June 1976. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1976__12_2_163_0.

Spokoiny:2017:PML

[Spo17] Vladimir Spokoiny. Penalized maximum likelihood estimation and effective dimension. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):389–429, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1486544896.

Sahu:2010:CUP

[SS10a] Lingaraj Sahu and Kalyan B. Sinha. Characterization of unitary processes with independent and stationary increments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):575–593, May 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1273584135;   http://www.numdam.org/
item?id=AIHPB_2010__46_2_575_0.

Sidoravicius:2010:CBV

[SS10b] Vladas Sidoravicius and Alain-Sol Sznitman. Connectivity bounds for the vacant set of random interlacements. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4): 976–990, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878333;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_976_0.

Shu:2018:CCC

[SS18] Yan Shu and Michał Strzelecki. A characterization of a class of convex log-Sobolev inequalities on the real line. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54 (4):2075–2091, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1539849793.

Shi:2021:USF

[SSS^{+}21] Zhan Shi, Vladas Sidoravicius, He Song, Longmin Wang, and Kainan Xiang. Uniform spanning forests on biased Euclidean lattices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1569–1582, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Uniform-spanning-forests-
on-biased-Euclidean-lattices/10.1214/20-AIHP1119.full.

Sanz-Sole:2003:EHS

[SSV03] Marta Sanz-Solé and Pierre-A. Vuillermot. Equivalence and Hölder–Sobolev regularity of solutions for a class of nonautonomous stochastic partial differential equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39 (4):703–742, July/August 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2003__39_4_703_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000153.

Sato:1994:MCR

[ST94] Hiroshi Sato and Masakazu Tamashiro. Multiplicative chaos and random translation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 30(2):245–264, ???? 1994. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1994_
_30_2_245_0.

Sabot:2011:RDE

[ST11] Christophe Sabot and Laurent Tournier. Reversed Dirichlet environment and directional transience of random walks in Dirichlet environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1):1–8, February 2011. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1294170226;   http://www.numdam.org/item?id=AIHPB_2011_
_47_1_1_0.

Savov:2020:SMP

[ST20a]  Mladen Savov and Bruno Toaldo.  Semi-Markov pro-

cesses, integro-differential equations and anomalous diffusionaggregation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2640–2671, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-56/issue-4/Semi-Markov-processes-
integro-differential-equations-and-anomalous-diffusion-
aggregation/10.1214/20-AIHP1053.full.

Sousi:2020:CRW

[ST20b] Perla Sousi and Sam Thomas. Cutoff for random walk on dynamical Erdős–Rényi graph. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(4):2745–2773, November 2020. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-56/issue-4/Cutoff-for-random-
walk-on-dynamical-Erd%c5%91sR%c3%a9nyi-graph/10.1214/
20-AIHP1057.full.

Stadje:1989:AES

[Sta89] W. Stadje. On the asymptotic equidistribution of sums of independent identically distributed random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2): 195–203, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1989__25_2_195_0.

Stadje:1990:PPR

[Sta90] Wolfgang Stadje. Bin-packing problems for a renewal process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(1):207–217, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1990__26_1_207_0.

Stephens:1974:CGF

[Ste74] M. A. Stephens. Components of goodness-of-fit statistics. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(1):37–54, January– March 1974. CODEN AHPBAR. ISSN 0020-2347. URL http:

//www.numdam.org/item?id=AIHPB_1974__10_1_37_0.

Steele:1989:KSE

[Ste89] J. Michael Steele. Kingman’s subadditive ergodic theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(1):93–98, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1989__25_1_93_0.

Stricker:1988:VCP

[Str88] C. Stricker. Variation conditionnelle des processus stochastiques. (French) [Conditional variation of stochastic processes]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):295–305, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1988__24_2_295_0.

Stricker:1990:ALM

[Str90] Christophe Stricker. Arbitrage et lois de martingale. (French) [Arbitration and martingale laws]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):451–460, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_3_451_0.

Stufler:2017:SLR

[Stu17] Benedikt Stufler. Scaling limits of random outerplanar maps with independent link-weights. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):900–915, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1491897750.

Stufler:2019:LLL

[Stu19] Benedikt Stufler. Local limits of large Galton–Watson trees rerooted at a random vertex. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(1):155–183, February 2019. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1547802398.

Soucaliuc:2000:RCB

[STW00] Florin Soucaliuc, Bálint Tóth, and Wendelin Werner. Reflection and coalescence between independent one-dimensional Brownian paths. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(4):509–545, July 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_4_509_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001369.

Suda:2021:FFD

[Sud21] Hayate Suda. A family of fractional diffusion equations derived from stochastic harmonic chains with long-range interactions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(4):2268–2314, November 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-4/A-family-of-fractional-
diffusion-equations-derived-from-stochastic-harmonic/
10.1214/20-AIHP1133.full.

Sued:2005:RPD

[Sue05] M. Sued. Regularity properties of the diffusion coefficient for a mean zero exclusion process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(1):1–33, January/ February 2005. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_2005__41_1_1_0;   http://www.sciencedirect.
com/science/article/pii/S0246020304000494.

Sunyach:1975:CCM

[Sun75] C. Sunyach. Une classe de chaînes de Markov récurrentes sur un espace métrique complet. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(4):325–343, October–December 1975. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1975__11_4_325_0.

Sunyach:1987:TRM

[Sun87] C. Sunyach. Sur la transience et la récurrence des marches aléatoires en milieu aléatoire. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(4):613–626, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1987__23_4_613_0.

Salminen:2007:MID

[SV07] Paavo Salminen and Pierre Vallois. On maximum increase and decrease of Brownian motion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(6):655–676, November/December 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2007__43_6_655_0;   http://
www.sciencedirect.com/science/article/pii/S0246020307000076.

Song:2008:SLP

[SV08] Renming Song and Zoran Vondraček. On suprema of Lévy processes and application in risk theory. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(5): 977–986, October 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1222261921;   http://www.numdam.org/
item?id=AIHPB_2008__44_5_977_0.

Stojkovic:2011:IMS

[SvG11] I. Stojkovic and O. van Gaans. Invariant measures and a stability theorem for locally Lipschitz stochastic delay equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1121–1146, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1317906504;   http://
www.numdam.org/item?id=AIHPB_2011__47_4_1121_0.

Schilling:2011:CPL

[SW11] René L. Schilling and Jian Wang. On the coupling property of Lévy processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(4):1147–1159, November 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1317906505;   http://www.numdam.org/item?id=AIHPB_2011_
_47_4_1147_0.

Sheffield:2017:SCD

[SWW17] Scott Sheffield, Samuel S. Watson, and Hao Wu. Simple CLE in doubly connected domains. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):594–615, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1491897737.

Salminen:2005:PPI

[SY05] Paavo Salminen and Marc Yor. Properties of perpetual integral functionals of Brownian motion with drift. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(3):335–347, ???? 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2005__41_3_335_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000269.

Schapira:2011:WPR

[SY11] Bruno Schapira and Robert Young. Windings of planar random walks and averaged Dehn function. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(1): 130–147, February 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1294170233;   http://www.numdam.org/
item?id=AIHPB_2011__47_1_130_0.

Schreiber:2013:LTG

[SY13] T. Schreiber and J. E. Yukich. Limit theorems for geometric functionals of Gibbs point processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1158–1182, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1380718742.

Stroock:1997:MCA

[SZ97] Daniel W. Stroock and Weian Zheng. Markov chain approximations to symmetric diffusions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 33(5):619–649, ???? 1997. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1997__33_5_619_0;   http://www.sciencedirect.
com/science/article/pii/S0246020397801070.

Storm:2016:TVD

[SZ16] Julia Storm and Dirk Zeindler. Total variation distance and the Erdős–Turán law for random permutations with polynomially growing cycle weights. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1614–1640, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1479373242.

Seppalainen:2017:HHP

[SZ17] Timo Seppäläinen and Yun Zhai. Hammersley’s harness process: Invariant distributions and height fluctuations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1): 287–321, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1486544893.

Sznitman:1980:PPE

[Szn80] Alain-Sol Sznitman. Perturbations ponctuelles d’évolutions: construction d’espaces stationnaires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(4):299–326, October–December 1980. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1980__16_4_299_0.

Sznitman:1982:MLM

[Szn82] Alain-Sol Sznitman. “minimal length” multi-channel. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 18(1):103–114, January–March 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1982__18_1_103_0.

Szpirglas:1978:EED

[Szp78] J. Szpirglas. Sur l’équivalence d’équations différentielles stochastiques à valeurs mesures intervenant dans le filtrage markovien non linéaire. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(1):33–59, January–March 1978. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_1_33_0.

Takacs:1970:DSS

[Tak70] Lajos Takács. On the distribution of the supremum for stochastic processes. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 6(3):237– 247, July–September 1970. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1970__
6_3_237_0.

Talagrand:1988:STS

[Tal88] Michel Talagrand. Small tails for the supremum of a Gaussian process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):307–315, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1988__24_2_307_0.

Talagrand:1989:SSP

[Tal89] M. Talagrand. On subsets of L^{p} and p-stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(2):153–166, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1989__25_2_153_0.

Talagrand:1992:AHF

[Tal92] Michel Talagrand. Approximating a helix in finitely many dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 28(3):355–363, ???? 1992. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1992__28_3_355_0.

Talagrand:1995:MFH

[Tal95] Michel Talagrand. The missing factor in Hoeffding’s inequalities. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(4):689–702, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1995__31_4_689_0.

Talagrand:1996:ATF

[Tal96] Michel Talagrand. Applying a theorem of Fernique. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(6): 779–799, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_6_779_0.

Telcs:2008:RWG

[Tel08] András Telcs. Random walk on graphs with regular resistance and volume growth. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1):143–169, February 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1203969872;   http://www.numdam.org/item?id=AIHPB_2008_
_44_1_143_0.

Testud:2006:MQB

[Tes06] Benoît Testud. Mesures quasi-Bernoulli au sens faible: résultats et exemples. (French) [Quasi-Bernoulli measures in the weak sense: results and examples]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1):1–35, January/ February 2006. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_2006__42_1_1_0;   http://www.sciencedirect.
com/science/article/pii/S0246020305000452.

Tessera:2018:SCF

[Tes18] Romain Tessera. Speed of convergence in first passage percolation and geodesicity of the average distance. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(2):569–586, May 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1524643222.

Thibault:1981:ECI

[Thi81] Lionel Thibault. Espérances conditionnelles d’intégrandes semicontinus. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(4):337–350, October–December 1981. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1981__17_4_337_0.

Thieullen:1991:TEE

[Thi91] Michèle Thieullen. Temps d’explosion d’équations différentielles stochastiques du type Doléans–Dade. (French) [Blow-up times for stochastic differential equations of Doléans–Dade type]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(4):549–557, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1991__27_4_549_0.

Thompson:2013:RER

[Tho13] Russ Thompson. The rate of escape for random walks on polycyclic and metabelian groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(1):270–287, February 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1359470135.

Thu:1981:NVD

[Thu81] Nguyen Van Thu. A new version of Doeblin’s theorem. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(2):213–217, April–June 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_2_213_0.

Tian:2017:NLE

[Tia17] Xueting Tian. Nonexistence of Lyapunov exponents for matrix cocycles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(1):493–502, February 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1486544900.

Timar:2014:SRG

[Tim14] Ádám Timár. A stationary random graph of no growth rate. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(4):1161–1164, November 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1413555495.

Tolli:2000:BET

[Tol00] Filippo Tolli. A Berry–Esseen theorem on semisimple Lie groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(3):275–290, May 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_
_36_3_275_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001345.

Tortrat:1965:LPE

[Tor65] A. Tortrat. Lois de probabilité sur un espace topologique complètement régulier et produits infinis à termes indépendants dans un groupe topologique. (French) [Probability distributions on a completely regular topological space and infinite products of independent terms in a topological group]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 1(3):217–237, January–March 1965. CODEN AHPBAR. ISSN 0020-2347. URL http://

www.numdam.org/item?id=AIHPB_1965__1_3_217_0.

Tortrat:1966:LTC

[Tor66] A. Tortrat. Lois tendues et convolutions dénombrables dans un groupe topologique X. (French) [Strained laws and convolutions in a countable topological group X]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 2(4):279–298, April–June 1966. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1966__2_4_0_0. See erratum [Ano66].

Tortrat:1969:MAG

[Tor69] A. Tortrat. Sur les mesures aléatoires dans les groupes non abéliens (compacité, mesure de Palm associée et représentation de Poisson). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(1):31–47, January–March 1969. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1969__5_1_31_0.

Tortrat:1971:MAF

[Tor71] A. Tortrat. Sur les mesures aléatoires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 7(1):1–8, January–March 1971. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1971__7_1_1_0.

Tortrat:1976:RLS

[Tor76] A. Tortrat. τ-Régularité des lois, séparation au sens de A. Tulcea et propriété de Radon–Nikodym. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 12(2):131–150, April–June 1976. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1976__12_2_131_0.

Tortrat:1977:CFS

[Tor77a] A. Tortrat. Complément sur ^{≪} le support des lois indéfiniment divisibles ^{≫}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(3):293–298, July–September 1977. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1977__13_3_293_0.

Tortrat:1977:SLI

[Tor77b] A. Tortrat. Sur le support des lois indéfiniment divisibles dans les espaces vectoriels localement convexes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(1):27–42, January–March 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1977__13_1_27_0.

Tortrat:1978:ALI

[Tor78a] A. Tortrat. Atomes et lois indéfiniment divisibles dans un espace vectoriel. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):343–347, July–September 1978. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_3_343_0.

Tortrat:1978:CME

[Tor78b] A. Tortrat. Sur la comparaison d’une mesure μ dans un espace vectoriel X avec ses translatées. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(1):61–83, January–March 1978. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1978__14_1_61_0.

Tortrat:1978:SCS

[Tor78c] Albert Tortrat. Second complément sur le support des lois indéfiniment divisibles. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(3):349–354, July–September 1978. CODEN AH-

PBAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1978__14_3_349_0.

Tortrat:1979:LID

[Tor79] Albert Tortrat. Lois indéfiniment divisibles et théorèmes de Itô–Nisio et Yuriskii. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 15(1):85–92, January–March 1979. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1979__15_1_85_0.

Tortrat:1981:LSG

[Tor81] A. Tortrat. Lois stables dans un groupe. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(1):51–61, January–March 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_1_51_0.

Tortrat:1982:SCL

[Tor82] A. Tortrat. Sur le support de certaines lois indéfiniment divisibles dans un espace vectoriel. (French) []. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(4): 395–410, October–December 1982. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1982__18_4_395_0.

Torrisi:2017:PAP

[Tor17] Giovanni Luca Torrisi. Poisson approximation of point processes with stochastic intensity, and application to nonlinear Hawkes processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53(2):679–700, May 2017. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1491897741.

Tournarie:1969:EOI

[Tou69] M. Tournarie. Exploitation d’une observation isolée pour l’estimation de plusieurs inconnues. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(1):49–67, January–March 1969. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1969__5_1_49_0.

Touati:1983:TLC

[Tou83] A. Touati. Théorèmes de limite centrale fonctionnels pour les processus de Markov. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):43–55, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1983_
_19_1_43_0.

Touati:1996:VCL

[Tou96] A. Touati. Vitesse de convergence en loi de l’estimateur des moindres carrés d’un modèle autorégressif (cas mixte). (French) [Speed of convergence in the law of the least squares estimator of an autoregressive model (mixed case)]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(2): 211–230, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1996__32_2_211_0.

Tournier:2015:ADR

[Tou15] Laurent Tournier. Asymptotic direction of random walks in Dirichlet environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):716–726, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://projecteuclid.org/euclid.aihp/1428672688.

Trashorras:2002:LDT

[Tra02] José Trashorras. Large deviations for a triangular array of exchangeable random variables. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(5):649–680, ???? 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_2002__38_5_649_0;   http://www.sciencedirect.
com/science/article/pii/S0246020301010949.

Trabs:2015:IBI

[Tra15] Mathias Trabs. Information bounds for inverse problems with application to deconvolution and Lévy models. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51 (4):1620–1650, November 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1445432055.

Trouve:1996:RLD

[Tro96] Alain Trouvé. Rough large deviation estimates for the optimal convergence speed exponent of generalized simulated annealing algorithms. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(3):299–348, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1996__32_3_299_0.

Tsirelson:2002:SDD

[Tsi02] Boris Tsirelson. Spectral densities describing off-white noises. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):1059–1069, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_2002__38_6_
1059_0;   http://www.sciencedirect.com/science/article/
pii/S0246020302011330.

Ton-That-Long:1968:RCP

[TTL68] Ton-That-Long. Représentations canoniques des processus gaussiens. (French) [Canonical representations of Gaussian processes]. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 4(3):179– 191, July–September 1968. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1968__
4_3_179_0.

Ton-That-Long:1969:CFC

[TTL69] Ton-That-Long. Sur le calcul fonctionnel d’une contraction complètement non unitaire II. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 5(3):213–231, July–September 1969. CO-DEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.

org/item?id=AIHPB_1969__5_2_113_0.

Ton-Than-Long:1974:PAV

[TTL74] Tôn-Thân-Long. Propriétés analytiques des valeurs et vecteurs propres des opérateurs de l’espace de Hilbert. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B.

Calcul des Probabilités et Statistique, 10(2):167–183, April– June 1974. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1974__10_2_167_0.

Tuyen:1981:ABS

[Tuy81] Dão Quang Tuyên. On the asymptotic behaviour of sequences of random variables and of their previsible compensators. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(1):63–73, January–March 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_1_63_0.

Truong-Van:1981:GDT

[TV81] B. Truong-Van. Une généralisation du théorème de Kolmogorov– Aronszajn ^{≪} processus V -bornés q-dimensionnels: domaine spectral § dilatations stationnaires ^{≫}. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 17(1):31–49, January–March 1981. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1981__17_1_31_0.

Tribouley:1998:ADE

[TV98] Karine Tribouley and Gabrielle Viennet. L_{p} adaptive density estimation in a β mixing framework. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(2):179–208, ???? 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=
AIHPB_1998__34_2_179_0;   http://www.sciencedirect.
com/science/article/pii/S0246020398800290.

Tsatsoulis:2018:SGS

[TW18] Pavlos Tsatsoulis and Hendrik Weber. Spectral gap for the stochastic quantization equation on the 2-dimensional torus. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(3):1204–1249, August 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1531296018.

Ueki:1999:AES

[Uek99] Naomasa Ueki. Asymptotic expansion of stochastic oscillatory integrals with rotation invariance. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(4):417–457, July 1999.

CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_4_417_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001016.

Ustunel:1982:SIN

[Ust82] S. Ustunel. Stochastic integration on nuclear spaces and its applications. Annales de l’Institut Henri Poincaré. Section B. Probabilités et Statistiques, 18(2):165–200, April–June 1982. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1982__18_2_165_0.

Vuolle-Apiala:1989:TCS

[VA89] J. Vuolle-Apiala. Time-changes of self-similar Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):581–587, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1989__25_4_581_0.

Vuolle-Apiala:1986:DTS

[VAG86] J. Vuolle-Apiala and S. E. Graversen. Duality theory for selfsimilar processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 22(3):323–332, ???? 1986. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1986__22_3_
323_0.

Valadier:1978:PCM

[Val78] Michel Valadier. Sur le plongement d’un champ mesurable d’espaces métriques dans un champ trivial. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 14(2):165–168, April–June 1978. CODEN AHPBAR. ISSN 0020-2347. URL http://www.

numdam.org/item?id=AIHPB_1978__14_2_165_0.

Valadier:1980:ECM

[Val80] Michel Valadier. Sur l’espérance conditionnelle multivoque non convexe. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16 (2):109–116, April–June 1980. CODEN AHPBAR. ISSN 0020-

2347. URL http://www.numdam.org/item?id=AIHPB_1980__
16_2_109_0.

Vallois:1989:PCM

[Val89] P. Vallois. Sur le passage de certaines marches aléatoires planes au-dessus d’une hyperbole équilatère. (French) [On the passage of certain random walks planar on an equilateral hyperbola]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 25(4):443–456, ???? 1989. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1989__25_4_443_0.

Valko:2006:HLP

[Val06]  Benedek Valkó.  Hydrodynamic limit for perturbation

of a hyperbolic equilibrium point in two-component systems. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(1):61–80, January/February 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_1_61_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000567.

Valentim:2012:HLD

[Val12] Fábio Júlio Valentim. Hydrodynamic limit of a d-dimensional exclusion process with conductances. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1):188–211, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1327328019.

VanCutsem:1972:MCF

[Van72] Bernard Van Cutsem. Martingales de convexes fermés aléatoires en dimension finie. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 8(4):365–385, October–December 1972. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1972__8_4_365_0.

Varadhan:1995:SDT

[Var95a] S. R. S. Varadhan. Self diffusion of a tagged particle in equilibrium for asymmetric mean zero random walk with simple exclusion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(1):273–285, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1995__31_1_273_0.

Varopoulos:1995:HLT

[Var95b] N. Th. Varopoulos. Hardy–Littlewood theory on unimodular groups. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 31(4):669–688, ???? 1995. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1995__31_4_669_0.

Vargas:2006:LLT

[Var06] Vincent Vargas. A local limit theorem for directed polymers in random media: the continuous and the discrete case. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 42(5):521–534, September/October 2006. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2006_
_42_5_521_0;   http://www.sciencedirect.com/science/
article/pii/S0246020305000920.

Vayatis:2003:ERV

[Vay03] Nicolas Vayatis. Exact rates in Vapnik–Chervonenkis bounds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(1):95–119, January/February 2003. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://www.numdam.org/item?id=AIHPB_2003__39_1_95_0;
 http://www.sciencedirect.com/science/article/pii/S0246020302000109.

vandenBerg:2007:HFB

[vdB07] M. van den Berg. Heat flow, Brownian motion and Newtonian capacity. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(2):193–214, March/April 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_2_193_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000707.

vandenBerg:2010:VIT

[vdB10] M. van den Berg. On the volume of intersection of three independent Wiener sausages. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):313–337, May 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1273584126;   http://www.numdam.org/item?id=AIHPB_2010_
_46_2_313_0.

vandenBerg:2008:RSG

[vdBPSV08] J. van den Berg, Y. Peres, V. Sidoravicius, and M. E. Vares. Random spatial growth with paralyzing obstacles. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(6):1173–1187, December 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1227287570;   http://
www.numdam.org/item?id=AIHPB_2008__44_6_1173_0.

vandenBerg:2022:UBT

[vdBvE22] J. van den Berg and D. G. P. van Engelenburg. An upper bound on the two-arms exponent for critical percolation on Z^{d}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):1–6, February 2022. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-58/issue-1/An-upper-bound-on-the-two-arms-exponent-
for-critical/10.1214/21-AIHP1153.full.

vanderHofstad:2007:SPC

[vdHdHS07] Remco van der Hofstad, Frank den Hollander, and Gordon Slade. The survival probability for critical spreadout oriented percolation above 4 + 1 dimensions. II. Expansion. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(5):509–570, September/October 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_5_509_0;   http://www.sciencedirect.com/science/
article/pii/S0246020306000653.

vanderHofstad:2003:CCO

[vdHS03] Remco van der Hofstad and Gordon Slade. Convergence of critical oriented percolation to super-Brownian motion above 4 + 1 dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(3):413–485, May/June 2003. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2003_
_39_3_413_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000086.

vanderHofstad:2014:CSP

[vdHS14] Remco van der Hofstad and Artëm Sapozhnikov. Cycle structure of percolation on high-dimensional tori. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3): 999–1027, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1403277006.

Vechambre:2019:EFS

[Véc19] Grégoire Véchambre. Exponential functionals of spectrally one-sided Lévy processes conditioned to stay positive. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(2):620–660, May 2019. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https://

projecteuclid.org/euclid.aihp/1557820826.

vanEnter:2012:GNG

[vEEIK12] Aernout C. D. van Enter, Victor N. Ermolaev, Giulio Iacobelli, and Christof Külske. Gibbs–non-Gibbs properties for evolving Ising models on trees. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(3):774–791, August 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1340714872.

Verzelen:2010:HDG

[Ver10] Nicolas Verzelen. High-dimensional Gaussian model selection on a Gaussian design. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):480–524, May 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1273584132;   http://www.numdam.org/item?id=AIHPB_2010_
_46_2_480_0.

Vigon:2003:CDC

[Vig03] Vincent Vigon. Comparaison des deux composantes d’un subordinateur bivarié, puis étude de l’enveloppe supérieure d’un processus de Lévy. (French) [Comparison of the two components of a bivariate subordinator, then study of the upper envelope of a Lévy process]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 39(6):993–1011, November/December 2003. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://www.numdam.org/item?id=

AIHPB_2003__39_6_993_0;   http://www.sciencedirect.
com/science/article/pii/S0246020303000323.

Vigon:2011:HMB

[Vig11] Vincent Vigon. (homogeneous) Markovian bridges. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 47(3): 875–916, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1308834862;   http://www.numdam.org/
item?id=AIHPB_2011__47_3_875_0.

Viveros:2021:DPS

[Viv21] Roberto Viveros. Directed polymer in γ-stable random environments. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):1081–1102, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-57/issue-2/Directed-polymer-in-%ce%b3-stable-
random-environments/10.1214/20-AIHP1108.full.

Vollering:2016:TII

[Völ16] Florian Völlering. Talagrand’s inequality for interacting particle systems satisfying a log-Sobolev inequality. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(1): 173–195, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1452089266.

Viklund:2012:SLA

[VST12] Fredrik Johansson Viklund, Alan Sola, and Amanda Turner. Scaling limits of anisotropic Hastings–Levitov clusters. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(1): 235–257, February 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1327328021.

vonWeizsacker:1983:EOT

[vW83] Heinrich von Weizsäcker. Exchanging the order of taking suprema and countable intersections of σ-algebras. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 19(1):91– 100, ???? 1983. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1983__19_1_91_0.

Volny:2004:C

[VW04] Dalibor Volný and Benjamin Weiss. Coboundaries in L_{0}^{∞}. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(6):771–778, November/December 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_6_771_0;   http://www.sciencedirect.com/science/
article/pii/S0246020304000482.

Veber:2015:SLF

[VW15] A. Véber and A. Wakolbinger. The spatial Lambda–Fleming– Viot process: An event-based construction and a lookdown representation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):570–598, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1428672683.

Vysotsky:2014:PIR

[Vys14] Vladislav Vysotsky. Positivity of integrated random walks. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(1):195–213, February 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1388545271.

Wang:2007:ASP

[Wan07] Wensheng Wang. Almost-sure path properties of fractional Brownian sheet. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(5):619–631, September/October 2007. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_5_619_0;   http://www.sciencedirect.com/science/
article/pii/S0246020307000040.

Wassmer:2015:PTV

[Was15] Tobias Wassmer. Phase transition for the vacant set left by random walk on the giant component of a random graph. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(2):756–780, May 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1428672690.

Watanabe:1987:CSP

[Wat87] Shinzo Watanabe. Construction of semimartingales from pieces by the method of excursion point processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 23(S2): 297–320, ???? 1987. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1987__23_S2_297_0.

Weber:1980:AAP

[Web80] Michel Weber. Analyse asymptotique des processus gaussiens stationnaires. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(2):117–176, April–June 1980. CODEN AHPBAR. ISSN

0020-2347. URL http://www.numdam.org/item?id=AIHPB_
1980__16_2_117_0.

Weber:2010:STA

[Web10] Hendrik Weber. On the short time asymptotic of the stochastic Allen–Cahn equation. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):965–975, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1288878332;   http://www.numdam.org/
item?id=AIHPB_2010__46_4_965_0.

Weinryb:1984:HPP

[Wei84] Sophie Weinryb. Homogénéisation pour des processus associés à des frontières perméables. (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 20(4): 373–407, ???? 1984. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.numdam.

org/item?id=AIHPB_1984__20_4_373_0.

Weron:1980:SOS

[Wer80] Aleksander Weron. Second order stochastic processes and the dilation theory in Banach spaces. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 16(1):29–38, January–March 1980. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1980__16_1_29_0.

Werner:1993:STL

[Wer93] Wendelin Werner. Sur les singularités des temps locaux d’intersection du mouvement brownien plan. (French) [On the local-time singularities of the intersection of planar Brownian motion]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(3):391–418, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_3_391_0.

Wood:2016:UEF

[Woo16] Philip Matchett Wood. Universality of the ESD for a fixed matrix plus small random noise: A stability approach. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1877–1896, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1479373252.

Wu:1991:GDP

[Wu91] L. M. Wu. Grandes déviations pour les mesures de Gibbs lorsque la température tend vers zéro. (French) [Large deviations for Gibbs measures when the temperature tends to zero]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(3):273–289, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1991__27_3_273_0.

Wu:1999:FBM

[Wu99] Liming Wu. Forward-backward martingale decomposition and compactness results for additive functionals of stationary ergodic Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(2):121–141, March/April 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_2_121_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399800089.

Wu:2000:DIN

[Wu00] Liming Wu. A deviation inequality for non-reversible Markov processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36(4):435–445, July 2000. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2000_

_36_4_435_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300001357.

Wu:2004:ESG

[Wu04] Liming Wu. Estimate of spectral gap for continuous gas. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(4):387–409, July/August 2004. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2004_
_40_4_387_0;   http://www.sciencedirect.com/science/
article/pii/S0246020303000700.

Wu:2010:TIS

[Wu10] Liming Wu. Transportation inequalities for stochastic differential equations of pure jumps. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(2):465–479, May 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1273584131;   http://www.numdam.org/item?id=AIHPB_2010_
_46_2_465_0.

Wuthrich:1998:FRB

[Wüt98] Mario V. Wüthrich. Fluctuation results for Brownian motion in a Poissonian potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 34(3):279–308, May/June 1998. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1998_
_34_3_279_0;   http://www.sciencedirect.com/science/
article/pii/S0246020398800137.

Wuthrich:1999:GCB

[Wüt99] Mario V. Wüthrich. Geodesics and crossing Brownian motion in a soft Poissonian potential. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 35(4):509–529, July 1999. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_1999_
_35_4_509_0;   http://www.sciencedirect.com/science/
article/pii/S0246020399001041.

Wuthrich:2004:BEP

[Wüt04] Mario V. Wüthrich. Bivariate extension of the Pickands– Balkema–de Haan theorem. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 40(1):33–41, January/February 2004. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2004__40_1_33_0;   http://
www.sciencedirect.com/science/article/pii/S0246020303000475.

Wang:2016:FIC

[WW16] Feng-Yu Wang and Jian Wang. Functional inequalities for convolution probability measures. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(2):898–914, May 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1462367899.

Wang:2016:MAS

[WY16] Qinwen Wang and Jianfeng Yao. Moment approach for singular values distribution of a large auto-covariance matrix. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(4):1641–1666, November 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1479373243.

Williams:1990:RBM

[WZ90] R. J. Williams and W. A. Zheng. On reflecting Brownian motion — a weak convergence approach. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 26(3):461–488, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1990__26_3_461_0.

Xu:2009:STA

[XZ09] Tiange Xu and Tusheng Zhang. On the small time asymptotics of the two-dimensional stochastic Navier–Stokes equations. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 45(4):1002–1019, November 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1257529889;   http://
www.numdam.org/item?id=AIHPB_2009__45_4_1002_0.

Xie:2020:ESD

[XZ20] Longjie Xie and Xicheng Zhang. Ergodicity of stochastic differential equations with jumps and singular coefficients. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 56(1):175–229, February 2020. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-56/issue-1/Ergodicity-of-stochastic-differential-
equations-with-jumps-and-singular-coefficients/10.1214/
19-AIHP959.full.

Yano:2013:RFM

[Yan13] Yuko Yano. A remarkable σ-finite measure unifying supremum penalisations for a stable Lévy process. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1014–1032, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1380718735.

Yang:2018:MJD

[Yan18] Xiaochuan Yang. Multifractality of jump diffusion processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):2042–2074, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL https:/

/projecteuclid.org/euclid.aihp/1539849792.

Yang:2021:REP

[Yan21a] Fan Yang. Rare event process and entry times distribution for arbitrary null sets on compact manifolds. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(2):1103–1135, May 2021. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
journals/annales-de-linstitut-henri-poincare-probabilites-
et-statistiques/volume-57/issue-2/Rare-event-process-
and-entry-times-distribution-for-arbitrary-null/10.
1214/20-AIHP1109.full.

Yang:2021:CPP

[Yan21b] Shangjie Yang. Cutoff for polymer pinning dynamics in the repulsive phase. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(3):1306–1335, August 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-3/Cutoff-for-polymer-pinning-
dynamics-in-the-repulsive-phase/10.1214/20-AIHP1127.
full.

Yao:2018:MAI

[Yao18] Chang-Long Yao. Multi-arm incipient infinite clusters in 2D: Scaling limits and winding numbers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 54(4):1848–1876, November 2018. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL https://projecteuclid.org/
euclid.aihp/1539849786.

Yaroslavtsev:2019:MDG

[Yar19] Ivan S. Yaroslavtsev. On the martingale decompositions of Gundy, Meyer, and Yoeurp in infinite dimensions. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 55(4):1988–2018, November 2019. CODEN AH-PBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-55/issue-4/On-the-martingale-decompositions-
of-Gundy-Meyer-and-Yoeurp-in/10.1214/18-AIHP940.full.

Yilmaz:2010:ALD

[Yil10] Atilla Yilmaz. Averaged large deviations for random walk in a random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3):853–868, August 2010. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://projecteuclid.org/euclid.aihp/
1281100401;   http://www.numdam.org/item?id=AIHPB_2010_
_46_3_853_0.

Ying:1996:BRM

[Yin96] Jiangang Ying. Bivariate Revuz measures and the Feynman– Kac formula. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 32(2):251–287, ???? 1996. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1996__32_2_251_0.

Yor:1974:EUD

[Yor74a] M. Yor. Existence et unicité de diffusions à valeurs dans un espace de Hilbert. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(1):55–88, January–March 1974. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1974__10_1_55_0.

Yor:1974:ISV

[Yor74b] M. Yor. Sur les intégrales stochastiques à valeurs dans un espace de Banach. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 10(1):31–36, January–March 1974. CODEN AHP-

BAR. ISSN 0020-2347. URL http://www.numdam.org/item?
id=AIHPB_1974__10_1_31_0.

Yor:1975:EMP

[Yor75] Marc Yor. Étude de mesures de probabilité sur C(R_{+}^{∗}; R) quasi invariantes sous les translations de D(R_{+}^{∗}; R). (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 11(2):127–171, April– June 1975. CODEN AHPBAR. ISSN 0020-2347. URL http:/

/www.numdam.org/item?id=AIHPB_1975__11_2_127_0.

Yor:1977:ECP

[Yor77] Marc Yor. Étude de certains processus (stochastiquement) différentiables ou holomorphes. (French) []. Annales de l’Institut Henri Poincaré. Nouvelle série. Section B. Calcul des Probabilités et Statistique, 13(1):1–25, January–March 1977. CODEN AHPBAR. ISSN 0020-2347. URL http://www.numdam.org/

item?id=AIHPB_1977__13_1_1_0.

Yor:1985:PID

[Yor85] M. Yor. À propos de l’inverse du mouvement brownien dans R^{n}(n ≥ 3). (French) []. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(1):27–38, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL http://www.numdam.org/item?id=AIHPB_1985__21_1_
27_0.

Yor:1991:EDT

[Yor91] Marc Yor. Une explication du théorème de Ciesielski–Taylor. (French) [An explanation of the Ciesielski–Taylor theorem]. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 27(2):201–213, ???? 1991. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1991__27_2_201_0.

Yor:1993:ILO

[Yor93] Marc Yor. On an identity in law obtained by A. Földes and P. Révész. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 29(2):321–324, ???? 1993. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_1993__29_2_321_0.

Yoshida:2001:ELS

[Yos01] Nobuo Yoshida. The equivalence of the log-Sobolev inequality and a mixing condition for unbounded spin systems on the lattice. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 37(2):223–243, March 2001. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2001_
_37_2_223_0;   http://www.sciencedirect.com/science/
article/pii/S0246020300010669.

Younes:1988:EAG

[You88] Laurent Younes. Estimation and annealing for Gibbsian fields. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 24(2):269–294, ???? 1988. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_1988__24_2_269_0.

Yurinsky:2008:LBP

[Yur08] V. V. Yurinsky. A lower bound for the principal eigenvalue of the Stokes operator in a random domain. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 44(1): 1–18, February 2008. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1203969865;   http://www.numdam.org/
item?id=AIHPB_2008__44_1_1_0.

Yano:2010:PSL

[YYY10] Kouji Yano, Yuko Yano, and Marc Yor. Penalisation of a stable Lévy process involving its one-sided supremum. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(4):1042–1054, November 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:/

/projecteuclid.org/euclid.aihp/1288878337;   http://
www.numdam.org/item?id=AIHPB_2010__46_4_1042_0.

Zaitsev:2002:ERA

[Zai02] Andrei Yu. Zaitsev. Estimates of the rate of approximation in a de-Poissonization lemma. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 38(6):1071–1086, November/December 2002. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2002__38_6_1071_0;   http:/
/www.sciencedirect.com/science/article/pii/S0246020302011408.

Zhu:2017:MIS

[ZBH17] Jiahui Zhu, Zdzislaw Brzeźniak, and Erika Hausenblas. Maximal inequalities for stochastic convolutions driven by compensated Poisson random measures in Banach spaces. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 53 (2):937–956, May 2017. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1491897752.

Zeng:2015:KLI

[Zen15] Qiang Zeng. Kolmogorov’s law of the iterated logarithm for noncommutative martingales. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 51(3):1124–1130, August 2015. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1435759242.

Zeng:2016:HVR

[Zen16] Xiaolin Zeng. How vertex reinforced jump process arises naturally. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 52(3):1061–1075, August 2016. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL

http://projecteuclid.org/euclid.aihp/1469723511.

Zerner:2000:VLE

[Zer00] Martin P. W. Zerner. Velocity and Lyapounov exponents of some random walks in random environment. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 36 (6):737–748, November 2000. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2000__36_6_737_0;   http://
www.sciencedirect.com/science/article/pii/S0246020300001448.

Zerner:2005:SPR

[Zer05] Martin P. W. Zerner. On the speed of a planar random walk avoiding its past convex hull. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 41(5):887–900, September/October 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://www.

numdam.org/item?id=AIHPB_2005__41_5_887_0;   http://
www.sciencedirect.com/science/article/pii/S0246020304000895.

Zhan:2010:DCS

[Zha10] Dapeng Zhan. Duality of chordal SLE, II. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(3): 740–759, August 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://projecteuclid.

org/euclid.aihp/1281100397;   http://www.numdam.org/
item?id=AIHPB_2010__46_3_740_0.

Zhang:2013:SDE

[Zha13] Xicheng Zhang. Stochastic differential equations with Sobolev drifts and driven by α-stable processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(4):1057–1079, November 2013. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://projecteuclid.org/
euclid.aihp/1380718737.

Zhan:2022:TRM

[Zha22] Dapeng Zhan. Time-reversal of multiple-force-point. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 58(1):489–523, February 2022. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

URL https://projecteuclid.org/journals/annales-de-
linstitut-henri-poincare-probabilites-et-statistiques/
volume-58/issue-1/Time-reversal-of-multiple-force-point-
SLE%ce%ba%cf%81_-with-all-force/10.1214/21-AIHP1170.
full.

Zheng:1985:TRL

[Zhe85] W. A. Zheng. Tightness results for laws of diffusion processes application to stochastic mechanics. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 21(2):103–124, ???? 1985. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic). URL http://www.numdam.org/item?
id=AIHPB_1985__21_2_103_0.

Zheng:2012:CLT

[Zhe12] Shurong Zheng. Central limit theorems for linear spectral statistics of large dimensional F -matrices. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 48(2):444–476, May 2012. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-

7017 (electronic). URL http://projecteuclid.org/euclid.
aihp/1334148207.

Zhou:2007:SIB

[Zho07] Xiaowen Zhou. A superprocess involving both branching and coalescing. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 43(5):599–618, September/October 2007. CO-DEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (elec-

tronic). URL http://www.numdam.org/item?id=AIHPB_2007_
_43_5_599_0;   http://www.sciencedirect.com/science/
article/pii/S0246020307000039.

Zhu:2014:PLL

[Zhu14] Lingjiong Zhu. Process-level large deviations for nonlinear Hawkes point processes. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 50(3):845–871, August 2014. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://projecteuclid.org/euclid.aihp/
1403277000.

Zhu:2021:CBR

[Zhu21] Qingsan Zhu. On the critical branching random walk III: The critical dimension. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 57(1):73–93, February 2021. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL https://projecteuclid.org/journals/
annales-de-linstitut-henri-poincare-probabilites-et-
statistiques/volume-57/issue-1/On-the-critical-branching-
random-walk-III--The-critical/10.1214/20-AIHP1071.full.

Zygouras:2013:SDS

[Zyg13] N. Zygouras. Strong disorder in semidirected random polymers. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 49(3):753–780, August 2013. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1372772643.